A Complete Bibliography of Publications in Fish and Fisheries

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

23 April 2025 Version 1.21

Title word cross-reference

$126.28 [1241]. $158.3 [1255]. ^{13} [863]. _{M}EY [456]. δ [863]. F_{msy} [1140].

. [183].

0-85238-252-9 [64]. 024 [1255]. 024-2080-7 [1255].

10th [42]. 16th [1185, 231]. 16th-century [1185]. 19 [1163, 1016]. 1990s [29]. 19th [231, 540].

2003 [164]. 2005 [228]. 2019 [959]. 20th [22]. 21st [538]. 247. [606]. 271pp [520]. 28 [959]. 2nd [364, 238, 328].

3rd [337, 79, 355].

4 [869]. 40 [106]. 4th [246].

5 [1324, 191].

6 [1241].

7 [1255].

8th [272].

978 [1241, 1324, 869]. 978-0-226-02234-5 [520]. 978-0-226-02248-2 [520]. 978-0-470-67024-8 [606]. 978-1-119-82191-5 [1324]. 978-1-61091-853-4 [869]. 978-3-030-98621-6 [1241].

= [363].

abalone [304]. Abascal [1284]. Abate [1255, 187]. abiotic [195]. Abundance [55, 937, 495, 1246, 594, 394, 1189, 880, 892, 1011, 231, 1021, 22, 804]. academic [1171]. Accelerated [1101]. access [613, 1109, 896]. accommodate [857]. Accomplishments [157]. accounting [514, 1010]. achieve [948, 626, 597]. achievements [617]. Achieving [1074, 847, 927, 516]. Achilles [528]. acid [1158]. acid-base [1158]. Acidification [423, 448, 687, 643, 609, 458]. Acipenseridae [191]. Acoustic [783, 1278, 508, 1070, 574]. acoustic-based [508]. Acoustics [238]. acquisition [551]. Across [155, 983, 1258, 1122, 631, 1289, 782, 1290, 892, 777, 462, 1144, 1095, 804, 1171]. actinopterygian [782]. action [604, 1286]. actions [245, 667]. Active [51]. activities [78, 1059, 1172, 1242]. Activity [150, 1258, 394, 41, 1037, 1145, 1099, 768, 949, 178]. ADApT [652]. adaptation [1088, 1098, 383, 315]. Adaptations [711, 226, 205]. adapting [1337]. adaptive [534, 186, 35, 234, 784, 1234]. Adding [1330]. additional [1327]. address [938]. adjacent [817]. adjustments [1024]. adult [843]. adults [525]. advance [900, 563]. advanced [865]. Advances [407, 319, 1264, 365, 221, 532, 106, 356]. advancing [1314]. advantages [234]. Adverse [245, 581, 1302]. advice [1210, 1220, 957, 1320, 766, 955]. advising [268]. advisory [368]. Aegean [728]. Aerobic [75]. affect [322]. affects [897]. Africa [940, 427, 613, 16, 516, 19, 48, 1173]. African [265, 1059, 415, 509]. AFS [107]. after [1194]. against [969]. age [1067, 1090, 1287, 857, 755, 1202, 297, 1161, 1042, 428, 1176]. age- [1176]. age-based [1287]. age-structured [1067, 1090]. ageing [755]. agency [1189]. agency-assessed [1189]. agendas [966]. agent [952]. agent-based [952]. aggregating [1313]. aggregation [1284, 656, 1276]. aggregations [667]. ago [41, 728]. agreement [1015]. AIDS [169]. aim [948]. al [269]. alalunga [1306]. alarm [131]. Alaska [520, 283, 958, 664, 1228, 898, 1228]. Alaskan [871]. albacore [1022, 1306]. Albulidae [527]. alcoholism [931].

Alexander [1131]. algal [603]. Alien [750]. align [1105]. Allee [976, 1123]. Allen [290]. allocation [1222, 551]. allowable [517, 1248]. almost [842]. along [917]. alpinus [653]. already [1113]. also [756]. alterations [875]. Alternative [541, 1090, 616, 295, 1243, 554, 384]. alternatives [500]. alters [1090, 784]. amarus [255]. Amazon [823, 770, 834, 1292]. ambush [1203]. America [906]. American [224, 165, 191, 146, 123, 166, 43, 760, 872, 600, 1124, 11, 591, 155]. Americas [101]. Ammodytes [950]. among [169, 1333, 828, 613, 974, 1123, 1267]. amphibious [205]. amphidromy [152, 247]. anadromous [619, 298, 902, 274]. Anadromy [35]. anaesthesia [1230]. Anaesthetic [337]. analyses [963, 1306, 317, 598]. Analysis [1140, 352, 518, 872, 663, 1225, 1343, 564, 610, 993, 1051, 1196, 464, 640, 1070, 892, 1340, 1283, 810, 860, 897, 915, 1041, 994, 473, 769, 910, 113, 885, 1335, 789, 474, 843, 179, 1211, 786, 585, 567, 436, 956, 622, 625, 630, 724, 804, 533, 854]. analytic [1293, 1230, 771]. analytical [831, 975, 933, 428]. Ancestry [152]. anchovy [1170]. Andersen [1078]. anecdotes [1005, 197]. Angler [696, 1049, 479, 506]. anglers [375]. Angling [604]. Anguilla [385]. anguillid [1121, 1053]. Animal [310, 145]. Animals [337, 1270, 460]. Announcement [218]. Announcing [398]. Annual [1038, 863]. anomalous [1172]. Antarctic [650, 495, 453, 373, 438]. Antarctica [495]. antecedent [1005]. Anthony [1256]. Anthropocene [1275]. anthropogenic [820, 1076, 984]. antipredator [130]. any [1194]. anyway [248]. apex [860]. appendage [907]. application [831, 794, 1332, 1243, 1240]. Applications [156, 235, 296, 991, 49, 511, 221, 928]. Applying [1135, 967]. appraisal [824, 907]. Approach [357, 378, 58, 250, 1134, 721, 440, 409, 975, 1219, 267, 1274, 1063, 275, 269, 957, 1030, 206, 408, 1166, 514, 1248, 1180, 941, 933, 1102, 568, 60, 1174, 560, 413, 893, 832, 802, 428, 734, 984]. Approaches [571, 900, 187, 1264, 941, 582, 32, 642]. approval [1086]. apps [696]. Aquaculture [126, 1225, 125, 71, 851, 1071, 1279, 913, 690, 1030, 640, 587, 1105, 505, 854, 1024, 617, 1100, 521, 392, 532, 1158, 1241, 44, 1324]. aquaculture-dominant [1071]. aquaculture-related [690]. Aquarium [102, 637, 697]. Aquatic [108, 337, 73, 299, 172, 640, 1270, 223, 493, 1185, 883, 556, 247, 1226, 4, 422]. Arabian [1012, 1170, 732, 817]. archaeological [546]. Archaeology [838]. archetypical [499]. architecture [628, 968]. architectures [1203]. archival [745]. arctic [1120, 118, 609, 1260, 653]. Arctica [863]. Arctico [1039]. Area [1129, 1229, 1188, 669, 428, 804, 702]. Area-based [1129]. areal [840]. areas [98, 1251, 975, 446, 718, 1232, 844, 1301, 552, 428, 9]. arena [439]. arid [534]. Aristotle [728]. Arlinghaus [269]. Art [486]. artefact [19]. article [1298]. artificial [1138, 792, 1308, 222]. artisanal [841, 491, 852]. Asia [1300, 402, 222, 617, 876]. Asian [787]. aspect [751]. Aspects [190]. assemblage [1250]. assemblages [1103, 1026, 267, 708]. assembly [484]. assess [967, 843]. assessed [1189, 957]. Assessing [975, 446, 698, 547, 1186, 1091, 382, 984, 171, 1313, 1063, 1243, 1193, 714, 1280].

Assessment [328, 71, 254, 463, 1210, 131, 826, 938, 1223, 117, 1293, 911, 170, 600, 632, 988, 1010, 1063, 809, 1142, 412, 957, 757, 1259, 514, 324, 904, 32, 494, 467, 1137, 39, 49, 1179, 1147, 1111, 1152, 1054, 478, 647]. assessments [1210, 836, 593, 1319, 889, 705, 1190, 903, 243]. Assigning [37]. Assimilation [124]. assist [459]. assisted [532]. associated [1063, 989, 1125, 1156]. association [1085]. associations [1063]. assumptions [415, 999, 585]. asymmetric [889]. Asymmetry [1095]. at-sea [843]. at-vessel [769]. Atlantic [191, 71, 950, 1178, 461, 385, 111, 747, 863, 1316, 213, 1089, 720, 988, 982, 201, 1101, 68, 1136, 1224, 976, 1260, 323, 410, 935, 1334, 840, 990, 93, 281, 1174, 848, 82, 812, 301, 168, 843, 625, 693, 630, 1128, 862]. Atlantis [411]. Atlas [824, 673]. attached [263]. attempt [941]. Attending [915]. attention [178]. attraction [1339]. Attributes [1132, 1198, 859]. Auchenipteridae [1007]. audiences [649]. Auditory [216]. augments [797]. auratus [602]. Austin [1241]. Australasia [343]. Australia [566, 920, 985, 634, 1076]. Australian [649, 602]. Author [148, 184, 217]. authorities [703]. automatic [1145]. Autopsy [189]. available [1142, 1055, 1282]. avoid [605]. avoidance [1311]. away [3, 114].

B [456]. background [1103]. Bacteriology [192]. bad [66, 497, 286, 325, 199]. Bahamas [1271]. Bailey [520]. Bait [686, 670]. Baja [745]. balance [832, 9]. Balanced [607, 819, 675, 655, 1204]. Balancing [964]. baleen [373, 749]. Balkans [750]. Baltic [673, 1028, 682, 967]. ban [581, 664, 557]. Bandt [259]. bans [664]. barefoot [139]. Barents [1119, 1174]. barrel [714]. Barrier [566, 327]. Barriers [294, 589]. base [861, 457, 463, 1158]. Based [424, 407, 673, 437, 398, 250, 900, 468, 177, 934, 952, 818, 1231, 795, 631, 729, 968, 863, 243, 1220, 470, 916, 1029, 354, 529, 834, 1287, 562, 149, 1045, 508, 1306, 1129, 939, 1140, 269, 300, 762, 718, 643, 368, 929, 410, 800, 677, 904, 1245, 258, 1165, 764, 1114, 415, 369, 850, 1048, 689, 1055, 1069, 766, 822, 827, 307, 1256]. baseline [350]. Baselines [944, 418]. Basin [1166, 1292, 1009]. Basin-scale [1166]. basis [133, 86, 68, 96]. bass [1095]. basses [1124]. batoid [191]. Bayesian [117, 1334, 967]. be [47, 358, 949, 325]. beaches [746]. beacons [1330]. beautiful [312]. beavers [447]. became [41]. because [427]. become [637]. becomes [517]. been [1023, 273]. Beginning [327]. begone [97]. Behavior [868, 225, 145, 1255, 299]. Behaviour [377, 102, 90, 406, 1034, 1134, 595, 482, 570, 1313, 517, 479, 399, 353, 1019, 464, 506, 951, 27, 168, 873]. Behavioural [1214, 973, 89, 1251, 446, 1199, 162, 885, 449]. behaviours [500, 1272, 1172]. Behnke [558]. being [931, 841, 1056, 734]. Beknown [985]. belief [967]. benchmarks [773]. beneficial [1182, 1173]. benefit [331, 179]. Benefits [1284, 1276, 417, 667, 806, 632, 17, 613, 1303, 1331, 695, 1173, 964, 522, 481]. benign [340]. benthic [789, 979, 224]. Benthos [251]. Beolens [1296]. Bering [835, 77, 730]. Bermuda [1271]. Bertalanffy [295]. BESA [1259]. Best [1125, 629, 626, 1259, 936, 608, 1055]. Best-practice [1125]. better [938, 1177, 388, 198, 295, 454, 1094, 316, 1054]. Between [174, 1246, 934, 713, 720, 640, 480, 94, 315, 1187, 1042, 349, 873, 138]. Beverton [356, 739]. Beyond [466, 890, 1026, 738, 525, 856, 580]. bias [1210, 1029]. biases [953]. Biennial [42]. big [667]. Bigelow [140]. bigger [1032]. billfish [659]. billfishes [331]. billion [478, 520]. Billion-Dollar [520]. binding [517]. bio [1259]. bio-economic [1259]. biochemical [1024]. Biodiversity [1027, 872, 339, 995, 1015, 1204, 294, 326]. Biogeographic [887]. Biogeographical [708]. biogeography [158, 523, 48]. Biological [144, 1279, 44, 82, 601, 1119, 953, 88, 659, 349, 1218, 333]. Biologist [253, 302]. Biology [108, 142, 435, 73, 328, 829, 253, 376, 387, 346, 223, 154, 578, 13, 76, 340, 1313, 314, 271, 106, 142, 355]. Biomass [1102, 522, 1003, 624, 773, 1344, 1048, 1111, 965]. biome [813]. Biomechanics [252]. biophysical [1224]. biota [1005, 936]. biotelemetry [285]. biotic [809]. biphenyl [430]. Birds [174, 1009, 775]. bird’s-eye [1009]. bitterling [255]. black [671, 520, 1326, 1281, 1139]. Blackwell [606]. Blennies [376]. blinded [1319]. blooms [603]. Blue [1081, 966, 1139, 747, 1071, 1129, 866, 1083]. bluefin [1281, 499, 213, 335, 990]. Bo [1296]. boat [23]. Boats [752]. Bob [414]. Body [782, 159, 892, 1118]. bonefishes [527]. Bony [191, 34]. Book [1241, 25, 30, 36, 105, 1297, 1256, 1324, 869, 520, 20, 606, 1228, 1240, 1255, 141, 1184, 64, 1296]. Books [26, 47, 167, 194, 64]. bootstrap [494]. borders [1095]. boreal [298, 1011]. both [320]. bottom [748, 699, 869, 626, 584, 789, 947, 540, 1173, 979]. bottom-up [584]. boundary [404, 1187]. boundless [1131]. box [1326, 64]. Breathing [460]. Breeders [83]. breeding [751]. Brian [1241]. Bridging [315]. Britain [519, 200]. British [1005, 540]. Broadbill [28]. broadcast [563]. broader [635]. Brown [461, 902]. BSc [228]. bubble [994]. Bucke [64]. budget [1273, 984]. buffers [1102]. build [678]. building [265]. Bump [107, 123]. burbot [360]. burdens [1323]. Burgess [64]. Buried [1008]. business [1135, 1081]. Buybacks [338, 390]. By-catch [97, 1082, 351, 419, 405, 564, 187, 219, 361, 632, 198, 215, 933, 778, 1060, 307]. Bycatch [985, 249, 772, 512, 529, 582, 1263, 1133, 1272, 1328]. bycaught [843].

C [863]. California [745, 197, 445, 504, 231]. call [666, 1216]. camera [928]. cameras [1278, 574]. Can [454, 1114, 515, 331, 1111, 358, 667, 198, 426, 1194, 393, 637, 709, 1197, 237, 10, 6, 675, 949, 325, 316, 516]. Canada [1298, 1249]. Canadian [72, 14, 1190, 812, 15]. Capacity [1191, 691, 1143, 374, 1219]. capacity-limited [1219]. capitalism [1139]. captive [592]. Capture [222, 851, 828, 564, 1341, 472, 609, 725, 693]. capturing [786]. Carangids [579]. Carbon [124]. card [762]. cardiorespiratory [1024]. care [992, 1025]. Caretta [843]. Caribbean [923, 995, 1271, 391, 909, 586]. Carlsberg [740]. Carp [13, 64, 617]. cartilaginous [34]. Case [1231, 1298, 283, 655, 972, 861, 995, 374, 600, 631, 531, 1249, 304, 549, 985, 580, 879, 459, 1028, 93, 516, 465, 197, 19, 509, 381, 730, 350, 1282]. case-studies [631]. case-study [465, 730]. cases [1135]. Caspian [21]. catalyse [694]. Catch [1339, 769, 948, 1222, 923, 351, 419, 1343, 285, 405, 517, 564, 906, 692, 466, 187, 219, 361, 632, 198, 215, 97, 492, 1259, 503, 933, 454, 778, 1137, 1060, 681, 1082, 602, 572, 822, 536, 1176, 307, 707, 165]. catch-and-release [285]. catch-at-length [1176]. catch-only [1137, 822]. catch-related [948]. catchability [714]. Catches [824, 663, 1192, 1248, 96, 1048, 1178, 1004, 199]. Catching [940, 259, 967, 590]. categorizing [267]. catfish [340]. catfishes [1247]. caught [849, 1340]. cautions [161]. CCAMLR [409, 668]. CD [46]. CD-ROM [46]. cell [1045]. cell-based [1045]. Celtic [673]. Cenozoic [16]. census [1116]. Central [1135, 191, 1263, 1178, 1229, 255]. centurial [576]. centuries [1281]. Century [1347, 1208, 1185, 720, 538, 168, 231, 540, 602, 1250, 22]. Century-scale [1347]. cephalopods [393]. certainly [325]. certification [1304, 1086, 1040, 436, 294]. Cetacean [104]. chain [375, 1340]. chains [876]. Challenge [1015, 250, 1313, 874]. Challenges [407, 157, 329, 423, 1195, 794, 928, 640, 1224, 1264, 1180, 617, 742, 113, 557, 1107, 221, 567, 696, 1036, 1191]. challenging [474]. Cham [1241]. Change [432, 334, 871, 1088, 652, 339, 344, 1139, 998, 805, 1182, 1098, 857, 1032, 989, 583, 1136, 1224, 643, 609, 1318, 905, 705, 1347, 305, 1109, 278, 429, 96, 784, 949, 660, 719, 1014, 986, 751, 875, 895, 1265, 1076, 954]. Changes [1022, 644, 1192, 118, 887, 790, 97, 371, 1125, 1028, 777, 1190, 967, 646, 1048, 714, 341]. Changing [941, 892, 166, 423, 690, 583, 427, 715, 1275, 1218]. Characterising [1345]. Characteristics [732]. Characterizing [230, 1257]. Charles [85, 228]. Charleston [123, 107]. charr [881, 1346, 653]. check [726]. checklist [1055]. Chelonia [745]. chemical [131]. chemically [474]. chemicals [621]. chemistry [771]. Chicago [520]. Chichester [606]. chimaeras [191, 817, 1193]. China [887, 523, 538, 942, 1211, 918]. Chinese [151]. Chinook [1253, 1246, 1235, 777, 1272, 1013]. Chionoecetes [1119]. choice [700]. choices [705]. choke [581]. chondrichthyan [1178, 1076]. Chondrichthyes [1282]. Choosing [936]. chronicle [958]. Chuck [228]. Cichla [1124]. Cichlid [158, 1255, 195]. cichlids [186, 90]. Ciliary [496]. circle [556, 769, 331]. Circumpolar [1039]. circumtropical [627]. cited [648]. CITES [787, 916, 543]. Claiming [683]. clam [622]. classical [948, 731]. classification [1164]. Classifying [595, 369]. Cleaner [704]. cleaning [704]. client [383]. Climate [1098, 1012, 201, 1002, 305, 784, 660, 1265, 334, 1214, 1074, 1088, 861, 339, 678, 1113, 1288, 1221, 1010, 982, 1170, 690, 989, 1336, 583, 1136, 1224, 643, 609, 917, 1318, 1334, 1309, 1132, 705, 278, 1335, 719, 1035, 1234, 1069, 875, 895, 1076, 166]. climate-adaptive [1234]. Climate-induced [1002, 1010, 982, 1334]. climate-resilient [1074]. climatic [1190, 22]. Close [1116]. Close-kin [1116]. closed [1319, 718]. closer [1194]. Closing [1251]. closures [1294]. cloth [520]. Clupea [863, 967, 22, 875]. Clupeidae [1205, 967, 22, 875]. Co [600, 157, 63, 512, 812, 1144]. Co-management [600, 157, 63, 812, 1144]. co-occurrence [512]. coast [1170, 1013, 1208, 1249, 1013]. coast-wide [1013]. Coastal [536, 1034, 1345, 1143, 1271, 361, 920, 1257, 568, 717, 719, 1173]. Cod [1039, 352, 118, 958, 1101, 68, 1224, 976, 1275, 323, 840, 1028, 1174, 8, 168, 18, 485]. codfish [958]. coexistence [607, 1004]. Cognition [299, 133]. cognitive [1029]. coherence [848]. cold [1214, 1346]. cold-water [1214, 1346]. collaboration [886]. collaborative [1274, 925]. Collapse [1163, 332, 303, 196, 956]. Collateral [969, 236]. collection [785, 222]. collective [619]. colonization [782]. Colonizing [775]. colour [371, 64]. Columbian [906]. combinations [933]. combining [816, 1048]. come [670]. comment [158]. commentary [1337]. commercial [845, 828, 257, 457, 1077, 369, 896, 1265]. commercially [463]. Common [831, 1110, 862, 585, 374, 889, 74, 401]. Commons [101, 400, 857]. Communication [487, 943, 783, 864]. communities [722, 1002, 613, 462, 248, 63, 924, 726, 586, 801]. Community [773, 1345, 784, 586]. Comparative [201, 721, 634, 474, 1295]. Comparing [995]. comparison [631, 263, 774]. comparisons [404, 284, 459]. Compensatory [120, 59]. competition [1318, 241]. Complex [545, 912]. complexes [541]. complexity [1142, 549, 511, 1320, 786, 903]. compliance [877, 561, 1251, 703, 318, 991]. complicated [170]. components [505, 1144]. composition [518]. Comprehensive [435, 831, 802, 1055]. comprehensively [565]. Computer [1164]. Computers [364]. concentrations [621, 430]. concept [929, 85, 710, 647]. concepts [86, 88, 84, 92]. conceptual [856, 725]. concerns [1284, 1276]. concise [270]. concluding [92]. condition [495, 1074, 1244, 159, 606]. conditions [708, 771, 1190, 1218]. Conference [98, 42]. Confidentiality [477]. conflict [1342, 798, 1058, 1004]. conflicts [632]. Confronting [424, 298]. Congress [232]. Connections [740, 278]. connectivity [525, 1091, 1141]. consciousness [176]. consensus [1051]. consequence [1032]. consequences [581, 483, 1233, 243, 611, 621, 989, 549, 992, 514, 507, 986, 874, 1234, 341]. Conservation [711, 451, 445, 288, 1072, 281, 465, 387, 56, 527, 629, 934, 76, 418, 1108, 604, 649, 667, 869, 1337, 563, 464, 556, 637, 1199, 320, 758, 60, 511, 552, 360, 950, 1331, 543, 294, 686, 760, 522, 242, 271, 271]. conservationist [959]. Considerations [1084, 662, 1083, 161]. Considering [243]. Consilience [308]. consistency [1154]. Consistent [1283, 1210, 631]. conspecifics [134]. constitutes [921]. constraints [1120, 1119, 248]. constructing [945]. consume [970]. Consumer [693, 635]. Consumption [1261, 588, 1016]. Contemporary [1333, 1180, 695]. content [518]. Contents [185]. context [1322, 727, 1231, 757, 431]. contexts [679, 1209]. Continental [1070, 388]. Continental-scale [1070]. continuous [676]. continuum [1310]. Contrasting [1113, 912, 976]. contribute [89, 718]. Contribution [475, 393, 513, 1156]. Contributions [440, 147]. Control [827, 1268, 631, 1221, 421, 1248, 774]. controls [374]. controversy [86, 59, 540]. conventional [298, 474]. cool [1113]. cooperatives [877]. coordinated [1151]. coordination [1009]. Coping [847]. coral [306, 544, 547, 332, 491, 773, 799, 489, 605, 305, 645, 890, 909, 271]. coral-feeding [306]. Cormorant [1041]. corporate [796]. Correction [1284, 1298]. corresponds [875]. corridors [1203]. Corrigendum [1253, 227, 712, 1079, 1148, 1252]. cortisol [815]. cosmopolitan [262, 1106]. cost [179]. costly [854]. costs [264, 446, 1182, 1303, 481]. Could [671, 819]. Council [1304]. counting [580]. Countries [462, 266, 95, 1071, 634]. coupled [508]. Coupling [1273]. Covelo [869]. cover [770, 64]. coverage [1040]. COVID [1163, 1016]. COVID-19 [1163, 1016]. crab [1119, 1340]. cracks [638]. crayfish [29, 1185]. CRC [142]. create [1108]. created [684]. creating [796]. credit [572]. Cren [444]. crimes [1083, 1081]. crisis [326, 395]. critical [1225, 824, 621, 907, 844, 555, 999, 724]. Critically [861]. critique [929]. croaker [303]. cross [632, 774]. cross-ecosystem [774]. cross-taxa [632]. crustacean [1293, 1142, 625]. Crustaceans [898, 191]. cryptic [264]. cucumber [420, 566, 403, 473]. Cues [672, 131]. culprit [998]. Culture [288, 13, 1162, 128, 222, 1267]. cultured [993]. culverts [951]. Cumulative [504]. Curious [869]. Current [1129, 146, 996, 415, 978, 1107, 176, 525, 491, 1264, 844, 284, 296, 588, 950, 221, 445, 504]. curves [441]. Cycle [647, 809]. cycles [177, 178]. Cyprinid [64].

D [1284, 183, 64]. daily [221]. damage [969]. dams [856]. danger [131]. Daniel [1150]. Darwin [85]. Data [309, 879, 254, 865, 1223, 1177, 821, 1057, 836, 1219, 791, 745, 1220, 335, 610, 547, 457, 1066, 963, 1186, 477, 1142, 1207, 492, 1259, 1161, 1155, 276, 785, 546, 1028, 1165, 1068, 897, 1094, 369, 842, 1147, 846, 1111, 932, 1093, 696, 822, 316, 804, 743]. data- [1219]. data-integrated [932]. Data-Limited [254, 1223, 821, 836, 1220, 547, 963, 822]. data-poor [745, 1066, 842]. data-rich [1220, 963]. Database [463]. databases [995]. David [444, 1255]. DC [869]. dead [317]. deadline [372]. deal [153]. Decadal [495, 177]. decades [935, 1052]. decapod [1293]. December [959]. decision [628, 652, 916, 1165]. decision-making [628, 916, 1165]. decisions [66, 675]. Declaration [927]. decline [1246, 111, 427, 546, 1013]. declined [614]. declines [540]. declining [199]. deep [997, 1027, 220, 1052, 1291, 1164, 999]. deep-sea [997]. deep-water [1052, 1291]. Defenses [443]. defined [710, 846]. Defining [735, 459]. definition [91]. delay [1243]. delay-differential [1243]. deliberate [245]. delivering [785, 413]. delivery [534]. delta [494]. demand [244]. demands [524]. demersal [490, 925]. Demographic [777, 909, 341]. demography [682]. Demystifying [391]. densities [974]. Density [953, 161, 894, 833, 701, 1233, 880, 840, 1154, 59, 888, 932, 622]. Density- [894]. Density-dependent [953, 161, 1233, 880, 888]. dependence [701, 1154, 59, 833, 932]. dependency [1029]. dependent [1322, 1233, 1207, 880, 739, 953, 161, 894, 888]. depends [826, 1321, 1094]. Depensation [33]. depleted [616, 1123]. depletion [809, 707]. deposition [234]. Depredation [1342, 1004]. depression [120]. depth [426, 1272]. depth-use [1272]. derived [603]. dermal [38]. Describing [679]. design [669, 597, 641, 49, 369, 978, 1328, 788]. Designation [668]. designing [785]. desires [1185]. despeciation [91]. despite [887]. dessert [919]. destined [716]. Destructive [1171]. Detecting [1172, 1291]. detection [530]. determinants [618, 725]. determination [611]. determine [1236, 1322, 426]. determined [584]. determines [550]. determining [297]. develop [1320, 734]. Developing [117, 266, 629, 636, 1119, 28, 856]. Development [493, 1139, 265, 1037, 792, 576, 929, 1160, 1003, 938]. Developmental [897, 541]. Developments [211, 438]. deviants [1315]. devices [1284, 1313, 564, 933, 1276, 1328]. diadromous [234]. Diadromy [52, 247]. Diagnosis [266, 593]. diagnostic [90]. Dictionary [1296, 46]. did [41, 942]. DIDSON [574]. diel [394, 1272]. diet [518, 698]. dietary [1111]. diets [963, 1306]. differences [1333]. different [1224]. differential [1243]. differentiated [1110]. differentiation [611, 195]. difficulties [483]. Digital [928, 1161]. dilemmas [455, 509]. diluted [651]. dimension [426]. dimensional [1181]. dimensions [952, 670]. diploid [284]. direct [724]. directions [1339, 408, 490, 1332, 996, 359, 113, 1004]. discard [581, 577, 557]. discarded [828, 215]. discards [911, 529, 198, 985, 582, 924, 199, 743, 307]. disciplinary [479]. Disciplines [155]. discomfort [368]. Discount [384]. discoveries [38]. discrete [700]. discrete-choice [700]. Disease [146, 1302]. diseases [1283, 64]. Disentangling [888]. dismantle [1267]. disparate [912]. disparity [907, 613]. dispersal [818, 775, 537]. displacement [343]. disproportionate [1323]. dispute [10]. disruption [790]. Dissecting [1144]. distant [538, 1064]. distant-water [538, 1064]. distinct [1333, 1224, 247]. Distribution [747, 200, 262, 1124, 1089, 982, 293, 1290, 431, 840, 530, 937, 1301, 1212, 949, 1014, 161, 730, 835, 1265]. distributions [938, 627, 1336, 1207, 1326]. divergence [52]. Diverse [1288, 962, 1306, 1196, 1283, 1006, 653]. diversification [1208, 871, 1160]. Diversity [306, 657, 1209, 1134, 1345, 728, 1292, 523, 52, 654, 704, 973]. divisions [1333]. DNA [1057, 457, 37, 794, 854, 318]. DNA-analysis [854]. Do [1143, 564, 1105, 396, 270, 322, 544, 1010, 1262, 198, 842, 896, 316]. documented [571]. Does [517, 243, 681, 1267, 701, 841, 670, 771, 1118, 942]. dollar [478, 520]. dolphins [1060]. domestic [931]. domesticated [998, 988]. Domestication [145, 521]. dominant [1071]. dominated [1346, 1340]. Dordrecht [1255]. down [584, 220]. Downstream [1000, 589]. Dr. [348]. drawing [668]. dredging [724]. Dried [1162, 151]. drifting [1284, 497, 1313, 1276]. drive [708, 1011, 979]. driven [1272, 714]. Drivers [1304, 622, 1122, 975, 779, 1346, 643, 1002, 1205, 803, 473, 1280, 586, 449]. drives [1049]. Drivin [1319]. drowned [651]. drug [981]. dry [775]. duration [385, 646]. during [75, 1336, 1266]. Dutch [925]. Dynamic [1289, 1331, 1273, 9, 1054]. dynamical [1048]. Dynamics [53, 867, 1217, 1236, 244, 594, 1103, 446, 213, 700, 545, 508, 1119, 1243, 1205, 1334, 362, 196, 810, 162, 161, 980, 1141, 585, 731, 1123, 1242, 862, 449].

e-book [520, 606]. E. [444]. Early [147, 1016, 283, 1089, 1336, 1213, 1275, 1266, 5, 894, 540, 1250, 16]. East [1136, 402, 516, 118, 717, 693, 862, 745, 1089]. eastern [77, 404, 1012, 1170, 613, 1192, 1028, 835]. eat [393, 391]. Eating [386]. EBFM [1044, 1218]. ebk [1241, 1255]. Eco [181, 1091, 1040]. eco-certification [1040]. eco-evolutionary [1091]. Eco-labelling [181]. Ecoepidemiology [5]. ecolabelling [436]. Ecological [321, 77, 309, 190, 645, 1076, 1126, 685, 1203, 744, 1322, 476, 687, 1008, 1022, 1289, 498, 776, 976, 1180, 480, 1155, 742, 915, 1303, 589, 1190, 429, 858, 980, 802, 1282, 541, 853]. Ecologically [893]. Ecologist [444, 139]. Ecology [109, 121, 143, 461, 711, 868, 73, 451, 422, 452, 239, 288, 81, 225, 807, 1202, 182, 183, 300, 591, 864, 253, 357, 1240, 1255, 210, 58, 301, 639, 56, 1078, 54, 753, 1281, 270, 1327, 213, 563, 160, 664, 904, 1199, 362, 746, 1039, 313, 846, 977, 862]. Economic [72, 190, 776, 662, 1126, 871, 744, 3, 1023, 687, 841, 1294, 887, 1259, 615, 480, 742, 641, 659, 1095, 395, 1173, 734, 1064, 485, 522]. Economics [288, 98, 42, 365]. economies [334]. economy [1162, 1139, 640]. Ecopath [1231]. ecophysiologist [348]. ecophysiology [1024]. Ecosim [1231]. Ecosystem [424, 596, 795, 661, 407, 824, 470, 834, 173, 1217, 929, 410, 378, 10, 814, 599, 749, 352, 468, 636, 1074, 688, 679, 440, 409, 1231, 998, 729, 968, 1029, 593, 354, 466, 529, 149, 1045, 508, 1019, 498, 275, 939, 412, 674, 206, 408, 762, 718, 453, 514, 323, 941, 677, 1002, 32, 258, 1347, 391, 605, 774, 413, 1320, 816, 179, 903, 875, 766, 1128, 1267, 827, 954, 307, 193, 950, 1256]. Ecosystem-Based [424, 407, 795, 470, 834, 929, 410, 468, 1231, 729, 968, 1029, 354, 529, 149, 939, 762, 718, 677, 258, 766, 827, 307, 1256]. ecosystem-informed [1320]. Ecosystems [124, 830, 432, 201, 202, 279, 322, 404, 493, 651, 497, 342, 649, 411, 920, 556, 882, 330, 774, 513, 6, 96, 1021, 1226, 4, 1054, 584]. Ecoviability [729]. ectotherms [644]. ed [142, 223]. eddies [499, 1242]. Eden [290]. Edition [246, 175, 140, 251, 259, 211, 868, 79, 141]. Editor [2, 31, 65, 203]. Editorial [233, 1097, 311]. editors [183]. edn [364, 337, 238, 328, 355]. eDNA [1047]. Eds [1324, 216]. Edward [228]. eel [427, 804, 175]. eels [805, 1121, 611, 1053]. Effect [1068, 877, 880, 976, 1248, 430, 1042, 697]. Effective [878, 968, 738, 785, 837, 921, 1116]. Effectiveness [359, 1075, 759, 859, 702]. Effects [1138, 224, 264, 1140, 1213, 659, 428, 77, 1308, 699, 603, 805, 1313, 661, 1113, 1196, 268, 447, 811, 480, 1205, 905, 882, 1072, 592, 278, 1041, 5, 888, 1123, 138, 724, 1016, 1294]. efficacy [1193]. Efficiency [674, 3, 1023, 575, 1051, 880, 467, 572]. efficient [1317]. Effort [689, 845, 691, 374, 928, 477, 757, 296]. Egg [1069, 983, 551, 234, 221, 888, 833]. Egg-production-based [1069]. eggs [813, 775]. Elasmobranch [732, 618]. elasmobranchs [632, 548]. electric [737]. electrofishing [394]. electromyogram [150]. Electronic [926].

Electroreception [216]. Electrotrawling [554]. elements [650, 597]. elephant [427]. eleven [555]. eliminating [1036]. Elopidae [527]. Elucidating [1157]. elusive [375]. embedded [1203]. emblematic [760]. embryos [505]. emerge [517, 675]. Emerging [407, 1161, 98]. eminent [959]. emphasis [95]. Empirical [531, 1210]. employment [1330, 475, 392]. Empowering [791]. Empowerment [455]. Empty [1060]. encourage [10]. End [327]. endangered [604]. endemic [465]. Ending [590]. endocrine [34, 505]. energetics [150, 528]. Energy [551, 1278, 1273, 984]. energy-budget [984]. enforcement [318]. Enforcing [74]. engineering [763]. England [531]. English [175]. enhance [663]. Enhancement [211, 304, 776, 1273, 18]. enhancements [76]. Enhancing [285, 649]. enough [1056]. enrichment [592]. ensemble [1301]. ensure [841, 1056]. entanglement [1005]. entire [840]. entrainment [856]. environment [416, 1026, 1244, 640, 230, 27, 530, 862]. Environmental [144, 1119, 1346, 592, 902, 82, 1107, 436, 534, 871, 340, 1057, 593, 118, 794, 708, 771, 643, 1243, 429, 1226, 586, 44]. Environmentalist [103]. environmentally [788]. Environments [172, 917, 592]. Epigenetics [1324]. epigenomic [505]. Eponym [1296]. Equation [280, 295]. equitable [738]. equity [3, 1023, 1314]. equivalence [1220]. era [913]. erode [695]. erosion [1026]. Errata [212]. Erratum [24, 40, 115, 240, 397, 471, 488, 1168]. Escalle [1284]. escaped [480]. escapees [988]. essential [1188, 1056]. established [28]. estimable [369]. estimate [1248, 799, 1200, 1116]. estimated [1028]. Estimates [910, 706, 1210, 880, 1099, 961, 1094]. Estimating [55, 236, 687, 692, 837, 39, 1263, 96, 1093, 1282, 707, 594, 512, 503, 1127]. estimation [483, 94, 1003]. estimators [542]. Estuaries [99, 711, 1236, 560, 1237]. estuarine [267, 708, 917]. estuary [906, 1063]. estuary-associated [1063]. Eternal [867]. Ethical [1083]. Ethics [72, 1225]. ethnographic [745, 573]. ethnohistoric [546]. EU [831]. eulachon [445]. Eumetopias [546]. EUR [606]. EUR120.99 [606]. Europe [95, 577, 861, 372, 757, 408, 427, 582, 510, 1265, 255]. European [29, 581, 340, 394, 374, 1087, 421, 593, 1037, 427, 615, 634, 459, 1060, 557, 1014, 1107, 255, 804]. evaluate [1319]. Evaluating [831, 1117, 1195, 552, 802, 1127, 1223, 726, 565, 586]. evaluation [881, 744, 1223, 176, 771, 844, 17, 1137, 555, 214, 608, 999, 1084, 925, 704, 734, 702, 742]. evaluations [368]. even [1032]. events [820, 1213]. Evidence [1188, 755, 848, 1021, 262, 671, 1333, 1005, 33, 1114, 231, 540, 702]. Evolution [452, 182, 1255, 942, 1078, 468, 577, 119, 782, 943, 819, 1240, 195, 660, 751, 695, 183]. Evolutionary [484, 514, 253, 721, 782, 976, 1091, 415, 555, 541]. evolving [470, 925]. exacerbate [1272]. Examining [787, 927, 463, 585, 861]. example [343, 1220, 368, 705, 986, 485]. Examples [876, 912]. exaptations [723]. exceeds [485]. exclusive [1064]. Excretion [69]. exercise [1194]. existing [408]. Expanding [647]. expansion [343, 560]. expansion-and-displacement [343]. expenditure [757]. Experience [157, 412, 411, 694]. experiences [668, 1175, 582, 634, 926]. experiencing [1309]. experiment [1319, 835]. experimental [789]. experiments [369]. experts [1147]. explain [923]. explained [949]. explaining [177]. explanation [891]. explicit [1117, 214, 1179]. Exploitation [1281, 893, 1078, 54, 420, 566, 607, 223]. Exploited [606, 245, 342, 1287, 297, 680, 463, 909, 1280, 622]. exploration [441, 554]. exploratory [967]. explore [858]. Exploring [728, 400, 1174, 1167, 1327]. exposed [621]. exposure [50]. extensions [739]. Extinction [112, 1178, 171, 1027, 817, 114, 1282, 478]. extinctions [256]. extraction [321]. extractions [1101]. extractivism [1173]. Extreme [620, 820]. extremely [842]. eye [1009]. Eyed [1020]. eyes [1319, 231].

F [1284, 1324, 1184]. face [1224]. facilities [359]. Factors [66, 82, 1236, 1289, 809, 459, 659, 384, 956]. Facts [918, 312]. FAD [1294]. fading [638]. FADs [497]. failed [857, 303]. failing [1060]. Fair [635]. Falling [638]. familiarity [138]. FAO [191]. Farmed [57, 720, 480, 617]. Farming [45, 29, 1279]. Farrell [183]. fast [1310, 1266]. fast-slow [1310]. faster [1262]. fauna [563, 1006, 568, 231, 789, 1076]. faunas [152]. Fay [216]. feasible [837]. features [1283, 333]. February [959, 414]. fecund [114]. Fecundity [1235, 779, 159, 1253]. federal [1221]. feed [709, 1004, 918]. feedback [245]. feedbacks [1308, 979]. feeding [306, 528]. Feel [396, 515]. feelings [250, 269]. feelings-based [250, 269]. feline [340]. female [188]. fertilization [373]. few [1188]. fewer [924]. Field [853, 1271, 823]. Fifth [175]. fight [969]. figures [64]. filter [883]. find [1317]. Finding [1117, 1343, 596, 697]. fine [938, 1093]. fine-scale [938, 1093]. finfish [806]. FIP [859]. first [941]. Fish [72, 261, 1, 47, 110, 108, 116, 1236, 250, 518, 347, 469, 55, 981, 57, 818, 53, 867, 443, 444, 358, 435, 667, 73, 46, 80, 1244, 450, 520, 377, 426, 174, 183, 192, 506, 434, 1318, 102, 672, 226, 486, 898, 1024, 1270, 259, 758, 460, 1000, 333, 223, 154, 944, 788, 948, 321, 1222, 1214, 628, 701, 1284, 685, 813, 534, 1162, 177, 1067, 872, 1210, 721, 997, 1074, 1050, 1322, 244, 416, 404, 826, 1122, 594, 972, 576, 716, 861, 1321, 501, 484, 117, 344, 1120, 1185, 699, 828, 150, 604, 603, 1278, 75]. fish [119, 1233, 998, 1103, 497, 1317, 482, 1026, 1085, 1313, 267, 517, 919, 815, 1230, 737, 372, 962, 472, 728, 1073, 370, 314, 1262, 286, 326, 388, 1287, 982, 887, 562, 208, 621, 37, 856, 1063, 575, 708, 1292, 1189, 775, 1194, 1336, 269, 797, 640, 490, 951, 295, 447, 669, 715, 643, 7, 1264, 297, 1213, 811, 345, 607, 922, 569, 615, 725, 1243, 431, 680, 1091, 1334, 637, 17, 430, 1311, 709, 1199, 574, 284, 551, 635, 152, 507, 1028, 803, 391, 229, 592, 60, 662, 467, 1068, 897, 132, 746, 1041, 723]. fish [39, 589, 513, 462, 750, 994, 1276, 1011, 560, 1200, 784, 539, 537, 878, 349, 1154, 359, 59, 515, 114, 1164, 885, 646, 1335, 1021, 205, 1014, 274, 1147, 598, 710, 999, 161, 763, 832, 873, 48, 894, 1141, 1107, 1146, 888, 833, 1069, 521, 730, 1250, 714, 1265, 1280, 586, 325, 129, 138, 665, 984, 977, 1191, 724, 1267, 50, 590, 862, 1284, 299, 396, 252, 1078, 1216, 299, 673, 79, 69, 252, 141]. Fish-Net [788]. Fisher [1034, 289, 1144, 1208, 595, 703, 23, 718, 197, 1302, 726, 714, 392, 713]. fisher-driven [714]. fisher-influenced [718]. fisherfolk [169]. fisherie [169, 1157]. Fisheries [261, 47, 107, 109, 116, 121, 143, 379, 122, 364, 224, 238, 356, 156, 232, 380, 780, 12, 424, 14, 867, 374, 366, 407, 824, 451, 470, 81, 157, 450, 852, 338, 328, 1186, 809, 329, 165, 181, 173, 190, 289, 74, 146, 378, 123, 254, 318, 210, 58, 166, 15, 387, 390, 154, 944, 401, 43, 1081, 753, 849, 948, 1222, 334, 437, 1126, 550, 420, 266, 1034, 1298, 95, 1062, 1104, 1268, 900, 1330, 871, 820, 981, 831, 744, 636, 1074, 934, 1050, 865, 648, 235, 748, 493, 1088, 249, 351, 688]. fisheries [679, 952, 483, 1188, 1345, 1143, 576, 735, 663, 1223, 1108, 727, 1256, 1177, 1308, 336, 1139, 851, 3, 417, 1023, 596, 476, 409, 828, 500, 257, 931, 1231, 119, 1206, 1219, 729, 405, 968, 661, 959, 667, 573, 564, 291, 687, 1027, 1220, 1341, 425, 906, 1277, 1221, 1249, 1274, 787, 421, 601, 692, 213, 335, 547, 399, 962, 1029, 908, 1098, 1175, 354, 466, 806, 187, 219, 361, 529, 1056, 545, 314, 1195, 834, 456, 1151, 1066, 1287, 562, 1015, 1323, 149, 794, 847, 1017, 319, 312, 265, 292, 273, 1129, 275, 939, 393, 943]. fisheries [732, 455, 674, 1030, 920, 206, 408, 640, 1166, 535, 383, 491, 506, 583, 1332, 762, 504, 1305, 542, 985, 580, 492, 1105, 368, 400, 160, 609, 514, 819, 458, 1161, 929, 1180, 410, 431, 941, 800, 1273, 664, 899, 32, 1228, 276, 494, 362, 571, 1245, 634, 854, 1132, 94, 553, 773, 961, 1075, 1125, 489, 905, 584, 454, 597, 1257, 1102, 697, 705, 1072, 1347, 196, 605, 810, 93, 1165, 1238, 645, 764, 742, 1114, 1109, 516, 1187, 278, 382, 1094, 555, 991, 1301, 367, 538, 1263, 214, 945, 49, 967, 207, 513, 10, 6]. fisheries [511, 1133, 369, 759, 921, 842, 473, 1304, 1020, 769, 349, 413, 1314, 1179, 1320, 1153, 1060, 1160, 949, 204, 308, 510, 893, 1083, 719, 1040, 1315, 694, 389, 986, 999, 924, 1193, 1092, 1100, 1110, 832, 509, 599, 189, 980, 765, 969, 798, 1058, 1209, 726, 876, 925, 802, 942, 1055, 384, 1211, 786, 886, 585, 1234, 475, 947, 1082, 567, 903, 1295, 1004, 350, 572, 428, 612, 955, 693, 696, 1331, 822, 28, 163, 536, 686, 392, 526, 1016, 42, 964, 1156, 481, 242, 50, 199, 743, 954, 1054, 1176, 926, 927, 365, 72, 1284]. Fisheries [165, 1216, 73, 166, 154]. fisheries-induced [943, 819]. fishermen [666, 1022, 677, 812]. fishers [923, 666, 841, 1056, 675, 162, 719, 1226, 947, 1267]. Fishery [770, 147, 191, 1217, 70, 978, 341, 307, 352, 877, 826, 151, 264, 836, 745, 517, 566, 761, 1288, 1022, 1289, 1159, 857, 1170, 215, 958, 1119, 829, 383, 1207, 880, 263, 303, 1155, 1340, 799, 438, 768, 675, 1095, 1212, 1272, 1167, 381, 731, 540, 350, 859, 695, 294, 686, 973, 485, 827, 478, 1184]. fishery-dependent [1207]. Fishery-induced [341]. Fishes [108, 116, 99, 246, 627, 260, 452, 433, 823, 519, 200, 657, 11, 591, 1255, 209, 100, 639, 43, 54, 120, 983, 1327, 1138, 1134, 1258, 133, 131, 136, 442, 1333, 176, 995, 111, 306, 394, 171, 649, 544, 779, 1271, 1310, 134, 989, 191, 1196, 523, 137, 130, 159, 1025, 783, 864, 1112, 128, 907, 34, 324, 27, 904, 1007, 62, 1118, 35, 52, 298, 234, 51, 882, 1347, 305, 16, 315, 195, 682, 1042, 1033, 1291, 733, 188, 850, 1266, 84, 660, 890, 814, 1084, 874, 654, 92, 909, 178, 1178, 835, 932, 1295]. fishes [704, 543, 241, 1237, 87, 548, 1156, 1038, 1158, 1296, 606, 64, 140, 355].

Fishing [14, 366, 593, 852, 1105, 1204, 220, 15, 838, 64, 918, 478, 437, 845, 1284, 250, 1339, 871, 940, 691, 595, 236, 699, 756, 1300, 41, 570, 1087, 405, 531, 1329, 343, 1182, 1277, 1005, 403, 1279, 633, 466, 670, 1294, 1045, 928, 1189, 477, 757, 268, 1037, 669, 97, 1145, 992, 1077, 1099, 514, 556, 1059, 722, 785, 320, 94, 1006, 613, 296, 936, 860, 1238, 318, 1198, 588, 462, 675, 1276, 139, 539, 910, 1344, 248, 1086, 63, 789, 896, 554, 969, 1048, 1211, 751, 895, 559, 586, 1036, 734, 1076, 502, 524, 4, 973]. fishing [1171, 1267, 801, 1064, 1242, 224]. fishing-induced [751]. fishmeal [716]. FishPath [1219]. fishway [1051, 788]. five [1021]. flag [1182]. flagship [649]. flatfish [1343, 1123]. flatfishes [283, 223]. fleet [594, 446, 700, 1294, 1119, 1172, 449]. fleets [1329, 1182, 588, 714, 524]. FLIRES [726]. floodplain [770]. Florida [664, 489, 909]. flow [534, 229]. flowing [818]. Fluctuation [118, 177]. fluctuations [797, 1309, 22, 875]. fluid [1203]. fluvial [234]. focus [1165, 163]. Focusing [1195]. Following [375]. follows [566, 1124]. Food [80, 979, 1104, 1162, 1143, 716, 440, 756, 1057, 908, 1244, 640, 1346, 715, 569, 386, 553, 1174, 1100, 241]. food-grade [716]. food-limiting [1244]. Forage [832, 797, 513, 1021]. Foraging [439, 129]. forcing [584]. Forecast [835]. Forecasting [1067, 39]. forecasts [1074, 795, 1334]. foreign [1173]. forensics [318]. Forestry [260]. Foreword [127]. form [247]. formal [500]. former [499]. Forthcoming [116]. fossil [38, 415]. Four [1052, 1135, 151, 631]. Fourth [259, 232]. Fragmentation [1141]. Framework [101, 534, 831, 934, 594, 938, 1223, 1220, 988, 725, 941, 1334, 764, 1020, 1193, 816, 1055, 565, 734, 652]. frameworks [856]. free [150, 161]. free-swimming [150]. frequencies [1090]. frequency [957, 1118]. frequentist [117]. fresh [908]. Freshwater [29, 753, 124, 444, 805, 519, 200, 11, 591, 100, 1214, 1327, 534, 998, 1103, 649, 286, 388, 887, 498, 1292, 563, 1196, 523, 1264, 811, 324, 27, 637, 152, 298, 758, 882, 1033, 733, 750, 878, 48, 1282, 325, 954]. Friday [164]. Frontiers [980]. fronts [1087, 656]. Fuel [588]. full [1047]. fully [571, 552]. fun [756]. function [75, 496, 1022, 646]. Functional [493, 1026, 314, 282, 1122, 306, 1063, 708, 790, 882, 974]. functions [345]. fundamental [362, 369]. fundamentally [1224]. funded [320]. further [554]. Future [852, 944, 1339, 1138, 736, 871, 244, 76, 851, 482, 988, 1017, 1336, 1030, 408, 490, 491, 1136, 917, 1260, 305, 996, 1033, 967, 858, 1314, 359, 660, 763, 642, 798, 1107, 221, 521, 1191, 926]. futures [1337]. fuzzy [1187].

G [1284]. gadid [1123]. Gadidae [1028]. gadids [1039]. Gadus [1028]. galaxiid [234]. Galaxiidae [234]. game [400]. gap [1251, 315]. gaps [1206, 847, 904, 882, 588]. gas [994]. Gasping [1270, 460]. GBP [869, 520, 606]. Gear [1108, 685, 236, 618, 1099, 664, 910]. gear-induced [685]. gears [933, 539]. Gender [580, 1245, 1135, 1314]. Gender-based [1245]. gendered [722, 392]. general [912, 883, 816]. generated [258]. generation [973]. Generic [836, 421]. Genetic [617, 195, 349, 344, 818, 720, 993, 68, 776, 315, 317, 316]. genetic-based [818]. genetically [1089]. Genetics [108, 156, 125, 83, 319, 312, 682, 555, 313, 281]. genome [532]. genomics [314]. genotype [315]. Geographic [122, 501, 882, 359]. Geography [761, 1246]. geolocation [1073]. Georgiy [606]. Geostatistics [55, 49]. get [942]. Getting [947]. ghost [1027]. Ghosts [1027]. Ghoti [277]. Giant [567, 1095, 622]. gigas [1095]. Gill [1229, 209, 1229]. gillnet [249]. gillnets [559]. Gills [1258, 1158]. gladius [1155]. glanis [340]. glauca [747]. Gleaner [392]. Global [527, 691, 727, 1308, 824, 1124, 1159, 1286, 432, 752, 329, 1065, 1142, 1196, 803, 768, 248, 731, 944, 1004, 502, 1152, 743, 29, 1222, 420, 1062, 820, 1162, 1050, 652, 1333, 663, 339, 336, 931, 651, 614, 805, 267, 566, 564, 690, 1306, 1063, 1017, 1032, 393, 757, 587, 1145, 492, 705, 810, 746, 1094, 1041, 588, 513, 921, 473, 769, 910, 1014, 789, 1193, 969, 1209, 886, 1082, 163, 536, 524, 623, 977, 801, 1242, 481, 199, 954, 478, 926, 981, 740]. globalization [857, 429]. globe [1006]. globule [813]. go [942]. goal [849, 481]. goalposts [1092]. goals [792, 10, 493, 265, 1037]. goes [139]. going [1010, 1224]. Golden [740]. Goldfish [83]. Gonorynchiformes [435]. Good [199, 66, 497, 830, 670, 320]. goods [400]. got [3, 114]. governability [380, 383, 381, 765]. governable [765]. Governance [261, 379, 876, 437, 1330, 688, 380, 791, 573, 1175, 529, 1286, 535, 921, 1144, 1209, 384, 1184]. Government [752]. Gozlan [326]. grade [716]. Grammatidae [191]. graphs [520]. Grayson [1296]. Great [760, 566, 327, 684]. greater [694, 995]. Greatest [838]. green [745]. Greenland [1011]. grey [197]. grim [164]. Grisel [1184]. groundfish [1065, 1181, 1290, 800, 1111]. group [840, 8]. grouper [847, 222, 1052, 197, 350]. groupers [478]. groups [840, 974]. grow [1262]. Growth [1270, 460, 120, 1258, 1229, 966, 336, 1233, 370, 1310, 1119, 68, 295, 297, 953, 1028, 803, 1266, 642, 932, 350, 1146]. guard [1286]. Guianas [823]. Guide [433, 807, 823, 191, 657, 629, 1271, 1218]. Guidelines [431, 390, 567]. Guiding [1165, 968]. guild [267, 560]. guilds [534, 1063]. Gulf [140, 197, 352, 283, 287, 197, 231]. gut [518, 919, 1283]. gut-limited [919]. Gyotaku [898]. gyre [1011].

H [1324]. Habitat [461, 790, 1281, 177, 1321, 1085, 782, 883, 490, 1264, 51, 605, 1068, 462, 161, 1093, 1265]. habitat-based [177]. Habitats [224, 711, 881, 826, 1343, 962, 78, 626, 1275, 1232, 936, 684, 1164, 890, 1107]. haddock [610]. hagfishes [191]. hake [581, 1272]. Half [720]. halibut [812]. Hall [1284]. halved [502]. Handbook [180, 216, 83, 270, 154]. Harbor [164]. Hard [64, 1176]. hard-to-age [1176]. harengus [967, 863, 22, 875]. harmful [603]. Hart [1184]. Harvest [512, 1221, 948, 1268, 631, 1341, 1277, 421, 457, 562, 822]. harvestable [1309]. harvesting [655, 1233, 819, 1204, 675]. hatcheries [1337]. Hatchery [141, 76]. Hawai’i [489, 912]. Hazards [912, 1076]. Hbk [606]. Health [849, 606, 801, 1062, 670, 1261, 498, 1030, 640, 1077, 1105]. Hearing [1247]. heatwave [1298, 1249]. heatwaves [1329]. heel [528]. help [1279, 597, 10, 6]. Hemisphere [749]. herbivore [605]. here [47, 942]. Heritage [838]. hermaphroditism [195, 282]. Herpetologists [191]. Herring [1228, 863, 1101, 840, 967, 22, 875, 956]. heterogeneity [1107]. Hidden [1291, 264]. hierarchy [934, 778]. High [1266, 678, 1206, 791, 1289, 453, 1248, 759, 974]. high-Antarctic [453]. highlight [1246]. Highly [1136, 648, 114]. highs [956]. Hilborn [1284]. hilsa [829]. hindcasts [1334]. histological [1200]. historic [1222, 602, 804]. Historical [418, 1005, 293, 321, 352, 576, 1343, 335, 1157, 1347, 317, 359, 749, 966]. Histories [461, 1321, 484, 1090, 1125, 1240]. History [310, 433, 1228, 209, 301, 1246, 644, 283, 1297, 821, 638, 1279, 1106, 1310, 782, 958, 883, 1140, 506, 297, 1275, 739, 773, 35, 902, 848, 415, 1137, 909, 932, 242, 61, 707, 965, 1127]. HIV [169]. Hoar [183]. hobby [637]. Holding [1182]. Hollingworth [228]. Holt [959, 739, 356]. homeward [619]. homing [35]. homology [907]. Hong [350]. hook [331]. hooks [725, 769]. Hoole [64]. Horizontal [509]. hot [232]. hotspot [1091]. hotspots [1052]. Human [399, 437, 644, 1251, 952, 417, 1124, 1289, 670, 1019, 1261, 640, 553, 6, 885, 1242]. human-influenced [644]. human-made [6]. Humanity [1106]. humans [1202]. Humboldt [404]. Hundred [111]. Hundred-year [111]. Hunting [41]. Hutchings [1240, 1131]. hybridisation [1291]. hydroacoustic [574]. Hydrobiologie [71]. hydrology [427]. hydropower [856, 662, 788]. hydrostatic [426]. hyperoxia [1024]. hypotheses [312, 159, 1069]. hypothesis [619, 177, 1121, 292, 569].

Iberian [1157]. ice [684, 846]. Iceland [634]. ICES [1220]. ichthyologist [728, 348]. Ichthyologists [191]. Ichthyology [180, 85]. iconic [638]. ideal [161]. Identification [191, 649, 790, 1145]. identify [1264, 568]. Identifying [1301, 846, 1058, 965]. If [653]. Ignore [23, 530]. IIFET [42]. ilisha [829]. ill [710]. ill-defined [710]. illegal [343, 1182, 1238, 96, 1036, 1172]. illuminate [1052]. illumination [292]. Illustrated [46, 433, 912]. imaging [1191]. imagining [1274]. IMBER [652]. IMBER-ADApT [652]. immediate [618]. Immunology [443]. Immunomodulation [922]. Impact [1298, 603, 1249, 688, 787, 633, 593, 319, 514, 922, 697, 705, 93, 749]. impacted [403]. Impacts [820, 824, 239, 44, 719, 630, 334, 861, 1256, 339, 497, 1113, 687, 466, 1012, 504, 776, 1136, 643, 458, 1006, 936, 641, 1033, 429, 1335, 557, 1167, 984, 1171, 918]. impairment [358]. imperative [204]. Imperfect [182, 530]. imperilled [563, 882, 733]. Implementation [366, 1088, 409, 742]. Implementing [468, 407, 1102, 1286, 802]. Implications [968, 174, 943, 669, 1109, 650, 352, 250, 685, 644, 499, 425, 1221, 208, 269, 757, 464, 505, 33, 1024, 52, 682, 1144, 162, 719, 978, 1092, 950, 521, 895, 91, 541, 686, 1267, 50, 918, 954]. imply [448]. importance [1256, 306, 1008, 1311, 656, 1291, 462, 59, 811, 758]. Important [1053, 618, 617, 90, 686]. imports [615]. imprint [1197]. improve [865, 931, 1029, 951, 915, 1133, 242]. improved [614]. improvement [617, 436, 859]. improvements [683, 859]. improves [1069].

Improving [706, 961, 1094, 772, 257, 318]. Incentive [877]. Incentivising [1182]. include [1074, 1277]. included [599]. Including [1287]. Inclusion [911, 698]. inclusive [1274]. income [448, 947]. Incorporating [1190, 594, 431]. increase [671, 405]. increased [575]. increasing [208, 674, 1320, 1040]. increments [863]. incursions [940, 1064]. index [148, 184, 217, 1229, 1244, 1069, 849]. India [579, 383]. Indian [747, 1012, 1179]. indicator [1287, 1244, 1032, 718]. Indicators [206, 606, 342, 153, 593, 1140, 1155, 645]. Indigenous [1332, 1020, 546, 1274]. Indirect [699, 278]. Individual [354, 300, 322, 1233, 517, 37, 597, 675, 466]. Individual-based [300]. individuals [671]. individuation [88]. induced [685, 1010, 982, 943, 819, 1002, 1334, 1021, 474, 751, 341]. industrial [940, 1099, 1059, 1155, 1133]. industrialized [95]. Industry [320, 1077, 785, 318, 1086, 925, 1228]. Industry-funded [320]. inequality [573]. infer [920]. inference [295, 362, 168]. inferences [492]. inferred [499]. Inferring [682]. influence [1050, 336, 1085, 1290, 1086, 730, 22, 586]. influenced [644, 718]. influences [501, 902, 195]. Influencing [82, 66, 1175, 459, 462]. inform [531, 1273]. Information [741, 754, 767, 781, 793, 808, 825, 839, 855, 870, 884, 901, 914, 930, 946, 960, 971, 987, 1001, 1018, 1031, 1046, 1061, 1080, 1096, 1115, 1130, 1149, 1169, 1183, 1201, 1215, 1227, 1239, 1254, 1269, 1285, 1299, 1312, 1325, 1338, 122, 847, 135, 542, 1282]. information-limited [542]. informative [276]. informed [1089, 1320]. Informing [916]. ingress [1237]. initiatives [812]. Inland [81, 927, 43, 1222, 95, 493, 1333, 735, 908, 1187]. Innate [1214]. Innovation [722]. Innovations [110]. innovative [865]. inshore [940]. Insights [1244, 310, 150, 1159, 640, 1326, 1070, 546, 1107, 722]. Instantaneous [840]. Institute [42]. institution [857]. Institutional [383]. institutions [500]. Intake [80]. Integrated [742, 1117, 1010, 932, 734]. Integrating [925, 1111, 316, 242, 744, 466, 1165]. integration [643]. Integrative [1134, 58]. integrity [498]. intensity [502]. inter [1210]. inter-assessment [1210]. interacting [1317, 607]. interaction [720]. Interactions [851, 260, 174, 1278, 640, 1180, 480, 1302, 876, 138]. Interactive [261]. interdisciplinary [874]. interest [693]. interface [1261]. interference [1124]. Intergenerational [179]. interim [957]. intermediate [511, 903]. International [42, 1286, 1319, 673, 616, 1095, 1058, 543, 945]. Internet [1161]. interruptive [245]. intersection [1162]. interspecies [1318]. Intertidal [1232]. intra [1318]. intra- [1318]. Intraspecific [241]. Intriguing [1203]. intrinsic [932, 1282]. Introduced [43, 340]. Introduction [286, 309, 191, 325, 326, 750, 84, 255]. introductions [1185]. Introgression [988]. intuitive [1147]. invader [340]. invasion [388]. invasions [388]. invasive [879]. inverse [1170, 739]. invertebrate [1210, 661]. investigate [873]. Investigating [404, 703, 335]. Investing [265]. invitation [595]. involvement [63]. ion [1158]. iron [373]. ISBN [1241, 1324, 869, 520, 606, 1255, 64]. ISBN-13 [520]. Island [107, 869, 462, 761, 152, 123]. islandica [863]. Islands [684, 775]. Isles [1005, 540]. isotopes [920, 1157, 1197]. Issue [116, 741, 754, 767, 781, 793, 808, 825, 839, 855, 870, 884, 901, 914, 930, 946, 960, 971, 987, 1001, 1018, 1031, 1046, 1061, 1080, 1096, 1115, 1130, 1149, 1169, 1183, 1201, 1215, 1227, 1239, 1254, 1269, 1285, 1299, 1312, 1325, 1338, 104, 47]. Issues [1216, 14, 15, 571, 978, 693]. iteroparous [188]. ITQs [336]. IUU [969].

J [959, 183, 558, 223, 769]. J-hooks [769]. Jamaican [332]. James [1228]. Japan [1250]. Japanese [304, 1205]. Jason [1256]. Jeffrey [1240, 1131]. Jersey [1324]. Johannes [740]. John [1324, 367]. joint [877]. Josep [606]. Journal [104]. Journey [658]. jubatus [546]. judgement [1147]. Junagadh [383]. Jurassic [782]. jurisdictional [985]. just [771, 278, 72]. juvenile [1321, 569, 999, 843, 350]. juveniles [671, 1237].

K. [290]. Kaplan [1284]. Keep [153]. Keeping [258]. Ken [1078]. Kevin [520]. Key [677, 859, 1298, 1249, 399, 453, 725, 1232]. key-habitats [1232]. kin [1116, 138]. kinetic [1203]. King [1185]. Kiribati [1345]. Knowledge [1274, 289, 861, 720, 1129, 23, 587, 951, 1332, 626, 904, 1007, 284, 996, 415, 588, 1190, 463, 950]. knowledge-needs [626]. Kong [350]. Krill [869, 373, 438, 73]. Kuroshio [1002].

L [1284, 142]. L.G [1255]. labelling [1261, 181]. Laboratories [164]. labour [1139]. ladyfishes [527]. Lake [121, 653, 186, 516, 382, 465, 381, 90]. lakes [881, 1032, 1346, 684]. Lampreys [209]. lances [950]. land [671, 770, 775, 1232]. Landings [257, 614, 208, 1012, 1344]. landlocked [465]. Landmark [406]. landscapes [774]. lane [1266]. Large [589, 948, 668, 638, 1341, 562, 303, 837, 1009, 759, 1141, 1004, 584]. large-scale [668, 1009]. largest [964]. Larkin [272, 2, 31, 65, 203]. larvae [1236, 385, 1085, 1237]. Larval [804, 528, 992, 263, 1146]. last [1023, 1250]. Late [782]. later [47, 336, 372, 498, 567]. lateral [1141]. Latin [600]. Latitudinal [544, 989]. Latitudinally [1224]. launch [1216]. lead [681]. Leadership [7]. leading [723]. Learned [134, 130, 700, 273, 505]. Learning [116, 131, 128, 558, 1315, 129, 136, 353, 132, 1164, 999, 1167, 1267]. Lecture [2, 272]. Lectures [31, 65, 203]. Led [752]. Leeches [209]. Legacy [463, 406, 887]. legislation [1088]. lekking [8]. length [948, 264, 562, 1140, 1176, 1003]. length-based [562, 1140]. length-structured [1176]. lens [1135, 1218, 1128, 526]. lenses [786]. lesions [34]. Leslie [367]. less [1337]. less-than-wild [1337]. Lessons [409, 335, 411, 332, 852, 1332, 1298, 76, 966, 1249, 304, 505, 617, 21, 567, 28, 926]. level [621, 674, 705, 874, 707]. Levels [521]. Levins [160]. liability [877]. Lice [209]. licences [896]. license [320]. Life [261, 461, 869, 147, 620, 82, 1266, 909, 701, 1150, 1246, 644, 283, 1321, 484, 821, 1090, 1279, 1310, 1202, 883, 809, 1140, 1336, 506, 453, 1213, 811, 1275, 1290, 739, 1307, 773, 1125, 35, 1240, 902, 848, 5, 205, 894, 932, 242, 61, 965, 1127, 647, 1297]. life-history [644, 1279, 1140, 739, 35, 932, 61, 965, 1127]. lifetime [1280]. light [1138, 1140, 1328]. like [988, 1030]. limit [948, 545, 1003]. Limitation [1229, 569]. Limited [254, 499, 1223, 821, 836, 1219, 919, 1220, 547, 963, 542, 822]. limiting [1244, 683]. limits [727, 264, 761, 601, 1184]. line [1, 748]. Link [1256, 1322, 525]. linkages [1246]. linked [1198]. Linking [1192, 288, 490, 850, 315]. links [265, 873, 862]. lionfishes [879]. lions [546]. listed [47]. listening [997]. listing [787, 916]. literature [698, 479, 904, 910, 197]. Litopenaeus [1283]. little [709, 1282]. live [715]. lived [1038]. livelihood [1160]. livelihoods [629, 1300, 908, 690, 1302, 726]. lives [1297]. Living [191, 1203, 189]. Lloret [606]. lobsters [528]. local [1104, 131, 663, 1087, 1332]. locally [1188]. locations [1113]. logbooks [352, 257]. loggerhead [843]. London [869, 520]. Long [1250, 600, 118, 1045, 530, 967, 875]. Long-term [1250, 600, 118, 967, 875]. longer [1054]. longline [249, 446, 564, 187, 219, 632, 769]. longlining [576]. look [908]. looks [988]. loopholes [177]. losers [609]. loss [1026, 1099, 1347, 910]. loudly [1112]. Love [606, 658]. low [499]. lower [957]. lowland [1107]. lows [956]. Lunar [178, 292]. lungfish [415]. Lutz [142].

M [1241, 1284, 520, 1193]. M-Risk [1193]. machine [1167]. Mackovjak [1228]. macroalgae [1156]. macroalgae-associated [1156]. Macroalgal [962]. macrofauna [979]. macrotidal [1237]. made [6, 765]. magnitudes [1263]. Maine [352, 140]. Majidae [1119]. Major [38, 243]. make [1274, 1030]. makes [648, 725, 276]. Making [366, 628, 916, 198, 549, 1165]. Malawi [186, 381, 90]. Malcolm [606]. male [992]. males [598]. Malthus [738]. mammalian [505]. mammals [191]. man [1026]. manage [568]. Managed [202, 845, 1257]. Management [109, 101, 121, 780, 424, 867, 366, 407, 470, 403, 104, 260, 982, 81, 328, 1186, 174, 71, 289, 281, 254, 279, 210, 58, 759, 608, 978, 650, 66, 266, 95, 644, 534, 744, 468, 865, 1088, 688, 483, 418, 1223, 1108, 1256, 678, 851, 417, 264, 440, 596, 500, 1231, 1233, 614, 600, 631, 1219, 729, 405, 968, 667, 668, 291, 479, 243, 1220, 713, 1221, 1274, 213, 399, 411, 830, 1029, 611, 354, 529, 834, 456, 1287, 1294, 1015, 149, 157, 1129, 939, 477, 829, 957, 732, 506, 762, 790, 718, 715, 368, 400]. management [514, 263, 1243, 929, 410, 431, 941, 582, 800, 1273, 677, 899, 32, 62, 1199, 362, 634, 94, 1075, 1125, 258, 454, 597, 1102, 93, 1165, 645, 764, 1114, 516, 915, 682, 1303, 555, 1301, 207, 812, 10, 511, 659, 1144, 473, 1095, 1020, 349, 1053, 949, 1218, 162, 893, 1040, 63, 694, 389, 1193, 1110, 832, 1084, 509, 599, 189, 1167, 689, 950, 942, 1055, 1234, 1178, 903, 565, 766, 955, 316, 686, 1128, 522, 827, 242, 954, 307, 1127, 862, 310, 108, 141]. management-strategy [368]. managerial [685]. Managing [792, 416, 573, 841, 857, 198, 273, 553, 413, 1153, 409, 912, 547, 1159, 263, 936, 842, 204, 572]. mangrove [1322]. Mangroves [663, 1322, 710, 986]. Manual [192, 45]. manuscripts [398]. many [298, 710]. Mapping [1151, 523, 886, 163, 532].

Margaret [348]. Maria [1255]. mariculture [303]. Marine [72, 106, 143, 124, 126, 142, 12, 566, 1329, 824, 470, 239, 310, 432, 450, 807, 165, 173, 191, 201, 606, 251, 609, 615, 853, 584, 279, 429, 740, 302, 639, 798, 895, 256, 983, 1034, 98, 1298, 644, 245, 561, 1251, 1210, 1074, 322, 652, 418, 1321, 339, 678, 344, 336, 1120, 851, 1023, 440, 699, 603, 1278, 651, 41, 1057, 497, 975, 446, 1085, 112, 171, 968, 1113, 1288, 1277, 1005, 1249, 916, 525, 1279, 411, 1073, 1089, 370, 545, 1310, 963, 1045, 794, 312, 1196, 1336, 393, 78, 669, 718, 776, 492, 643]. marine [160, 1326, 569, 230, 680, 1091, 844, 837, 1118, 905, 1257, 697, 778, 530, 1347, 391, 330, 1165, 568, 60, 774, 848, 641, 1303, 513, 6, 96, 858, 349, 463, 1154, 850, 1179, 1212, 113, 1218, 197, 231, 552, 313, 161, 67, 925, 942, 179, 585, 833, 475, 947, 835, 1004, 1280, 428, 1331, 695, 543, 630, 316, 294, 9, 502, 1237, 702, 522, 743, 1202, 1304, 838]. maritime [1139, 1172]. marker [532]. marker-assisted [532]. markers [1293, 37]. Market [641, 671, 151, 1114, 876]. market-based [1114]. marlin [860]. Marseille [857]. Marshak [1256]. masculinized [598]. Masked [651]. Mass [1042, 1213, 892, 9]. mass-balance [9]. Mass-effect [1042]. Maternal [815, 159]. matter [813]. matters [1195, 715]. maturation [297]. maxima [622]. Maximizing [806]. Maximum [1330, 395, 847, 94, 1153, 841]. may [47, 545]. mbuna [186]. McDowall [414]. McGill [228]. MCMC [494]. McMurdo [495]. meadow [962]. mean [243]. means [1139, 897]. measures [518, 351, 632, 1294, 473, 360]. Measuring [103, 561, 494]. mechanism [1203]. mechanisms [737, 953, 459, 1069, 827]. Mechanistic [357, 919]. mechanistically [861]. media [970, 1171]. mediate [1289]. medieval [857]. Mediterranean [1184, 1281, 671, 1188, 419, 440, 476, 1155, 937, 625]. Meet [423, 119, 372]. meets [1232]. megafauna [576, 778, 1179]. Megalopidae [527]. melanostictus [1205]. membras [967]. memory [128]. menhaden [795]. Mental [1062, 1077]. mesh [62]. Mesohabitat [977]. Mesoscale [1087, 1242]. Mesozoic [16]. messages [970]. Meta [1293, 993, 860, 872, 663, 975, 564, 1230, 610, 1051, 1196, 771, 1283, 933, 810, 897, 1041, 994, 769, 910, 885, 789, 598, 1211, 585, 567, 956, 622, 428, 625, 630, 533]. meta-analyses [598]. Meta-analysis [993, 860, 872, 663, 564, 610, 1051, 1196, 1283, 810, 897, 1041, 994, 769, 910, 885, 789, 1211, 585, 567, 956, 622, 625, 630, 533]. Meta-analytic [1293, 1230, 771]. meta-analytical [975, 933, 428]. metabarcoding [1047, 1057]. metabolic [550]. Metabolism [1310]. Metapopulation [160, 274, 67]. Method [1223, 1259, 494, 503, 858, 221, 1003]. methodological [953, 705]. methodologies [466, 1307]. Methodology [807, 985]. Methods [171, 253, 27, 259, 117, 821, 170, 1103, 512, 1180, 94, 296, 1295, 1116, 1127, 853, 251]. metric [362]. metrics [856, 586, 955]. Mexico [745, 197]. MEY [363]. Michael [1296]. microarrays [314]. microbiome [873]. microbiota [1283]. microchemistry [501]. Microplastics [1033, 1196]. microsatellite [37]. microtidal [1237]. midlatitude [1242]. Migration [1000, 100, 639, 619, 385, 1316, 669, 846, 751]. Migrations [935, 1121, 1292]. Migratory [168, 574, 662, 589]. Millennium [493, 265]. million [782, 924]. mine [702]. minimal [330]. Minimizing [633]. minimum [948, 417]. minimum-length [948]. mining [343]. minke [457]. minnows [563, 760]. Miocene [48]. mis [1261]. Misconstrued [1005]. misdiagnosis [498]. mislabelling [459]. Mismanagement [450]. mismatch [1117, 349]. miss [23]. missing [525]. Mississippi [1009]. Misspecification [889, 1117]. Mitigating [361, 559, 828]. mitigation [934, 351, 632, 778, 1133]. mixed [1305, 1136, 263, 1072, 891, 1301, 843]. mixed-species [891]. mixed-stock [1305, 1072]. mode [618]. Model [275, 404, 446, 919, 1106, 186, 883, 1119, 295, 400, 160, 1243, 480, 330, 1174, 676, 978, 932, 9, 1176]. Modeling [280, 300]. Modelling [479, 295, 643, 1212, 952, 440, 342, 411, 700, 508, 1142, 904, 530, 980, 1328, 749, 977]. Models [309, 4, 706, 1067, 1117, 679, 596, 911, 1231, 512, 1073, 700, 1010, 1019, 33, 1273, 742, 1068, 1137, 1301, 214, 49, 511, 1320, 873, 816, 1048, 731, 903, 822, 973, 788, 449]. moderation [67]. Modern [1077, 913]. Modes [1237]. Modification [78]. modifications [1321]. modulate [1242]. Molecular [262, 1343, 319, 555]. Molidae [191]. molluscs [448, 191]. monitor [854]. Monitoring [867, 1145, 279, 1191, 561, 235, 457, 508, 928, 757, 1161, 574, 878, 359, 788, 926]. Monograph [579]. monoterpene [1230]. moratoria [616]. Moreno [1284]. morhua [1028]. Morphological [371, 850]. morphology [1327]. mortality [236, 264, 618, 358, 370, 856, 797, 542, 669, 992, 1213, 739, 860, 539, 769, 1344, 1021, 1048, 894, 559, 4, 1127, 1146]. Mortality/Growth [1146]. mosaics [917]. mosquitofish [1106]. Most [716, 931, 1195, 653]. motivation [1311]. motives [459]. move [531]. move-on [531]. movement [818, 1010, 1091, 1039, 1093]. movement- [818]. movements [935, 589]. Movescapes [1091]. Moving [738, 525, 1092, 856, 1304]. MPA [243]. MPAs [668]. MSc [228]. MSY [363, 1317, 32, 503]. much [198]. mud [1340]. Multi [1188, 1343, 1203, 479, 295, 1174, 1301, 1335, 1048]. multi-area [1188]. multi-disciplinary [479]. multi-mechanism [1203]. multi-model [295, 1174]. Multi-molecular [1343]. multi-scenario [1335]. multi-species [1301, 1048]. Multi-specific [1188]. multibeam [235]. Multicentury [489]. Multicriteria [799]. Multidecadal [1028]. multidisciplinary [629]. Multimodel [642]. Multiple [811, 723, 948, 1195, 1205, 516, 1093]. Multiple-stressor [811]. multiscale [1172]. Multisectoral [1030]. Multispecies [511, 974, 1317, 893, 1211]. multistressor [917]. multitracking [482]. multivoltine [1038]. Munro [367]. muscle [75]. muscles [229]. mussel [758]. mussels [758]. mutual [632]. mutually [1173]. mutually-beneficial [1173]. mydas [745]. Myers [1333]. myth [391]. Myths [67].

N [216]. namaycush [653]. Namibia [190]. Narco [981]. Narco-Fish [981]. narrative [945]. narratives [602]. Narrow [786, 1045]. National [423, 334, 587, 1105, 705]. national-level [705]. nations [613, 1286]. nationwide [1211]. native [416, 340, 286, 326, 498, 1260, 324, 684, 325]. Natural [280, 487, 433, 292, 202, 797, 542, 226, 209, 820, 1297, 370, 27, 88, 915, 1021, 509, 1127, 301]. naturalist [197]. navigation [619, 426]. Near [114, 1136]. near-future [1136]. nearly [114]. Necessary [650]. need [792, 1009, 846]. needed [1166]. needs [527, 1129, 587, 626]. Neglected [1155, 899, 1295, 751]. negotiations [722]. Nemo [697]. Neogene [48]. Neotropical [1026, 1007, 733, 1282, 589]. nesting [1134]. Net [788, 361, 664]. network [872, 1070, 886]. networks [1157, 967, 548]. neural [133]. Newfoundland [485]. News [64, 199]. next [873, 942, 973]. nexus [756, 945]. Nicol [869]. night [1138]. Nineteenth [602]. Nitrogen [69]. No [72, 919, 191, 561, 372, 263, 428, 1173]. no-take [561, 263, 428]. Noakes [1255]. noise [810]. Non [1205, 416, 517, 528, 286, 326, 498, 324, 1002, 991, 890, 1123, 325]. non- [991]. non-binding [517]. non-feeding [528]. non-native [416, 286, 326, 498, 324, 325]. non-reef [890]. Non-stationary [1205, 1002, 1123]. Nonindigenous [43]. Nonlinear [810]. nonlinearity [1002]. nonstationarity [1309]. Nordic [1114]. Norms [1184]. North [673, 118, 19, 843, 1265, 693, 862, 77, 576, 111, 863, 1089, 201, 1101, 293, 11, 591, 1136, 546, 164, 93, 278, 848, 924, 1111, 155, 760]. north-east [118, 693, 862, 1089]. north-west [843]. north-western [19]. northeast [412, 1035, 982, 1181, 1334, 1174, 777]. northern [404, 1335]. Northwest [1022, 427, 410, 1235, 950, 1253]. Norwegian [457, 193, 22, 875]. note [2, 31, 65, 203]. Nothing [1131, 842]. notothenioid [1025]. novel [938, 1045, 1260, 568]. novelty [1126]. Nudging [1100]. numerical [1063]. nurseries [568]. nursery [568, 1107]. nutrient [1153]. Nutrition [448, 756, 1105].

Obituary [1131, 228, 290]. objectives [948, 744, 934, 597, 516]. obligate [992]. obligation [671]. observation [508, 1099]. observations [1052, 197]. observed [1192]. observer [257]. Observing [1278]. occur [701]. Occurrence [1033, 512]. Ocean [283, 499, 867, 433, 752, 1017, 1202, 82, 387, 423, 1064, 1150, 1203, 448, 863, 687, 508, 563, 643, 609, 458, 746, 1291, 798, 875, 1242, 849, 321, 77, 373, 777, 555, 1263, 1179, 1035, 1178]. oceanic [1206, 656]. oceanographic [1289, 1190]. Oceanography [107, 122, 123, 58]. Oceans [310, 852, 343, 1260, 1272, 747]. octavolateralis [345]. off [177, 830, 915]. offs [934, 748, 1108, 424, 713, 507, 1303, 832, 822, 964, 1054]. offshore [283]. offshore-spawning [283]. ogives [1344]. Oil [813, 1012]. old [966, 1342, 562]. older [1287]. olfactory [496]. Omnidirectional [235]. omnivory [814]. on-going [1224]. once [427]. Oncorhynchus [881, 1013]. one [3, 1030, 114, 964]. online [1161]. only [1137, 822]. Ontogenetic [1120, 1010, 1181]. ontogenies [541]. ontogeny [1290]. Oocyte [779]. Open [433, 387, 1182, 508, 1109]. Opening [1326]. Operating [735, 978, 1055]. operational [1231, 941, 809]. Operationalizing [381]. opilio [1119]. Opistognathidae [191]. Opportunities [865, 952, 1029, 211, 1104, 1087, 1198, 696, 760, 926]. opportunity [931, 446, 1015]. optic [951]. Optimal [1316, 506, 933]. optimist [794].

Optimizing [999]. Options [744, 828, 291, 713]. organisation [135]. organisms [247]. Organizations [978, 529, 1015, 762, 812, 1186]. organs [496]. orientation [132]. Origins [540, 425, 1197]. Orinoco [823]. ornamental [547, 615]. Ostariophysan [435]. Other [64, 1294, 643, 634, 274, 586, 1076]. Otolith [771, 501, 1042]. Our [852, 393, 163]. ourselves [1337]. outcome [705]. outcomes [748, 119, 1160]. Outlook [164]. outperforms [948]. overcome [400]. overcoming [1029]. Overfishing [752, 401, 614, 738, 153, 605, 473, 681, 590]. overlap [1289]. Overlooked [557]. Overturning [1187]. overview [646]. ownership [1182]. Oxford [1256, 1240, 64]. oxidative [721]. Oxygen [1229, 1270, 460, 1192, 1197]. oyster [806].

P [1324, 183, 64, 1184]. Pacific [77, 868, 777, 1263, 1253, 1135, 736, 1192, 1300, 745, 425, 1022, 1289, 225, 958, 412, 293, 1326, 1275, 1260, 1181, 1235, 546, 1307, 164, 278, 462, 812, 1272, 1035, 876, 1111]. paddlefish [207]. Pagrus [602]. pain [176, 515, 396]. Palaeozoic [16]. paleo [684]. paleo-ice [684]. Paleocene [782]. Pan [388]. Pan-continental [388]. pandemic [1030]. pandemic-like [1030]. Panting [1270, 460]. Papers [72, 406]. Paradigm [312, 476, 525, 680, 1266, 986]. paradigms [563]. paradox [404]. Parallel [91]. paralytic [603]. parameter [909]. parameters [993, 369, 999, 965, 1127]. Parasites [443, 11, 209, 262, 469, 630]. Parental [1025]. Park [566]. part [479, 191]. partially [902, 552]. Participation [757, 871, 1144]. participatory [1151, 455, 858]. particular [178]. Parties [916]. partnerships [925]. party [1086]. parva [388]. Passage [110, 856, 951, 507, 467, 359, 763]. Passive [685, 361, 775]. Past [310, 736, 1231, 863, 335, 988, 332, 660, 231]. path [1029, 1045]. pathobiology [469]. Pathogens [443]. Pathology [79]. paths [283]. pathways [1288, 1323, 750, 1314]. patient} [1131]. patron [383]. pattern [887]. Patterns [394, 537, 1236, 747, 566, 403, 1306, 989, 464, 490, 359, 1004, 956, 61, 707]. Pauly [1150]. Payments [636]. payoff [499]. pays [830]. peacock [1124]. Peggy [348]. pelagic [177, 404, 972, 1192, 497, 1329, 219, 632, 1189, 563, 1243, 1002, 860, 1291, 759, 769, 1014, 1084]. pelagic-broadcast [563]. Peninsula [1157]. People [14, 15, 931, 273, 546]. per-recruit [965]. Perception [697]. Perceptions [613, 571, 905]. Percid [54]. perfect [1117]. Performance [529, 1229, 1085, 718, 1068, 1304, 384]. Perspective [201, 29, 89, 573, 470, 408, 587, 1275, 615, 929, 677, 656, 274, 1191, 801, 307]. Perspectives [155, 966, 176, 482, 794, 844, 1033]. Peru [177]. Peter [142]. Petrarch [6]. pharmacological [333]. PhD [228]. phenology [883]. phenomena [1309]. phenotype [315]. phenotypic [897]. Pheromones [672]. philosophy [97]. photographs [520]. photoprotective [922]. phylogenetic [1295]. physical [1002, 592]. physiological [1293]. Physiology [183, 1158, 951, 458, 284, 646, 242, 252, 69]. Phytoplankton [124]. Piferrer [1324]. pillars [1159]. pilot [151]. pink [1260]. pioneer [367, 348]. piscine [89]. Pitcher [223]. Pitfalls [211, 572]. pits [499]. place [243]. place-based [243]. placed [667]. Places [14, 15]. plaice [1343]. Plains [760]. plan [1286]. Plankton [807, 357, 1335]. planning [1177, 1331, 788]. plans [565]. Plastic [1032, 1099]. plasticity [541]. plate [715]. plate-sized [715]. platyhelminth [262]. plays [783]. Pleistocene [48]. Plenty [665]. plight [63]. Poeciliid [452]. point [394, 1287, 1092]. Points [71, 550, 821, 692, 939, 1102, 965, 1127, 1003]. poisoning [603]. Policies [366, 1105, 733]. Policy [1088, 1206, 1323, 945, 113, 577, 662, 214, 784, 510, 1171, 831, 374, 74, 401]. political [664]. pollock [1111, 730, 520]. pollution [1032, 5]. polychlorinated [430]. polymorphisms [881]. polypteriforms [38]. ‘Pons [1284]. pool [263]. pooled [912]. poor [745, 1066, 842, 947]. poorly [710, 986, 731]. Popper [216]. Population [108, 156, 772, 867, 621, 874, 706, 1067, 1117, 1321, 747, 525, 213, 1010, 1310, 943, 797, 1273, 840, 362, 837, 574, 317, 682, 555, 369, 349, 441, 1141, 1116]. Population-level [621, 874]. Populations [310, 701, 685, 177, 1143, 344, 1090, 112, 37, 643, 297, 1213, 837, 1307, 60, 1041, 59, 313, 1035, 843, 1123]. porpoises [1060]. portfolio [1180, 1072]. portfolios [923]. posed [324]. position [814]. positive [945, 1315, 307]. positively [688]. Possible [345, 535, 234]. post [618, 528, 860]. post-larval [528]. post-release [618, 860]. potadromous [856]. Potamotrygoninae [1282]. Potential [695, 685, 900, 997, 1047, 624, 458, 774, 1226, 1167, 612]. potentials [572]. poverty [382]. pp [1241, 1324, 606, 1255, 64, 1240]. pp. [1256, 1296]. Practical [144, 125, 807, 571]. Practice [238, 629, 76, 505, 1132, 1125, 991, 978, 763]. practices [626, 936, 608]. practitioners [1245]. prawns [617]. Pre [906, 1155, 51]. Pre-Columbian [906]. pre-industrial [1155]. pre-settlement [51]. precautionary [650, 291, 1274, 1102]. Predation [1112, 774, 1041, 539, 974, 4]. predator [499, 638, 963, 130, 774]. Predators [443, 279, 1341, 939, 1004]. predatory [111, 919, 860]. predict [1111, 1064]. predicted [358, 989, 731]. Predicting [983, 933, 932, 508, 835]. prediction [1229, 545]. predictors [537, 1058]. Preference [1311]. preferences [1281, 713, 137, 1093]. Preliminary [717]. prescription [842]. presence [885]. Press [1256, 869, 520, 1240]. pressure [1005, 1279, 1045, 1189, 426]. pressures [871, 151]. prey [131, 919, 974]. Price [869, 606, 1259]. prices [1163, 237]. Primary [417, 524]. primer [1240]. Principle [6]. Principles [156, 187, 677, 1165]. Prints [898]. Prionace [747]. priori [295]. Priorities [1314, 1138, 491]. Prioritization [587, 626]. prioritizing [60]. priority [1301]. Pristiophoridae [996]. private [1184]. probabilistic [512, 788]. Probability [1344]. Probiotics [1241]. Problem [243, 89, 604, 912, 383, 739, 1238]. problems [258]. procedure [957]. Procedures [192]. Process [239, 398, 643]. process-based [643]. processes [1236, 779, 52, 599]. Processing [172]. processor [392]. produce [999]. production [1203, 716, 761, 993, 992, 1112, 551, 584, 723, 676, 221, 888, 833, 1069, 827].

Productivity [124, 1321, 699, 1087, 847, 1136, 892, 1334, 1154, 524]. products [727, 151, 457, 1261]. professional [734]. Professionals [487]. Professor [1131]. Profile [303]. profit [964]. profitability [583]. profits [320]. prognosis [78, 59]. Programme [21]. programming [505]. programs [1337]. Progress [1256, 1286, 785, 1036, 1023, 198, 1186, 939, 477, 408, 927, 1053]. project [861, 583, 859]. Projecting [244, 339, 1336, 1207]. projections [1265]. promiscuous [8]. promises [413, 1060]. promising [1327, 554]. promote [612]. promoting [656, 859]. pronounced [1334, 1309]. properties [118, 1155]. proposal [189]. proposals [1216]. proposed [696]. Propositions [1339]. Prospects [157, 1260]. protect [1060]. protected [98, 1251, 975, 446, 669, 844, 6, 552, 428, 9, 702]. Protecting [101, 562]. Protection [1000, 469]. Protein [343]. protocols [257]. Protopterus [415]. Provenance [623]. provide [1143, 966, 1005, 766]. provision [417]. proxies [1140, 1136]. proximate [628]. proximity [1322, 1145]. proxy [624]. Pseudorasbora [388]. public [400, 10]. Publication [191, 43]. Publishing [1296, 1184]. puffins [1111]. Puget [892]. Purposes [191]. purse [1294, 714]. purse-seine [1294]. Putting [1222]. puzzle [1170].

quahog [863]. Qualitative [342, 447, 1347]. Quality [57, 66, 703, 879, 646]. quantifications [1103]. quantify [749]. Quantifying [849, 1126, 670, 1007, 296, 1048, 953]. Quantitative [53, 253, 792, 1225, 412, 447, 904, 467, 842]. quantity [879]. quarter [1208]. question [1044]. questions [655, 720, 1053, 1267]. Quota [1305, 517, 597]. quotas [322, 354, 466].

R [1284, 1256, 216, 606]. R. [326]. rabbitfishes [178]. race [517]. Racial [1139]. radiation [186, 922, 50]. Radway [290]. rails [1286]. Rajidae [846]. RAM [463]. Ranching [211, 21, 776]. Randall [183]. random [880]. range [917, 168]. ranges [340]. Rapid [954, 1057, 186, 887, 905]. Rapidly [1040]. rarely [599]. rate [769, 932, 1127]. Rates [893, 757, 1028, 910, 1048, 1093, 602]. ratio [1146]. ray [614, 755, 642]. Rays [578, 496, 1206, 564, 1008, 787, 817, 464, 1232, 844, 1193]. Re [1216, 1274, 949, 704]. re-distribution [949]. re-evaluation [704]. Re-imagining [1274]. Re-launch [1216]. reaction [885]. Real [103, 291, 582, 497]. Real-time [582]. realities [1022]. reality [697]. Realizing [900]. really [515]. rearing [881]. reassess [197]. Rebuilding [476, 372, 481, 847, 616]. Received [26, 167, 194]. Reception [226]. recognition [134, 130]. recognizing [1029]. Recommendations [925]. reconciliation [86]. Reconciling [95, 3, 1023, 1142, 934, 556]. reconstructed [1178, 743]. Reconstructing [745, 738, 765]. Reconstruction [863, 1057, 906, 6]. record [415]. records [546]. recovering [748]. recovery [1163, 581, 403, 1045, 1017, 943, 976, 717, 956]. Recreational [1049, 1341, 329, 165, 757, 250, 1339, 735, 1308, 500, 1277, 1195, 670, 928, 506, 1161, 1109, 1187, 921, 1226, 980, 895, 696, 756]. recruit [680, 965]. Recruitment [563, 71, 1309, 1236, 1067, 972, 1090, 795, 863, 779, 810, 585, 888, 1069, 932, 1123, 875, 1280, 862]. Redefining [1066]. Redemption [996]. redistribute [1329]. Redistribution [1035]. redoubles [502]. reduce [405, 564, 1279, 1323, 1045, 819, 582]. reduced [1125, 974, 875]. Reducing [249, 219, 351]. reduction [727, 564, 1318, 320, 933, 1328]. Reef [566, 271, 327, 792, 1322, 995, 544, 806, 989, 491, 715, 62, 773, 799, 961, 1075, 489, 51, 803, 305, 645, 1052, 462, 660, 204, 890, 909, 178, 586, 1156, 327]. reef-associated [989]. reefs [1308, 306, 547, 332, 605, 6]. Reference [939, 71, 550, 821, 692, 1287, 1102, 617, 1092, 87, 965, 1127, 1003]. references [648]. refinement [560]. refining [943, 779]. reflect [772, 779, 771]. reflected [576]. reflex [358]. Refocusing [943]. Reforming [374]. Reframing [1238]. regarding [662]. regime [77, 1322, 230, 1309, 1250]. region [708, 732]. Regional [1015, 978, 529, 762, 1186]. Regions [155, 1006]. register [457]. registers [1182]. regulated [465, 1128]. Regulation [972, 703, 455, 401, 1158]. regulations [500, 562]. regulatory [871, 877]. Rehabilitation [56]. reintroduced [447]. reinvigorate [10]. relate [896]. Related [672, 948, 118, 690, 34, 859]. relation [626, 551]. relations [383]. relationship [94, 1042]. Relationships [435, 972, 1002]. Relative [462, 1290, 840, 237, 552, 598, 730, 586]. Release [165, 618, 285, 472, 860]. released [539]. Relevance [1245, 1062]. relevant [649]. reliable [970]. Reliance [881]. religion [899]. relocation [446]. rely [1188]. Remaining [655]. remarkable [1150, 654]. remarks [92]. remote [1099, 1264]. repercussions [321]. reply [269]. Report [762, 945]. Reporting [268, 1186, 985]. Representing [1019, 1243]. Reproduction [434, 990, 815, 5]. Reproductive [680, 362, 1280, 499, 813, 404, 624, 1025, 551, 654, 178, 341, 485]. required [861]. requirement [417]. requirements [369, 1107]. requires [1074, 832]. requiring [784]. rescue [976]. Research [104, 216, 1095, 527, 1339, 1138, 736, 745, 673, 1129, 491, 1332, 1247, 1161, 991, 1020, 1314, 1083, 1100, 873, 925, 1171, 364]. research-vessel [673]. reserves [561, 641, 113, 695, 522]. Reservoir [121]. reservoirs [1026, 1009, 589]. resilience [1214, 678, 1288, 1221, 692, 680, 503, 1132, 248, 726, 1280]. resilient [1074, 1030]. resolution [10]. Resource [101, 487, 1203, 440, 915, 509]. Resources [108, 73, 299, 191, 223, 418, 151, 215, 809, 1212, 1218, 316]. respeciation [91]. Respiratory [618]. Response [326, 789, 120, 871, 652, 1321, 1318, 1283, 1035, 979]. responses [703, 1122, 828, 130, 429, 974]. responsibilities [1110]. responsibility [796]. Responsible [126, 173, 927, 169, 1182]. Responsive [957]. Restoration [21, 806, 1166, 179]. Restoring [202]. Restrepo [1284]. restrictions [1108, 613]. Results [1319, 739, 764, 401]. results-based [764]. Rethinking [562]. retrospective [835]. return [1316]. reveal [1022, 1306, 1070, 1334, 974, 540, 602]. revealed [887, 935, 1347]. reveals [1343, 1196]. reversal [474]. Review [1241, 1256, 1324, 1296, 435, 869, 260, 520, 606, 1228, 1255, 141, 1078, 1184, 64, 1236, 721, 136, 1297, 158, 1225, 340, 394, 38, 267, 1230, 1073, 621, 1332, 718, 776, 1264, 725, 634, 574, 1307, 1240, 742, 746, 415, 921, 910, 598, 980, 390, 1158]. reviewed [47]. Reviews [25, 30, 36, 105, 20, 146, 702]. revisited [1333, 815, 627, 256]. Revisiting [1229, 601, 498, 1266, 1069]. Reviving [1226]. revolution [468]. rheophilic [1107]. Rhodeus [255]. rich [1220, 963]. richness [394]. right [1343, 293, 104]. Rights [437, 466, 1261, 800, 1110, 689]. Rights-based [437, 800, 689]. Risk [570, 1220, 389, 499, 934, 131, 1223, 631, 171, 1027, 988, 1066, 817, 1179, 197, 1178, 1082, 1193]. risk-based [934]. risks [889, 1077, 324, 17, 1072, 683, 478]. River [109, 202, 1128, 770, 1166]. riverine [534]. Rivers [658, 1032, 465, 537, 1141, 1107, 1128]. Robert [558]. robust [518, 117, 624]. Rockfish [164, 164]. Role [279, 796, 1322, 815, 908, 472, 353, 1337, 297, 783, 1326, 323, 899, 32, 1174, 132, 555, 539, 950, 1295, 543]. roles [1008]. ROM [46]. Rome [927]. room [919]. rule [631, 627, 710]. Rules [715, 550, 1268, 484, 531, 1221, 421, 1184]. run [1246]. rural [248].

S [1241, 1256, 869, 183]. safe [601, 735]. safety [640]. saints [637]. salar [480, 967]. Salmo [881, 480, 902, 967]. Salmon [461, 868, 225, 664, 281, 278, 82, 56, 155, 1253, 881, 1246, 736, 1316, 1089, 720, 988, 1337, 1326, 1260, 480, 892, 1235, 1307, 777, 848, 967, 1272, 274, 1035, 630, 1128, 1013, 71]. Salmonid [45, 35, 108]. Salmonidae [902, 967, 1013]. Salmonids [711, 619, 953, 465, 684]. Salvelinus [881, 653]. same [41]. samples [317]. sampling [394, 257, 880, 19, 999]. sand [1008, 950]. sardine [1012, 1170, 1205, 1069]. Sardinella [19]. Sardinops [1205]. satellite [791, 935]. satisfaction [1049]. save [303]. Saving [948]. Sawshark [996]. sawsharks [996]. say [830]. Scale [1327, 1081, 923, 266, 1104, 177, 1188, 1143, 938, 417, 14, 600, 761, 668, 573, 1175, 1151, 690, 1166, 580, 1070, 722, 1118, 1009, 1347, 1238, 1114, 1068, 382, 991, 675, 1314, 910, 1160, 15, 1083, 969, 876, 1082, 1093, 526, 1156]. scales [490]. scaling [550, 458]. scapegoats [326]. Scavenger [924]. scenario [967, 858, 1335]. scenarios [339, 1279, 1336, 583, 1014]. sceptical [794]. scheme [499]. Schmidt [740]. ScholarOne [398]. scholarship [1225]. school [575]. schooling [345]. Schroeder [140]. Science [356, 780, 147, 289, 945, 66, 900, 1162, 235, 285, 916, 477, 563, 943, 455, 1332, 368, 160, 785, 367, 1053, 308, 986, 763, 189, 925, 1055, 786, 886, 1295, 740]. science-based [916, 1055]. science-industry [925]. Science-Policy [945]. Sciences [44, 1251]. Scientific [487, 324, 616, 362, 766]. Scientist [658, 959, 368]. Scientists [156, 144]. scope [66]. scorecard [718]. SE [222]. Sea [419, 450, 673, 620, 193, 473, 211, 21, 420, 997, 1139, 1342, 566, 1005, 403, 219, 361, 191, 549, 757, 776, 160, 1232, 546, 1095, 1302, 843, 665, 1281, 77, 576, 440, 863, 1271, 1012, 1170, 1119, 817, 937, 1174, 682, 924, 730, 835, 625]. seabed [626, 1006, 936]. Seabird [1021, 249, 351, 187, 1111].

Seabird-induced [1021]. seabirds [924]. seafloor [1050]. Seafood [459, 1163, 796, 678, 151, 651, 698, 970, 1045, 1197, 635, 237, 436, 683, 623, 1016, 647]. seagrass [768]. Seamounts [451, 1145, 656]. search [1071]. Seas [732, 302, 678, 1206, 791, 1289, 593, 1248, 759, 1335, 625, 1087]. Seascape [313, 568]. seascapes [962, 1064]. seasonal [1294]. Seatrout [142]. Second [868, 211, 141, 128]. secondary [41]. Section [379]. sector [705]. security [1143, 908, 640, 553]. Sedative [337]. Seeing [1020]. seine [1294, 714]. seismic [874]. Selection [751, 919, 610, 51, 315, 441, 161, 532]. selective [992]. selectively [1341]. Selectivity [715, 955, 405, 62, 612]. semi [534]. semi-arid [534]. Senses [226]. sensing [1264]. sensitive [1003]. sensitivities [915]. sensitivity [570, 1073]. Sensory [172]. Sensuous [302]. September [414, 164]. sequences [245]. sequential [195]. Serial [420, 566]. Series [108, 142, 73, 299, 271, 183, 252, 223, 216, 71, 1344, 1003]. serpent [1005]. services [636, 1173, 1128]. sets [1057]. settlement [51]. Seussian [796]. seventeen [462]. several [877]. severe [1194]. severity [828]. Sex [344, 611, 430, 715, 195, 474]. sex-changing [715]. shad [829]. shallow [574]. shapes [441, 814]. shaping [1175]. share [454]. Sharifuzzaman [1241]. shark [934, 1297, 747, 1206, 614, 353, 755, 1261, 937, 278, 642, 1152, 1127]. Sharks [844, 387, 578, 934, 1297, 821, 496, 564, 1027, 787, 1286, 191, 817, 464, 402, 1232, 860, 717, 1193]. sharpening [163]. Sheffield [44]. shell [863]. Shellfish [46, 192, 861, 603, 60]. shellfisheries [600]. shelters [222]. shift [544, 680]. shifted [1281]. Shifting [968, 905, 944, 1022]. Shifts [871, 77, 544, 982, 312, 917, 230, 1181, 431, 1309, 730, 835, 1250]. Shipwreck [838]. shoal [135]. shoaling [891]. shoals [119, 7]. Shooting [714]. Shorelines [301]. Short [1038, 641]. Short-lived [1038]. short-term [641]. shortfalls [1007]. Should [666, 530]. shoulders [567]. shout [1112]. show [1113]. shrimp [993, 1283, 704]. Shrinking [350]. Shulman [606]. Shut [1112]. Sidney [959]. Sifting [1057]. sightings [1005]. signals [651, 737, 1250]. significance [768, 333]. Silurus [340]. silverfish [453]. simple [836, 153, 503, 842, 586]. simplify [534]. simplifying [549]. Simulating [1303]. simulation [938, 1319, 368, 214]. simulation-based [368]. single [679, 32]. single-species [679, 32]. sinners [637]. site [263]. site-attached [263]. situation [1030]. Situations [254]. six [722]. Sixty [406]. Size [550, 370, 495, 983, 1246, 813, 610, 782, 992, 159, 297, 607, 840, 739, 837, 1118, 1200, 784, 850, 1048, 894, 1146, 730, 428, 1116]. size-based [850, 1048]. size-dependent [739, 894]. size-selection [610]. size-selective [992]. size-spectrum [1146]. size-structure [784, 730]. size-structured [607]. sized [715]. Skating [846]. skeleton [38]. Skeptical [103]. skill [835]. skills [129]. Skin [209, 371]. Skipped [188]. slain [312]. slippery [1043]. slope [1043, 1146]. slots [948, 562]. Slow [1109, 1310]. Small [1104, 14, 15, 526, 1081, 923, 266, 177, 404, 1188, 972, 1143, 417, 394, 600, 761, 573, 1175, 1151, 1189, 580, 1243, 722, 1238, 1114, 382, 991, 675, 1314, 1160, 1083, 1014, 1084, 876, 1082, 1156, 969]. Small-scale [1104, 14, 15, 526, 923, 266, 177, 1188, 1143, 417, 600, 761, 573, 1175, 1151, 580, 722, 1238, 1114, 382, 991, 675, 1314, 1160, 1083, 876, 1082, 1156, 969]. snapper [847, 602]. snapshot [1173]. snow [1119]. Social [72, 923, 1268, 136, 135, 190, 464, 548, 1126, 796, 744, 687, 1022, 1180, 1109, 915, 195, 1303, 429, 858, 873, 980, 802, 693, 964, 1267]. social-ecological [1126, 1022, 1303, 858]. societal [1086]. Society [224, 165, 191, 146, 123, 166, 43]. socio [1126, 887, 734]. socio-economic [1126, 887, 734]. socioecological [1308]. solar [922, 50]. Solutions [326, 1284, 836, 738, 640, 78, 258, 641, 1276]. some [648, 544, 1262, 234]. Something [966]. sonar [235, 1191]. Sons [1324]. Sound [495, 864, 1112, 723, 892]. source [399, 843, 696]. Sources [202, 1155, 1347, 1093]. south [1012, 1170, 613, 717, 872, 906, 1124, 402]. south-east [717]. south-eastern [1012, 1170, 613]. Southeast [787, 898]. southern [1106, 1076, 321, 373, 749]. Southward [949]. southwest [427]. sp [950]. space [1208, 940, 1059, 1048, 735]. Spain [1185]. Spatial [133, 594, 418, 1177, 1120, 1319, 982, 582, 800, 27, 568, 1068, 915, 848, 682, 1303, 1301, 313, 161, 903, 1331]. Spatially [214, 1179, 1117, 1010]. Spatio [405, 963, 903]. Spatio-temporal [405, 963, 903]. spatiotemporal [938]. spawner [1246, 680, 8, 833, 965]. spawner-recruit [680]. spawning [881, 283, 1121, 667, 563, 624, 840, 1118, 1200, 188, 684, 1111, 22, 875, 804, 341]. Special [116, 379, 1216, 104, 191, 43]. speciation [186, 52, 84, 92, 91, 87]. Species [1208, 86, 200, 191, 917, 92, 89, 581, 1122, 679, 1188, 938, 1120, 772, 394, 912, 795, 970, 544, 916, 993, 498, 490, 1207, 992, 453, 607, 263, 879, 32, 88, 1307, 530, 617, 891, 1301, 85, 750, 463, 84, 197, 846, 1048, 91, 541, 87]. species-specific [846]. specific [1188, 1089, 846]. spectrum [1146]. speed [1304]. Sperm [598]. spillover [975]. spined [346]. spiny [528]. Sportfisheries [629]. spot [596]. Spotted [142]. spp [262, 1124, 846]. spring [22, 875]. spring-spawning [22, 875]. Springer [1241, 216, 1255]. Squaring [556]. squid [393]. Squids [1270, 460]. stability [1195, 774, 441]. Stable [920, 1157]. stage [528, 811]. Stages [147, 283, 1336, 1290, 1307, 1266, 5, 894]. stakeholder [571, 613, 1165]. stakeholder-based [1165]. Standard [1200]. standardized [61]. standards [696]. start [1200]. State [103, 1257, 1094, 1048, 42]. state-managed [1257]. state-space [1048]. states [1182, 1257, 582, 1257, 717, 1256, 43]. stationary [1002, 1205, 1123]. statistical [1176]. Statistics [144, 151, 947]. Status [720, 207, 527, 706, 826, 1088, 688, 336, 772, 821, 593, 445, 1244, 1065, 1006, 237, 659, 463, 360, 28, 104]. steady [75, 533]. Steepness [1043, 362]. Steller [546]. step [873]. Steps [927]. Stereolepis [1095]. stereotypes [1187]. stewards [1226]. Stewardship [1304]. Stickleback [346, 406]. Stimuli [226]. stingrays [1282]. Stochastic [1090, 676]. Stock [821, 463, 211, 581, 1210, 244, 76, 594, 972, 938, 117, 336, 911, 170, 512, 593, 1089, 118, 1319, 1244, 847, 1142, 889, 1305, 624, 776, 1136, 1259, 1070, 1243, 1273, 1334, 32, 494, 1072, 1137, 39, 1190, 49, 659, 1344, 1147, 843, 1111, 1054, 707, 1003, 71]. stock-specific [1089]. stocking [1337]. stocks [1210, 972, 1233, 1317, 573, 403, 372, 1287, 982, 304, 1065, 1189, 616, 957, 1224, 1309, 1125, 237, 473, 1335, 18, 625, 1176, 862]. Story [520, 869, 1228]. storylines [967]. straddling [982]. strandings [1152]. strategic [362]. Strategies [202, 1208, 881, 404, 1223, 531, 1277, 453, 1025, 922, 1243, 1091, 551, 902, 822, 1038]. Strategy [423, 813, 1151, 368, 608, 1084]. stratified [1010]. Stratifying [605]. Stream [107, 123, 290, 447, 684, 287]. streams [394, 684]. Strength [1154]. Strengthening [916, 945]. strengths [1264, 642]. stress [721, 358, 815, 646]. stressor [811]. stressors [828, 1302, 1076, 984]. Strong [688]. Structural [280]. structure [550, 1117, 344, 747, 1022, 135, 989, 1070, 682, 784, 349, 730, 1267]. structured [1067, 1090, 607, 1176]. structures [1203, 592]. student [558]. studies [1298, 89, 818, 1231, 600, 631, 1249, 304, 755, 505, 459, 642, 541]. Study [251, 283, 1188, 861, 995, 151, 531, 1119, 985, 879, 1028, 516, 465, 381, 730, 350]. studying [27, 536]. Sturgeon [21, 207]. sturgeons [21]. styles [595]. sub [1120]. sub-arctic [1120]. subject [543]. subjective [670]. sublethal [621, 984]. submission [398]. Subpolar [1011]. Subsidies [752, 93]. subsistence [1187]. substrates [592]. subtropical [906]. succeed [516]. success [1304, 694, 1280, 341]. successful [401]. Sudden [875]. suitable [1265]. summary [98]. Summit [481]. superensemble [706]. supplies [244]. supply [678]. support [652, 1231, 962, 206, 635, 10, 924, 788]. suppositions [474]. surf [746]. surplus [676]. Surprises [1298, 1249]. survey [1162, 49, 1164]. surveys [257, 880, 999, 874, 673]. Survival [82, 892, 953, 1266, 5, 888, 1069, 630, 1013]. survive [1114]. surviving [205]. susceptibility [1282]. Sussex [606]. Sustainability [945, 95, 796, 3, 1023, 476, 1159, 1261, 1180, 799, 489, 635, 93, 1086, 1092, 859, 612, 693]. Sustainable [867, 366, 841, 1056, 70, 1160, 893, 1268, 629, 1330, 1074, 727, 417, 1233, 698, 1277, 847, 1017, 587, 992, 94, 961, 1198, 1053, 510, 802, 521, 683, 788, 647, 1037]. Sustaining [12, 155]. swamps [1106]. Sweden [465]. Swedish [576]. sweepstakes [1280]. sweet [596]. swimming [150, 75, 1085, 575, 229, 533]. Switzerland [1241, 1255]. swordfish [1155, 28]. symbioses [704]. symbol [805]. Symmetry [182]. Symposium [224, 165, 146, 164, 123, 945, 166, 107]. synchrony [682]. Syndemics [1302]. syndrome [685]. synergies [35]. synthesis [1138, 1050, 679, 1308, 818, 828, 1103, 482, 1089, 1292, 430, 296, 991, 980, 950, 909, 977, 1013, 760, 743, 1078]. Synthesizing [1122, 479]. synthetic [1248]. system [479, 508, 345, 34, 505]. systematic [1225, 1332, 776, 1264, 1307, 598, 702]. systematically [1200]. Systematics [54, 48]. systemic [755]. Systems [122, 280, 239, 202, 70, 1126, 1104, 245, 1203, 534, 792, 831, 865, 652, 1308, 631, 1288, 243, 1159, 963, 1145, 680, 1303, 858, 1100, 802, 384, 572, 1237, 1274, 422].

T [223, 354]. tables [520, 64]. taboos [425]. TAC [401]. Tackling [931, 1180]. tactical [511, 599]. tactically [1231]. tagging [935, 1028, 369, 168]. tailored [832]. take [561, 263, 428]. tale [625]. Tales [302]. Tanzania [382]. target [772, 649, 456]. targets [454, 784, 1234, 522]. tarpon [935]. tarpons [527]. Taste [137]. taxa [632]. Taxonomic [501, 882]. Taxonomy [807, 1327, 811]. technique [554]. Techniques [337, 192, 1200]. technological [258, 949]. Technologies [1000, 1133]. Technology [110, 29, 865, 319, 97, 574]. telemetry [150, 1070]. teleost [983, 133, 1085, 654, 548]. Teleostean [435, 505]. Teleostei [234]. teleosts [371, 5, 282, 474]. tell [237]. tells [736]. temperate [1236, 1121, 1113, 989, 490, 1347, 774]. Temperature [1270, 460, 1192, 472, 370, 1326, 1290, 905, 897, 1011, 730, 783]. temperatures [983]. Template [461]. Temporal [1103, 405, 963, 903]. Temporary [1123, 883]. Ten [927]. Tenualosa [829]. Tenure [425]. term [600, 118, 641, 967, 1250, 875, 1171]. terms [779]. terrain [1050]. terrestrial [234, 778]. territories [1257, 462]. testimonies [540]. Testing [159, 1320, 938]. tests [1073]. their [727, 336, 828, 1103, 1182, 908, 806, 319, 200, 657, 1311, 296, 530, 862, 279]. them [666, 393]. Theoretical [253, 679, 1078]. Theories [449, 991]. Theory [238, 439, 1132, 436, 1229]. there [1023, 68, 866]. Thermal [1307, 1275, 1272]. Thermodynamics [182]. thermoregulation [1214]. thin [846]. think [1010]. Third [175, 140, 251, 1086]. third-party [1086]. Thirty [871, 336, 700]. those [670, 1105]. Threat [838, 169, 1112]. threatens [477]. Threats [817, 1104, 872, 760, 478]. Three [1181, 346, 153, 1159, 128, 1260, 369]. Three-dimensional [1181]. three-second [128]. Three-spined [346]. throughout [22]. thumb [715]. Thunnus [262, 1306]. Tidal [1322]. tier [631]. tier-based [631]. ties [923]. tilapia [617]. Time [327, 1208, 1163, 1231, 1343, 582, 848, 676, 1344, 1003]. time-series [1003]. timidity [685]. timing [1316, 751, 341]. Tinbergen [406]. tips [1009]. today [448]. together [427]. tolerance [1307]. tomorrow [448]. Too [854, 1135, 170, 393]. Tool [442, 1327, 652, 1317, 726, 90, 903, 536]. Toolbox [253]. toolkit [1199]. Tools [407, 412, 263, 1161, 800, 574, 318, 694]. toothed [1082]. toothfish [650, 495]. Top [279, 499, 584]. top-down [584]. topic [1062, 904, 786]. Topics [146]. total [1246, 517, 1248, 840, 833, 199]. towed [236]. toxins [603]. traceability [318]. Tracing [1147]. track [1197, 454]. Tracking [1059, 791, 1073]. Trade [748, 424, 822, 42, 934, 1108, 151, 651, 713, 457, 1261, 732, 715, 635, 507, 915, 1303, 832, 543, 1173, 964, 1054]. trade-off [915]. Trade-offs [748, 424, 822, 934, 1108, 713, 507, 1303, 832, 964, 1054]. trader [392]. traditional [95, 1077, 263]. Traditions [14, 15]. trafficking [981]. Tragedy [400]. Trait [1306, 904, 900, 1333, 1026, 790, 751]. Trait-based [1306, 904, 900]. traits [993, 708, 1140, 35, 848, 850, 598]. transboundary [1289, 982]. transects [536]. transfer [135]. transferable [322, 517, 354]. Transferrable [466]. Transform [678, 1020]. Transformations [1104]. transforming [1040]. Transgenic [17]. transition [866]. transitions [1071]. Translating [778]. translational [873]. Transmission [737]. transparency [1206]. transport [283, 646]. Trap [62, 1294]. traps [559]. trauma [994]. travellers [231]. Trawl [1006, 249, 351, 936, 933]. trawling [78, 626, 540, 1173, 979]. Treatise [913]. Trees [1295]. trend [706]. Trends [1246, 644, 491, 1253, 495, 772, 1189, 892, 904, 1235, 489, 848, 1263, 207, 1095, 1344, 359, 862]. tribute [558]. Tricky [1081]. Tridacna [622]. triple [748]. triple-bottom-line [748]. triploid [284]. Trophic [323, 1281, 881, 968, 989, 814, 974]. trophodynamics [850]. Tropical [1192, 829, 1322, 306, 1313, 962, 961, 1075, 367, 949, 113, 714, 1156]. Trouble [204]. Troubled [817]. Trout [461, 225, 558, 56, 881, 1346, 902, 868]. trout-dominated [1346]. trust [1167]. trutta [902]. tshawytscha [1013]. Tumours [34]. Tuna [1135, 964, 1281, 499, 1206, 1313, 213, 335, 1294, 1323, 1306, 1186, 616, 762, 990, 1263, 1133, 659, 1179, 949, 1040, 978, 1110, 714, 183]. tunas [262, 1284, 1276]. Turbine [856]. turbines [1278]. turtle [419, 745, 219, 361]. turtles [191, 843]. two [912, 795, 1259, 664, 935, 739, 35, 459, 625, 1020]. Two-Eyed [1020]. type [618, 779, 814, 598]. Types [1209, 369]. typology [694].

ubiquitously [1171]. ugly [66, 312]. UK [1256, 687, 785]. ultimate [891]. ultraviolet [922]. unaccounted [559]. unanswered [720]. uncertain [594]. uncertainties [889, 1077]. Uncertainty [202, 849, 1339, 1210, 399, 275, 324, 1243, 494, 1102, 39, 1048]. Uncovering [1008]. under-estimate [1200]. under-utilized [970]. underappreciated [998, 539]. underestimation [755]. underlying [779, 449]. undersized [671]. understand [828]. Understanding [151, 1042, 1086, 572, 1128, 388, 1194, 491, 680, 882, 588, 59, 1069, 392]. understudied [1062, 539]. underwater [1164, 885]. underway [866]. Undulatory [229, 533]. unexploited [997]. unfinished [1106]. unfulfilled [612]. unifying [1121]. Unintended [507, 1234, 549]. Union [1037, 615, 634, 1060]. Unique [147, 1226]. United [1256, 43, 1286, 582, 1257, 717]. units [349]. universal [1197]. University [1256, 520, 1240, 164]. unknown [402]. unmanaged [402, 1257]. unmonitored [402]. unreported [96, 681]. unselective [1211]. unsustainable [733]. Untangling [1180]. Untold [520]. Unveiling [1211]. unwatchable [258]. Updating [1051]. upon [1321, 93]. upwelling [404]. USA [164, 215, 1165]. Usage [1171]. USD [869, 520]. use [442, 1177, 1231, 497, 883, 426, 1305, 800, 574, 318, 560, 1272, 890, 331, 710, 1328]. used [1221, 1171]. Using [534, 583, 951, 985, 492, 1259, 1218, 197, 873, 843, 734, 244, 264, 821, 1047, 1278, 1231, 358, 446, 700, 37, 1207, 295, 1145, 1161, 1157, 904, 1174, 1301, 511, 694, 1328, 1055, 932, 1093, 822, 1191, 1127, 647, 999]. utility [586]. utilized [970]. UV [50].

V [1284]. vague [1171]. validated [710, 986]. validation [1073]. valuable [204, 686]. Value [1340, 562, 920, 317, 1095, 231, 552, 986, 485]. Vänern [465]. Vanishing [944]. vannamei [1283]. Vanuatu [1345]. Variability [202, 774, 483, 208, 1170, 292, 330, 897, 278, 19, 1069, 1250]. Variable [483]. variance [880]. variances [49]. Variation [201, 22, 880, 315]. varied [974]. varies [1290, 1021]. Varieties [83]. Varley [348]. vary [770]. Vectors [750]. vegetated [920]. velocity [1113]. venom [333]. versus [769]. vertebrate [653]. Vertical [1203, 426, 509]. vessel [791, 673, 769, 1064]. vessels [570]. Victoria [516, 382]. Video [482, 999]. view [982, 1009]. Views [1075]. violence [931, 1245]. vision [1164, 1328]. visitors [656]. Visualisation [61]. Vol [106, 69, 183, 44]. Volume [148, 184, 443, 271, 191, 252, 154, 217]. Voluntary [500]. Voracious [340]. vs [586]. Vulnerability [334, 690, 448, 112, 403, 705, 1179, 1076]. vulnerable [931, 725, 204].

W [183]. Waking [317]. Wales [228]. Walleye [730]. Wang [1324]. Warm [1272, 1113]. warmer [1113]. warming [1214, 1113, 687, 643, 1039]. warning [1250]. warranted [245]. warranting [554]. wars [664]. was [232]. Washington [869, 164]. waste [924]. Wasted [215]. watch [258]. Water [460, 1214, 1129, 775, 1346, 538, 1052, 1291, 646, 205, 1064, 883]. Water-Breathing [460]. Waters [43, 940, 818, 1113, 817, 1037, 1059, 574, 1011]. Watershed [260]. Watkins [1296]. Ways [1000, 560]. weakens [651]. weaknesses [642]. Wealth [516]. web [398, 440, 386, 1174, 979]. web-based [398]. webs [1057, 1346]. weight [1318, 1042]. Welfare [347, 250, 1293, 269, 640, 1311, 516]. well [931, 667, 841, 1056, 413, 734]. well-being [931, 841, 1056, 734]. well-placed [667]. wellbeing [526]. Wellby [64]. were [427, 684]. West [1208, 606, 843, 255, 1249, 1173, 1013]. WestCoast [1298]. Western [191, 459, 19, 1265, 1135, 1263, 1178]. wetlands [834]. whale [457]. Whales [104, 293, 373, 1082, 749, 391]. whaling [749]. Where [1010, 427, 1232, 1212, 667, 942]. Which [597, 1198, 560]. while [590]. Whirling [146]. whistleblower [1150]. white [520, 937, 860]. whitefish [693]. Whittles [1296]. who [670]. wholesale [1163]. whose [248]. wicked [1238]. wide [1013]. widespread [1309, 890]. Wild [102, 849, 685, 1330, 1279, 720, 621, 1337, 480, 1340, 555, 598]. wild-caught [849, 1340]. wild-type [598]. Wiley [1324, 606]. will [372, 939]. winners [609, 1173]. wisdom [298]. within [1117, 584, 974, 714]. without [972, 63]. witness [540]. Women [1175, 1135, 580, 1340]. women-dominated [1340]. Work [366, 1150]. World [103, 711, 246, 796, 636, 1219, 343, 627, 1056, 492, 1105, 709, 88, 196, 1039, 1040, 964, 232, 481, 259, 869]. worlds [1259]. Worldwide [104, 260, 360, 708, 634, 585, 475, 932]. worms [686]. worse [1113].

Xiphias [1155]. Xiphiidae [1155]. xiv [1296, 64].

Yangtze [1166, 954]. year [111]. Years [356, 406, 871, 336, 1023, 394, 41, 745, 863, 728, 700, 782, 498, 1247, 892, 567]. yellow [303]. yield [667, 847, 992, 94, 961, 1153, 395, 502, 841]. yields [770, 428]. young [394]. you’re [1010].

Zealand [968, 634]. zone [976]. zones [263, 746, 1064]. zooplankton [1146].

References

Anonymous:2000:FL

[1] Anonymous. Fish on the line. Fish and Fisheries, 1(1):1, March 2000. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2000:LLE

[2] Anonymous. The Larkin Lecture — Editor’s note. Fish and Fisheries, 1 (1):2, March 2000. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cochrane:2000:RSE

[3] Kevern L. Cochrane. Reconciling sustainability, economic efficiency and equity in fisheries: the one that got away? Fish and Fisheries, 1(1): 3–21, March 2000. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Whipple:2000:MPF

[4] Stuart J. Whipple, Jason S. Link, Lance P. Garrison, and Michael J. Fogarty. Models of predation and fishing mortality in aquatic ecosystems. Fish and Fisheries, 1(1):22–40, March 2000. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Rolland:2000:EEP

[5] Rosalind Murl Rolland. Ecoepidemiology of the effects of pollution on reproduction and survival of early life stages in teleosts. Fish and Fisheries, 1(1):41–72, March 2000. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Pitcher:2000:PPH

[6] Tony J. Pitcher and William Seaman, Jr. Petrarch’s Principle: how protected human-made reefs can help the reconstruction of fisheries and marine ecosystems. Fish and Fisheries, 1(1):73–81, March 2000. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Krause:2000:LFS

[7] J. Krause, D. Hoare, S. Krause, C. K. Hemelrijk, and D. I. Rubenstein. Leadership in fish shoals. Fish and Fisheries, 1(1):82–89, March 2000. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Nordeide:2000:CLP

[8] Nordeide and Folstad. Is cod lekking or a promiscuous group spawner? Fish and Fisheries, 1(1):90–93, March 2000. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Watson:2000:DMB

[9] Watson, Alder, and Walters. A dynamic mass-balance model for marine protected areas. Fish and Fisheries, 1(1):94–98, March 2000. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Pitcher:2000:EGC

[10] Tony J. Pitcher. Ecosystem goals can reinvigorate fisheries management, help dispute resolution and encourage public support. Fish and Fisheries, 1(1):99–103, March 2000. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kennedy:2000:PNA

[11] Kennedy. Parasites of North American freshwater fishes. Fish and Fisheries, 1(1):104–105, March 2000. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cochrane:2000:SMF

[12] Cochrane. Sustaining marine fisheries. Fish and Fisheries, 1(1):105– 107, March 2000. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Winfield:2000:CBC

[13] Ian J. Winfield. Carp: Biology and culture. Fish and Fisheries, 1(1): 107, March 2000. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Colgan:2000:FPF

[14] Patrick Colgan. Fishing places, fishing people: Traditions and issues in Canadian small-scale fisheries. Fish and Fisheries, 1(1):108–109, March 2000. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ruttan:2000:FPF

[15] Lore M. Ruttan. Fishing places, fishing people: Traditions and issues in Canadian small-scale fisheries. Fish and Fisheries, 1(1):109–110, March 2000. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Murray:2000:PME

[16] A. M. Murray. The Palaeozoic, Mesozoic and Early Cenozoic fishes of Africa. Fish and Fisheries, 1(2):111–145, June 2000. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Maclean:2000:TFE

[17] N. Maclean and R. J. Laight. Transgenic fish: an evaluation of benefits and risks. Fish and Fisheries, 1(2):146–172, June 2000. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Svaasand:2000:ECS

[18] T. Svåsand, T. S. Kristiansen, T. Pedersen, A. G. V. Salvanes, R. Engelsen, G. Nævdal, and M. Nødtvedt. The enhancement of cod stocks. Fish and Fisheries, 1(2):173–205, June 2000. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Samb:2000:VSA

[19] B. Samb and D. Pauly. On ‘variability’ as a sampling artefact: the case of Sardinella in north-western Africa. Fish and Fisheries, 1(2):206– 210, June 2000. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hollingworth:2000:BR

[20] Chuck Hollingworth. Book reviews. Fish and Fisheries, 1(2):211–214, June 2000. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Secor:2000:RSL

[21] D. H. Secor, V. Arefjev, A. Nikolaev, and A. Sharov. Restoration of sturgeons: lessons from the Caspian Sea Sturgeon Ranching Programme. Fish and Fisheries, 1(3):215–230, September 2000. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Toresen:2000:VAN

[22] R. Toresen and O. J. Østvedt. Variation in abundance of Norwegian spring-spawning herring (Clupea harengus, Clupeidae) throughout the 20th century and the influence of climatic fluctuations. Fish and Fisheries, 1(3):231–256, September 2000. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Johannes:2000:IFK

[23] R. E. Johannes, M. M. R. Freeman, and R. J. Hamilton. Ignore fishers’ knowledge and miss the boat. Fish and Fisheries, 1(3):257–271, September 2000. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2000:E

[24] Anonymous. Erratum. Fish and Fisheries, 1(3):272–273, September 2000. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2000:BRa

[25] Anonymous. Book reviews. Fish and Fisheries, 1(3):274–281, September 2000. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2000:BRb

[26] Anonymous. Books received. Fish and Fisheries, 1(3):282, September 2000. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Lucas:2000:MSS

[27] M. C. Lucas and E. Baras. Methods for studying spatial behaviour of freshwater fishes in the natural environment. Fish and Fisheries, 1(4): 283–316, December 2000. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ward:2000:BSS

[28] P. Ward, J. M. Porter, and S. Elscot. Broadbill swordfish: status of established fisheries and lessons for developing fisheries. Fish and Fisheries, 1(4):317–336, December 2000. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ackefors:2000:FCF

[29] H. E. G. Ackefors. Freshwater crayfish farming technology in the 1990s: a European and global perspective. Fish and Fisheries, 1(4):337–359, December 2000. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2000:BRc

[30] Anonymous. Book reviews. Fish and Fisheries, 1(4):360–364, December 2000. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Pitcher:2001:LLE

[31] Pitcher. The Larkin Lectures — Editor’s note. Fish and Fisheries, 2(1): 1, March 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Mace:2001:NRM

[32] Mace. A new role for MSY in single-species and ecosystem approaches to fisheries stock assessment and management. Fish and Fisheries, 2(1): 2–32, March 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Liermann:2001:DEM

[33] Liermann and Hilborn. Depensation: evidence, models and implications. Fish and Fisheries, 2(1):33–58, March 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Leatherland:2001:TRL

[34] Leatherland and Down. Tumours and related lesions of the endocrine system of bony and cartilaginous fishes. Fish and Fisheries, 2(1):59– 77, March 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

McDowall:2001:AHT

[35] McDowall. Anadromy and homing: two life-history traits with adaptive synergies in salmonid fishes? Fish and Fisheries, 2(1):78–85, March 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2001:BR

[36] Anonymous. Book reviews. Fish and Fisheries, 2(1):86–90, March 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hansen:2001:AIF

[37] Michael M. Hansen, Ellen Kenchington, and Einar E. Nielsen. Assigning individual fish to populations using microsatellite DNA markers. Fish and Fisheries, 2(2):93–112, June 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Daget:2001:MDD

[38] Jacques Daget, Mireille Gayet, François J. Meunier, and Jean-Yves Sire. Major discoveries on the dermal skeleton of fossil and recent polypteriforms: a review. Fish and Fisheries, 2(2):113–124, June 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Patterson:2001:EUF

[39] Kenneth Patterson, Robin Cook, Chris Darby, Stratis Gavaris, Laurence Kell, Peter Lewy, Benoît Mesnil, André Punt, Victor Restrepo, Dankert W. Skagen, and Gunnar Stefánsson. Estimating uncertainty in fish stock assessment and forecasting. Fish and Fisheries, 2(2):125–157, June 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2001:E

[40] Anonymous. Erratum. Fish and Fisheries, 2(2):158–161, June 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cury:2001:HBS

[41] Philippe Cury and Patrice Cayré. Hunting became a secondary activity 2000 years ago: marine fishing did the same in 2021. Fish and Fisheries, 2(2):162–169, June 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Whitmarsh:2001:SFE

[42] David Whitmarsh and Anthony Charles. The state of fisheries economics: the 10th Biennial Conference of the International Institute of Fisheries Economics and Trade (IIFET). Fish and Fisheries, 2(2):170–171, June 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Winfield:2001:NFI

[43] Ian J. Winfield and Chuck Hollingworth. Nonindigenous Fishes Introduced into Inland Waters of the United States (American Fisheries Society, Special Publication 27). Fish and Fisheries, 2(2):172–173, June 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Jobling:2001:EIA

[44] Malcolm Jobling. Environmental Impacts of Aquaculture (Sheffield Biological Sciences, Vol. 5). Fish and Fisheries, 2(2):173, June 2001. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Purser:2001:MSF

[45] John Purser. Manual of salmonid farming. Fish and Fisheries, 2(2): 174, June 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Froese:2001:IDF

[46] Rainer Froese. An illustrated dictionary of fish and shellfish CD-ROM. Fish and Fisheries, 2(2):174–175, June 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2001:BLH

[47] Anonymous. Books listed here may be reviewed in a later issue of fish and fisheries. Fish and Fisheries, 2(2):176, June 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Stewart:2001:FFN

[48] Kathlyn M. Stewart. The freshwater fish of Neogene Africa (Miocene– Pleistocene): systematics and biogeography. Fish and Fisheries, 2(3): 177–230, September 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Petitgas:2001:GFS

[49] Pierre Petitgas. Geostatistics in fisheries survey design and stock assessment: models, variances and applications. Fish and Fisheries, 2(3): 231–249, September 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Zagarese:2001:ISU

[50] Horacio E. Zagarese and Craig E. Williamson. The implications of solar UV radiation exposure for fish and fisheries. Fish and Fisheries, 2(3): 250–260, September 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Montgomery:2001:AHS

[51] John Charles Montgomery, Nicholas Tolimieri, and Olivia Sophie Haine. Active habitat selection by pre-settlement reef fishes. Fish and Fisheries, 2(3):261–277, September 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

McDowall:2001:DDD

[52] R. M. McDowall. Diadromy, diversity and divergence: implications for speciation processes in fishes. Fish and Fisheries, 2(3):278–285, September 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cook:2001:QFD

[53] Robin Cook and Chuck Hollingworth. Quantitative fish dynamics. Fish and Fisheries, 2(3):286–287, September 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Winfield:2001:PFS

[54] Lan J. Winfield. Percid fishes: Systematics, ecology and exploitation. Fish and Fisheries, 2(3):287, September 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bartlett:2001:GEF

[55] Darius Bartlett. Geostatistics for estimating fish abundance. Fish and Fisheries, 2(3):287–288, September 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Thorpe:2001:TSE

[56] John Thorpe. Trout and salmon: Ecology, conservation and rehabilitation. Fish and Fisheries, 2(3):289–290, September 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Carter:2001:FFQ

[57] Chris Carter. Farmed fish quality. Fish and Fisheries, 2(3):289–290, September 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Platt:2001:FOI

[58] Trevor Platt. Fisheries oceanography: An integrative approach to fisheries ecology and management. Fish and Fisheries, 2(3):290–291, September 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Rose:2001:CDD

[59] Kenneth A. Rose, James H. Cowan, Jr., Kirk O. Winemiller, Ransom A. Myers, and Ray Hilborn. Compensatory density dependence in fish populations: importance, controversy, understanding and prognosis. Fish and Fisheries, 2(4):293–327, December 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Nielsen:2001:NAP

[60] Einar Eg Nielsen and Ellen Kenchington. A new approach to prioritizing marine fish and shellfish populations for conservation. Fish and Fisheries, 2(4):328–343, December 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Zeller:2001:VSL

[61] Dirk Zeller and Daniel Pauly. Visualisation of standardized life-history patterns. Fish and Fisheries, 2(4):344–355, December 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Mahon:2001:TMS

[62] Robin Mahon and Wayne Hunte. Trap mesh selectivity and the management of reef fishes. Fish and Fisheries, 2(4):356–375, December 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Schreiber:2001:CMI

[63] Dorothee K. Schreiber. Co-management without involvement: the plight of fishing communities. Fish and Fisheries, 2(4):376–384, December 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Winfield:2001:BRD

[64] Ian J. Winfield. Book review: Diseases of Carp and Other Cyprinid Fishes, Hoole, D., Bucke, D., Burgess, P. and Wellby, I. Fishing News Books, Oxford, 2001. ISBN 0-85238-252-9, £49.50. Hard cover, xiv + 264 pp., 5 tables, 1 box, 158 figures (82 in colour). Fish and Fisheries, 2(4):388, December 2001. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Pitcher:2002:LLE

[65] Tony J. Pitcher. The Larkin Lectures — Editor’s note. Fish and Fisheries, 3(1):1–2, March 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Alverson:2002:FIS

[66] Dayton L. Alverson. Factors influencing the scope and quality of science and management decisions: (the good, the bad and the ugly). Fish and Fisheries, 3(1):3–19, March 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Smedbol:2002:MMM

[67] R. Kent Smedbol, Arran McPherson, Michael M. Hansen, and Ellen Kenchington. Myths and moderation in marine ‘metapopulations’? Fish and Fisheries, 3(1):20–35, March 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Imsland:2002:TGB

[68] Albert K. Imsland and Ólöf Dóra Bartels Jónsdóttir. Is there a genetic basis to growth in Atlantic cod? Fish and Fisheries, 3(1):36–52, March 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hollingworth:2002:NEF

[69] Chuck Hollingworth. Nitrogen Excretion (Fish Physiology, vol. 20). Fish and Fisheries, 3(1):53–54, March 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Mace:2002:SFS

[70] Pamela M. Mace. Sustainable fishery systems. Fish and Fisheries, 3(1): 54–55, March 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hutchings:2002:SRR

[71] Jeffrey A. Hutchings. Stock, Recruitment and Reference Points: Assessment and Management of Atlantic Salmon (Hydrobiologie et Aquaculture series). Fish and Fisheries, 3(1):55–56, March 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Aas:2002:JFE

[72] Øystein Aas. Just Fish: Ethics and Canadian Marine Fisheries (Social and Economic Papers No. 23). Fish and Fisheries, 3(1):57–58, March 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Fevolden:2002:KBE

[73] Svein-Erik Fevolden. Krill: Biology, Ecology and Fisheries (fish and aquatic resources series 6). Fish and Fisheries, 3(1):57, March 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kaye:2002:ECF

[74] Stuart Kaye. Enforcing the Common Fisheries Policy. Fish and Fisheries, 3(1):58–59, March 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Coughlin:2002:AMF

[75] David J. Coughlin. Aerobic muscle function during steady swimming in fish. Fish and Fisheries, 3(2):63–78, June 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Brown:2002:FSE

[76] Culum Brown and Rachel L. Day. The future of stock enhancements: lessons for hatchery practice from conservation biology. Fish and Fisheries, 3(2):79–94, June 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Benson:2002:EER

[77] Ashleen J. Benson and Andrew W. Trites. Ecological effects of regime shifts in the Bering Sea and eastern North Pacific Ocean. Fish and Fisheries, 3(2):95–113, June 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kaiser:2002:MMH

[78] Michel J. Kaiser, Jeremy S. Collie, Stephen J. Hall, Simon Jennings, and Ian R. Poiner. Modification of marine habitats by trawling activities: prognosis and solutions. Fish and Fisheries, 3(2):114–136, June 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hollingworth:2002:FPE

[79] Chuck Hollingworth. Fish Pathology, 3rd Edition. Fish and Fisheries, 3 (2):137, June 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Grove:2002:FIF

[80] Dave J. Grove. Food intake in fish. Fish and Fisheries, 3(2):138–139, June 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hart:2002:IFE

[81] Paul J. B. Hart. Inland fisheries: Ecology and management. Fish and Fisheries, 3(2):139–140, June 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Pearcy:2002:OLA

[82] William G. Pearcy. The ocean life of Atlantic salmon: Environmental and biological factors influencing survival. Fish and Fisheries, 3(2):140– 141, June 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Purdom:2002:GVG

[83] Colin Purdom. Goldfish varieties and genetics: A handbook for breeders. Fish and Fisheries, 3(2):141, June 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ruffing:2002:ISC

[84] Renea A. Ruffing, Patrick M. Kocovsky, and Jay R. Stauffer, Jr. An introduction to species concepts and speciation of fishes. Fish and Fisheries, 3(3):143–145, September 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Pauly:2002:CDI

[85] Daniel Pauly. Charles Darwin, ichthyology and the species concept. Fish and Fisheries, 3(3):146–150, September 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ghiselin:2002:SCB

[86] Michael T. Ghiselin. Species concepts: the basis for controversy and reconciliation. Fish and Fisheries, 3(3):151–160, September 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Wiley:2002:SSR

[87] E. O. Wiley. On species and speciation with reference to the fishes. Fish and Fisheries, 3(3):161–170, September 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Mayden:2002:BSS

[88] Richard L. Mayden. On biological species, species concepts and individuation in the natural world. Fish and Fisheries, 3(3):171–196, September 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Barlow:2002:HBS

[89] George W. Barlow. How behavioural studies contribute to the species problem: a piscine perspective. Fish and Fisheries, 3(3):197–212, September 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Stauffer:2002:BID

[90] Jay R. Stauffer, Jr., Kenneth R. McKaye, and Ad F. Konings. Behaviour: an important diagnostic tool for Lake Malawi cichlids. Fish and Fisheries, 3(3):213–224, September 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Turner:2002:PSD

[91] George F. Turner. Parallel speciation, despeciation and respeciation: implications for species definition. Fish and Fisheries, 3(3):225–229, September 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Stauffer:2002:SCS

[92] Jay R. Stauffer, Jr., Patrick M. Kocovsky, and Renea A. Ruffing. Species concepts and speciation of fishes: concluding remarks. Fish and Fisheries, 3(3):230–232, September 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Munro:2002:ISU

[93] Gordon Munro and Ussif R. Sumaila. The impact of subsidies upon fisheries management and sustainability: the case of the North Atlantic. Fish and Fisheries, 3(4):233–250, December 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Maunder:2002:RBF

[94] Mark N. Maunder. The relationship between fishing methods, fisheries management and the estimation of maximum sustainable yield. Fish and Fisheries, 3(4):251–260, December 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Arlinghaus:2002:RTI

[95] Robert Arlinghaus, Thomas Mehner, and Ian G. Cowx. Reconciling traditional inland fisheries management and sustainability in industrialized countries, with emphasis on Europe. Fish and Fisheries, 3(4):261–316, December 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Pitcher:2002:EIU

[96] Tony J. Pitcher, Reg Watson, Robyn Forrest, Hreidar Tór Valtýsson, and Sylvie Guénette. Estimating illegal and unreported catches from marine ecosystems: a basis for change. Fish and Fisheries, 3(4):317–339, December 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kennelly:2002:CBC

[97] Steven J. Kennelly and Matt K. Broadhurst. By-catch begone: changes in the philosophy of fishing technology. Fish and Fisheries, 3(4):340–355, December 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2002:CSN

[98] Anonymous. Conference summary: the ‘new’ emerging economics of marine protected areas. Fish and Fisheries, 3(4):356–359, December 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Blaber:2002:FE

[99] Stephen J. M. Blaber. Fishes in estuaries. Fish and Fisheries, 3(4): 360–361, December 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Rose:2002:MFF

[100] George Rose. Migration of freshwater fishes. Fish and Fisheries, 3(4): 361–362, December 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2002:PCF

[101] Anonymous. Protecting the commons: A framework for resource management in the Americas. Fish and Fisheries, 3(4):362–363, December 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Magurran:2002:FBA

[102] Anne Magurran. Fish behaviour in the aquarium and in the wild. Fish and Fisheries, 3(4):362, December 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2002:SEM

[103] Anonymous. The skeptical environmentalist: Measuring the real state of the world. Fish and Fisheries, 3(4):364–365, December 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Goold:2002:RWW

[104] John C. Goold. Right Whales: Worldwide Status (The Journal of Cetacean Research and Management, Special Issue 2). Fish and Fisheries, 3(4):364, December 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2002:BR

[105] Anonymous. Book reviews. Fish and Fisheries, 3(4):366, December 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2002:AMB

[106] Anonymous. Advances in Marine Biology, Vol. 40. Fish and Fisheries, 3(4):366, December 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2002:ISO

[107] Anonymous. Island in the stream: Oceanography and fisheries of the Charleston Bump (AFS Symposium 25). Fish and Fisheries, 3(4):366, December 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2002:SFP

[108] Anonymous. Salmonid Fishes: Population Biology, Genetics and Management (Fish and Aquatic Resources Series 2). Fish and Fisheries, 3 (4):366, December 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2002:MER

[109] Anonymous. Management and ecology of river fisheries. Fish and Fisheries, 3(4):366, December 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2002:IFP

[110] Anonymous. Innovations in fish passage technology. Fish and Fisheries, 3(4):366, December 2002. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Christensen:2003:HYD

[111] Villy Christensen, Sylvie Guénette, Johanna J. Heymans, Carl J. Walters, Reginald Watson, Dirk Zeller, and Daniel Pauly. Hundred-year decline of North Atlantic predatory fishes. Fish and Fisheries, 4(1):1– 24, March 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Dulvy:2003:EVM

[112] Nicholas K. Dulvy, Yvonne Sadovy, and John D. Reynolds. Extinction vulnerability in marine populations. Fish and Fisheries, 4(1):25–64, March 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Rudd:2003:PAT

[113] Murray A. Rudd, Mark H. Tupper, Henk Folmer, and G. C. Van Kooten. Policy analysis for tropical marine reserves: challenges and directions. Fish and Fisheries, 4(1):65–85, March 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Sadovy:2003:NEH

[114] Yvonne Sadovy and Wai Lung Cheung. Near extinction of a highly fecund fish: the one that nearly got away. Fish and Fisheries, 4(1):86–99, March 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2003:E

[115] Anonymous. Erratum. Fish and Fisheries, 4(1):100–101, March 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2003:LFF

[116] Anonymous. Learning in fishes: Forthcoming fish and fisheries special issue. Fish and Fisheries, 4(1):102, March 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Chen:2003:DRF

[117] Yong Chen, Yan Jiao, and Liqiao Chen. Developing robust frequentist and Bayesian fish stock assessment methods. Fish and Fisheries, 4(2): 105–120, June 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Godo:2003:FSP

[118] O. R. Godø. Fluctuation in stock properties of north-east Arctic cod related to long-term environmental changes. Fish and Fisheries, 4(2): 121–137, June 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Croft:2003:WFS

[119] Darren P. Croft, Jens Krause, Iain D. Couzin, and Tony J. Pitcher. When fish shoals meet: outcomes for evolution and fisheries. Fish and Fisheries, 4(2):138–146, June 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ali:2003:CGF

[120] M. Ali, A. Nicieza, and R. J. Wootton. Compensatory growth in fishes: a response to growth depression. Fish and Fisheries, 4(2):147–190, June 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2003:MEL

[121] Anonymous. Management and ecology of lake and reservoir fisheries. Fish and Fisheries, 4(2):191–192, June 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bartlett:2003:GIS

[122] Darius Bartlett. Geographic information systems in oceanography and fisheries. Fish and Fisheries, 4(2):192–193, June 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Mitchelson–Jacob:2003:ISO

[123] E. G. Mitchelson-Jacob. Island in the Stream: Oceanography and Fisheries of the Charleston Bump (American Fisheries Society Symposium 25). Fish and Fisheries, 4(2):193–194, June 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2003:PPC

[124] Anonymous. Phytoplankton productivity: Carbon assimilation in marine and freshwater ecosystems. Fish and Fisheries, 4(2):194–195, June 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cross:2003:PGA

[125] Tom F. Cross. Practical genetics for aquaculture. Fish and Fisheries, 4(2):195–196, June 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2003:RMA

[126] Anonymous. Responsible marine aquaculture. Fish and Fisheries, 4(2): 196, June 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Huntingford:2003:F

[127] Felicity Huntingford. Foreword. Fish and Fisheries, 4(3):197–198, September 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Laland:2003:LFT

[128] Keven N. Laland, Culum Brown, and Jens Krause. Learning in fishes: from three-second memory to culture. Fish and Fisheries, 4(3):199–202, September 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Warburton:2003:LFS

[129] Kevin Warburton. Learning of foraging skills by fish. Fish and Fisheries, 4(3):203–215, September 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kelley:2003:LPR

[130] Jennifer L. Kelley and Anne E. Magurran. Learned predator recognition and antipredator responses in fishes. Fish and Fisheries, 4(3):216–226, September 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Brown:2003:LAD

[131] Grant E. Brown. Learning about danger: chemical alarm cues and local risk assessment in prey fishes. Fish and Fisheries, 4(3):227–234, September 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Odling-Smee:2003:RLF

[132] Lucy Odling-Smee and Victoria A. Braithwaite. The role of learning in fish orientation. Fish and Fisheries, 4(3):235–246, September 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Broglio:2003:SCN

[133] Cristina Broglio, Fernando Rodríguez, and Cosme Salas. Spatial cognition and its neural basis in teleost fishes. Fish and Fisheries, 4(3): 247–255, September 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Griffiths:2003:LRC

[134] Siân W. Griffiths. Learned recognition of conspecifics by fishes. Fish and Fisheries, 4(3):256–268, September 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hoare:2003:SOS

[135] D. J. Hoare and J. Krause. Social organisation, shoal structure and information transfer. Fish and Fisheries, 4(3):269–279, September 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Brown:2003:SLF

[136] Culum Brown and Kevin N. Laland. Social learning in fishes: a review. Fish and Fisheries, 4(3):280–288, September 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kasumyan:2003:TPF

[137] Alexander O. Kasumyan and Kjell B. Døving. Taste preferences in fishes. Fish and Fisheries, 4(4):289–347, December 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ward:2003:EKF

[138] Ashley J. W. Ward and Paul J. B. Hart. The effects of kin and familiarity on interactions between fish. Fish and Fisheries, 4(4):348–358, December 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Prince:2003:BEG

[139] Jeremy D. Prince. The barefoot ecologist goes fishing. Fish and Fisheries, 4(4):359–371, December 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kaiser:2003:BSF

[140] Michel J. Kaiser. Bigelow and Schroeder’s Fishes of the Gulf of Maine, third edition. Fish and Fisheries, 4(4):372, December 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Stefansson:2003:BRF

[141] Sigurd O. Stefansson. Book review: Fish Hatchery Management, Second Edition. Fish and Fisheries, 4(4):372–373, December 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Blaber:2003:BSS

[142] Stephen J. M. Blaber. Biology of the Spotted Seatrout (CRC Marine Biology Series, Peter L. Lutz, ed.). Fish and Fisheries, 4(4):374, December 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2003:MFE

[143] Anonymous. Marine fisheries ecology. Fish and Fisheries, 4(4):374–375, December 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Clarke:2003:PSE

[144] Graham Clarke. Practical statistics for environmental and biological scientists. Fish and Fisheries, 4(4):375–376, December 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Lorenzen:2003:ADB

[145] Kai Lorenzen. Animal domestication and behavior. Fish and Fisheries, 4(4):376, December 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Lorenzen:2003:WDR

[146] Kai Lorenzen. Whirling Disease: Reviews and Current Topics (American Fisheries Society Symposium 29). Fish and Fisheries, 4(4):377, December 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hart:2003:FSU

[147] Paul J. B. Hart. Fishery science: The unique contributions of early life stages. Fish and Fisheries, 4(4):377–378, December 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2003:AIV

[148] Anonymous. Author index to volume 4 (2003). Fish and Fisheries, 4 (4):379, December 2003. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hall:2004:TEB

[149] Stephen J. Hall and B. Mainprize. Towards ecosystem-based fisheries management. Fish and Fisheries, 5(1):1–20, March 2004. CODEN FFI-IAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cooke:2004:AEF

[150] Steven J. Cooke, Eva B. Thorstad, and Scott G. Hinch. Activity and energetics of free-swimming fish: insights from electromyogram telemetry. Fish and Fisheries, 5(1):21–52, March 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Clarke:2004:UPF

[151] Shelley Clarke. Understanding pressures on fishery resources through trade statistics: a pilot study of four products in the Chinese dried seafood market. Fish and Fisheries, 5(1):53–74, March 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

McDowall:2004:AAI

[152] R. M. McDowall. Ancestry and amphidromy in island freshwater fish faunas. Fish and Fisheries, 5(1):75–85, March 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Froese:2004:KIS

[153] Rainer Froese. Keep it simple: three indicators to deal with overfishing. Fish and Fisheries, 5(1):86–91, March 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Sullivan:2004:HFB

[154] Michael G. Sullivan. Handbook of Fish Biology and Fisheries. Volume 1: Fish Biology. Volume 2: Fisheries. Fish and Fisheries, 5(1):92–93, March 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Thorpe:2004:SNA

[155] John Thorpe. Sustaining North American salmon: Perspectives across regions and disciplines. Fish and Fisheries, 5(1):93, March 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Carvalho:2004:PGP

[156] Gary R. Carvalho. Population genetics: Principles and applications for fisheries scientists. Fish and Fisheries, 5(1):94–95, March 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hart:2004:FCM

[157] Paul J. B. Hart. The fisheries co-management experience: Accomplishments, challenges and prospects. Fish and Fisheries, 5(1):95–96, March 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Chakrabarty:2004:CBC

[158] Prosanta Chakrabarty. Cichlid biogeography: comment and review. Fish and Fisheries, 5(2):97–119, June 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Koops:2004:THA

[159] Marten A. Koops, Jeffrey A. Hutchings, and Tara M. McIntyre. Testing hypotheses about fecundity, body size and maternal condition in fishes. Fish and Fisheries, 5(2):120–130, June 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kritzer:2004:MES

[160] Jacob P. Kritzer and Peter F. Sale. Metapopulation ecology in the sea: from Levins’ model to marine ecology and fisheries science. Fish and Fisheries, 5(2):131–140, June 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Shepherd:2004:DDH

[161] Travis D. Shepherd and Matthew K. Litvak. Density-dependent habitat selection and the ideal free distribution in marine fish spatial dynamics: considerations and cautions. Fish and Fisheries, 5(2):141–152, June 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Salas:2004:BDF

[162] Silvia Salas and Daniel Gaertner. The behavioural dynamics of fishers: management implications. Fish and Fisheries, 5(2):153–167, June 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Watson:2004:MGF

[163] Reg Watson, Adrian Kitchingman, Ahmed Gelchu, and Daniel Pauly. Mapping global fisheries: sharpening our focus. Fish and Fisheries, 5(2): 168–177, June 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Mills:2004:OGN

[164] Claudia E. Mills and Kit Rawson. Outlook grim for North Pacific rockfish: Rockfish Symposium, Friday Harbor Laboratories, University of Washington, USA, September 25–26, 2003. Fish and Fisheries, 5(2): 178–180, June 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hollingworth:2004:CRM

[165] Chuck Hollingworth. Catch and Release in Marine Recreational Fisheries (American Fisheries Society Symposium 30). Fish and Fisheries, 5(2): 181–182, June 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Rose:2004:FCC

[166] George A. Rose. Fisheries in a Changing Climate (American Fisheries Society Symposium 32). Fish and Fisheries, 5(2):182–183, June 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2004:BR

[167] Anonymous. Books received. Fish and Fisheries, 5(2):184, June 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Robichaud:2004:MBR

[168] David Robichaud and George A. Rose. Migratory behaviour and range in Atlantic cod: inference from a century of tagging. Fish and Fisheries, 5 (3):185–214, September 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Allison:2004:HAA

[169] Edward H. Allison and Janet A. Seeley. HIV and AIDS among fisherfolk: a threat to ‘responsible fisheries’? Fish and Fisheries, 5(3):215–234, September 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cotter:2004:SAM

[170] A. J. R. Cotter, L. Burt, C. G. M. Paxton, C. Fernandez, S. T. Buckland, and J-X. Pan. Are stock assessment methods too complicated? Fish and Fisheries, 5(3):235–254, September 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Dulvy:2004:MAE

[171] Nicholas K. Dulvy, Jim R. Ellis, Nicholas B. Goodwin, Alastair Grant, John D. Reynolds, and Simon Jennings. Methods of assessing extinction risk in marine fishes. Fish and Fisheries, 5(3):255–276, September 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hollingworth:2004:SPA

[172] Chuck Hollingworth. Sensory processing in aquatic environments. Fish and Fisheries, 5(3):277, September 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hollingworth:2004:RFM

[173] Chuck Hollingworth. Responsible fisheries in the marine ecosystem. Fish and Fisheries, 5(3):277–278, September 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hollingworth:2004:IBF

[174] Chuck Hollingworth. Interactions between fish and birds: Implications for management. Fish and Fisheries, 5(3):278–279, September 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hollingworth:2004:ETE

[175] Chuck Hollingworth. The Eel, Third Edition (English Fifth Edition). Fish and Fisheries, 5(3):279–280, September 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Chandroo:2004:ECP

[176] Kristopher Paul Chandroo, Stephanie Yue, and Richard David Moccia. An evaluation of current perspectives on consciousness and pain in fishes. Fish and Fisheries, 5(4):281–295, December 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bertrand:2004:SSH

[177] Arnaud Bertrand, Marceliano Segura, Mariano Gutiérrez, and Luis Vásquez. From small-scale habitat loopholes to decadal cycles: a habitatbased hypothesis explaining fluctuation in pelagic fish populations off Peru. Fish and Fisheries, 5(4):296–316, December 2004. CODEN FFI-IAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Takemura:2004:LCR

[178] Akihiro Takemura, Md. Saydur Rahman, Sigeo Nakamura, Young Ju Park, and Kazunori Takano. Lunar cycles and reproductive activity in reef fishes with particular attention to rabbitfishes. Fish and Fisheries, 5 (4):317–328, December 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Sumaila:2004:ICB

[179] Ussif Rashid Sumaila. Intergenerational cost–benefit analysis and marine ecosystem restoration. Fish and Fisheries, 5(4):329–343, December 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hollingworth:2004:IH

[180] Chuck Hollingworth. Ichthyology handbook. Fish and Fisheries, 5(4): 344, December 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hollingworth:2004:ELF

[181] Chuck Hollingworth. Eco-labelling in fisheries: What is it all about? Fish and Fisheries, 5(4):344–346, December 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hollingworth:2004:IST

[182] Chuck Hollingworth. Imperfect symmetry: Thermodynamics in ecology and evolution. Fish and Fisheries, 5(4):346–347, December 2004. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hollingworth:2004:TPE

[183] Chuck Hollingworth. Tuna: Physiology, Ecology, and Evolution. Fish Physiology Series Vol. 19 (Series editors: W. S. Hoar, D. J. Randall and A. P. Farrell). Fish and Fisheries, 5(4):348, December 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2004:AIV

[184] Anonymous. Author index to volume 5 (2004). Fish and Fisheries, 5 (4):349, December 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2004:C

[185] Anonymous. Contents. Fish and Fisheries, 5(4):iii–iv, December 2004. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Genner:2005:MCL

[186] Martin J. Genner and George F. Turner. The mbuna cichlids of Lake Malawi: a model for rapid speciation and adaptive radiation. Fish and Fisheries, 6(1):1–34, March 2005. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Gilman:2005:PAA

[187] Eric Gilman, Nigel Brothers, and Donald R. Kobayashi. Principles and approaches to abate seabird by-catch in longline fisheries. Fish and Fisheries, 6(1):35–49, March 2005. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Rideout:2005:SSF

[188] Rick M. Rideout, George A. Rose, and Margaret P. M. Burton. Skipped spawning in female iteroparous fishes. Fish and Fisheries, 6(1):50–72, March 2005. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Smith:2005:AYD

[189] Tim D. Smith and Jason S. Link. Autopsy your dead… and living: a proposal for fisheries science, fisheries management and fisheries. Fish and Fisheries, 6(1):73–87, March 2005. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hollingworth:2005:NFE

[190] Chuck Hollingworth. Namibia’s fisheries: Ecological, economic and social aspects. Fish and Fisheries, 6(1):88, March 2005. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hollingworth:2005:LMR

[191] Chuck Hollingworth. The Living Marine Resources of the Western Central Atlantic. Volume 1: Introduction, molluscs, crustaceans, hagfishes, sharks, batoid fishes, and chimaeras. Volume 2: Bony fishes part 1 (Acipenseridae to Grammatidae). Volume 3: Bony fishes part 2 (Opistognathidae to Molidae), sea turtles and marine mammals. FAO Species Identification Guide for Fishery Purposes and American Society of Ichthyologists and Herpetologists Special Publication No. 5. Fish and Fisheries, 6(1):89–90, March 2005. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hollingworth:2005:FSB

[192] Chuck Hollingworth. Fish and shellfish bacteriology manual: Techniques and procedures. Fish and Fisheries, 6(1):90, March 2005. CODEN FFI-IAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hollingworth:2005:NSE

[193] Chuck Hollingworth. The Norwegian Sea Ecosystem. Fish and Fisheries, 6(1):90–91, March 2005. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2005:BR

[194] Anonymous. Books received. Fish and Fisheries, 6(1):92, March 2005. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Oldfield:2005:GAS

[195] Ronald George Oldfield. Genetic, abiotic and social influences on sex differentiation in cichlid fishes and the evolution of sequential hermaphroditism. Fish and Fisheries, 6(2):93–110, June 2005. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Mullon:2005:DCW

[196] Christian Mullon, Pierre Fréon, and Philippe Cury. The dynamics of collapse in world fisheries. Fish and Fisheries, 6(2):111–120, June 2005. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Saenz-Arroyo:2005:UFA

[197] Andrea Sáenz-Arroyo, Callum M. Roberts, Jorge Torre, and Micheline Cariño-Olvera. Using fishers’ anecdotes, naturalists’ observations and grey literature to reassess marine species at risk: the case of the gulf grouper in the Gulf of California, Mexico. Fish and Fisheries, 6(2):121– 133, June 2005. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hall:2005:MCD

[198] Stephen J. Hall and Brooke M. Mainprize. Managing by-catch and discards: how much progress are we making and how can we do better? Fish and Fisheries, 6(2):134–155, June 2005. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Zeller:2005:GNB

[199] Dirk Zeller and Daniel Pauly. Good news, bad news: global fisheries discards are declining, but so are total catches. Fish and Fisheries, 6(2): 156–159, June 2005. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hollingworth:2005:FFB

[200] Chuck Hollingworth. Freshwater fishes in Britain: The species and their distribution. Fish and Fisheries, 6(2):160–161, June 2005. CODEN FFI-IAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hollingworth:2005:MEC

[201] Chuck Hollingworth. Marine ecosystems and climate variation. The North Atlantic: A comparative perspective. Fish and Fisheries, 6(2): 161–162, June 2005. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hollingworth:2005:SRR

[202] Chuck Hollingworth. Strategies for restoring river ecosystems: Sources of variability and uncertainty in natural and managed systems. Fish and Fisheries, 6(2):162–164, June 2005. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Pitcher:2005:LLE

[203] Tony J. Pitcher. The Larkin Lectures — Editor’s note. Fish and Fisheries, 6(3):165, September 2005. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Sadovy:2005:TRI

[204] Yvonne Sadovy. Trouble on the reef: the imperative for managing vulnerable and valuable fisheries. Fish and Fisheries, 6(3):167–185, September 2005. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Sayer:2005:AAF

[205] Martin D. J. Sayer. Adaptations of amphibious fish for surviving life out of water. Fish and Fisheries, 6(3):186–211, September 2005. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Jennings:2005:ISE

[206] Simon Jennings. Indicators to support an ecosystem approach to fisheries. Fish and Fisheries, 6(3):212–232, September 2005. CODEN FFI-IAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Pikitch:2005:STM

[207] Ellen K. Pikitch, Phaedra Doukakis, Liz Lauck, Prosanta Chakrabarty, and Daniel L. Erickson. Status, trends and management of sturgeon and paddlefish fisheries. Fish and Fisheries, 6(3):233–265, September 2005. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Halley:2005:IIV

[208] John M. Halley and Konstantinos I. Stergiou. The implications of increasing variability of fish landings. Fish and Fisheries, 6(3):266–276, September 2005. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Poulin:2005:LLL

[209] Robert Poulin. Leeches, lice and lampreys: A natural history of skin and gill parasites of fishes. Fish and Fisheries, 6(3):277, September 2005. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Pepin:2005:FEM

[210] Pierre Pepin. Fisheries ecology and management. Fish and Fisheries, 6 (3):278–279, September 2005. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Rose:2005:SES

[211] Kenneth A. Rose. Stock enhancement and sea ranching: Developments, pitfalls and opportunities, second edition. Fish and Fisheries, 6(3):279, September 2005. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2005:E

[212] Anonymous. Errata. Fish and Fisheries, 6(3):280, September 2005. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Fromentin:2005:ABT

[213] Jean-Marc Fromentin and Joseph E. Powers. Atlantic bluefin tuna: population dynamics, ecology, fisheries and management. Fish and Fisheries, 6(4):281–306, December 2005. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Pelletier:2005:SEF

[214] Dominique Pelletier and Stéphanie Mahévas. Spatially explicit fisheries simulation models for policy evaluation. Fish and Fisheries, 6(4):307– 349, December 2005. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Harrington:2005:WFR

[215] Jennie M. Harrington, Ransom A. Myers, and Andrew A. Rosenberg. Wasted fishery resources: discarded by-catch in the USA. Fish and Fisheries, 6(4):350–361, December 2005. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hollingworth:2005:ESH

[216] Chuck Hollingworth. Electroreception (Springer Handbook of Auditory Research 21, series eds R. R. Fay and A. N. Popper). Fish and Fisheries, 6(4):362–363, December 2005. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2005:AIV

[217] Anonymous. Author index to Volume 6 (2005). Fish and Fisheries, 6 (4):364, December 2005. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2006:A

[218] Anonymous. Announcement. Fish and Fisheries, 7(1):1, March 2006. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Gilman:2006:RST

[219] Eric Gilman, Erika Zollett, Stephen Beverly, Hideki Nakano, Kimberly Davis, Daisuke Shiode, Paul Dalzell, and Irene Kinan. Reducing sea turtle by-catch in pelagic longline fisheries. Fish and Fisheries, 7(1):2– 23, March 2006. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Morato:2006:FD

[220] Telmo Morato, Reg Watson, Tony J. Pitcher, and Daniel Pauly. Fishing down the deep. Fish and Fisheries, 7(1):24–34, March 2006. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Stratoudakis:2006:DEP

[221] Yorgos Stratoudakis, Miguel Bernal, Konstantinos Ganias, and Andrés Uriarte. The daily egg production method: recent advances, current applications and future challenges. Fish and Fisheries, 7(1):35–57, March 2006. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Mous:2006:CCA

[222] Peter J. Mous, Yvonne Sadovy, Abdul Halim, and Jos S. Pet. Capture for culture: artificial shelters for grouper collection in SE Asia. Fish and Fisheries, 7(1):58–72, March 2006. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Stiassny:2006:FBE

[223] Melanie L. J. Stiassny. Flatfishes: Biology and Exploitation (Fish and Aquatic Resources Series 9, series ed. T. J. Pitcher). Fish and Fisheries, 7(1):73–74, March 2006. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Blaber:2006:BHE

[224] Stephen J. M. Blaber. Benthic Habitats and the Effects of Fishing (American Fisheries Society Symposium 41). Fish and Fisheries, 7(1):74–75, March 2006. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hart:2006:BEP

[225] Paul J. B. Hart. The behavior and ecology of Pacific salmon and trout. Fish and Fisheries, 7(1):75–76, March 2006. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Marshall:2006:SFA

[226] Justin Marshall. The senses of fish: Adaptations for the reception of natural stimuli. Fish and Fisheries, 7(1):76–77, March 2006. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2006:C

[227] Anonymous. Corrigendum. Fish and Fisheries, 7(1):78–79, March 2006. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Pitcher:2006:OCC

[228] Tony J. Pitcher and Paul J. B. Hart. Obituary for Charles (Chuck) Edward Hollingworth BSc, MSc (McGill), PhD (Wales) 1950–2005. Fish and Fisheries, 7(2):81–83, June 2006. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Muller:2006:UFS

[229] Ulrike K. Müller and Johan L. Van Leeuwen. Undulatory fish swimming: from muscles to flow. Fish and Fisheries, 7(2):84–103, June 2006. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Lees:2006:CRS

[230] Kathryn Lees, Sophie Pitois, Catherine Scott, Chris Frid, and Steven Mackinson. Characterizing regime shifts in the marine environment. Fish and Fisheries, 7(2):104–127, June 2006. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Saenz-Arroyo:2006:VEA

[231] Andrea Sáenz-Arroyo, Callum M. Roberts, Jorge Torre, Micheline Cariño-Olvera, and Julie P. Hawkins. The value of evidence about past abundance: marine fauna of the Gulf of California through the eyes of 16th to 19th century travellers. Fish and Fisheries, 7(2):128–146, June 2006. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Chuenpagdee:2006:WWH

[232] Ratana Chuenpagdee and Alida Bundy. What was hot at the fourth World Fisheries Congress? Fish and Fisheries, 7(3):147–150, June 2006. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2006:Ea

[233] Anonymous. Editorial. Fish and Fisheries, 7(3):151–152, September 2006. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Mcdowall:2006:PAA

[234] Rm Mcdowall and Sc Charteris. The possible adaptive advantages of terrestrial egg deposition in some fluvial diadromous galaxiid fishes (Teleostei: Galaxiidae). Fish and Fisheries, 7(3):153–164, September 2006. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Brehmer:2006:OMS

[235] Patrice Brehmer, Thierry Lafont, Stratis Georgakarakos, Erwan Josse, François Gerlotto, and Christophe Collet. Omnidirectional multibeam sonar monitoring: applications in fisheries science. Fish and Fisheries, 7 (3):165–179, September 2006. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Broadhurst:2006:ECM

[236] Matt K. Broadhurst, Petri Suuronen, and Alex Hulme. Estimating collateral mortality from towed fishing gear. Fish and Fisheries, 7(3):180–218, September 2006. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Pinnegar:2006:WRS

[237] John K. Pinnegar, Trevor P. Hutton, and Vincenzo Placenti. What relative seafood prices can tell us about the status of stocks. Fish and Fisheries, 7(3):219–226, September 2006. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Brehmer:2006:FAT

[238] Patrice A. J-P. Brehmer. Fisheries acoustics: Theory and practice, 2nd edn. Fish and Fisheries, 7(3):227–228, September 2006. CODEN FFI-IAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hall-Spencer:2006:MEP

[239] Jason Hall-Spencer. Marine ecology. process, systems, and impacts. Fish and Fisheries, 7(3):228, September 2006. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2006:Eb

[240] Anonymous. Erratum. Fish and Fisheries, 7(3):229, September 2006. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ward:2006:IFC

[241] Ashley J. W. Ward, Michael M. Webster, and Paul J. B. Hart. Intraspecific food competition in fishes. Fish and Fisheries, 7(4):231–261, December 2006. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Young:2006:IPL

[242] Jeffery L. Young, Zosia B. Bornik, Michelle L. Marcotte, Kim N. Charlie, Glenn N. Wagner, Scott G. Hinch, and Steven J. Cooke. Integrating physiology and life history to improve fisheries management and conservation. Fish and Fisheries, 7(4):262–283, December 2006. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Field:2006:DMM

[243] John C. Field, André E. Punt, Richard D. Methot, Jr., and Cynthia J. Thomson. Does MPA mean ‘Major Problem for Assessments’? Considering the consequences of place-based management systems. Fish and Fisheries, 7(4):284–302, December 2006. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Briones:2006:PFF

[244] Roehlano M. Briones. Projecting future fish supplies using stock dynamics and demand. Fish and Fisheries, 7(4):303–315, December 2006. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bakun:2006:AFS

[245] Andrew Bakun and Scarla J. Weeks. Adverse feedback sequences in exploited marine systems: are deliberate interruptive actions warranted? Fish and Fisheries, 7(4):316–333, December 2006. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Day:2006:FWE

[246] Julia J. Day. Fishes of the world, 4th edition. Fish and Fisheries, 7 (4):334, December 2006. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

McDowall:2007:ADF

[247] R. M. McDowall. On amphidromy, a distinct form of diadromy in aquatic organisms. Fish and Fisheries, 8(1):1–13, March 2007. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Robards:2007:GCR

[248] Martin D. Robards and Joshua A. Greenberg. Global constraints on rural fishing communities: whose resilience is it anyway? Fish and Fisheries, 8(1):14–30, March 2007. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bull:2007:RSB

[249] Leigh S. Bull. Reducing seabird bycatch in longline, trawl and gillnet fisheries. Fish and Fisheries, 8(1):31–56, March 2007. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Arlinghaus:2007:FWC

[250] Robert Arlinghaus, Steven J. Cooke, Alexander Schwab, and Ian G. Cowx. Fish welfare: a challenge to the feelings-based approach, with implications for recreational fishing. Fish and Fisheries, 8(1):57–71, March 2007. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kaiser:2007:MSM

[251] Michel Kaiser. Methods for the Study of Marine Benthos (Third Edition). Fish and Fisheries, 8(1):72–73, March 2007. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Muller:2007:FBV

[252] Ulrike Müller. Fish Biomechanics. volume 23 in the series ‘Fish Physiology’. Fish and Fisheries, 8(1):73–74, March 2007. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Law:2007:TBT

[253] Richard Law. The theoretical biologist’s toolbox: Quantitative methods for ecology and evolutionary biology. Fish and Fisheries, 8(1):74–75, March 2007. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Needle:2007:FAM

[254] Coby L. Needle. Fisheries assessment and management in data-limited situations. Fish and Fisheries, 8(1):75–76, March 2007. CODEN FFI-IAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

VanDamme:2007:IEB

[255] Dirk Van Damme, Nina Bogutskaya, Richard C. Hoffmann, and Carl Smith. The introduction of the European bitterling (Rhodeus amarus) to west and central Europe. Fish and Fisheries, 8(2):79–106, June 2007. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

delMonte-Luna:2007:MER

[256] Pablo del Monte-Luna, Daniel Lluch-Belda, Elisa Serviere-Zaragoza, Roberto Carmona, Héctor Reyes-Bonilla, David Aurioles-Gamboa, José Luis Castro-Aguirre, Sergio A. Guzmán del Próo, Oscar Trujillo-Millán, and Barry W. Brook. Marine extinctions revisited. Fish and Fisheries, 8(2):107–122, June 2007. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cotter:2007:LLO

[257] A. J. R. Cotter and G. M. Pilling. Landings, logbooks and observer surveys: improving the protocols for sampling commercial fisheries. Fish and Fisheries, 8(2):123–152, June 2007. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

McShane:2007:KWU

[258] Paul E. McShane, Matt K. Broadhurst, and Alan Williams. Keeping watch on the unwatchable: technological solutions for the problems generated by ecosystem-based management. Fish and Fisheries, 8(2):153– 161, June 2007. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Misund:2007:BFC

[259] Ole Arve Misund. Von Bandt’s Fish Catching Methods of the World (Fourth Edition). Fish and Fisheries, 8(2):162–163, June 2007. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Greenberg:2007:RFF

[260] Larry Greenberg. Review of fishes and forestry, worldwide watershed interactions and management. Fish and Fisheries, 8(2):164, June 2007. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Annala:2007:FLI

[261] John Annala. Fish for life: Interactive governance for fisheries. Fish and Fisheries, 8(2):164–165, June 2007. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Aiken:2007:MEC

[262] Hamish M. Aiken, Nathan J. Bott, Ivona Mladineo, Francisco E. Montero, Barbara F. Nowak, and Craig J. Hayward. Molecular evidence for cosmopolitan distribution of platyhelminth parasites of tunas (Thunnus spp.). Fish and Fisheries, 8(3):167–180, September 2007. CODEN FFI-IAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

LeQuesne:2007:CNT

[263] William J. F. Le Quesne, Stephen J. Hawkins, and John G. Shepherd. A comparison of no-take zones and traditional fishery management tools for managing site-attached species with a mixed larval pool. Fish and Fisheries, 8(3):181–195, September 2007. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Coggins:2007:ECM

[264] Lewis G. Coggins, Jr., Matthew J. Catalano, Micheal S. Allen, William E. Pine III, and Carl J. Walters. Effects of cryptic mortality and the hidden costs of using length limits in fishery management. Fish and Fisheries, 8 (3):196–210, September 2007. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Heck:2007:IAF

[265] Simon Heck, Christophe Béné, and Roberto Reyes-Gaskin. Investing in African fisheries: building links to the Millennium Development Goals. Fish and Fisheries, 8(3):211–226, September 2007. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Andrew:2007:DMS

[266] Neil L. Andrew, Christophe Béné, Stephen J. Hall, Edward H. Allison, Simon Heck, and Blake D. Ratner. Diagnosis and management of smallscale fisheries in developing countries. Fish and Fisheries, 8(3):227–240, September 2007. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Elliott:2007:GAC

[267] Michael Elliott, Alan K. Whitfield, Ian C. Potter, Stephen J. M. Blaber, Digby P. Cyrus, Frank G. Nordlie, and Trevor D. Harrison. The guild approach to categorizing estuarine fish assemblages: a global review. Fish and Fisheries, 8(3):241–268, September 2007. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Jennings:2007:RAE

[268] Simon Jennings. Reporting and advising on the effects of fishing. Fish and Fisheries, 8(3):269–276, September 2007. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Huntingford:2007:IFB

[269] Felicity Huntingford, Colin Adams, Victoria A. Braithwaite, Sunil Kadri, Tom G. Pottinger, Peter Sandoe, and James F. Turnbull. The implications of a feelings-based approach to fish welfare: a reply to Arlinghaus et al. Fish and Fisheries, 8(3):277–280, September 2007. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Baines:2007:HDE

[270] David Baines. How to do ecology — a concise handbook. Fish and Fisheries, 8(3):281–282, September 2007. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Goodwin:2007:CRC

[271] Edward Goodwin. Coral Reef Conservation. volume 13 in the series ‘Conservation Biology’. Fish and Fisheries, 8(3):282–283, September 2007. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2007:LL

[272] Anonymous. The 8th Larkin Lecture. Fish and Fisheries, 8(4):285, December 2007. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hilborn:2007:MFM

[273] Ray Hilborn. Managing fisheries is managing people: what has been learned? Fish and Fisheries, 8(4):285–296, December 2007. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Schtickzelle:2007:MPS

[274] Nicolas Schtickzelle and Thomas P. Quinn. A metapopulation perspective for salmon and other anadromous fish. Fish and Fisheries, 8(4): 297–314, December 2007. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hill:2007:MUE

[275] Simeon L. Hill, George M. Watters, André E. Punt, Murdoch K. McAllister, Corinne Le Quéré, and John Turner. Model uncertainty in the ecosystem approach to fisheries. Fish and Fisheries, 8(4):315–336, December 2007. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Magnusson:2007:WMF

[276] Arni Magnusson and Ray Hilborn. What makes fisheries data informative? Fish and Fisheries, 8(4):337–358, December 2007. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2007:G

[277] Anonymous. Ghoti. Fish and Fisheries, 8(4):359, December 2007. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Okey:2007:SSC

[278] Thomas A. Okey, Bruce A. Wright, and Michael Y. Brubaker. Salmon shark connections: North Pacific climate change, indirect fisheries effects, or just variability? Fish and Fisheries, 8(4):359–366, December 2007. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Nottestad:2007:TPM

[279] Leif Nøttestad. Top predators in marine ecosystems. Their role in monitoring and management. Fish and Fisheries, 8(4):367–368, December 2007. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Doncaster:2007:SEM

[280] C. Patrick Doncaster. Structural equation modeling and natural systems. Fish and Fisheries, 8(4):368–369, December 2007. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Naevdal:2007:ASG

[281] Gunnar Nævdal. The Atlantic salmon. Genetics, conservation and management. Fish and Fisheries, 8(4):369–370, December 2007. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

SadovyDeMitcheson:2008:FHT

[282] Yvonne Sadovy De Mitcheson and Min Liu. Functional hermaphroditism in teleosts. Fish and Fisheries, 9(1):1–43, March 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bailey:2008:OTP

[283] Kevin M. Bailey, Alisa A. Abookire, and Janet T. Duffy-Anderson. Ocean transport paths for the early life history stages of offshorespawning flatfishes: a case study in the Gulf of Alaska. Fish and Fisheries, 9(1):44–66, March 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Maxime:2008:PTF

[284] Valérie Maxime. The physiology of triploid fish: current knowledge and comparisons with diploid fish. Fish and Fisheries, 9(1):67–78, March 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Donaldson:2008:ECR

[285] Michael R. Donaldson, Robert Arlinghaus, Kyle C. Hanson, and Steven J. Cooke. Enhancing catch-and-release science with biotelemetry. Fish and Fisheries, 9(1):79–105, March 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Gozlan:2008:INN

[286] Rodolphe Elie Gozlan. Introduction of non-native freshwater fish: is it all bad? Fish and Fisheries, 9(1):106–115, March 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Reid:2008:GS

[287] Philip C. Reid. The Gulf Stream. Fish and Fisheries, 9(1):116–117, March 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Harper:2008:CLE

[288] David Harper. Conservation: Linking ecology, economics and culture. Fish and Fisheries, 9(1):117–118, March 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Johnson:2008:FKF

[289] Teresa R. Johnson. Fishers’ knowledge in fisheries science and management. Fish and Fisheries, 9(1):118, March 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Pitcher:2008:OSE

[290] Tony J. Pitcher. Obituary: Stream out of Eden: K. Radway Allen 1911– 2008. Fish and Fisheries, 9(2):119–120, June 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Fenichel:2008:ROP

[291] Eli P. Fenichel, Jean I. Tsao, Michael L. Jones, and Graham J. Hickling. Real options for precautionary fisheries management. Fish and Fisheries, 9(2):121–137, June 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hernandez-Leon:2008:NVF

[292] Santiago Hernández-León. Natural variability of fisheries and lunar illumination: a hypothesis. Fish and Fisheries, 9(2):138–154, June 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Josephson:2008:HDR

[293] Elizabeth Josephson, Tim D. Smith, and Randall R. Reeves. Historical distribution of right whales in the North Pacific. Fish and Fisheries, 9(2):155–168, June 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ward:2008:BBC

[294] Trevor J. Ward. Barriers to biodiversity conservation in marine fishery certification. Fish and Fisheries, 9(2):169–177, June 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Katsanevakis:2008:MFG

[295] Stelios Katsanevakis and Christos D. Maravelias. Modelling fish growth: multi-model inference as a better alternative to a priori using von Bertalanffy equation. Fish and Fisheries, 9(2):178–187, June 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

McCluskey:2008:QFE

[296] Shannon M. McCluskey and Rebecca L. Lewison. Quantifying fishing effort: a synthesis of current methods and their applications. Fish and Fisheries, 9(2):188–200, June 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kuparinen:2008:RGH

[297] Anna Kuparinen, Robert B. O’Hara, and Juha Merilä. The role of growth history in determining age and size at maturation in exploited fish populations. Fish and Fisheries, 9(2):201–207, June 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

McDowall:2008:WMB

[298] Robert M. McDowall. Why are so many boreal freshwater fishes anadromous? Confronting ‘conventional wisdom’. Fish and Fisheries, 9(2): 208–213, June 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Filer:2008:FCB

[299] Joshua Filer. Fish Cognition and Behavior. Fish and aquatic resources series 11. Fish and Fisheries, 9(2):214–215, June 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Huse:2008:IBM

[300] Geir Huse. Individual-based modeling and ecology. Fish and Fisheries, 9(2):215–216, June 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Raffaelli:2008:ASN

[301] Dave Raffaelli. Atlantic shorelines. Natural history and ecology. Fish and Fisheries, 9(2):216–217, June 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Smith:2008:SST

[302] Carl Smith. Sensuous seas: Tales of a marine biologist. Fish and Fisheries, 9(2):217–218, June 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Liu:2008:PFC

[303] Min Liu and Yvonne Sadovy De Mitcheson. Profile of a fishery collapse: why mariculture failed to save the large yellow croaker. Fish and Fisheries, 9(3):219–242, September 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hamasaki:2008:EAS

[304] Katsuyuki Hamasaki and Shuichi Kitada. The enhancement of abalone stocks: lessons from Japanese case studies. Fish and Fisheries, 9(3): 243–260, September 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Munday:2008:CCF

[305] Philip L. Munday, Geoffrey P. Jones, Morgan S. Pratchett, and Ashley J. Williams. Climate change and the future for coral reef fishes. Fish and Fisheries, 9(3):261–285, September 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cole:2008:DFI

[306] Andrew J. Cole, Morgan S. Pratchett, and Geoffrey P. Jones. Diversity and functional importance of coral-feeding fishes on tropical coral reefs. Fish and Fisheries, 9(3):286–307, September 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Zhou:2008:FCD

[307] Shijie Zhou. Fishery by-catch and discards: a positive perspective from ecosystem-based fishery management. Fish and Fisheries, 9(3):308–315, September 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Saenz-Arroyo:2008:CFS

[308] Andrea Sáenz-Arroyo and Callum M. Roberts. Consilience in fisheries science. Fish and Fisheries, 9(3):316–327, September 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Haydon:2008:MED

[309] Daniel Haydon. Models for ecological data: An introduction. Fish and Fisheries, 9(3):328–329, September 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hart:2008:OPM

[310] Paul J. B. Hart. Oceans past. Management insights from the history of marine animal populations. Fish and Fisheries, 9(3):329–330, September 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hauser:2008:E

[311] Lorenz Hauser, Robin S. Waples, and Gary R. Carvalho. Editorial. Fish and Fisheries, 9(4):331–332, December 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hauser:2008:PSM

[312] Lorenz Hauser and Gary R. Carvalho. Paradigm shifts in marine fisheries genetics: ugly hypotheses slain by beautiful facts. Fish and Fisheries, 9 (4):333–362, December 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Selkoe:2008:SGS

[313] Kimberly A. Selkoe, Christine M. Henzler, and Steven D. Gaines. Seascape genetics and the spatial ecology of marine populations. Fish and Fisheries, 9(4):363–377, December 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Goetz:2008:FGM

[314] Frederick William Goetz and Simon MacKenzie. Functional genomics with microarrays in fish biology and fisheries. Fish and Fisheries, 9(4): 378–395, December 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Naish:2008:BGB

[315] Kerry A. Naish and Jeffrey J. Hard. Bridging the gap between the genotype and the phenotype: linking genetic variation, selection and adaptation in fishes. Fish and Fisheries, 9(4):396–422, December 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Waples:2008:IGD

[316] Robin S. Waples, André E. Punt, and Jason M. Cope. Integrating genetic data into management of marine resources: how can we do it better? Fish and Fisheries, 9(4):423–449, December 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Nielsen:2008:WDV

[317] Einar Eg Nielsen and Michael Møller Hansen. Waking the dead: the value of population genetic analyses of historical samples. Fish and Fisheries, 9 (4):450–461, December 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ogden:2008:FFU

[318] Rob Ogden. Fisheries forensics: the use of DNA tools for improving compliance, traceability and enforcement in the fishing industry. Fish and Fisheries, 9(4):462–472, December 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hauser:2008:AMT

[319] Lorenz Hauser and James E. Seeb. Advances in molecular technology and their impact on fisheries genetics. Fish and Fisheries, 9(4):473–486, December 2008. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Martell:2009:IFF

[320] Steven J. Martell, Carl Walters, and Ussif Rashid Sumaila. Industryfunded fishing license reduction good for both profits and conservation. Fish and Fisheries, 10(1):1–12, March 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ainley:2009:ERH

[321] David G. Ainley and Louise K. Blight. Ecological repercussions of historical fish extraction from the Southern Ocean. Fish and Fisheries, 10(1):13–38, March 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Branch:2009:HDI

[322] Trevor A. Branch. How do individual transferable quotas affect marine ecosystems? Fish and Fisheries, 10(1):39–57, March 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Link:2009:TRA

[323] Jason S. Link, Bjarte Bogstad, Henrik Sparholt, and George R. Lilly. Trophic role of Atlantic cod in the ecosystem. Fish and Fisheries, 10(1): 58–87, March 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Leprieur:2009:SUA

[324] F. Leprieur, S. Brosse, E. García-Berthou, T. Oberdorff, J. D. Olden, and C. R. Townsend. Scientific uncertainty and the assessment of risks posed by non-native freshwater fishes. Fish and Fisheries, 10(1):88–97, March 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Vitule:2009:INN

[325] Jean Ricardo Simões Vitule, Carolina Arruda Freire, and Daniel Simberloff. Introduction of non-native freshwater fish can certainly be bad. Fish and Fisheries, 10(1):98–108, March 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Gozlan:2009:RRG

[326] R. E. Gozlan. Response by R. Gozlan, Biodiversity crisis and the introduction of non-native fish: Solutions, not scapegoats. Fish and Fisheries, 10(1):109–110, March 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Nevill:2009:RTG

[327] C. Jonathan Nevill. A reef in time: The Great Barrier Reef from beginning to end. Fish and Fisheries, 10(1):111–112, March 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hatfield:2009:FBA

[328] Emma Hatfield. Fisheries biology, assessment and management, 2nd edn. Fish and Fisheries, 10(1):112–113, March 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hickley:2009:GCR

[329] Phil Hickley. Global challenges in recreational fisheries. Fish and Fisheries, 10(1):113–114, March 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Mullon:2009:MMV

[330] Christian Mullon, Pierre Fréon, Philippe Cury, Lynne Shannon, and Claude Roy. A minimal model of the variability of marine ecosystems.

Fish and Fisheries, 10(2):115–131, June 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Serafy:2009:CCH

[331] Joseph E. Serafy, David W. Kerstetter, and Patrick H. Rice. Can circle hook use benefit billfishes? Fish and Fisheries, 10(2):132–142, June 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hardt:2009:LPC

[332] Marah J. Hardt. Lessons from the past: the collapse of Jamaican coral reefs. Fish and Fisheries, 10(2):143–158, June 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Sivan:2009:FVP

[333] Gisha Sivan. Fish venom: pharmacological features and biological significance. Fish and Fisheries, 10(2):159–172, June 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Allison:2009:VNE

[334] Edward H. Allison, Allison L. Perry, Marie-Caroline Badjeck, W. Neil Adger, Katrina Brown, Declan Conway, Ashley S. Halls, Graham M. Pilling, John D. Reynolds, Neil L. Andrew, and Nicholas K. Dulvy. Vulnerability of national economies to the impacts of climate change on fisheries. Fish and Fisheries, 10(2):173–196, June 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Fromentin:2009:LPI

[335] Jean-Marc Fromentin. Lessons from the past: investigating historical data from bluefin tuna fisheries. Fish and Fisheries, 10(2):197–216, June 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Chu:2009:TYL

[336] Cindy Chu. Thirty years later: the global growth of ITQs and their influence on stock status in marine fisheries. Fish and Fisheries, 10(2): 217–230, June 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Betts:2009:AST

[337] Tim Betts. Anaesthetic and sedative techniques for aquatic animals, 3rd edn. Fish and Fisheries, 10(2):231–232, June 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hatcher:2009:FB

[338] Aaron Hatcher. Fisheries buybacks. Fish and Fisheries, 10(2):232–233, June 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cheung:2009:PGM

[339] William W. L. Cheung, Vicky W. Y. Lam, Jorge L. Sarmiento, Kelly Kearney, Reg Watson, and Daniel Pauly. Projecting global marine biodiversity impacts under climate change scenarios. Fish and Fisheries, 10 (3):235–251, September 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Copp:2009:VIB

[340] Gordon H. Copp, J. Robert Britton, Julien Cucherousset, Emili García-Berthou, Ruth Kirk, Edmund Peeler, and Saulius Stakėnas. Voracious invader or benign feline? A review of the environmental biology of European catfish Silurus glanis in its native and introduced ranges. Fish and Fisheries, 10(3):252–282, September 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Wright:2009:FID

[341] Peter J. Wright and Edward A. Trippel. Fishery-induced demographic changes in the timing of spawning: consequences for reproductive success. Fish and Fisheries, 10(3):283–304, September 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Dambacher:2009:QMI

[342] Jeffrey M. Dambacher, Daniel J. Gaughan, Marie-Joëlle Rochet, Philippe A. Rossignol, and Verena M. Trenkel. Qualitative modelling and indicators of exploited ecosystems. Fish and Fisheries, 10(3):305– 322, September 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Field:2009:PMW

[343] Iain C. Field, Mark G. Meekan, Rik C. Buckworth, and Corey J. A. Bradshaw. Protein mining the world’s oceans: Australasia as an example of illegal expansion-and-displacement fishing. Fish and Fisheries, 10(3): 323–328, September 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Chopelet:2009:SCG

[344] Julien Chopelet, Robin S. Waples, and Stefano Mariani. Sex change and the genetic structure of marine fish populations. Fish and Fisheries, 10 (3):329–343, September 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Larsson:2009:PFO

[345] Matz Larsson. Possible functions of the octavolateralis system in fish schooling. Fish and Fisheries, 10(3):344–353, September 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Smith:2009:BTS

[346] Carl Smith. Biology of the three-spined stickleback. Fish and Fisheries, 10(3):354–355, September 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Barber:2009:FW

[347] Iain Barber. Fish welfare. Fish and Fisheries, 10(3):355–357, September 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Pond:2009:DMP

[348] Caroline Pond. Dr. Margaret (Peggy) Varley, pioneer ecophysiologist and ichthyologist. Fish and Fisheries, 10(4):359–360, December 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Reiss:2009:GPS

[349] Henning Reiss, Galice Hoarau, Mark Dickey-Collas, and Wim J. Wolff. Genetic population structure of marine fish: mismatch between biological and fisheries management units. Fish and Fisheries, 10(4):361–395, December 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

To:2009:SBG

[350] Allen W. L. To and Yvonne Sadovy De Mitcheson. Shrinking baseline: the growth in juvenile fisheries, with the Hong Kong grouper fishery as a case study. Fish and Fisheries, 10(4):396–407, December 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bull:2009:NMM

[351] Leigh S. Bull. New mitigation measures reducing seabird by-catch in trawl fisheries. Fish and Fisheries, 10(4):408–427, December 2009. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Alexander:2009:GMC

[352] Karen E. Alexander, William B. Leavenworth, Jamie Cournane, Andrew B. Cooper, Stefan Claesson, Stephen Brennan, Gwynna Smith, Lesley Rains, Katherine Magness, Reneé Dunn, Tristan K. Law, Robert Gee, W. Jeffrey Bolster, and Andrew A. Rosenberg. Gulf of Maine cod in 1861: historical analysis of fishery logbooks, with ecosystem implications. Fish and Fisheries, 10(4):428–449, December 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Guttridge:2009:RLS

[353] Tristan L. Guttridge, Arthur A. Myrberg, Ila F. Porcher, David W. Sims, and Jens Krause. The role of learning in shark behaviour. Fish and Fisheries, 10(4):450–469, December 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Gibbs:2009:ITQ

[354] Mark T. Gibbs. Individual transferable quotas and ecosystem-based fisheries management: it’s all in the T. Fish and Fisheries, 10(4):470–474, December 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Metcalfe:2009:BFE

[355] Neil Metcalfe. Biology of Fishes, 3rd edn. Fish and Fisheries, 10(4): 475–476, December 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bundy:2009:AFS

[356] Alida Bundy. Advances in Fisheries Science: 50 Years on from Beverton and Holt. Fish and Fisheries, 10(4):476–477, December 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Lindley:2009:MAP

[357] J. A. Lindley. A mechanistic approach to plankton ecology. Fish and Fisheries, 10(4):477–478, December 2009. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Davis:2010:FSM

[358] Michael W. Davis. Fish stress and mortality can be predicted using reflex impairment. Fish and Fisheries, 11(1):1–11, March 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Roscoe:2010:EMF

[359] David W. Roscoe and Scott G. Hinch. Effectiveness monitoring of fish passage facilities: historical trends, geographic patterns and future directions. Fish and Fisheries, 11(1):12–33, March 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Stapanian:2010:WSB

[360] Martin A. Stapanian, Vaughn L. Paragamian, Charles P. Madenjian, James R. Jackson, Jyrki Lappalainen, Matthew J. Evenson, and Matthew D. Neufeld. Worldwide status of burbot and conservation measures. Fish and Fisheries, 11(1):34–56, March 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Gilman:2010:MST

[361] Eric Gilman, Jeff Gearhart, Blake Price, Scott Eckert, Henry Milliken, John Wang, Yonat Swimmer, Daisuke Shiode, Osamu Abe, S. Hoyt Peckham, Milani Chaloupka, Martin Hall, Jeff Mangel, Joanna Alfaro-Shigueto, Paul Dalzell, and Asuka Ishizaki. Mitigating sea turtle by-catch in coastal passive net fisheries. Fish and Fisheries, 11(1):57–88, March 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Mangel:2010:RES

[362] Marc Mangel, Jon Brodziak, and Gerard DiNardo. Reproductive ecology and scientific inference of steepness: a fundamental metric of population dynamics and strategic fisheries management. Fish and Fisheries, 11(1): 89–104, March 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Christensen:2010:MM

[363] Villy Christensen. MEY = MSY. Fish and Fisheries, 11(1):105–110, March 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Beecham:2010:CFR

[364] Jonathan Beecham. Computers in Fisheries Research, 2nd edn. Fish and Fisheries, 11(1):111–112, March 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Squires:2010:AFE

[365] Dale Squires. Advances in Fisheries Economics. Fish and Fisheries, 11 (1):112–114, March 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Dimech:2010:MFM

[366] Mark Dimech. Making fisheries management work: Implementation of policies for sustainable fishing. Fish and Fisheries, 11(1):114–115, March 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Pauly:2010:JLM

[367] Daniel Pauly. John Leslie Munro (1938–2009) — a pioneer of tropical fisheries science. Fish and Fisheries, 11(2):117–118, June 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kraak:2010:SDF

[368] Sarah B. M. Kraak, Ciaran J. Kelly, Edward A. Codling, and Emer Rogan. On scientists’ discomfort in fisheries advisory science: the example of simulation-based fisheries management-strategy evaluations. Fish and Fisheries, 11(2):119–132, June 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Polacheck:2010:CTE

[369] Tom Polacheck, J. Paige Eveson, and Geoff M. Laslett. Classifying tagging experiments for commercial fisheries into three fundamental types based on design, data requirements and estimable population parameters. Fish and Fisheries, 11(2):133–148, June 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Gislason:2010:SGT

[370] Henrik Gislason, Niels Daan, Jake C. Rice, and John G. Pope. Size, growth, temperature and the natural mortality of marine fish. Fish and Fisheries, 11(2):149–158, June 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Leclercq:2010:MSC

[371] Eric Leclercq, John F. Taylor, and Hervé Migaud. Morphological skin colour changes in teleosts. Fish and Fisheries, 11(2):159–193, June 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Froese:2010:RFS

[372] Rainer Froese and Alexander Proelß. Rebuilding fish stocks no later than 2015: will Europe meet the deadline? Fish and Fisheries, 11(2): 194–202, June 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Nicol:2010:SOI

[373] Stephen Nicol, Andrew Bowie, Simon Jarman, Delphine Lannuzel, Klaus M. Meiners, and Pier Van Der Merwe. Southern Ocean iron fertilization by baleen whales and Antarctic krill. Fish and Fisheries, 11(2): 203–209, June 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cotter:2010:REC

[374] John Cotter. Reforming the European Common Fisheries Policy: a case for capacity and effort controls. Fish and Fisheries, 11(2):210–219, June 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Griffiths:2010:FCE

[375] Shane P. Griffiths, Kenneth H. Pollock, Jeremy M. Lyle, Julian G. Pepperell, Mark L. Tonks, and William Sawynok. Following the chain to elusive anglers. Fish and Fisheries, 11(2):220–228, June 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Miller:2010:BB

[376] Peter J. Miller. The biology of blennies. Fish and Fisheries, 11(2):229– 231, June 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Head:2010:FB

[377] Megan Head. Fish behaviour. Fish and Fisheries, 11(2):231–232, June 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Mackinson:2010:EAF

[378] Steve Mackinson. The ecosystem approach to fisheries. Fish and Fisheries, 11(2):232, June 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2010:FGS

[379] Anonymous. Fisheries governance — special section. Fish and Fisheries, 11(3):233, September 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Chuenpagdee:2010:FGG

[380] Guest editors Ratana Chuenpagdee and Rashid Sumaila. Fisheries governance and governability. Fish and Fisheries, 11(3):234, September 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Song:2010:OGC

[381] Andrew Moonseok Song and Ratana Chuenpagdee. Operationalizing governability: a case study of a Lake Malawi fishery. Fish and Fisheries, 11(3):235–249, September 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Onyango:2010:APS

[382] Paul Onyango and Svein Jentoft. Assessing poverty in small-scale fisheries in Lake Victoria, Tanzania. Fish and Fisheries, 11(3):250–263, September 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Johnson:2010:IAG

[383] Derek S. Johnson. Institutional adaptation as a governability problem in fisheries: patron–client relations in the Junagadh fishery, India. Fish and Fisheries, 11(3):264–277, September 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Sumaila:2010:DFP

[384] Ussif Rashid Sumaila and Marcos Domínguez-Torreiro. Discount factors and the performance of alternative fisheries governance systems. Fish and Fisheries, 11(3):278–287, September 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bonhommeau:2010:DMA

[385] Sylvain Bonhommeau, Martin Castonguay, Etienne Rivot, Richard Sabatié, and Olivier Le Pape. The duration of migration of Atlantic Anguilla larvae. Fish and Fisheries, 11(3):289–306, September 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Levin:2010:EFW

[386] Phillip S. Levin and Aaron Dufault. Eating up the food web. Fish and Fisheries, 11(3):307–312, September 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Sims:2010:SOO

[387] David Sims. Sharks of the open ocean: Biology, fisheries and conservation. Fish and Fisheries, 11(3):313–314, September 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Gozlan:2010:PCI

[388] Rodolphe E. Gozlan, Demetra Andreou, Takashi Asaeda, Kathleen Beyer, Rachid Bouhadad, Dean Burnard, Nuno Caiola, Predrag Cakic, Vesna Djikanovic, Hamid R. Esmaeili, Istvan Falka, Duncan Golicher, Akos Harka, Galina Jeney, Vladimír Kováč, Jiří Musil, Annamaria Nocita, Meta Povz, Nicolas Poulet, Tomas Virbickas, Christian Wolter, A. Serhan Tarkan, Elena Tricarico, Teodora Trichkova, Hugo Verreycken, Andrzej Witkowski, Chun Guang Zhang, Irene Zweimueller, and J. Robert Britton. Pan-continental invasion of Pseudorasbora parva: towards a better understanding of freshwater fish invasions. Fish and Fisheries, 11(4):315–340, December 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Sethi:2010:RMF

[389] Suresh A. Sethi. Risk management for fisheries. Fish and Fisheries, 11 (4):341–365, December 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Squires:2010:FBR

[390] Dale Squires. Fisheries buybacks: a review and guidelines. Fish and Fisheries, 11(4):366–387, December 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Morissette:2010:WEF

[391] Lyne Morissette, Kristin Kaschner, and Leah R. Gerber. ‘Whales eat fish’? Demystifying the myth in the Caribbean marine ecosystem. Fish and Fisheries, 11(4):388–404, December 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Weeratunge:2010:GFT

[392] Nireka Weeratunge, Katherine A. Snyder, and Choo Poh Sze. Gleaner, fisher, trader, processor: understanding gendered employment in fisheries and aquaculture. Fish and Fisheries, 11(4):405–420, December 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hunsicker:2010:CCG

[393] Mary E. Hunsicker, Timothy E. Essington, Reg Watson, and Ussif R. Sumaila. The contribution of cephalopods to global marine fisheries: can we have our squid and eat them too? Fish and Fisheries, 11(4): 421–438, December 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Copp:2010:PDA

[394] Gordon H. Copp. Patterns of diel activity and species richness in young and small fishes of European streams: a review of 20 years of point abundance sampling by electrofishing. Fish and Fisheries, 11(4):439–460, December 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Sumaila:2010:MEY

[395] U. R. Sumaila and Rögnvaldur Hannesson. Maximum economic yield in crisis? Fish and Fisheries, 11(4):461–465, December 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Metcalfe:2010:DFF

[396] Julian Metcalfe. Do Fish Feel Pain? Fish and Fisheries, 11(4):466–467, December 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2010:E

[397] Anonymous. Erratum. Fish and Fisheries, 11(4):468, December 2010. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2011:AWB

[398] Anonymous. Announcing web-based submission process: ScholarOne manuscripts. Fish and Fisheries, 12(1):1, March 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Fulton:2011:HBK

[399] Elizabeth A. Fulton, Anthony D. M. Smith, David C. Smith, and Ingrid E. van Putten. Human behaviour: the key source of uncertainty in fisheries management. Fish and Fisheries, 12(1):2–17, March 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kraak:2011:EPG

[400] Sarah B. M. Kraak. Exploring the ‘public goods game’ model to overcome the tragedy of the commons in fisheries management. Fish and Fisheries, 12(1):18–33, March 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Villasante:2011:OCF

[401] Sebastián Villasante, María do Carme García-Negro, Fernando González-Laxe, and Gonzalo Rodríguez Rodríguez. Overfishing and the Common Fisheries Policy: (un)successful results from TAC regulation? Fish and Fisheries, 12(1):34–50, March 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Lam:2011:SSE

[402] Vivian Y. Y. Lam and Yvonne Sadovy de Mitcheson. The sharks of South East Asia — unknown, unmonitored and unmanaged. Fish and Fisheries, 12(1):51–74, March 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Friedman:2011:MSC

[403] Kim Friedman, Hampus Eriksson, Emmanuel Tardy, and Kalo Pakoa. Management of sea cucumber stocks: patterns of vulnerability and recovery of sea cucumber stocks impacted by fishing. Fish and Fisheries, 12(1):75–93, March 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Brochier:2011:INH

[404] Timothée Brochier, Christophe Lett, and Pierre Fréon. Investigating the ‘northern Humboldt paradox’ from model comparisons of small pelagic fish reproductive strategies in eastern boundary upwelling ecosystems. Fish and Fisheries, 12(1):94–109, March 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Dunn:2011:STM

[405] Daniel C. Dunn, André M. Boustany, and Patrick N. Halpin. Spatiotemporal management of fisheries to reduce by-catch and increase fishing selectivity. Fish and Fisheries, 12(1):110–119, March 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ward:2011:TLB

[406] Ashley Ward. Tinbergen’s legacy in behaviour: Sixty years of landmark stickleback papers. Fish and Fisheries, 12(1):120–121, March 2011. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Essington:2011:IEB

[407] Timothy E. Essington and André E. Punt. Implementing ecosystembased fisheries management: Advances, challenges and emerging tools. Fish and Fisheries, 12(2):123–124, June 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Jennings:2011:TEA

[408] Simon Jennings and Jake Rice. Towards an ecosystem approach to fisheries in Europe: a perspective on existing progress and future directions. Fish and Fisheries, 12(2):125–137, June 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Constable:2011:LCI

[409] Andrew J. Constable. Lessons from CCAMLR on the implementation of the ecosystem approach to managing fisheries. Fish and Fisheries, 12(2):138–151, June 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Link:2011:EBF

[410] Jason S. Link, Alida Bundy, William J. Overholtz, Nancy Shackell, John Manderson, Daniel Duplisea, Jon Hare, Mariano Koen-Alonso, and Kevin D. Friedland. Ecosystem-based fisheries management in the Northwest Atlantic. Fish and Fisheries, 12(2):152–170, June 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Fulton:2011:LMM

[411] Elizabeth A. Fulton, Jason S. Link, Isaac C. Kaplan, Marie Savina-Rolland, Penelope Johnson, Cameron Ainsworth, Peter Horne, Rebecca Gorton, Robert J. Gamble, Anthony D. M. Smith, and David C. Smith. Lessons in modelling and management of marine ecosystems: the Atlantis experience. Fish and Fisheries, 12(2):171–188, June 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hollowed:2011:EQE

[412] Anne B. Hollowed, Kerim Y. Aydin, Timothy E. Essington, James N. Ianelli, Bernard A. Megrey, André E. Punt, and Anthony D. M. Smith. Experience with quantitative ecosystem assessment tools in the northeast Pacific. Fish and Fisheries, 12(2):189–208, June 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Rice:2011:MFW

[413] Jake Rice. Managing fisheries well: delivering the promises of an ecosystem approach. Fish and Fisheries, 12(2):209–231, June 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Jellyman:2011:BMS

[414] Donald Jellyman. Bob McDowall (15 September 1939–20 February 2011). Fish and Fisheries, 12(3):233–234, September 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Otero:2011:CKN

[415] Olga Otero. Current knowledge and new assumptions on the evolutionary history of the African lungfish, Protopterus, based on a review of its fossil record. Fish and Fisheries, 12(3):235–255, September 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Britton:2011:MNN

[416] J. R. Britton, Rodolphe E. Gozlan, and Gordon H. Copp. Managing non-native fish in the environment. Fish and Fisheries, 12(3):256–274, September 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cochrane:2011:PFM

[417] Kevern L. Cochrane, Neil L. Andrew, and Ana M. Parma. Primary fisheries management: a minimum requirement for provision of sustainable human benefits in small-scale fisheries. Fish and Fisheries, 12(3): 275–288, September 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cardinale:2011:HSB

[418] Massimiliano Cardinale, Valerio Bartolino, Marcos Llope, Luigi Maiorano, Mattias Sköld, and Jacob Hagberg. Historical spatial baselines in conservation and management of marine resources. Fish and Fisheries, 12 (3):289–298, September 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Casale:2011:STC

[419] Paolo Casale. Sea turtle by-catch in the Mediterranean. Fish and Fisheries, 12(3):299–316, September 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anderson:2011:SEG

[420] Sean C. Anderson, Joanna Mills Flemming, Reg Watson, and Heike K. Lotze. Serial exploitation of global sea cucumber fisheries. Fish and Fisheries, 12(3):317–339, September 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Froese:2011:GHC

[421] Rainer Froese, Trevor A. Branch, Alexander Proelß, Martin Quaas, Keith Sainsbury, and Christopher Zimmermann. Generic harvest control rules for European fisheries. Fish and Fisheries, 12(3):340–351, September 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Gill:2011:EAS

[422] Andrew B. Gill. Ecology of Aquatic Systems. Fish and Fisheries, 12 (3):352, September 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Turley:2011:OAN

[423] Carol Turley. Ocean acidification. A national strategy to meet the challenges of a changing ocean. Fish and Fisheries, 12(3):352–354, September 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cochrane:2011:EBF

[424] Kevern Cochrane. Ecosystem-based fisheries management: Confronting trade-offs. Fish and Fisheries, 12(3):354–356, September 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Foale:2011:TTO

[425] Simon Foale, Philippa Cohen, Stephanie Januchowski-Hartley, Amelia Wenger, and Martha Macintyre. Tenure and taboos: origins and implications for fisheries in the Pacific. Fish and Fisheries, 12(4):357–369, December 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Holbrook:2011:FNV

[426] Robert I. Holbrook and Theresa Burt de Perera. Fish navigation in the vertical dimension: can fish use hydrostatic pressure to determine depth? Fish and Fisheries, 12(4):370–379, December 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kettle:2011:WOE

[427] A. James Kettle, L. Asbjørn Vøllestad, and Joanna Wibig. Where once the eel and the elephant were together: decline of the European eel because of changing hydrology in southwest Europe and northwest Africa? Fish and Fisheries, 12(4):380–411, December 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Vandeperre:2011:ENT

[428] Frederic Vandeperre, Ruth M. Higgins, Julio Sánchez-Meca, Francesc Maynou, Raquel Goñi, Pablo Martín-Sosa, Angel Pérez-Ruzafa, Pedro Afonso, Iacopo Bertocci, Romain Crec’hriou, Giovanni D’Anna, Mark Dimech, Carmelo Dorta, Oscar Esparza, Jesús M. Falcón, Aitor Forcada, Ivan Guala, Laurence Le Direach, Concepción Marcos, Celia Ojeda-Martínez, Carlo Pipitone, Patrick J. Schembri, Vanessa Stelzenmüller, Ben Stobart, and Ricardo S. Santos. Effects of no-take area size and age of marine protected areas on fisheries yields: a meta-analytical approach. Fish and Fisheries, 12(4):412–426, December 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Perry:2011:MSE

[429] R. Ian Perry, Rosemary E. Ommer, Manuel Barange, Svein Jentoft, Barbara Neis, and U. Rashid Sumaila. Marine social–ecological responses to environmental change and the impacts of globalization. Fish and Fisheries, 12(4):427–450, December 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Madenjian:2011:SEP

[430] Charles P. Madenjian. Sex effect on polychlorinated biphenyl concentrations in fish: a synthesis. Fish and Fisheries, 12(4):451–460, December 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Link:2011:GIF

[431] Jason S. Link, Janet A. Nye, and Jonathan A. Hare. Guidelines for incorporating fish distribution shifts into a fisheries management context. Fish and Fisheries, 12(4):461–469, December 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hart:2011:MEG

[432] Paul J. B. Hart. Marine ecosystems and global change. Fish and Fisheries, 12(4):470–471, December 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hart:2011:FOO

[433] Paul J. B. Hart. Fishes of the open ocean: A natural history and illustrated guide. Fish and Fisheries, 12(4):471–472, December 2011. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kjesbu:2011:FR

[434] Olav S. Kjesbu. Fish reproduction. Fish and Fisheries, 12(4):472–473, December 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Day:2011:GOR

[435] Julia J. Day. Gonorynchiformes and Ostariophysan relationships: A comprehensive review, teleostean fish biology. Fish and Fisheries, 12(4): 473–474, December 2011. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Tlusty:2012:EIS

[436] Michael F. Tlusty. Environmental improvement of seafood through certification and ecolabelling: theory and analysis. Fish and Fisheries, 13(1): 1–13, March 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Allison:2012:RBF

[437] Edward H. Allison, Blake D. Ratner, Björn Åsgård, Rolf Willmann, Robert Pomeroy, and John Kurien. Rights-based fisheries governance: from fishing rights to human rights. Fish and Fisheries, 13(1):14–29, March 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Nicol:2012:FAK

[438] Stephen Nicol, Jacqueline Foster, and So Kawaguchi. The fishery for Antarctic krill — recent developments. Fish and Fisheries, 13(1):30– 40, March 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ahrens:2012:FAT

[439] Robert N. M. Ahrens, Carl J. Walters, and Villy Christensen. Foraging arena theory. Fish and Fisheries, 13(1):41–59, March 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Coll:2012:CFW

[440] Marta Coll and Simone Libralato. Contributions of food web modelling to the ecosystem approach to marine resource management in the Mediterranean Sea. Fish and Fisheries, 13(1):60–88, March 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Sampson:2012:ESS

[441] David B. Sampson and Robert D. Scott. An exploration of the shapes and stability of population–selection curves. Fish and Fisheries, 13(1): 89–104, March 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Brown:2012:TUF

[442] Culum Brown. Tool use in fishes. Fish and Fisheries, 13(1):105–115, March 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Costa:2012:FDV

[443] Graça Costa. Fish defenses, volume 1: Immunology; volume 2: Pathogens, parasites and predators. Fish and Fisheries, 13(1):116–117, March 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Craig:2012:DCF

[444] John F. Craig. E. David Le Cren, 1922–2011 (freshwater fish ecologist). Fish and Fisheries, 13(1):118–120, March 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Gustafson:2012:CSE

[445] Richard G. Gustafson, Michael J. Ford, Peter B. Adams, Jonathan S. Drake, Robert L. Emmett, Kurt L. Fresh, Mindy Rowse, Elizabeth A. K. Spangler, Robert E. Spangler, David J. Teel, and Matthew T. Wilson. Conservation status of eulachon in the California Current. Fish and Fisheries, 13(2):121–138, June 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Dowling:2012:AOR

[446] Natalie A. Dowling, Chris Wilcox, Marc Mangel, and Sean Pascoe. Assessing opportunity and relocation costs of marine protected areas using a behavioural model of longline fleet dynamics. Fish and Fisheries, 13 (2):139–157, June 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kemp:2012:QQE

[447] Paul S. Kemp, Tom A. Worthington, Terence E. L. Langford, Angus R. J. Tree, and Martin J. Gaywood. Qualitative and quantitative effects of reintroduced beavers on stream fish. Fish and Fisheries, 13(2):158– 181, June 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cooley:2012:NIM

[448] Sarah R. Cooley, Noelle Lucey, Hauke Kite-Powell, and Scott C. Doney. Nutrition and income from molluscs today imply vulnerability to ocean acidification tomorrow. Fish and Fisheries, 13(2):182–215, June 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

vanPutten:2012:TBD

[449] Ingrid E. van Putten, Soile Kulmala, Olivier Thébaud, Natalie Dowling, Katell G. Hamon, Trevor Hutton, and Sean Pascoe. Theories and behavioural drivers underlying fleet dynamics models. Fish and Fisheries, 13(2):216–235, June 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hart:2012:MMF

[450] Paul J. B. Hart. Mismanagement of marine fisheries; all the fish in the sea. Fish and Fisheries, 13(2):236–237, June 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Fock:2012:SEF

[451] Heino O. Fock. Seamounts: Ecology, fisheries & conservation. Fish and Fisheries, 13(2):238–239, June 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Griffiths:2012:EEP

[452] Siân W. Griffiths. Ecology and evolution of poeciliid fishes. Fish and Fisheries, 13(2):239–240, June 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

LaMesa:2012:ASL

[453] Mario La Mesa and Joseph T. Eastman. Antarctic silverfish: life strategies of a key species in the high-Antarctic ecosystem. Fish and Fisheries, 13(3):241–266, September 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Melnychuk:2012:CCS

[454] Michael C. Melnychuk, Timothy E. Essington, Trevor A. Branch, Selina S. Heppell, Olaf P. Jensen, Jason S. Link, Steven J. D. Martell, Ana M. Parma, John G. Pope, and Anthony D. M. Smith. Can catch share fisheries better track management targets? Fish and Fisheries, 13 (3):267–290, September 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Jacobsen:2012:ERD

[455] Rikke B. Jacobsen, Douglas C. K. Wilson, and Paulina Ramirez-Monsalve. Empowerment and regulation — dilemmas in participatory fisheries science. Fish and Fisheries, 13(3):291–302, September 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Grafton:2012:BFM

[456] Quentin R. Grafton, Tom Kompas, Tuong N. Che, Long Chu, and Ray Hilborn. B_{M}EY as a fisheries management target. Fish and Fisheries, 13 (3):303–312, September 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Glover:2012:NMW

[457] Kevin A. Glover, Tore Haug, Nils Øien, Lars Walløe, Lotta Lindblom, Bjørghild B. Seliussen, and Hans J. Skaug. The Norwegian minke whale DNA register: a data base monitoring commercial harvest and trade of whale products. Fish and Fisheries, 13(3):313–332, September 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

LeQuesne:2012:PIO

[458] Will J. F. Le Quesne and John K. Pinnegar. The potential impacts of ocean acidification: scaling from physiology to fisheries. Fish and Fisheries, 13(3):333–344, September 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Miller:2012:SMC

[459] Dana Miller, Adam Jessel, and Stefano Mariani. Seafood mislabelling: comparisons of two western European case studies assist in defining influencing factors, mechanisms and motives. Fish and Fisheries, 13(3): 345–358, September 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Piccolo:2012:GFP

[460] John Piccolo. Gasping fish and panting squids: Oxygen, temperature and the growth of water-breathing animals. Fish and Fisheries, 13(3): 359, September 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bergman:2012:EAS

[461] Eva Bergman. Ecology of Atlantic salmon and brown trout — habitat as a template for life histories. Fish and Fisheries, 13(3):360, September 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Pinca:2012:RIH

[462] S. Pinca, M. Kronen, F. Magron, B. McArdle, L. Vigliola, M. Kulbicki, and S. Andréfouët. Relative importance of habitat and fishing in influencing reef fish communities across seventeen Pacific Island Countries and Territories. Fish and Fisheries, 13(4):361–379, December 2012. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ricard:2012:EKB

[463] Daniel Ricard, Cóilín Minto, Olaf P. Jensen, and Julia K. Baum. Examining the knowledge base and status of commercially exploited marine species with the RAM Legacy Stock Assessment Database. Fish and Fisheries, 13(4):380–398, December 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Jacoby:2012:SBS

[464] David M. P. Jacoby, Darren P. Croft, and David W. Sims. Social behaviour in sharks and rays: analysis, patterns and implications for conservation. Fish and Fisheries, 13(4):399–417, December 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Piccolo:2012:CEL

[465] John J. Piccolo, Johnny R. Norrgård, Larry A. Greenberg, Monika Schmitz, and Eva Bergman. Conservation of endemic landlocked salmonids in regulated rivers: a case-study from Lake Vänern, Sweden. Fish and Fisheries, 13(4):418–433, December 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Gibbs:2012:BIT

[466] Mark T. Gibbs and Olivier Thébaud. Beyond Individual Transferrable Quotas: methodologies for integrating ecosystem impacts of fishing into fisheries catch rights. Fish and Fisheries, 13(4):434–449, December 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Noonan:2012:QAF

[467] Michael J. Noonan, James W. A. Grant, and Christopher D. Jackson. A quantitative assessment of fish passage efficiency. Fish and Fisheries, 13 (4):450–464, December 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Berkes:2012:IEB

[468] Fikret Berkes. Implementing ecosystem-based management: evolution or revolution? Fish and Fisheries, 13(4):465–476, December 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Barber:2012:FPP

[469] Iain Barber. Fish parasites: pathobiology and protection. Fish and Fisheries, 13(4):477, December 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Frid:2012:EBM

[470] Chris Frid. Ecosystem-based management for marine fisheries: an evolving perspective. Fish and Fisheries, 13(4):478, December 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2012:E

[471] Anonymous. Erratum. Fish and Fisheries, 13(4):479–480, December 2012. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Gale:2013:RTC

[472] Marika Kirstin Gale, Scott G. Hinch, and Michael R. Donaldson. The role of temperature in the capture and release of fish. Fish and Fisheries, 14 (1):1–33, March 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Purcell:2013:SCF

[473] Steven W. Purcell, Annie Mercier, Chantal Conand, Jean-François Hamel, M. Verónica Toral-Granda, Alessandro Lovatelli, and Sven Uthicke. Sea cucumber fisheries: global analysis of stocks, management measures and drivers of overfishing. Fish and Fisheries, 14(1):34–59, March 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Senior:2013:CAC

[474] Alistair M. Senior and Shinichi Nakagawa. A comparative analysis of chemically induced sex reversal in teleosts: challenging conventional suppositions. Fish and Fisheries, 14(1):60–76, March 2013. CODEN FFI-IAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Teh:2013:CMF

[475] Lydia C. L. Teh and U. R. Sumaila. Contribution of marine fisheries to worldwide employment. Fish and Fisheries, 14(1):77–88, March 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Colloca:2013:RMF

[476] Francesco Colloca, Massimiliano Cardinale, Francesc Maynou, Marianna Giannoulaki, Giuseppe Scarcella, Klavdija Jenko, Josè Maria Bellido, and Fabio Fiorentino. Rebuilding Mediterranean fisheries: a new paradigm for ecological sustainability. Fish and Fisheries, 14(1):89–109, March 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hinz:2013:CFE

[477] H. Hinz, L. G. Murray, G. I. Lambert, J. G. Hiddink, and M. J. Kaiser. Confidentiality over fishing effort data threatens science and management progress. Fish and Fisheries, 14(1):110–117, March 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

deMitcheson:2013:FGT

[478] Yvonne Sadovy de Mitcheson, Matthew T. Craig, Athila A. Bertoncini, Kent E. Carpenter, William W. L. Cheung, John H. Choat, Andrew S. Cornish, Sean T. Fennessy, Beatrice P. Ferreira, Philip C. Heemstra, Min Liu, Robert F. Myers, David A. Pollard, Kevin L. Rhodes, Luiz A. Rocha, Barry C. Russell, Melita A. Samoilys, and Jonnell Sanciangco. Fishing groupers towards extinction: a global assessment of threats and extinction risks in a billion dollar fishery. Fish and Fisheries, 14(2):119– 136, June 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Fenichel:2013:MAB

[479] Eli P. Fenichel, Joshua K. Abbott, and Biao Huang. Modelling angler behaviour as a part of the management system: synthesizing a multidisciplinary literature. Fish and Fisheries, 14(2):137–157, June 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Liu:2013:EEM

[480] Yajie Liu, Ola H. Diserud, Kjetil Hindar, and Anders Skonhoft. An ecological–economic model on the effects of interactions between escaped farmed and wild salmon (Salmo salar). Fish and Fisheries, 14(2):158– 173, June 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ye:2013:RGF

[481] Yimin Ye, Kevern Cochrane, Gabriella Bianchi, Rolf Willmann, Jacek Majkowski, Merete Tandstad, and Fabio Carocci. Rebuilding global fisheries: the World Summit Goal, costs and benefits. Fish and Fisheries, 14(2):174–185, June 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Delcourt:2013:VMF

[482] Johann Delcourt, Mathieu Denoël, Marc Ylieff, and Pascal Poncin. Video multitracking of fish behaviour: a synthesis and future perspectives. Fish and Fisheries, 14(2):186–204, June 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Burrow:2013:VVD

[483] Jennifer F. Burrow, Joe W. Horwood, and Jonathan W. Pitchford. Variable variability: difficulties in estimation and consequences for fisheries management. Fish and Fisheries, 14(2):205–212, June 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Charnov:2013:EAR

[484] Eric L. Charnov, Henrik Gislason, and John G. Pope. Evolutionary assembly rules for fish life histories. Fish and Fisheries, 14(2):213–224, June 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Xu:2013:WRV

[485] Cailin Xu, David C. Schneider, and Cassandra Rideout. When reproductive value exceeds economic value: an example from the Newfoundland cod fishery. Fish and Fisheries, 14(2):225–233, June 2013. CODEN FFI-IAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Mason:2013:FA

[486] Debby Mason. Fish in Art. Fish and Fisheries, 14(2):234–235, June 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Freshwater:2013:SCN

[487] Cameron Freshwater, James P. W. Robinson, Cayla Naumann, Eric Hertz, David Stormer, Amy K. Teffer, and Francis Juanes. Scientific communication for natural resource professionals. Fish and Fisheries, 14(2):235–236, June 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2013:E

[488] Anonymous. Erratum. Fish and Fisheries, 14(2):237, June 2013. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

McClenachan:2013:MTS

[489] Loren McClenachan and John N. Kittinger. Multicentury trends and the sustainability of coral reef fisheries in Hawai’i and Florida. Fish and Fisheries, 14(3):239–255, September 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Johnson:2013:LTD

[490] Andrew F. Johnson, Stuart R. Jenkins, Jan G. Hiddink, and Hilmar Hinz. Linking temperate demersal fish species to habitat: scales, patterns and future directions. Fish and Fisheries, 14(3):256–280, September 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Johnson:2013:TCU

[491] Ayana E. Johnson, Joshua E. Cinner, Marah J. Hardt, Jennifer Jacquet, Tim R. McClanahan, and James N. Sanchirico. Trends, current understanding and future research priorities for artisanal coral reef fisheries research. Fish and Fisheries, 14(3):281–292, September 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kleisner:2013:UGC

[492] K. Kleisner, D. Zeller, R. Froese, and D. Pauly. Using global catch data for inferences on the world’s marine fisheries. Fish and Fisheries, 14(3): 293–311, September 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Brummett:2013:FAE

[493] Randall E. Brummett, Malcolm C. M. Beveridge, and Ian G. Cowx. Functional aquatic ecosystems, inland fisheries and the Millennium Development Goals. Fish and Fisheries, 14(3):312–324, September 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Magnusson:2013:MUF

[494] Arni Magnusson, André E. Punt, and Ray Hilborn. Measuring uncertainty in fisheries stock assessment: the delta method, bootstrap, and MCMC. Fish and Fisheries, 14(3):325–342, September 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ainley:2013:DTA

[495] David G. Ainley, Nadav Nur, Joseph T. Eastman, Grant Ballard, Claire L. Parkinson, Clive W. Evans, and Arthur L. DeVries. Decadal trends in abundance, size and condition of Antarctic toothfish in Mc-Murdo Sound, Antarctica, 1972–2011. Fish and Fisheries, 14(3):343–363, September 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cox:2013:CFO

[496] Jonathan P. L. Cox. Ciliary function in the olfactory organs of sharks and rays. Fish and Fisheries, 14(3):364–390, September 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Dagorn:2013:IGB

[497] Laurent Dagorn, Kim N. Holland, Victor Restrepo, and Gala Moreno. Is it good or bad to fish with FADs? What are the real impacts of the use of drifting FADs on pelagic marine ecosystems? Fish and Fisheries, 14(3): 391–415, September 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hermoso:2013:REI

[498] Virgilio Hermoso and Miguel Clavero. Revisiting ecological integrity 30 years later: non-native species and the misdiagnosis of freshwater ecosystem health. Fish and Fisheries, 14(3):416–423, September 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bakun:2013:OEP

[499] Andrew Bakun. Ocean eddies, predator pits and bluefin tuna: implications of an inferred ‘low risk–limited payoff’ reproductive scheme of a (former) archetypical top predator. Fish and Fisheries, 14(3):424–438, September 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cooke:2013:VIB

[500] Steven J. Cooke, Cory D. Suski, Robert Arlinghaus, and Andy J. Danylchuk. Voluntary institutions and behaviours as alternatives to formal regulations in recreational fisheries management. Fish and Fisheries, 14(4):439–457, December 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Chang:2013:TGI

[501] Mei-Yu Chang and Audrey J. Geffen. Taxonomic and geographic influences on fish otolith microchemistry. Fish and Fisheries, 14(4):458–492, December 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Watson:2013:GMY

[502] Reg A. Watson, William W. L. Cheung, Jonathan A. Anticamara, Rashid U. Sumaila, Dirk Zeller, and Daniel Pauly. Global marine yield halved as fishing intensity redoubles. Fish and Fisheries, 14(4):493–503, December 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Martell:2013:SME

[503] Steven Martell and Rainer Froese. A simple method for estimating MSY from catch and resilience. Fish and Fisheries, 14(4):504–514, December 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kaplan:2013:CIF

[504] Isaac C. Kaplan, Iris A. Gray, and Phillip S. Levin. Cumulative impacts of fisheries in the California Current. Fish and Fisheries, 14(4):515–527, December 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Li:2013:IAP

[505] Mao Li and John F. Leatherland. The implications for aquaculture practice of epigenomic programming of components of the endocrine system of teleostean embryos: lessons learned from mammalian studies. Fish and Fisheries, 14(4):528–553, December 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Johnston:2013:FLH

[506] Fiona D. Johnston, Robert Arlinghaus, and Ulf Dieckmann. Fish life history, angler behaviour and optimal management of recreational fisheries. Fish and Fisheries, 14(4):554–579, December 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

McLaughlin:2013:UCT

[507] Robert L. McLaughlin, Eric R. B. Smyth, Theodore Castro-Santos, Michael L. Jones, Marten A. Koops, Thomas C. Pratt, and Luis-Antonio Vélez-Espino. Unintended consequences and trade-offs of fish passage.

Fish and Fisheries, 14(4):580–604, December 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Handegard:2013:TAB

[508] Nils Olav Handegard, Louis du Buisson, Patrice Brehmer, Stewart J. Chalmers, Alex De Robertis, Geir Huse, Rudy Kloser, Gavin Macaulay, Olivier Maury, Patrick H. Ressler, Nils Chr Stenseth, and Olav Rune Godø. Towards an acoustic-based coupled observation and modelling system for monitoring and predicting ecosystem dynamics of the open ocean. Fish and Fisheries, 14(4):605–615, December 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Sjostedt:2013:HVR

[509] Martin Sjöstedt. Horizontal and vertical resource dilemmas in natural resource management: the case of African fisheries. Fish and Fisheries, 14 (4):616–624, December 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Salomon:2013:TSF

[510] Markus Salomon and Karin Holm-Müller. Towards a sustainable fisheries policy in Europe. Fish and Fisheries, 14(4):625–638, December 2013. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Plaganyi:2014:MFM

[511] Éva E. Plagányi, André E. Punt, Richard Hillary, Elisabetta B. Morello, Olivier Thébaud, Trevor Hutton, Richard D. Pillans, James T. Thorson, Elizabeth A. Fulton, Anthony D. M. Smith, Franz Smith, Peter Bayliss, Michael Haywood, Vincent Lyne, and Peter C. Rothlisberg. Multispecies fisheries management and conservation: tactical applications using models of intermediate complexity. Fish and Fisheries, 15(1):1–22, March 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Drake:2014:HMS

[512] D. Andrew R. Drake and Nicholas E. Mandrak. Harvest models and stock co-occurrence: probabilistic methods for estimating bycatch. Fish and Fisheries, 15(1):23–42, March 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Pikitch:2014:GCF

[513] Ellen K. Pikitch, Konstantine J. Rountos, Timothy E. Essington, Christine Santora, Daniel Pauly, Reg Watson, Ussif R. Sumaila, P. Dee Boersma, Ian L. Boyd, David O. Conover, Philippe Cury, Selina S. Heppell, Edward D. Houde, Marc Mangel, Éva Plagányi, Keith Sainsbury, Robert S. Steneck, Tess M. Geers, Natasha Gownaris, and Stephan B. Munch. The global contribution of forage fish to marine fisheries and ecosystems. Fish and Fisheries, 15(1):43–64, March 2014. CODEN FFI-IAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Laugen:2014:EIA

[514] Ane T. Laugen, Georg H. Engelhard, Rebecca Whitlock, Robert Arlinghaus, Dorothy J. Dankel, Erin S. Dunlop, Anne M. Eikeset, Katja Enberg, Christian Jørgensen, Shuichi Matsumura, Sébastien Nusslé, Davnah Urbach, Loïc Baulier, David S. Boukal, Bruno Ernande, Fiona D. Johnston, Fabian Mollet, Heidi Pardoe, Nina O. Therkildsen, Silva Uusi-Heikkilä, Anssi Vainikka, Mikko Heino, Adriaan D. Rijnsdorp, and Ulf Dieckmann. Evolutionary impact assessment: accounting for evolutionary consequences of fishing in an ecosystem approach to fisheries management. Fish and Fisheries, 15(1):65–96, March 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Rose:2014:CFR

[515] J. D. Rose, R. Arlinghaus, S. J. Cooke, B. K. Diggles, W. Sawynok, E. D. Stevens, and C. D. L. Wynne. Can fish really feel pain? Fish and Fisheries, 15(1):97–133, March 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Nunan:2014:WWC

[516] Fiona Nunan. Wealth and welfare? Can fisheries management succeed in achieving multiple objectives? A case study of Lake Victoria, East Africa. Fish and Fisheries, 15(1):134–150, March 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Emery:2014:DRF

[517] Timothy J. Emery, Klaas Hartmann, Bridget S. Green, Caleb Gardner, and John Tisdell. Does ‘race to fish’ behaviour emerge in an individual transferable quota fishery when the total allowable catch becomes nonbinding? Fish and Fisheries, 15(1):151–169, March 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Baker:2014:FGC

[518] Ronald Baker, Amanda Buckland, and Marcus Sheaves. Fish gut content analysis: robust measures of diet composition. Fish and Fisheries, 15(1): 170–177, March 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hickley:2014:BFF

[519] Phil Hickley. Britain’s freshwater fishes. Fish and Fisheries, 15(1):178, March 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hart:2014:BRB

[520] Paul J. B. Hart. Book review: Billion-Dollar Fish. The Untold Story of Alaska Pollock. by Bailey, Kevin M., University of Chicago Press, Chicago and London, 2013, GBP 17.50, USD 25.00 ISBN-13: 978-0-226-02234-5 (cloth); 978-0-226-02248-2 (e-book). 271pp., graphs, tables and 19 black and white photographs. Fish and Fisheries, 15(1):179– 180, March 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Teletchea:2014:LDF

[521] Fabrice Teletchea and Pascal Fontaine. Levels of domestication in fish: implications for the sustainable future of aquaculture. Fish and Fisheries, 15(2):181–195, June 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Yamazaki:2014:BMT

[522] Satoshi Yamazaki, Quentin R. Grafton, Tom Kompas, and Sarah Jennings. Biomass management targets and the conservation and economic benefits of marine reserves. Fish and Fisheries, 15(2):196–208, June 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kang:2014:MCF

[523] Bin Kang, Junming Deng, Yunfei Wu, Liqiang Chen, Jie Zhang, Haiyuan Qiu, Yin Lu, and Daming He. Mapping China’s freshwater fishes: diversity and biogeography. Fish and Fisheries, 15(2):209–230, June 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Watson:2014:PPD

[524] Reg Watson, Dirk Zeller, and Daniel Pauly. Primary productivity demands of global fishing fleets. Fish and Fisheries, 15(2):231–241, June 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Frisk:2014:MBC

[525] Michael G. Frisk, Adrian Jordaan, and Thomas J. Miller. Moving beyond the current paradigm in marine population connectivity: are adults the missing link? Fish and Fisheries, 15(2):242–254, June 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Weeratunge:2014:SSF

[526] Nireka Weeratunge, Christophe Béné, Rapti Siriwardane, Anthony Charles, Derek Johnson, Edward H. Allison, Prateep K. Nayak, and Marie-Caroline Badjeck. Small-scale fisheries through the wellbeing lens. Fish and Fisheries, 15(2):255–279, June 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Adams:2014:GCS

[527] Aaron J. Adams, Andrij Z. Horodysky, Richard S. McBride, Kathryn Guindon, Jonathan Shenker, Timothy C. MacDonald, Heather D. Harwell, Rocky Ward, and Kent Carpenter. Global conservation status and research needs for tarpons (Megalopidae), ladyfishes (Elopidae) and bonefishes (Albulidae). Fish and Fisheries, 15(2):280–311, June 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Fitzgibbon:2014:AHS

[528] Quinn P. Fitzgibbon, Andrew G. Jeffs, and Stephen C. Battaglene. The Achilles heel for spiny lobsters: the energetics of the non-feeding postlarval stage. Fish and Fisheries, 15(2):312–326, June 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Gilman:2014:PRF

[529] Eric Gilman, Kelvin Passfield, and Katrina Nakamura. Performance of regional fisheries management organizations: ecosystem-based governance of bycatch and discards. Fish and Fisheries, 15(2):327–351, June 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Monk:2014:HLS

[530] Jacquomo Monk. How long should we ignore imperfect detection of species in the marine environment when modelling their distribution? Fish and Fisheries, 15(2):352–358, June 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Dunn:2014:EMR

[531] Daniel C. Dunn, André M. Boustany, Jason J. Roberts, Eric Brazer, Melissa Sanderson, Beth Gardner, and Patrick N. Halpin. Empirical move-on rules to inform fishing strategies: a New England case study. Fish and Fisheries, 15(3):359–375, September 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Yue:2014:RAG

[532] Gen Hua Yue. Recent advances of genome mapping and marker-assisted selection in aquaculture. Fish and Fisheries, 15(3):376–396, September 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

vanWeerden:2014:MAS

[533] J. Fransje van Weerden, Daniel A. P. Reid, and Charlotte K. Hemelrijk. A meta-analysis of steady undulatory swimming. Fish and Fisheries, 15 (3):397–409, September 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Baumgartner:2014:UFG

[534] Lee J. Baumgartner, John Conallin, Ian Wooden, Bruce Campbell, Rebecca Gee, Wayne A. Robinson, and Martin Mallen-Cooper. Using flow guilds of freshwater fish in an adaptive management framework to simplify environmental flow delivery for semi-arid riverine systems. Fish and Fisheries, 15(3):410–427, September 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Johnsen:2014:FGP

[535] Jahn P. Johnsen. Is fisheries governance possible? Fish and Fisheries, 15 (3):428–444, September 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Watson:2014:CCT

[536] Reg Watson and Daniel Pauly. Coastal catch transects as a tool for studying global fisheries. Fish and Fisheries, 15(3):445–455, September 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Radinger:2014:PPF

[537] Johannes Radinger and Christian Wolter. Patterns and predictors of fish dispersal in rivers. Fish and Fisheries, 15(3):456–473, September 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Pauly:2014:CDW

[538] Daniel Pauly, Dyhia Belhabib, Roland Blomeyer, William W. W. L. Cheung, Andrés M. Cisneros-Montemayor, Duncan Copeland, Sarah Harper, Vicky W. Y. Lam, Yining Mai, Frédéric Le Manach, Henrik Österblom, Ka Man Mok, Liesbeth van der Meer, Antonio Sanz, Soohyun Shon, U. Rashid Sumaila, Wilf Swartz, Reg Watson, Yunlei Zhai, and Dirk Zeller. China’s distant-water fisheries in the 21st century. Fish and Fisheries, 15(3):474–488, September 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Raby:2014:UUR

[539] Graham D. Raby, Jessica R. Packer, Andy J. Danylchuk, and Steven J. Cooke. The understudied and underappreciated role of predation in the mortality of fish released from fishing gears. Fish and Fisheries, 15(3): 489–505, September 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Thurstan:2014:OBT

[540] Ruth H. Thurstan, Julie P. Hawkins, and Callum M. Roberts. Origins of the bottom trawling controversy in the British Isles: 19th century witness testimonies reveal evidence of early fishery declines. Fish and Fisheries, 15(3):506–522, September 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Vilizzi:2014:AOD

[541] Lorenzo Vilizzi and Vladimir Kováč. Alternative ontogenies and developmental plasticity: implications for ecological and evolutionary studies on species complexes. Fish and Fisheries, 15(3):523–531, September 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kenchington:2014:NME

[542] Trevor J. Kenchington. Natural mortality estimators for informationlimited fisheries. Fish and Fisheries, 15(4):533–562, December 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Vincent:2014:RCC

[543] Amanda C. J. Vincent, Yvonne J. Sadovy de Mitcheson, Sarah L. Fowler, and Susan Lieberman. The role of CITES in the conservation of marine fishes subject to international trade. Fish and Fisheries, 15(4):563–592, December 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Feary:2014:LSC

[544] David A. Feary, Morgan S. Pratchett, Micheal J. Emslie, Ashley M. Fowler, Will F. Figueira, Osmar J. Luiz, Yohei Nakamura, and David J. Booth. Latitudinal shifts in coral reef fishes: why some species do and others do not shift. Fish and Fisheries, 15(4):593–615, December 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Glaser:2014:CDM

[545] Sarah M. Glaser, Michael J. Fogarty, Hui Liu, Irit Altman, Chih-Hao Hsieh, Les Kaufman, Alec D. MacCall, Andrew A. Rosenberg, Hao Ye, and George Sugihara. Complex dynamics may limit prediction in marine fisheries. Fish and Fisheries, 15(4):616–633, December 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Maschner:2014:DSS

[546] Herbert D. G. Maschner, Andrew W. Trites, Katherine L. Reedy-Maschner, and Matthew Betts. The decline of Steller sea lions (Eumetopias jubatus) in the North Pacific: insights from indigenous people, ethnohistoric records and archaeological data. Fish and Fisheries, 15(4): 634–660, December 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Fujita:2014:AMD

[547] Rod Fujita, Daniel J. Thornhill, Kendra Karr, Cara H. Cooper, and Laura E. Dee. Assessing and managing data-limited ornamental fisheries in coral reefs. Fish and Fisheries, 15(4):661–675, December 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Wilson:2014:SNE

[548] A. D. M. Wilson, D. P. Croft, and J. Krause. Social networks in elasmobranchs and teleost fishes. Fish and Fisheries, 15(4):676–689, December 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Howarth:2014:UCS

[549] Leigh M. Howarth, Callum M. Roberts, Ruth H. Thurstan, and Bryce D. Stewart. The unintended consequences of simplifying the sea: making the case for complexity. Fish and Fisheries, 15(4):690–711, December 2014. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Andersen:2015:SSM

[550] Ken H. Andersen and Jan E. Beyer. Size structure, not metabolic scaling rules, determines fisheries reference points. Fish and Fisheries, 16(1):1– 22, March 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

McBride:2015:EAA

[551] Richard S. McBride, Stylianos Somarakis, Gary R. Fitzhugh, Anu Albert, Nathalia A. Yaragina, Mark J. Wuenschel, Alexandre Alonso-Fernández, and Gualtiero Basilone. Energy acquisition and allocation to egg production in relation to fish reproductive strategies. Fish and Fisheries, 16(1):23–57, March 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Sciberras:2015:ERC

[552] Marija Sciberras, Stuart R. Jenkins, Rebecca Mant, Michel J. Kaiser, Stephen J. Hawkins, and Andrew S. Pullin. Evaluating the relative conservation value of fully and partially protected marine areas. Fish and Fisheries, 16(1):58–77, March 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

McClanahan:2015:MFH

[553] Tim McClanahan, Edward H. Allison, and Joshua E. Cinner. Managing fisheries for human and food security. Fish and Fisheries, 16(1):78– 103, March 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Soetaert:2015:EPA

[554] Maarten Soetaert, Annemie Decostere, Hans Polet, Bart Verschueren, and Koen Chiers. Electrotrawling: a promising alternative fishing technique warranting further exploration. Fish and Fisheries, 16(1):104–124, March 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ovenden:2015:OEC

[555] Jennifer R. Ovenden, Oliver Berry, David J. Welch, Rik C. Buckworth, and Catherine M. Dichmont. Ocean’s eleven: a critical evaluation of the role of population, evolutionary and molecular genetics in the management of wild fisheries. Fish and Fisheries, 16(1):125–159, March 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Law:2015:SCR

[556] Richard Law, Jeppe Kolding, and Michael J. Plank. Squaring the circle: reconciling fishing and conservation of aquatic ecosystems. Fish and Fisheries, 16(1):160–174, March 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Sarda:2015:OIC

[557] Francesc Sardà, Marta Coll, Johanna J. Heymans, and Konstantinos I. Stergiou. Overlooked impacts and challenges of the new European discard ban. Fish and Fisheries, 16(1):175–180, March 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Piccolo:2015:LAT

[558] John J. Piccolo. Learning ‘about trout’: a student’s tribute to Robert J. Behnke. Fish and Fisheries, 16(1):181–182, March 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Uhlmann:2015:MUF

[559] Sebastian S. Uhlmann and Matt K. Broadhurst. Mitigating unaccounted fishing mortality from gillnets and traps. Fish and Fisheries, 16(2):183– 229, June 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Potter:2015:WWF

[560] Ian C. Potter, James R. Tweedley, Michael Elliott, and Alan K. Whitfield. The ways in which fish use estuaries: a refinement and expansion of the guild approach. Fish and Fisheries, 16(2):230–239, June 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bergseth:2015:MMC

[561] Brock J. Bergseth, Garry R. Russ, and Joshua E. Cinner. Measuring and monitoring compliance in no-take marine reserves. Fish and Fisheries, 16(2):240–258, June 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Gwinn:2015:RLB

[562] Daniel C. Gwinn, Micheal S. Allen, Fiona D. Johnston, Paul Brown, Charles R. Todd, and Robert Arlinghaus. Rethinking length-based fisheries regulations: the value of protecting old and large fish with harvest slots. Fish and Fisheries, 16(2):259–281, June 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hoagstrom:2015:REP

[563] Christopher W. Hoagstrom and Thomas F. Turner. Recruitment ecology of pelagic-broadcast spawning minnows: paradigms from the ocean advance science and conservation of an imperilled freshwater fauna. Fish and Fisheries, 16(2):282–299, June 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Favaro:2015:DCR

[564] Brett Favaro and Isabelle M. Côté. Do by-catch reduction devices in longline fisheries reduce capture of sharks and rays? A global meta-analysis. Fish and Fisheries, 16(2):300–309, June 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Trenkel:2015:FEM

[565] Verena M. Trenkel, Marie-Joëlle Rochet, and Jake C. Rice. A framework for evaluating management plans comprehensively. Fish and Fisheries, 16(2):310–328, June 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Eriksson:2015:SCF

[566] Hampus Eriksson and Maria Byrne. The sea cucumber fishery in Australia’s Great Barrier Reef Marine Park follows global patterns of serial exploitation. Fish and Fisheries, 16(2):329–341, June 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Thorson:2015:GSY

[567] James T. Thorson, Jason M. Cope, Kristin M. Kleisner, Jameal F. Samhouri, Andrew O. Shelton, and Eric J. Ward. Giants’ shoulders 15 years later: lessons, challenges and guidelines in fisheries meta-analysis. Fish and Fisheries, 16(2):342–361, June 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Nagelkerken:2015:SNN

[568] Ivan Nagelkerken, Marcus Sheaves, Ronald Baker, and Rod M. Connolly. The seascape nursery: a novel spatial approach to identify and manage nurseries for coastal marine fauna. Fish and Fisheries, 16(2):362–371, June 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

LePape:2015:FLH

[569] Olivier Le Pape and Sylvain Bonhommeau. The food limitation hypothesis for juvenile marine fish. Fish and Fisheries, 16(3):373–398, September 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Dowling:2015:RSB

[570] Natalie A. Dowling, Chris Wilcox, and Marc Mangel. Risk sensitivity and the behaviour of fishing vessels. Fish and Fisheries, 16(3):399–425, September 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Mangi:2015:AFD

[571] Stephen C. Mangi, Paul J. Dolder, Thomas L. Catchpole, Dale Rodmell, and Nathan de Rozarieux. Approaches to fully documented fisheries: practical issues and stakeholder perceptions. Fish and Fisheries, 16(3):

426–452, September 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

VanRiel:2015:UFC

[572] Maria C. Van Riel, Simon R. Bush, Paul A. M. van Zwieten, and Arthur P. J. Mol. Understanding fisheries credit systems: potentials and pitfalls of managing catch efficiency. Fish and Fisheries, 16(3):453–470, September 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Fabinyi:2015:MIM

[573] Michael Fabinyi, Simon Foale, and Martha Macintyre. Managing inequality or managing stocks? An ethnographic perspective on the governance of small-scale fisheries. Fish and Fisheries, 16(3):471–485, September 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Martignac:2015:UAC

[574] François Martignac, Aurélie Daroux, Jean-Luc Bagliniere, Dominique Ombredane, and Jean Guillard. The use of acoustic cameras in shallow waters: new hydroacoustic tools for monitoring migratory fish population. A review of DIDSON technology. Fish and Fisheries, 16(3):486–510, September 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hemelrijk:2015:IEF

[575] Ck Hemelrijk, Dap Reid, H. Hildenbrandt, and Jt Padding. The increased efficiency of fish swimming in a school. Fish and Fisheries, 16(3):511–521, September 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cardinale:2015:CDN

[576] M. Cardinale, V. Bartolino, H. Svedäng, A. Sundelöf, R. T. Poulsen, and M. Casini. A centurial development of the North Sea fish megafauna as reflected by the historical Swedish longlining fisheries. Fish and Fisheries, 16(3):522–533, September 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Borges:2015:EDP

[577] Lisa Borges. The evolution of a discard policy in Europe. Fish and Fisheries, 16(3):534–540, September 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Wearmouth:2015:BSR

[578] Victoria J. Wearmouth. The biology of sharks and rays. Fish and Fisheries, 16(3):541–542, September 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Smith-Vaniz:2015:CIM

[579] William F. Smith-Vaniz and Kent E. Carpenter. The Carangids of India — a Monograph. Fish and Fisheries, 16(3):543–546, September 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kleiber:2015:GSS

[580] Danika Kleiber, Leila M. Harris, and Amanda C. J. Vincent. Gender and small-scale fisheries: a case for counting women and beyond. Fish and Fisheries, 16(4):547–562, December 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Baudron:2015:ACS

[581] Alan R. Baudron and Paul G. Fernandes. Adverse consequences of stock recovery: European hake, a new “choke” species under a discard ban? Fish and Fisheries, 16(4):563–575, December 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Little:2015:RTS

[582] Alyson S. Little, Coby L. Needle, Ray Hilborn, Daniel S. Holland, and C. Tara Marshall. Real-time spatial management approaches to reduce bycatch and discards: experiences from Europe and the United States. Fish and Fisheries, 16(4):576–602, December 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Jones:2015:USP

[583] Miranda C. Jones, Stephen R. Dye, John K. Pinnegar, Rachel Warren, and William W. L. Cheung. Using scenarios to project the changing profitability of fisheries under climate change. Fish and Fisheries, 16(4): 603–622, December 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Mcowen:2015:FPW

[584] Chris J. Mcowen, William W. L. Cheung, Ryan R. Rykaczewski, Reg A. Watson, and Louisa J. Wood. Is fisheries production within Large Marine Ecosystems determined by bottom-up or top-down forcing? Fish and Fisheries, 16(4):623–632, December 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Szuwalski:2015:ECA

[585] Cody S. Szuwalski, Katyana A. Vert-Pre, André E. Punt, Trevor A. Branch, and Ray Hilborn. Examining common assumptions about recruitment: a meta-analysis of recruitment dynamics for worldwide marine fisheries. Fish and Fisheries, 16(4):633–648, December 2015. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Valles:2015:USF

[586] Henri Vallès and Hazel A. Oxenford. The utility of simple fish community metrics for evaluating the relative influence of fishing vs. other environmental drivers on Caribbean reef fish communities. Fish and Fisheries, 16(4):649–667, December 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Jones:2015:PKN

[587] Aisla C. Jones, Angela Mead, Michel J. Kaiser, Melanie C. V. Austen, Alex W. Adrian, Neil A. Auchterlonie, Kenneth D. Black, Lucy R. Blow, Charlotte Bury, Janet H. Brown, Gavin M. Burnell, Elaine Connolly, Alastair Dingwall, Simon Derrick, N. Clare Eno, Dominique J. H. Gautier, Karen A. Green, Matthew Gubbins, Piers R. Hart, John M. Holmyard, Anton J. Immink, David L. Jarrad, Emi Katoh, Jeremy C. R. Langley, Daniel O’C. Lee, Lewis Le Vay, Chris P. Leftwich, Mike Mitchell, Andrew Moore, Alexander G. Murray, Emma M. R. McLaren, Hannah Norbury, David Parker, Stephen O. Parry, Dawn Purchase, Amanna Rahman, Feyza Sanver, Melanie Siggs, Stephen D. Simpson, Richard J. Slaski, Katie Smith, Martin Le Q.Syvret, Claire Tibbott, Phil C. Thomas, Jimmy Turnbull, Robert Whiteley, Matthew Whittles, Mary J. Wilcockson, James Wilson, Lynn V. Dicks, and William J. Sutherland. Prioritization of knowledge needs for sustainable aquaculture: a national and global perspective. Fish and Fisheries, 16(4):668– 683, December 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Parker:2015:FCG

[588] Robert W. R. Parker and Peter H. Tyedmers. Fuel consumption of global fishing fleets: current understanding and knowledge gaps. Fish and Fisheries, 16(4):684–696, December 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Pelicice:2015:LRE

[589] Fernando M. Pelicice, Paulo S. Pompeu, and Angelo A. Agostinho. Large reservoirs as ecological barriers to downstream movements of Neotropical migratory fish. Fish and Fisheries, 16(4):697–715, December 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Zhou:2015:EOW

[590] Shijie Zhou, Anthony Dm Smith, and E. Eric Knudsen. Ending overfishing while catching more fish. Fish and Fisheries, 16(4):716–722, December 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kennedy:2015:ENA

[591] Laura A. Kennedy, Aharon G. Fleury, Jessica J. Holden, Alison E. Macnaughton, Piatã S. Marques, Geoffrey J. Osgood, and Francis Juanes. The ecology of North American freshwater fishes. Fish and Fisheries, 16 (4):723–725, December 2015. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Naslund:2016:EEF

[592] Joacim Näslund and Jörgen I. Johnsson. Environmental enrichment for fish in captive environments: effects of physical structures and substrates. Fish and Fisheries, 17(1):1–30, March 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Gascuel:2016:FIE

[593] Didier Gascuel, Marta Coll, Clive Fox, Sylvie Guénette, Jérome Guitton, Andrew Kenny, Leyla Knittweis, J. Rasmus Nielsen, Gerjan Piet, Tiit Raid, Morgane Travers-Trolet, and Samuel Shephard. Fishing impact and environmental status in European seas: a diagnosis from stock assessments and ecosystem indicators. Fish and Fisheries, 17(1):31–55, March 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Campbell:2016:NSF

[594] Robert A. Campbell. A new spatial framework incorporating uncertain stock and fleet dynamics for estimating fish abundance. Fish and Fisheries, 17(1):56–77, March 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Boonstra:2016:CFB

[595] Wiebren J. Boonstra and Jonas Hentati-Sundberg. Classifying fishers’ behaviour. An invitation to fishing styles. Fish and Fisheries, 17(1):78– 100, March 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Collie:2016:EMF

[596] Jeremy S. Collie, Louis W. Botsford, Alan Hastings, Isaac C. Kaplan, John L. Largier, Patricia A. Livingston, Éva Plagányi, Kenneth A. Rose, Brian K. Wells, and Francisco E. Werner. Ecosystem models for fisheries management: finding the sweet spot. Fish and Fisheries, 17(1):101– 125, March 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Melnychuk:2016:WDE

[597] Michael C. Melnychuk, Timothy E. Essington, Trevor A. Branch, Selina S. Heppell, Olaf P. Jensen, Jason S. Link, Steven J. D. Martell, Ana M. Parma, and Anthony D. M. Smith. Which design elements of individual quota fisheries help to achieve management objectives? Fish and Fisheries, 17(1):126–142, March 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Senior:2016:STM

[598] Alistair M. Senior, Sheri L. Johnson, and Shinichi Nakagawa. Sperm traits of masculinized fish relative to wild-type males: a systematic review and meta-analyses. Fish and Fisheries, 17(1):143–164, March 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Skern-Mauritzen:2016:EPR

[599] Mette Skern-Mauritzen, Geir Ottersen, Nils Olav Handegard, Geir Huse, Gjert E. Dingsør, Nils C. Stenseth, and Olav S. Kjesbu. Ecosystem processes are rarely included in tactical fisheries management. Fish and Fisheries, 17(1):165–175, March 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Defeo:2016:CML

[600] Omar Defeo, Mauricio Castrejón, Roberto Pérez-Castañeda, Juan C. Castilla, Nicolás L. Gutiérrez, Timothy E. Essington, and Carl Folke. Co-management in Latin American small-scale shellfisheries: assessment from long-term case studies. Fish and Fisheries, 17(1):176–192, March 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Froese:2016:RSB

[601] Rainer Froese, Gianpaolo Coro, Kristin Kleisner, and Nazli Demirel. Revisiting safe biological limits in fisheries. Fish and Fisheries, 17(1):193– 209, March 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Thurstan:2016:NCN

[602] Ruth H. Thurstan, Alexander B. Campbell, and John M. Pandolfi. Nineteenth century narratives reveal historic catch rates for Australian snapper (Pagrus auratus). Fish and Fisheries, 17(1):210–225, March 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Costa:2016:IEP

[603] Pedro Reis Costa. Impact and effects of paralytic shellfish poisoning toxins derived from harmful algal blooms to marine fish. Fish and Fisheries, 17(1):226–248, March 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cooke:2016:AEF

[604] Steven J. Cooke, Zeb S. Hogan, Paul A. Butcher, Michael J. W. Stokesbury, Rajeev Raghavan, Austin J. Gallagher, Neil Hammerschlag, and Andy J. Danylchuk. Angling for endangered fish: conservation problem or conservation action? Fish and Fisheries, 17(1):249–265, March 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Mumby:2016:SHF

[605] Peter J. Mumby. Stratifying herbivore fisheries by habitat to avoid ecosystem overfishing of coral reefs. Fish and Fisheries, 17(1):266–278, March 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Johnston:2016:BRC

[606] Ian A. Johnston. Book review: Condition and Health Indicators of Exploited Marine Fishes, by Josep Lloret, Georgiy Shulman, and R. Malcolm Love, Wiley Blackwell, Chichester, West Sussex, 2014, ISBN: 978-0-470-67024-8, Price: GBP 130.00; EUR 150.80 Hbk. GBP 104.99; EUR120.99 e-book, pp. 247. Fish and Fisheries, 17(1):279, March 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Law:2016:BEC

[607] Richard Law, Michael J. Plank, and Jeppe Kolding. Balanced exploitation and coexistence of interacting, size-structured, fish species. Fish and Fisheries, 17(2):281–302, June 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Punt:2016:MSE

[608] André E. Punt, Doug S. Butterworth, Carryn L. de Moor, José A. A. De Oliveira, and Malcolm Haddon. Management strategy evaluation: best practices. Fish and Fisheries, 17(2):303–334, June 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Lam:2016:MCF

[609] Vicky W. Y. Lam, William W. L. Cheung, and U. Rashid Sumaila. Marine capture fisheries in the Arctic: winners or losers under climate change and ocean acidification? Fish and Fisheries, 17(2):335–357, June 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Fryer:2016:MAH

[610] Robert John Fryer, Finbarr Gerard O’Neill, and Alexius Edridge. A meta-analysis of haddock size-selection data. Fish and Fisheries, 17(2): 358–374, June 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Geffroy:2016:SDS

[611] Benjamin Geffroy and Agnès Bardonnet. Sex differentiation and sex determination in eels: consequences for management. Fish and Fisheries, 17(2):375–398, June 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Vasilakopoulos:2016:UPF

[612] Paraskevas Vasilakopoulos, Finbarr G. O’Neill, and C. Tara Marshall. The unfulfilled potential of fisheries selectivity to promote sustainability. Fish and Fisheries, 17(2):399–416, June 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

McClanahan:2016:PFA

[613] Tim R. McClanahan and Caroline A. Abunge. Perceptions of fishing access restrictions and the disparity of benefits among stakeholder communities and nations of south-eastern Africa. Fish and Fisheries, 17(2): 417–437, June 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Davidson:2016:WGS

[614] Lindsay N. K. Davidson, Meg A. Krawchuk, and Nicholas K. Dulvy. Why have global shark and ray landings declined: improved management or overfishing? Fish and Fisheries, 17(2):438–458, June 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Leal:2016:MOF

[615] Miguel Costa Leal, Marcela Carraro Melo Vaz, João Puga, Rui Jorge Miranda Rocha, Chris Brown, Rui Rosa, and Ricardo Calado. Marine ornamental fish imports in the European Union: an economic perspective. Fish and Fisheries, 17(2):459–468, June 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hillary:2016:SAM

[616] Richard M. Hillary, Ann L. Preece, Campbell R. Davies, Hiroyuki Kurota, Osamu Sakai, Tomoyuki Itoh, Ana M. Parma, Doug S. Butterworth, James Ianelli, and Trevor A. Branch. A scientific alternative to moratoria for rebuilding depleted international tuna stocks. Fish and Fisheries, 17(2):469–482, June 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Nguyen:2016:GII

[617] Nguyen Hong Nguyen. Genetic improvement for important farmed aquaculture species with a reference to carp, tilapia and prawns in Asia: achievements, lessons and challenges. Fish and Fisheries, 17(2):483–506, June 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Dapp:2016:RMG

[618] Derek R. Dapp, Terence I. Walker, Charlie Huveneers, and Richard D. Reina. Respiratory mode and gear type are important determinants of elasmobranch immediate and post-release mortality. Fish and Fisheries, 17(2):507–524, June 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Berdahl:2016:CNH

[619] Andrew Berdahl, Peter A. H. Westley, Simon A. Levin, Iain D. Couzin, and Thomas P. Quinn. A collective navigation hypothesis for homeward migration in anadromous salmonids. Fish and Fisheries, 17(2):525–542, June 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hill:2016:ELS

[620] Simeon L. Hill. The extreme life of the sea. Fish and Fisheries, 17(2): 543–544, June 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hamilton:2016:PLC

[621] Patrick B. Hamilton, Ian G. Cowx, Marjorie F. Oleksiak, Andrew M. Griffiths, Mats Grahn, Jamie R. Stevens, Gary R. Carvalho, Elizabeth Nicol, and Charles R. Tyler. Population-level consequences for wild fish exposed to sublethal concentrations of chemicals — a critical review. Fish and Fisheries, 17(3):545–566, September 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

VanWynsberge:2016:DDE

[622] Simon Van Wynsberge, Serge Andréfouët, Nabila Gaertner-Mazouni, Colette C. C. Wabnitz, Antoine Gilbert, Georges Remoissenet, Claude Payri, and Cécile Fauvelot. Drivers of density for the exploited giant clam Tridacna maxima: a meta-analysis. Fish and Fisheries, 17(3):567–584, September 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Watson:2016:PGS

[623] Reg A. Watson, Bridget S. Green, Sean R. Tracey, Anna Farmery, and Tony J. Pitcher. Provenance of global seafood. Fish and Fisheries, 17(3): 585–595, September 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kell:2016:SSB

[624] Laurence T. Kell, Richard D. M. Nash, Mark Dickey-Collas, Iago Mosqueira, and Cody Szuwalski. Is spawning stock biomass a robust proxy for reproductive potential? Fish and Fisheries, 17(3):596–616, September 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Vasilakopoulos:2016:TTS

[625] Paraskevas Vasilakopoulos and Christos D. Maravelias. A tale of two seas: a meta-analysis of crustacean stocks in the NE Atlantic and the Mediterranean Sea. Fish and Fisheries, 17(3):617–636, September 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kaiser:2016:PKN

[626] Michel J. Kaiser, Ray Hilborn, Simon Jennings, Ricky Amaroso, Michael Andersen, Kris Balliet, Eric Barratt, Odd A. Bergstad, Stephen Bishop, Jodi L. Bostrom, Catherine Boyd, Eduardo A. Bruce, Merrick Burden, Chris Carey, Jason Clermont, Jeremy S. Collie, Antony Delahunty, Jacqui Dixon, Steve Eayrs, Nigel Edwards, Rod Fujita, John Gauvin, Mary Gleason, Brad Harris, Pingguo He, Jan G. Hiddink, Kathryn M. Hughes, Mario Inostroza, Andrew Kenny, Jake Kritzer, Volker Kuntzsch, Mario Lasta, Ivan Lopez, Craig Loveridge, Don Lynch, Jim Masters, Tessa Mazor, Robert A. McConnaughey, Marcel Moenne, Francis, Aileen M. Nimick, Alex Olsen, David Parker, Ana Parma, Christine Penney, David Pierce, Roland Pitcher, Michael Pol, Ed Richardson, Adriaan D. Rijnsdorp, Simon Rilatt, Dale P. Rodmell, Craig Rose, Suresh A. Sethi, Katherine Short, Petri Suuronen, Erin Taylor, Scott Wallace, Lisa Webb, Eric Wickham, Sam R. Wilding, Ashley Wilson, Paul Winger, and William J. Sutherland. Prioritization of knowledge-needs to achieve best practices for bottom trawling in relation to seabed habitats. Fish and Fisheries, 17(3):637–663, September 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Gaither:2016:FRW

[627] Michelle R. Gaither, Brian W. Bowen, Luiz A. Rocha, and John C. Briggs. Fishes that rule the world: circumtropical distributions revisited. Fish and Fisheries, 17(3):664–679, September 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Andersen:2016:PAD

[628] Bjørn Snorre Andersen, Christian Jørgensen, Sigrunn Eliassen, and Jarl Giske. The proximate architecture for decision-making in fish. Fish and Fisheries, 17(3):680–695, September 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Barnett:2016:SCS

[629] Adam Barnett, Kátya G. Abrantes, Ronald Baker, Amy S. Diedrich, Marina Farr, Alf Kuilboer, Tracey Mahony, Ian McLeod, Gianna Moscardo, Murray Prideaux, Natalie Stoeckl, Ariella van Luyn, and Marcus Sheaves. Sportfisheries, conservation and sustainable livelihoods: a multidisciplinary guide to developing best practice. Fish and Fisheries, 17(3):696–713, September 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Vollset:2016:IPM

[630] Knut Wiik Vollset, Randi Ingebjørg Krontveit, Peder A. Jansen, Bengt Finstad, Bjørn Torgeir Barlaup, Ove Tommy Skilbrei, Martin Krkošek, Pål Romunstad, Arnfinn Aunsmo, Arne J. Jensen, and Ian Dohoo. Impacts of parasites on marine survival of Atlantic salmon: a meta-analysis. Fish and Fisheries, 17(3):714–730, September 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Dichmont:2016:RCA

[631] Catherine M. Dichmont, André E. Punt, Natalie Dowling, José A. A. De Oliveira, Lorne R. Little, Miriana Sporcic, Elizabeth Fulton, Rebecca Gorton, Neil Klaer, Malcolm Haddon, and David C. Smith. Is risk consistent across tier-based harvest control rule management systems? A comparison of four case-studies. Fish and Fisheries, 17(3):731–747, September 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Gilman:2016:CTA

[632] Eric Gilman, Milani Chaloupka, Yonat Swimmer, and Susanna Piovano. A cross-taxa assessment of pelagic longline by-catch mitigation measures: conflicts and mutual benefits to elasmobranchs. Fish and Fisheries, 17 (3):748–784, September 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Froese:2016:MIF

[633] Rainer Froese, Henning Winker, Didier Gascuel, U. Rashid Sumaila, and Daniel Pauly. Minimizing the impact of fishing. Fish and Fisheries, 17 (3):785–802, September 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Marchal:2016:CRF

[634] Paul Marchal, Jesper Levring Andersen, Martin Aranda, Mike Fitzpatrick, Leyre Goti, Olivier Guyader, Gunnar Haraldsson, Aaron Hatcher, Troels Jacob Hegland, Pascal Le Floc’h, Claire Macher, Loretta Malvarosa, Christos D. Maravelias, Simon Mardle, Arantza Murillas, J. Rasmus Nielsen, Rosaria Sabatella, Anthony D. M. Smith, Kevin Stokes, Thomas Thoegersen, and Clara Ulrich. A comparative review of fisheries management experiences in the European Union and in other countries worldwide: Iceland, Australia, and New Zealand. Fish and Fisheries, 17(3):803–824, September 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

McClenachan:2016:FTF

[635] Loren McClenachan, Sahan T. M. Dissanayake, and Xiaojie Chen. Fair trade fish: consumer support for broader seafood sustainability. Fish and Fisheries, 17(3):825–838, September 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bladon:2016:PES

[636] Annabelle J. Bladon, Katherine M. Short, Essam Yassin Mohammed, and E. J. Milner-Gulland. Payments for ecosystem services in developing world fisheries. Fish and Fisheries, 17(3):839–859, September 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Maceda-Veiga:2016:AHC

[637] Alberto Maceda-Veiga, Omar Domínguez-Domínguez, Josep Escribano-Alacid, and John Lyons. The aquarium hobby: can sinners become saints in freshwater fish conservation? Fish and Fisheries, 17(3):860–874, September 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ferretti:2016:FTC

[638] Francesco Ferretti, Gabriel Morey Verd, Bernard Seret, Jadranka Sulić Šprem, and Fiorenza Micheli. Falling through the cracks: the fading history of a large iconic predator. Fish and Fisheries, 17(3):875–889, September 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Smith:2016:MEM

[639] Wade D. Smith. Migration ecology of marine fishes. Fish and Fisheries, 17(3):890–891, September 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Jennings:2016:AFS

[640] Simon Jennings, Grant D. Stentiford, Ana M. Leocadio, Keith R. Jeffery, Julian D. Metcalfe, Ioanna Katsiadaki, Neil A. Auchterlonie, Stephen C. Mangi, John K. Pinnegar, Tim Ellis, Edmund J. Peeler, Tiziana Luisetti, Craig Baker-Austin, Mary Brown, Thomas L. Catchpole, Fiona J. Clyne, Stephen R. Dye, Nathan J. Edmonds, Kieran Hyder, Janette Lee, David N. Lees, Owen C. Morgan, Carl M. O’Brien, Birgit Oidtmann, Paulette E. Posen, Ana Ribeiro Santos, Nick G. H. Taylor, Andrew D. Turner, Bryony L. Townhill, and David W. Verner-Jeffreys. Aquatic food security: insights into challenges and solutions from an analysis of interactions between fisheries, aquaculture, food safety, human health, fish and human welfare, economy and environment. Fish and Fisheries, 17 (4):893–938, December 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ovando:2016:MDS

[641] Daniel Ovando, Dawn Dougherty, and Jono R. Wilson. Market and design solutions to the short-term economic impacts of marine reserves. Fish and Fisheries, 17(4):939–954, December 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Smart:2016:MAS

[642] Jonathan J. Smart, Andrew Chin, Andrew J. Tobin, and Colin A. Simpfendorfer. Multimodel approaches in shark and ray growth studies: strengths, weaknesses and the future. Fish and Fisheries, 17(4): 955–971, December 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Koenigstein:2016:MCC

[643] Stefan Koenigstein, Felix C. Mark, Stefan Gößling-Reisemann, Hauke Reuter, and Hans-Otto Poertner. Modelling climate change impacts on marine fish populations: process-based integration of ocean warming, acidification and other environmental drivers. Fish and Fisheries, 17(4): 972–1004, December 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Audzijonyte:2016:TMI

[644] Asta Audzijonyte, Elizabeth Fulton, Malcolm Haddon, Fay Helidoniotis, Alistair J. Hobday, Anna Kuparinen, John Morrongiello, Anthony Dm Smith, Judy Upston, and Robin S. Waples. Trends and management implications of human-influenced life-history changes in marine ectotherms. Fish and Fisheries, 17(4):1005–1028, December 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Nash:2016:EIC

[645] Kirsty L. Nash and Nicholas A. J. Graham. Ecological indicators for coral reef fisheries management. Fish and Fisheries, 17(4):1029–1054, December 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Sampaio:2016:OSP

[646] Flávia D. F. Sampaio and Carolina A. Freire. An overview of stress physiology of fish transport: changes in water quality as a function of transport duration. Fish and Fisheries, 17(4):1055–1072, December 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ziegler:2016:ECS

[647] Friederike Ziegler, Sara Hornborg, Bridget S. Green, Ole Ritzau Eigaard, Anna K. Farmery, Linus Hammar, Klaas Hartmann, Sverker Molander, Robert W. R. Parker, Erik Skontorp Hognes, Ian Vázquez-Rowe, and Anthony D. M. Smith. Expanding the concept of sustainable seafood using Life Cycle Assessment. Fish and Fisheries, 17(4):1073–1093, December 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Branch:2016:WMS

[648] Trevor A. Branch and Allison E. Linnell. What makes some fisheries references highly cited? Fish and Fisheries, 17(4):1094–1133, December 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ebner:2016:ECA

[649] Brendan C. Ebner, David L. Morgan, Adam Kerezsy, Scott Hardie, Stephen J. Beatty, Jamie E. Seymour, James A. Donaldson, Simon Linke, Stirling Peverell, David Roberts, Thomas Espinoza, Nadine Marshall, Frederieke J. Kroon, Damien W. Burrows, and Ryan R. J. McAllister. Enhancing conservation of Australian freshwater ecosystems: identification of freshwater flagship fishes and relevant target audiences. Fish and Fisheries, 17(4):1134–1151, December 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Abrams:2016:NEP

[650] Peter A. Abrams, David G. Ainley, Louise K. Blight, Paul K. Dayton, Joseph T. Eastman, and Jennifer L. Jacquet. Necessary elements of precautionary management: implications for the Antarctic toothfish. Fish and Fisheries, 17(4):1152–1174, December 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Crona:2016:MDD

[651] Beatrice I. Crona, Tim M. Daw, Wilf Swartz, Albert V. Norström, Magnus Nyström, Matilda Thyresson, Carl Folke, Jonas Hentati-Sundberg, Henrik Österblom, Lisa Deutsch, and Max Troell. Masked, diluted and drowned out: how global seafood trade weakens signals from marine ecosystems. Fish and Fisheries, 17(4):1175–1182, December 2016. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bundy:2016:DST

[652] Alida Bundy, Ratana Chuenpagdee, Sarah R. Cooley, Omar Defeo, Bernhard Glaeser, Patrice Guillotreau, Moenieba Isaacs, Makino Mitsutaku, and R. Ian Perry. A decision support tool for response to global change in marine systems: the IMBER-ADApT Framework. Fish and Fisheries, 17(4):1183–1193, December 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Muir:2016:IAC

[653] Andrew Michael Muir, Michael Jay Hansen, Charles Richard Bronte, and Charles Conrad Krueger. If Arctic charr Salvelinus alpinus is ‘the most diverse vertebrate’, what is the lake charr Salvelinus namaycush? Fish and Fisheries, 17(4):1194–1207, December 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Smith:2016:RRD

[654] Carl Smith and Robert J. Wootton. The remarkable reproductive diversity of teleost fishes. Fish and Fisheries, 17(4):1208–1215, December 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Burgess:2016:RQC

[655] Matthew G. Burgess, Florian K. Diekert, Nis S. Jacobsen, Ken H. Andersen, and Steven D. Gaines. Remaining questions in the case for balanced harvesting. Fish and Fisheries, 17(4):1216–1226, December 2016. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Morato:2016:PIO

[656] Telmo Morato, Peter I. Miller, Daniel C. Dunn, Simon J. Nicol, James Bowcott, and Patrick N. Halpin. A perspective on the importance of oceanic fronts in promoting aggregation of visitors to seamounts. Fish and Fisheries, 17(4):1227–1233, December 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Johnson:2016:FGT

[657] Andrew Frederick Johnson. Fishes: A guide to their diversity. Fish and Fisheries, 17(4):1234–1235, December 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Piccolo:2016:LRS

[658] John J. Piccolo. For the love of rivers: A scientist’s journey. Fish and Fisheries, 17(4):1236–1237, December 2016. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Pons:2017:EBE

[659] Maite Pons, Trevor A. Branch, Michael C. Melnychuk, Olaf P. Jensen, Jon Brodziak, Jean M. Fromentin, Shelton J. Harley, Alan C. Haynie, Laurie T. Kell, Mark N. Maunder, Ana M. Parma, Victor R. Restrepo, Rishi Sharma, Robert Ahrens, and Ray Hilborn. Effects of biological, economic and management factors on tuna and billfish stock status. Fish and Fisheries, 18(1):1–21, January 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Rummer:2017:CCE

[660] Jodie L. Rummer and Philip L. Munday. Climate change and the evolution of reef fishes: past and future. Fish and Fisheries, 18(1):22–39, January 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Eddy:2017:EEI

[661] Tyler D. Eddy, Heike K. Lotze, Elizabeth A. Fulton, Marta Coll, Cameron H. Ainsworth, Júlio Neves de Araújo, Catherine M. Bulman, Alida Bundy, Villy Christensen, John C. Field, Neil A. Gribble, Mejs Hasan, Steve Mackinson, and Howard Townsend. Ecosystem effects of invertebrate fisheries. Fish and Fisheries, 18(1):40–53, January 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Nieminen:2017:EPC

[662] Emmi Nieminen, Kari Hyytiäinen, and Marko Lindroos. Economic and policy considerations regarding hydropower and migratory fish. Fish and Fisheries, 18(1):54–78, January 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Carrasquilla-Henao:2017:MEL

[663] Mauricio Carrasquilla-Henao and Francis Juanes. Mangroves enhance local fisheries catches: a global meta-analysis. Fish and Fisheries, 18 (1):79–93, January 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Loring:2017:PEG

[664] Philip A. Loring. The political ecology of gear bans in two fisheries: Florida’s net ban and Alaska’s salmon wars. Fish and Fisheries, 18(1): 94–104, January 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Watson:2017:PMF

[665] Reg A. Watson, Tony J. Pitcher, and Simon Jennings. Plenty more fish in the sea? Fish and Fisheries, 18(1):105–113, January 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Branch:2017:SWC

[666] Trevor A. Branch and Danika Kleiber. Should we call them fishers or fishermen? Fish and Fisheries, 18(1):114–127, January 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Erisman:2017:FSA

[667] Brad Erisman, William Heyman, Shinichi Kobara, Tal Ezer, Simon Pittman, Octavio Aburto-Oropeza, and Richard S. Nemeth. Fish spawning aggregations: where well-placed management actions can yield big benefits for fisheries and conservation. Fish and Fisheries, 18(1):128– 144, January 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Everson:2017:DML

[668] Inigo Everson. Designation and management of large-scale MPAs drawing on the experiences of CCAMLR. Fish and Fisheries, 18(1):145–159, January 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kenchington:2017:IFM

[669] Trevor J. Kenchington. Implications of fish migration and fishing mortality for marine protected area design. Fish and Fisheries, 18(1):160–170, January 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Griffiths:2017:QSH

[670] Shane P. Griffiths, Joanne Bryant, Henry F. Raymond, and Peter A. Newcombe. Quantifying subjective human dimensions of recreational fishing: does good health come to those who bait? Fish and Fisheries, 18(1):171–184, January 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bellido:2017:COL

[671] Jose M. Bellido, Mariano García-Rodriguez, Teresa García-Jiménez, María González-Aguilar, and Ana Carbonell-Quetglas. Could the obligation to land undersized individuals increase the black market for juveniles: evidence from the Mediterranean? Fish and Fisheries, 18(1): 185–194, January 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Magurran:2017:FPR

[672] Anne Magurran. Fish pheromones and related cues. Fish and Fisheries, 18(1):195–196, January 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hart:2017:FAC

[673] Paul J. B. Hart. Fish atlas of the Celtic Sea, North Sea, and Baltic Sea: Based on international research-vessel surveys. Fish and Fisheries, 18 (1):197–198, January 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Jacobsen:2017:EFI

[674] Nis S. Jacobsen, Matthew G. Burgess, and Ken H. Andersen. Efficiency of fisheries is increasing at the ecosystem level. Fish and Fisheries, 18 (2):199–211, March 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Plank:2017:BHC

[675] Michael J. Plank, Jeppe Kolding, Richard Law, Hans D. Gerritsen, and David Reid. Balanced harvesting can emerge from fishing decisions by individual fishers in a small-scale fishery. Fish and Fisheries, 18(2):212– 225, March 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Pedersen:2017:SSP

[676] Martin W. Pedersen and Casper W. Berg. A stochastic surplus production model in continuous time. Fish and Fisheries, 18(2):226–243, March 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Long:2017:KPE

[677] Rachel D. Long, Anthony Charles, and Robert L. Stephenson. Key principles of ecosystem-based management: the fishermen’s perspective. Fish and Fisheries, 18(2):244–253, March 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cheung:2017:THS

[678] William W. L. Cheung, Miranda C. Jones, Vicky W. Y. Lam, Dana D. Miller, Yoshitaka Ota, Louise Teh, and Ussif R. Sumaila. Transform high seas management to build climate resilience in marine seafood supply. Fish and Fisheries, 18(2):254–263, March 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Burgess:2017:DEC

[679] Matthew G. Burgess, Henrique C. Giacomini, Cody S. Szuwalski, Christopher Costello, and Steven D. Gaines. Describing ecosystem contexts with single-species models: a theoretical synthesis for fisheries. Fish and Fisheries, 18(2):264–284, March 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Lowerre-Barbieri:2017:RRP

[680] Susan Lowerre-Barbieri, Greg DeCelles, Pierre Pepin, Ignacio A. Catalán, Barbara Muhling, Brad Erisman, Steven X. Cadrin, Josep Alós, Andres Ospina-Alvarez, Megan M. Stachura, Michael D. Tringali, Sarah Walters Burnsed, and Claire B. Paris. Reproductive resilience: a paradigm shift in understanding spawner-recruit systems in exploited marine fish. Fish and Fisheries, 18(2):285–312, March 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Rudd:2017:DUC

[681] Merrill B. Rudd and Trevor A. Branch. Does unreported catch lead to overfishing? Fish and Fisheries, 18(2):313–323, March 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ostman:2017:ISS

[682] Örjan Östman, Jens Olsson, Johan Dannewitz, Stefan Palm, and Ann-Britt Florin. Inferring spatial structure from population genetics and spatial synchrony in demography of Baltic Sea fishes: implications for management. Fish and Fisheries, 18(2):324–339, March 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Tlusty:2017:CSS

[683] Michael F. Tlusty and Øistein Thorsen. Claiming seafood is ‘sustainable’ risks limiting improvements. Fish and Fisheries, 18(2):340–346, March 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Riley:2017:IIS

[684] Stephen C. Riley, Thomas R. Binder, Taaja R. Tucker, John Menzies, Nick Eyles, John Janssen, Andrew M. Muir, Peter C. Esselman, Nigel J. Wattrus, and Charles C. Krueger. Islands in the ice stream: were spawning habitats for native salmonids in the Great Lakes created by paleo-ice streams? Fish and Fisheries, 18(2):347–359, March 2017. CODEN FFI-IAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Arlinghaus:2017:PGI

[685] Robert Arlinghaus, Kate L. Laskowski, Josep Alós, Thomas Klefoth, Christopher T. Monk, Shinnosuke Nakayama, and Arne Schröder. Passive gear-induced timidity syndrome in wild fish populations and its potential ecological and managerial implications. Fish and Fisheries, 18 (2):360–373, March 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Watson:2017:BWV

[686] Gordon J. Watson, Joanna M. Murray, Martin Schaefer, and Adam Bonner. Bait worms: a valuable and important fishery with implications for fisheries and conservation management. Fish and Fisheries, 18(2):374– 388, March 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Fernandes:2017:EEE

[687] Jose A. Fernandes, Eleni Papathanasopoulou, Caroline Hattam, Ana M. Queirós, William W. W. L. Cheung, Andrew Yool, Yuri Artioli, Edward C. Pope, Kevin J. Flynn, Gorka Merino, Piero Calosi, Nicola Beaumont, Melanie C. Austen, Stephen Widdicombe, and Manuel Barange. Estimating the ecological, economic and social impacts of ocean acidification and warming on UK fisheries. Fish and Fisheries, 18(3):389–411, May 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bundy:2017:SFM

[688] Alida Bundy, Ratana Chuenpagdee, Jennifer L. Boldt, Maria de Fatima Borges, Mohamed Lamine Camara, Marta Coll, Ibrahima Diallo, Clive Fox, Elizabeth A. Fulton, Ayse Gazihan, Astrid Jarre, Didier Jouffre, Kristin M. Kleisner, Ben Knight, Jason Link, Patroba P. Matiku, Hicham Masski, Dimitrios K. Moutopoulos, Chiara Piroddi, Tiit Raid, Ignacio Sobrino, Jorge Tam, Djiga Thiao, Maria Angeles Torres, Konstantinos Tsagarakis, Gro I. van der Meeren, and Yunne-Jai Shin. Strong fisheries management and governance positively impact ecosystem status. Fish and Fisheries, 18(3):412–439, May 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Squires:2017:ERB

[689] Dale Squires, Mark Maunder, Robin Allen, Peder Andersen, Kepa Astorkiza, Douglas Butterworth, Gonzalo Caballero, Raymond Clarke, Hans Ellefsen, Patrice Guillotreau, John Hampton, Rögnvaldur Hannesson, Elizabeth Havice, Mark Helvey, Samuel Herrick, Jr., Kjartan Hoydal, Vishwanie Maharaj, Rebecca Metzner, Iago Mosqueira, Ana Parma, Ivan Prieto-Bowen, Victor Restrepo, Shaufique Fahmi Sidique, Stein Ivar Steinsham, Eric Thunberg, Ikerne del Valle, and Niels Vestergaard. Effort rights-based management. Fish and Fisheries, 18(3):440–465, May 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Handisyde:2017:VAR

[690] Neil Handisyde, Trevor C. Telfer, and Lindsay G. Ross. Vulnerability of aquaculture-related livelihoods to changing climate at the global scale. Fish and Fisheries, 18(3):466–488, May 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bell:2017:GFC

[691] Justin D. Bell, Reg A. Watson, and Yimin Ye. Global fishing capacity and fishing effort from 1950 to 2012. Fish and Fisheries, 18(3):489– 505, May 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Froese:2017:EFR

[692] Rainer Froese, Nazli Demirel, Gianpaolo Coro, Kristin M. Kleisner, and Henning Winker. Estimating fisheries reference points from catch and resilience. Fish and Fisheries, 18(4):506–526, May 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Veldhuizen:2017:CIS

[693] Linda J. L. Veldhuizen, Ivo A. van der Lans, Paul Bm Berentsen, Imke J. M. de Boer, and Eddie A. M. Bokkers. Consumer interest in social sustainability issues of whitefish from capture fisheries in the north-east Atlantic. Fish and Fisheries, 18(4):527–542, May 2017. CODEN FFI-IAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Selig:2017:TFM

[694] Elizabeth R. Selig, Kristin M. Kleisner, Oren Ahoobim, Freddy Arocha, Annabelle Cruz-Trinidad, Rod Fujita, Mafaniso Hara, Laure Katz, Patrick McConney, Blake D. Ratner, Lina M. Saavedra-Díaz, Anne-Maree Schwarz, Djiga Thiao, Elin Torell, Sebastian Troëng, and Sebastian Villasante. A typology of fisheries management tools: using experience to catalyse greater success. Fish and Fisheries, 18(4):543–570, May 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Villegas-Rios:2017:PCE

[695] David Villegas-Ríos, Even Moland, and Esben Moland Olsen. Potential of contemporary evolution to erode fishery benefits from marine reserves. Fish and Fisheries, 18(4):571–577, May 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Venturelli:2017:AAS

[696] Paul A. Venturelli, Kieran Hyder, and Christian Skov. Angler apps as a source of recreational fisheries data: opportunities, challenges and proposed standards. Fish and Fisheries, 18(4):578–595, May 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Militz:2017:NEP

[697] Thane A. Militz and Simon Foale. The “Nemo Effect”: Perception and reality of Finding Nemo’s impact on marine aquarium fisheries. Fish and Fisheries, 18(4):596–606, May 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Farmery:2017:AIS

[698] Anna K. Farmery, Caleb Gardner, Sarah Jennings, Bridget S. Green, and Reg A. Watson. Assessing the inclusion of seafood in the sustainable diet literature. Fish and Fisheries, 18(4):607–618, May 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Collie:2017:IEB

[699] Jeremy Collie, Jan Geert Hiddink, Tobias van Kooten, Adriaan D. Rijnsdorp, Michel J. Kaiser, Simon Jennings, and Ray Hilborn. Indirect effects of bottom fishing on the productivity of marine fish. Fish and Fisheries, 18(4):619–637, July 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Girardin:2017:TYF

[700] Raphaël Girardin, Katell G. Hamon, John Pinnegar, Jan Jaap Poos, Olivier Thébaud, Alex Tidd, Youen Vermard, and Paul Marchal. Thirty years of fleet dynamics modelling using discrete-choice models: What have we learned? Fish and Fisheries, 18(4):638–655, July 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Andersen:2017:WLD

[701] Ken H. Andersen, Nis S. Jacobsen, Teunis Jansen, and Jan E. Beyer. When in life does density dependence occur in fish populations? Fish and Fisheries, 18(4):656–667, July 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Woodcock:2017:YEM

[702] Paul Woodcock, Bethan C. O’Leary, Michel J. Kaiser, and Andrew S. Pullin. Your evidence or mine? Systematic evaluation of reviews of marine protected area effectiveness. Fish and Fisheries, 18(4):668–681, July 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Boonstra:2017:QCI

[703] Wiebren J. Boonstra, Simon Birnbaum, and Emma Björkvik. The quality of compliance: investigating fishers’ responses towards regulation and authorities. Fish and Fisheries, 18(4):682–697, July 2017. CODEN FFI-IAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Vaughan:2017:CFS

[704] David Brendan Vaughan, Alexandra Sara Grutter, Mark John Costello, and Kate Suzanne Hutson. Cleaner fishes and shrimp diversity and a reevaluation of cleaning symbioses. Fish and Fisheries, 18(4):698–716, July 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Monnereau:2017:IMC

[705] Iris Monnereau, Robin Mahon, Patrick McConney, Leonard Nurse, Rachel Turner, and Henri Vallès. The impact of methodological choices on the outcome of national-level climate change vulnerability assessments: an example from the global fisheries sector. Fish and Fisheries, 18(4):717–731, July 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anderson:2017:IEP

[706] Sean C. Anderson, Andrew B. Cooper, Olaf P. Jensen, Cóilín Minto, James T. Thorson, Jessica C. Walsh, Jamie Afflerbach, Mark Dickey-Collas, Kristin M. Kleisner, Catherine Longo, Giacomo Chato Osio, Daniel Ovando, Iago Mosqueira, Andrew A. Rosenberg, and Elizabeth R. Selig. Improving estimates of population status and trend with superensemble models. Fish and Fisheries, 18(4):732–741, July 2017. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Zhou:2017:ESD

[707] Shijie Zhou, André E. Punt, Yimin Ye, Nick Ellis, Cathy M. Dichmont, Malcolm Haddon, David C. Smith, and Anthony Dm Smith. Estimating stock depletion level from patterns of catch history. Fish and Fisheries, 18(4):742–751, July 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Henriques:2017:BRE

[708] Sofia Henriques, François Guilhaumon, Sébastien Villéger, Sandra Amoroso, Susana França, Stéphanie Pasquaud, Henrique N. Cabral, and Rita P. Vasconcelos. Biogeographical region and environmental conditions drive functional traits of estuarine fish assemblages worldwide. Fish and Fisheries, 18(4):752–771, July 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Majluf:2017:LFC

[709] Patricia Majluf, Santiago De la Puente, and Villy Christensen. The little fish that can feed the world. Fish and Fisheries, 18(4):772–777, July 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Sheaves:2017:HMF

[710] Marcus Sheaves. How many fish use mangroves? The 75% rule an illdefined and poorly validated concept. Fish and Fisheries, 18(4):778– 789, July 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Chalifour:2017:ESE

[711] Lia Chalifour, Morgan J. Black, Kieran Cox, Sarah Schroeder, and Francis Juanes. Ecology of salmonids in estuaries around the world: Adaptations, habitats and conservation. Fish and Fisheries, 18(4):790–791, July 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2017:C

[712] Anonymous. Corrigendum. Fish and Fisheries, 18(4):792–793, July 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Fitzpatrick:2017:FPT

[713] Mike Fitzpatrick, Christos D. Maravelias, Ole Ritzau Eigaard, Stephen Hynes, and David Reid. Fisher’s preferences and trade-offs between management options. Fish and Fisheries, 18(5):795–807, September 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Tidd:2017:SFB

[714] Alex Tidd, Stephen Brouwer, and Graham Pilling. Shooting fish in a barrel? Assessing fisher-driven changes in catchability within tropical tuna purse seine fleets. Fish and Fisheries, 18(5):808–820, September 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kindsvater:2017:SMR

[715] Holly K. Kindsvater, John D. Reynolds, Yvonne Sadovy de Mitcheson, and Marc Mangel. Selectivity matters: Rules of thumb for management of plate-sized, sex-changing fish in the live reef food fish trade. Fish and Fisheries, 18(5):821–836, September 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cashion:2017:MFD

[716] Tim Cashion, Frédéric Le Manach, Dirk Zeller, and Daniel Pauly. Most fish destined for fishmeal production are food-grade fish. Fish and Fisheries, 18(5):837–844, September 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Peterson:2017:PRC

[717] Cassidy D. Peterson, Carolyn N. Belcher, Dana M. Bethea, William B. Driggers III, Bryan S. Frazier, and Robert J. Latour. Preliminary recovery of coastal sharks in the south-east United States. Fish and Fisheries, 18(5):845–859, September 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kincaid:2017:HFI

[718] Kate Kincaid, George Rose, and Rodolphe Devillers. How fisherinfluenced marine closed areas contribute to ecosystem-based management: a review and performance indicator scorecard. Fish and Fisheries, 18(5):860–876, September 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Savo:2017:ICC

[719] Valentina Savo, Cedar Morton, and Dana Lepofsky. Impacts of climate change for coastal fishers and implications for fisheries. Fish and Fisheries, 18(5):877–889, September 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Glover:2017:HCG

[720] Kevin A. Glover, Monica F. Solberg, Phil McGinnity, Kjetil Hindar, Eric Verspoor, Mark W. Coulson, Michael M. Hansen, Hitoshi Araki, Øystein Skaala, and Terje Svåsand. Half a century of genetic interaction between farmed and wild Atlantic salmon: Status of knowledge and unanswered questions. Fish and Fisheries, 18(5):890–927, September 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Birnie-Gauvin:2017:CEA

[721] Kim Birnie-Gauvin, David Costantini, Steven J. Cooke, and William G. Willmore. A comparative and evolutionary approach to oxidative stress in fish: a review. Fish and Fisheries, 18(5):928–942, September 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Locke:2017:IGN

[722] Catherine Locke, Paramita Muljono, Cynthia McDougall, and Miranda Morgan. Innovation and gendered negotiations: Insights from six smallscale fishing communities. Fish and Fisheries, 18(5):943–957, September 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Parmentier:2017:MEL

[723] Eric Parmentier, Rui Diogo, and Michael L. Fine. Multiple exaptations leading to fish sound production. Fish and Fisheries, 18(5):958–966, September 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Wenger:2017:CAD

[724] Amelia S. Wenger, Euan Harvey, Shaun Wilson, Chris Rawson, Stephen J. Newman, Douglas Clarke, Benjamin J. Saunders, Nicola Browne, Michael J. Travers, Jennifer L. Mcilwain, Paul L. A. Erftemeijer, Jean-Paul A. Hobbs, Dianne Mclean, Martial Depczynski, and Richard D. Evans. A critical analysis of the direct effects of dredging on fish. Fish and Fisheries, 18(5):967–985, September 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Lennox:2017:WMF

[725] Robert J. Lennox, Josep Alós, Robert Arlinghaus, Andrij Horodysky, Thomas Klefoth, Christopher T. Monk, and Steven J. Cooke. What makes fish vulnerable to capture by hooks? A conceptual framework and a review of key determinants. Fish and Fisheries, 18(5):986–1010, September 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Stanford:2017:FLR

[726] Richard J. Stanford, Budy Wiryawan, Dietriech G. Bengen, Rudi Febriamansyah, and John Haluan. The fisheries livelihoods resilience check (FLIRES check): a tool for evaluating resilience in fisher communities. Fish and Fisheries, 18(6):1011–1025, November 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cashion:2017:GRF

[727] Tim Cashion, Peter Tyedmers, and Robert W. R. Parker. Global reduction fisheries and their products in the context of sustainable limits. Fish and Fisheries, 18(6):1026–1037, November 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ganias:2017:AIE

[728] Kostas Ganias, Charikleia Mezarli, and Eleni Voultsiadou. Aristotle as an ichthyologist: Exploring Aegean fish diversity 2,400 years ago. Fish and Fisheries, 18(6):1038–1055, November 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Doyen:2017:EEB

[729] Luc Doyen, Christophe Béné, Michel Bertignac, Fabian Blanchard, Abdoul Ahad Cissé, Catherine Dichmont, Sophie Gourguet, Olivier Guyader, Pierre-Yves Hardy, Sarah Jennings, Lorne Richard Little, Claire Macher, David Jonathan Mills, Ahmed Noussair, Sean Pascoe, Jean-Christophe Pereau, Nicolas Sanz, Anne-Maree Schwarz, Tony Smith, and Olivier Thébaud. Ecoviability for ecosystem-based fisheries management. Fish and Fisheries, 18(6):1056–1072, November 2017. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Thorson:2017:RIT

[730] James T. Thorson, James N. Ianelli, and Stan Kotwicki. The relative influence of temperature and size-structure on fish distribution shifts: a case-study on walleye pollock in the Bering Sea. Fish and Fisheries, 18(6):1073–1084, November 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Szuwalski:2017:GFD

[731] Cody S. Szuwalski and James T. Thorson. Global fishery dynamics are poorly predicted by classical models. Fish and Fisheries, 18(6):1085– 1095, November 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Jabado:2017:EFA

[732] Rima W. Jabado and Julia L. Y. Spaet. Elasmobranch fisheries in the Arabian Seas Region: Characteristics, trade and management. Fish and Fisheries, 18(6):1096–1118, November 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Pelicice:2017:NFF

[733] Fernando M. Pelicice, Valter M. Azevedo-Santos, Jean R. S. Vitule, Mário L. Orsi, Dilermando P. Lima Junior, André L. B. Magalhães, Paulo S. Pompeu, Miguel Petrere, Jr., and Angelo A. Agostinho. Neotropical freshwater fishes imperilled by unsustainable policies. Fish and Fisheries, 18(6):1119–1133, November 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Voyer:2017:UWB

[734] Michelle Voyer, Kate Barclay, Alistair McIlgorm, and Nicole Mazur. Using a well-being approach to develop a framework for an integrated socioeconomic evaluation of professional fishing. Fish and Fisheries, 18(6): 1134–1149, November 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Carpenter:2017:DSO

[735] Stephen R. Carpenter, William A. Brock, Gretchen J. A. Hansen, Jonathan F. Hansen, Joseph M. Hennessy, Daniel A. Isermann, Eric J. Pedersen, K. Martin Perales, Andrew L. Rypel, Greg G. Sass, Tyler D. Tunney, and M. Jake Vander Zanden. Defining a Safe Operating Space for inland recreational fisheries. Fish and Fisheries, 18(6):1150–1160, November 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Beamish:2017:WPT

[736] Richard J. Beamish. What the past tells us about the future of Pacific salmon research. Fish and Fisheries, 18(6):1161–1175, November 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Fine:2017:TMF

[737] Michael L. Fine, Lemont B. Kier, and Julio C. Alvarez. Transmission mechanisms of fish electric signals. Fish and Fisheries, 18(6):1176–1179, November 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Finkbeiner:2017:ROM

[738] Elena M. Finkbeiner, Nathan J. Bennett, Timothy H. Frawley, Julia G. Mason, Dana K. Briscoe, Cassandra M. Brooks, Crystal A. Ng, Rosana Ourens, Katherine Seto, Shannon Switzer Swanson, Jose Urteaga, and Larry B. Crowder. Reconstructing overfishing: Moving beyond Malthus for effective and equitable solutions. Fish and Fisheries, 18(6):1180–1191, November 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Mangel:2017:ILH

[739] Marc Mangel. The inverse life-history problem, size-dependent mortality and two extensions of results of Holt and Beverton. Fish and Fisheries, 18(6):1192–1200, November 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Righton:2017:GMS

[740] David Righton. Global Marine Science and Carlsberg: The golden connections of Johannes Schmidt (1877–1933). Fish and Fisheries, 18(6): 1201, November 2017. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2018:IIa

[741] Anonymous. Issue information. Fish and Fisheries, 19(1):i–iii, January 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Nielsen:2018:IEE

[742] J. Rasmus Nielsen, Eric Thunberg, Daniel S. Holland, Jorn O. Schmidt, Elizabeth A. Fulton, Francois Bastardie, André E. Punt, Icarus Allen, Heleen Bartelings, Michel Bertignac, Eckhard Bethke, Sieme Bossier, Rik Buckworth, Griffin Carpenter, Asbjørn Christensen, Villy Christensen, José M. Da-Rocha, Roy Deng, Catherine Dichmont, Ralf Doering, Aniol Esteban, Jose A. Fernandes, Hans Frost, Dorleta Garcia, Loic Gasche, Didier Gascuel, Sophie Gourguet, Rolf A. Groeneveld, Jordi Guillén, Olivier Guyader, Katell G. Hamon, Ayoe Hoff, Jan Horbowy, Trevor Hutton, Sigrid Lehuta, L. Richard Little, Jordi Lleonart, Claire Macher, Steven Mackinson, Stephanie Mahevas, Paul Marchal, Rosa Mato-Amboage, Bruce Mapstone, Francesc Maynou, Mathieu Merzéréaud, Artur Palacz, Sean Pascoe, Anton Paulrud, Eva Plaganyi, Raul Prellezo, Elizabeth I. van Putten, Martin Quaas, Lars Ravn-Jonsen, Sonia Sanchez, Sarah Simons, Olivier Thébaud, Maciej T. Tomczak, Clara Ulrich, Diana van Dijk, Youen Vermard, Rudi Voss, and Staffan Waldo. Integrated ecological–economic fisheries models — Evaluation, review and challenges for implementation. Fish and Fisheries, 19(1):1–29, January 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Zeller:2018:GMF

[743] Dirk Zeller, Tim Cashion, Maria Palomares, and Daniel Pauly. Global marine fisheries discards: a synthesis of reconstructed data. Fish and Fisheries, 19(1):30–39, January 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Benson:2018:OIE

[744] Ashleen J. Benson and Robert L. Stephenson. Options for integrating ecological, economic, and social objectives in evaluation and management of fisheries. Fish and Fisheries, 19(1):40–56, January 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Early-Capistran:2018:RYD

[745] Michelle-María Early-Capistrán, Andrea Sáenz-Arroyo, José-Gilberto Cardoso-Mohedano, Gerardo Garibay-Melo, S. Hoyt Peckham, and Volker Koch. Reconstructing 290 years of a data-poor fishery through ethnographic and archival research: the East Pacific green turtle (Chelonia mydas) in Baja California, Mexico. Fish and Fisheries, 19(1):57–77, January 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Olds:2018:EFS

[746] Andrew D. Olds, Elena Vargas-Fonseca, Rod M. Connolly, Ben L. Gilby, Chantal M. Huijbers, Glenn A. Hyndes, Craig A. Layman, Alan K. Whitfield, and Thomas A. Schlacher. The ecology of fish in the surf zones of ocean beaches: a global review. Fish and Fisheries, 19(1):78–89, January 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Coelho:2018:DPP

[747] Rui Coelho, Jaime Mejuto, Andrés Domingo, Kotaro Yokawa, Kwang-Ming Liu, Enric Cortés, Evgeny V. Romanov, Charlene da Silva, Fábio Hazin, Freddy Arocha, Aldrin Masawbi Mwilima, Pascal Bach, Victoria Ortiz de Zárate, William Roche, Pedro G. Lino, Blanca García-Cortés, Ana M. Ramos-Cartelle, Rodrigo Forselledo, Federico Mas, Seiji Ohshimo, Dean Courtney, Philippe S. Sabarros, Bernardo Perez, Ciara Wogerbauer, Wen-Pei Tsai, Felipe Carvalho, and Miguel N. Santos. Distribution patterns and population structure of the blue shark (Prionace glauca) in the Atlantic and Indian Oceans. Fish and Fisheries, 19(1):90– 106, January 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Brown:2018:TOT

[748] Christopher J. Brown, Glenn Althor, Benjamin S. Halpern, Md Sayed Iftekhar, Carissa J. Klein, Simon Linke, Elizabeth C. Pryde, Steven Schilizzi, James E. M. Watson, Becky Twohey, and Hugh P. Possingham. Trade-offs in triple-bottom-line outcomes when recovering fisheries. Fish and Fisheries, 19(1):107–116, January 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Tulloch:2018:EMQ

[749] Vivitskaia J. D. Tulloch, Éva E. Plagányi, Richard Matear, Christopher J. Brown, and Anthony J. Richardson. Ecosystem modelling to quantify the impact of historical whaling on Southern Hemisphere baleen whales. Fish and Fisheries, 19(1):117–137, January 2018. CODEN FFI-IAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Piria:2018:AFF

[750] Marina Piria, Predrag Simonović, Eleni Kalogianni, Leonidas Vardakas, Nicholas Koutsikos, Davor Zanella, Milica Ristovska, Apostolos Apostolou, Avdul Adrović, Danilo Mrdak, Ali Serhan Tarkan, Dragana Milošević, Linda N. Zanella, Rigers Bakiu, F. Güler Ekmekçi, Metka Povž, Kastriot Korro, Vera Nikolić, Rifat Škrijelj, Vasil Kostov, Andrej Gregori, and Michael K. Joy. Alien freshwater fish species in the Balkans — vectors and pathways of introduction. Fish and Fisheries, 19(1):138– 169, January 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Tillotson:2018:STM

[751] Michael D. Tillotson and Thomas P. Quinn. Selection on the timing of migration and breeding: a neglected aspect of fishing-induced evolution and trait change. Fish and Fisheries, 19(1):170–181, January 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hart:2018:ABO

[752] Paul J. B. Hart. All the boats on the ocean. How government subsidies led to global overfishing. Fish and Fisheries, 19(1):182, January 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Adams:2018:FFE

[753] Colin E. Adams. Freshwater fisheries ecology. Fish and Fisheries, 19 (1):183–184, January 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2018:IIb

[754] Anonymous. Issue information. Fish and Fisheries, 19(2):i–iii, March 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Harry:2018:ESA

[755] Alastair Varley Harry. Evidence for systemic age underestimation in shark and ray ageing studies. Fish and Fisheries, 19(2):185–200, March 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cooke:2018:NFN

[756] Steven J. Cooke, William M. Twardek, Robert J. Lennox, Aaron J. Zolderdo, Shannon D. Bower, Lee F. G. Gutowsky, Andy J. Danylchuk, Robert Arlinghaus, and Douglas Beard. The nexus of fun and nutrition: Recreational fishing is also about food. Fish and Fisheries, 19 (2):201–224, March 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hyder:2018:RSF

[757] Kieran Hyder, Marc Simon Weltersbach, Mike Armstrong, Keno Ferter, Bryony Townhill, Anssi Ahvonen, Robert Arlinghaus, Andrei Baikov, Manuel Bellanger, Janis Birzaks, Trude Borch, Giulia Cambie, Martin de Graaf, Hugo M. C. Diogo, Łukasz Dziemian, Ana Gordoa, Ryszard Grzebielec, Bruce Hartill, Anders Kagervall, Kostas Kapiris, Martin Karlsson, Alf Ring Kleiven, Adam M. Lejk, Harold Levrel, Sabrina Lovell, Jeremy Lyle, Pentti Moilanen, Graham Monkman, Beatriz Morales-Nin, Estanis Mugerza, Roi Martinez, Paul O’Reilly, Hans Jakob Olesen, Anastasios Papadopoulos, Pablo Pita, Zachary Radford, Krzysztof Radtke, William Roche, Delphine Rocklin, Jon Ruiz, Callum Scougal, Roberto Silvestri, Christian Skov, Scott Steinback, Andreas Sundelöf, Arvydas Svagzdys, David Turnbull, Tessa van der Hammen, David van Voorhees, Frankwin van Winsen, Thomas Verleye, Pedro Veiga, Jon-Helge Vølstad, Lucia Zarauz, Tomas Zolubas, and Harry V. Strehlow. Recreational sea fishing in Europe in a global context — participation rates, fishing effort, expenditure, and implications for monitoring and assessment. Fish and Fisheries, 19(2):225–243, March 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Modesto:2018:FMI

[758] Vanessa Modesto, Martina Ilarri, Allan T. Souza, Manuel Lopes-Lima, Karel Douda, Miguel Clavero, and Ronaldo Sousa. Fish and mussels: Importance of fish for freshwater mussel conservation. Fish and Fisheries, 19(2):244–259, March 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Pons:2018:MEL

[759] Maite Pons, Michael C. Melnychuk, and Ray Hilborn. Management effectiveness of large pelagic fisheries in the high seas. Fish and Fisheries, 19(2):260–270, March 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Worthington:2018:EMN

[760] Thomas A. Worthington, Anthony A. Echelle, Joshuah S. Perkin, Robert Mollenhauer, Nicole Farless, Joseph J. Dyer, Daniel Logue, and Shannon K. Brewer. The emblematic minnows of the North American Great Plains: a synthesis of threats and conservation opportunities. Fish and Fisheries, 19(2):271–307, March 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Eriksson:2018:GLI

[761] Hampus Eriksson, Kim Friedman, Moses Amos, Ian Bertram, Kalo Pakoa, Rebecca Fisher, and Neil Andrew. Geography limits island smallscale fishery production. Fish and Fisheries, 19(2):308–320, March 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Juan-Jorda:2018:RCE

[762] Maria José Juan-Jordá, Hilario Murua, Haritz Arrizabalaga, Nicholas K. Dulvy, and Victor Restrepo. Report card on ecosystem-based fisheries management in tuna regional fisheries management organizations. Fish and Fisheries, 19(2):321–339, March 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Silva:2018:FFP

[763] Ana T. Silva, Martyn C. Lucas, Theodore Castro-Santos, Christos Katopodis, Lee J. Baumgartner, Jason D. Thiem, Kim Aarestrup, Paulo S. Pompeu, Gordon C. O’Brien, Douglas C. Braun, Nicholas J. Burnett, David Z. Zhu, Hans-Petter Fjeldstad, Torbjørn Forseth, Nallamuthu Rajaratnam, John G. Williams, and Steven J. Cooke. The future of fish passage science, engineering, and practice. Fish and Fisheries, 19(2):340–362, March 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Nielsen:2018:FRB

[764] Kåre Nolde Nielsen, Michaela Maria Aschan, Sveinn Agnarsson, Marta Ballesteros, Alan Baudron, Maria de Fátima Borges, Aida Campos, Rosa Chapela, Anna Kristín Daníelsdóttir, Karim Erzini, Ólavur Gregersen, Petter Holm, Alessandro Lucchetti, Sveinn Margeirsson, Hugo Vilela Mendes, Petter Olsen, Mafalda Rangel, Antonello Sala, José Luis Santiago, Sigrídur Sigurdardóttir, Cristina Silva, Daryl Sykes, Jónas Rúnar Vidarsson, Massimo Virgili, Laura Wise, and Paul George Fernandes. A framework for results-based management in fisheries. Fish and Fisheries, 19(2):363–376, March 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Song:2018:RGH

[765] Andrew M. Song, Jahn P. Johnsen, and Tiffany H. Morrison. Reconstructing governability: How fisheries are made governable. Fish and Fisheries, 19(2):377–389, March 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Trenkel:2018:HPS

[766] Verena M. Trenkel. How to provide scientific advice for ecosystem-based management now. Fish and Fisheries, 19(2):390–398, March 2018. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2018:IIc

[767] Anonymous. Issue information. Fish and Fisheries, 19(3):i–iii, May 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Nordlund:2018:GSS

[768] Lina M. Nordlund, Richard K. F. Unsworth, Martin Gullström, and Leanne C. Cullen-Unsworth. Global significance of seagrass fishery activity. Fish and Fisheries, 19(3):399–412, May 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Reinhardt:2018:CRV

[769] James F. Reinhardt, Jennifer Weaver, Pamela J. Latham, Andrea Dell’Apa, Joseph E. Serafy, Joan A. Browder, Mary Christman, Daniel G. Foster, and David R. Blankinship. Catch rate and at-vessel mortality of circle hooks versus J-hooks in pelagic longline fisheries: a global meta-analysis. Fish and Fisheries, 19(3):413–430, May 2018. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Castello:2018:FYV

[770] Leandro Castello, Laura L. Hess, Ram Thapa, David G. McGrath, Caroline C. Arantes, Vivian F. Renó, and Victoria J. Isaac. Fishery yields vary with land cover on the Amazon River floodplain. Fish and Fisheries, 19(3):431–440, May 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Izzo:2018:OCD

[771] Christopher Izzo, Patrick Reis-Santos, and Bronwyn M. Gillanders. Otolith chemistry does not just reflect environmental conditions: a metaanalytic evaluation. Fish and Fisheries, 19(3):441–454, May 2018. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cook:2018:PTB

[772] Robin M. Cook and Michael R. Heath. Population trends of bycatch species reflect improving status of target species. Fish and Fisheries, 19(3):455–470, May 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

McClanahan:2018:CBL

[773] Timothy R. McClanahan. Community biomass and life history benchmarks for coral reef fisheries. Fish and Fisheries, 19(3):471–488, May 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Oken:2018:VSP

[774] Kiva L. Oken, Timothy E. Essington, and Caihong Fu. Variability and stability in predation landscapes: a cross-ecosystem comparison on the potential for predator control in temperate marine ecosystems. Fish and Fisheries, 19(3):489–501, May 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hirsch:2018:CIW

[775] Philipp Emanuel Hirsch, Anouk N’Guyen, Roxane Muller, Irene Adrian-Kalchhauser, and Patricia Burkhardt-Holm. Colonizing Islands of water on dry land — on the passive dispersal of fish eggs by birds. Fish and Fisheries, 19(3):502–510, May 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kitada:2018:EEG

[776] Shuichi Kitada. Economic, ecological and genetic impacts of marine stock enhancement and sea ranching: a systematic review. Fish and Fisheries, 19(3):511–532, May 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ohlberger:2018:DCC

[777] Jan Ohlberger, Eric J. Ward, Daniel E. Schindler, and Bert Lewis. Demographic changes in Chinook salmon across the Northeast Pacific Ocean. Fish and Fisheries, 19(3):533–546, May 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Milner-Gulland:2018:TTM

[778] E. J. Milner-Gulland, Serge Garcia, William Arlidge, Joseph Bull, Anthony Charles, Laurent Dagorn, Sonya Fordham, Joshua Graff Zivin, Martin Hall, Jeffrey Shrader, Niels Vestergaard, Chris Wilcox, and Dale Squires. Translating the terrestrial mitigation hierarchy to marine megafauna by-catch. Fish and Fisheries, 19(3):547–561, May 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ganias:2018:ORF

[779] Kostas Ganias and Susan Lowerre-Barbieri. Oocyte recruitment and fecundity type in fishes: Refining terms to reflect underlying processes and drivers. Fish and Fisheries, 19(3):562–572, May 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Clausen:2018:MSF

[780] Lotte W. Clausen. Management science in fisheries. Fish and Fisheries, 19(3):573–574, May 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2018:IId

[781] Anonymous. Issue information. Fish and Fisheries, 19(4):i–ii, July 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Guinot:2018:BSE

[782] Guillaume Guinot and Lionel Cavin. Body size evolution and habitat colonization across 100 million years (Late Jurassic–Paleocene) of the actinopterygian evolutionary history. Fish and Fisheries, 19(4):577–597, July 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ladich:2018:ACF

[783] Friedrich Ladich. Acoustic communication in fishes: Temperature plays a role. Fish and Fisheries, 19(4):598–612, July 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Queiros:2018:CCA

[784] Ana M. Queirós, José Fernandes, Lily Genevier, and Christopher P. Lynam. Climate change alters fish community size-structure, requiring adaptive policy targets. Fish and Fisheries, 19(4):613–621, July 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Mangi:2018:PDD

[785] Stephen C. Mangi, Sven Kupschus, Steven Mackinson, Dale Rodmell, Alexandra Lee, Elizabeth Bourke, Tom Rossiter, Jim Masters, Stuart Hetherington, Thomas Catchpole, and David Righton. Progress in designing and delivering effective fishing industry–science data collection in the UK. Fish and Fisheries, 19(4):622–642, July 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Syed:2018:NLC

[786] Shaheen Syed, Melania Borit, and Marco Spruit. Narrow lenses for capturing the complexity of fisheries: a topic analysis of fisheries science from 1990 to 2016. Fish and Fisheries, 19(4):643–661, July 2018. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Friedman:2018:EIC

[787] K. Friedman, S. Gabriel, O. Abe, A. Adnan Nuruddin, A. Ali, R. Bidin Raja Hassan, S. X. Cadrin, A. Cornish, T. De Meulenaer, Dharmadi, Fahmi, L. Huu Tuan Anh, D. Kachelriess, L. Kissol, Jr., T. Krajangdara, A. Rahman Wahab, W. Tanoue, C. Tharith, F. Torres, Jr., W. Wanchana, S. Win, K. Yokawa, and Y. Ye. Examining the impact of CITES listing of sharks and rays in Southeast Asian fisheries. Fish and Fisheries, 19(4):662–676, July 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Wilkes:2018:FNP

[788] Martin Wilkes, Lee Baumgartner, Craig Boys, Luiz G. M. Silva, Justin O’Connor, Matthew Jones, Ivor Stuart, Evelyn Habit, Oscar Link, and J. Angus Webb. Fish-net: Probabilistic models for fishway planning, design and monitoring to support environmentally sustainable hydropower. Fish and Fisheries, 19(4):677–697, July 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Sciberras:2018:RBF

[789] Marija Sciberras, Jan Geert Hiddink, Simon Jennings, Claire L. Szostek, Kathryn M. Hughes, Brian Kneafsey, Leo J. Clarke, Nick Ellis, Adriaan D. Rijnsdorp, Robert A. McConnaughey, Ray Hilborn, Jeremy S. Collie, C. Roland Pitcher, Ricardo O. Amoroso, Ana M. Parma, Petri Suuronen, and Michel J. Kaiser. Response of benthic fauna to experimental bottom fishing: a global meta-analysis. Fish and Fisheries, 19 (4):698–715, July 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kelley:2018:HDI

[790] Jennifer L. Kelley, Pauline F. Grierson, Shaun P. Collin, and Peter M. Davies. Habitat disruption and the identification and management of functional trait changes. Fish and Fisheries, 19(4):716–728, July 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Dunn:2018:EHS

[791] Daniel C. Dunn, Caroline Jablonicky, Guillermo O. Crespo, Douglas J. McCauley, David A. Kroodsma, Kristina Boerder, Kristina M. Gjerde, and Patrick N. Halpin. Empowering high seas governance with satellite vessel tracking data. Fish and Fisheries, 19(4):729–739, July 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Becker:2018:MDA

[792] Alistair Becker, Matthew D. Taylor, Heath Folpp, and Michael B. Lowry. Managing the development of artificial reef systems: the need for quantitative goals. Fish and Fisheries, 19(4):740–752, July 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2018:IIe

[793] Anonymous. Issue information. Fish and Fisheries, 19(5):i–ii, September 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hansen:2018:SOC

[794] Brian Klitgaard Hansen, Dorte Bekkevold, Lotte Worsøe Clausen, and Einar Eg Nielsen. The sceptical optimist: challenges and perspectives for the application of environmental DNA in marine fisheries. Fish and Fisheries, 19(5):751–768, September 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Deyle:2018:EBF

[795] Ethan Deyle, Amy M. Schueller, Hao Ye, Gerald M. Pao, and George Sugihara. Ecosystem-based forecasts of recruitment in two menhaden species. Fish and Fisheries, 19(5):769–781, September 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bailey:2018:RCS

[796] Megan Bailey, Helen Packer, Laurenne Schiller, Michael Tlusty, and Wilf Swartz. The role of corporate social responsibility in creating a Seussian world of seafood sustainability. Fish and Fisheries, 19(5):782– 790, September 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Jacobsen:2018:NMA

[797] Nis S. Jacobsen and Timothy E. Essington. Natural mortality augments population fluctuations of forage fish. Fish and Fisheries, 19(5):791–797, September 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Spijkers:2018:MFF

[798] Jessica Spijkers, Tiffany H. Morrison, Robert Blasiak, Graeme S. Cumming, Matthew Osborne, James Watson, and Henrik Österblom. Marine fisheries and future ocean conflict. Fish and Fisheries, 19(5):798–806, September 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

McClanahan:2018:MEC

[799] Timothy R. McClanahan. Multicriteria estimate of coral reef fishery sustainability. Fish and Fisheries, 19(5):807–820, September 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Liu:2018:USM

[800] Owen R. Liu, Kristin M. Kleisner, Sarah L. Smith, and Jacob P. Kritzer. The use of spatial management tools in rights-based groundfish fisheries. Fish and Fisheries, 19(5):821–838, September 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Woodhead:2018:HFC

[801] Anna J. Woodhead, Kirsten E. Abernethy, Lucy Szaboova, and Rachel A. Turner. Health in fishing communities: a global perspective. Fish and Fisheries, 19(5):839–852, September 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Stephenson:2018:EIS

[802] Robert L. Stephenson, Stacey Paul, Melanie Wiber, Eric Angel, Ashleen J. Benson, Anthony Charles, Omer Chouinard, Marc Clemens, Dan Edwards, Paul Foley, Lindsay Jennings, Owen Jones, Dan Lane, Jim McIsaac, Claire Mussells, Barbara Neis, Bethany Nordstrom, Courtenay Parlee, Evelyn Pinkerton, Mark Saunders, Kevin Squires, and U. Rashid Sumaila. Evaluating and implementing social–ecological systems: a comprehensive approach to sustainable fisheries. Fish and Fisheries, 19(5): 853–873, September 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Morais:2018:GDR

[803] Renato A. Morais and David R. Bellwood. Global drivers of reef fish growth. Fish and Fisheries, 19(5):874–889, September 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Westerberg:2018:LAA

[804] Håkan Westerberg, Michael J. Miller, Klaus Wysujack, Lasse Marohn, Marko Freese, Jan-Dag Pohlmann, Shun Watanabe, Katsumi Tsukamoto, and Reinhold Hanel. Larval abundance across the European eel spawning area: an analysis of recent and historic data. Fish and Fisheries, 19 (5):890–902, September 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Drouineau:2018:FES

[805] Hilaire Drouineau, Caroline Durif, Martin Castonguay, Maria Mateo, Eric Rochard, Guy Verreault, Kazuki Yokouchi, and Patrick Lambert. Freshwater eels: a symbol of the effects of global change. Fish and Fisheries, 19(5):903–930, September 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Gilby:2018:MBO

[806] Ben L. Gilby, Andrew D. Olds, Charles H. Peterson, Rod M. Connolly, Christine M. Voss, Melanie J. Bishop, Michael Elliott, Jonathan H. Grabowski, Nicholas L. Ortodossi, and Thomas A. Schlacher. Maximizing the benefits of oyster reef restoration for finfish and their fisheries. Fish and Fisheries, 19(5):931–947, September 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hart:2018:MPP

[807] Paul J. B. Hart. Marine plankton: A practical guide to ecology, methodology and taxonomy. Fish and Fisheries, 19(5):948–949, September 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2018:IIf

[808] Anonymous. Issue information. Fish and Fisheries, 19(6):i–ii, November 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Helias:2018:FLC

[809] Arnaud Hélias, Juliette Langlois, and Pierre Fréon. Fisheries in life cycle assessment: Operational factors for biotic resources depletion. Fish and Fisheries, 19(6):951–963, November 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Munch:2018:NDN

[810] Stephan B. Munch, Alfredo Giron-Nava, and George Sugihara. Nonlinear dynamics and noise in fisheries recruitment: a global meta-analysis. Fish and Fisheries, 19(6):964–973, November 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Lange:2018:MSE

[811] Katharina Lange, Andreas Bruder, Christoph D. Matthaei, Jakob Brodersen, and Rachel A. Paterson. Multiple-stressor effects on freshwater fish: Importance of taxonomy and life stage. Fish and Fisheries, 19 (6):974–983, November 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Pinkerton:2018:APH

[812] Evelyn Pinkerton, Marc Allain, David Decker, and Ken Carew. Atlantic and Pacific halibut co-management initiatives by Canadian fishermen’s organizations. Fish and Fisheries, 19(6):984–995, November 2018. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Baras:2018:OGS

[813] Etienne Baras, Otong Zenal Arifin, Jacques Slembrouck, Jojo Subagja, Anang Hari Kristanto, and Marc Legendre. Oil globule size in fish eggs: a matter of biome and reproductive strategy. Fish and Fisheries, 19(6): 996–1002, November 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Sanchez-Hernandez:2018:ETS

[814] Javier Sánchez-Hernández and Per-Arne Amundsen. Ecosystem type shapes trophic position and omnivory in fishes. Fish and Fisheries, 19(6): 1003–1015, November 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Faught:2018:MSF

[815] Erin Faught and Mathilakath M. Vijayan. Maternal stress and fish reproduction: the role of cortisol revisited. Fish and Fisheries, 19(6): 1016–1030, November 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Spence:2018:GFC

[816] Michael A. Spence, Julia L. Blanchard, Axel G. Rossberg, Michael R. Heath, Johanna J. Heymans, Steven Mackinson, Natalia Serpetti, Douglas C. Speirs, Robert B. Thorpe, and Paul G. Blackwell. A general framework for combining ecosystem models. Fish and Fisheries, 19(6): 1031–1042, November 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Jabado:2018:TWT

[817] Rima W. Jabado, Peter M. Kyne, Riley A. Pollom, David A. Ebert, Colin A. Simpfendorfer, Gina M. Ralph, Shaikha S. Al Dhaheri, K. V. Akhilesh, Khadeeja Ali, Mohamud Hassan Ali, Tariq M. S. Al Mamari, K. K. Bineesh, Igbal S. El Hassan, Daniel Fernando, Edwin M. Grandcourt, Muhammad Moazzam Khan, Alec B. M. Moore, Fereidoon Owfi, David P. Robinson, Evgeny Romanov, Ana-Lucia Soares, Julia L. Y. Spaet, Dawit Tesfamichael, Tooraj Valinassab, and Nicholas K. Dulvy. Troubled waters: Threats and extinction risk of the sharks, rays and chimaeras of the Arabian Sea and adjacent waters. Fish and Fisheries, 19(6):1043–1062, November 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Comte:2018:FDF

[818] Lise Comte and Julian D. Olden. Fish dispersal in flowing waters: a synthesis of movement- and genetic-based studies. Fish and Fisheries, 19(6):1063–1077, November 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Law:2018:BHC

[819] Richard Law and Michael J. Plank. Balanced harvesting could reduce fisheries-induced evolution. Fish and Fisheries, 19(6):1078–1091, November 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Belhabib:2018:IAN

[820] Dyhia Belhabib, Raouf Dridi, Allan Padilla, Melanie Ang, and Philippe Le Billon. Impacts of anthropogenic and natural “extreme events” on global fisheries. Fish and Fisheries, 19(6):1092–1109, November 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cortes:2018:SSR

[821] Enric Cortés and Elizabeth N. Brooks. Stock status and reference points for sharks using data-limited methods and life history. Fish and Fisheries, 19(6):1110–1129, November 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Walsh:2018:TOD

[822] Jessica C. Walsh, Cóilín Minto, Ernesto Jardim, Sean C. Anderson, Olaf P. Jensen, Jamie Afflerbach, Mark Dickey-Collas, Kristin M. Kleisner, Catherine Longo, Giacomo Chato Osio, Elizabeth R. Selig, James T. Thorson, Merrill B. Rudd, Katherine J. Papacostas, John N. Kittinger, Andrew A. Rosenberg, and Andrew B. Cooper. Trade-offs for datalimited fisheries when using harvest strategies based on catch-only models. Fish and Fisheries, 19(6):1130–1146, November 2018. CODEN FFI-IAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Henderson:2018:FGF

[823] Peter A. Henderson. Field Guide to the Fishes of the Amazon, Orinoco Guianas. Fish and Fisheries, 19(6):1147, November 2018. CODEN FFI-IAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Fernandes:2018:GAM

[824] Paul G. Fernandes. Global atlas of marine fisheries: a critical appraisal of catches and ecosystem impacts. Fish and Fisheries, 19(6):1148–1149, November 2018. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2019:IIa

[825] Anonymous. Issue information. Fish and Fisheries, 20(1):i–ii, January 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Brown:2019:AFS

[826] Christopher J. Brown, Andrew Broadley, Maria F. Adame, Trevor A. Branch, Mischa P. Turschwell, and Rod M. Connolly. The assessment of fishery status depends on fish habitats. Fish and Fisheries, 20(1):1– 14, January 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ye:2019:CME

[827] Yimin Ye and Fabio Carocci. Control mechanisms and ecosystem-based management of fishery production. Fish and Fisheries, 20(1):15–24, January 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cook:2019:SUR

[828] Katrina V. Cook, Andrea J. Reid, David A. Patterson, Kendra A. Robinson, Jacqueline M. Chapman, Scott G. Hinch, and Steven J. Cooke. A synthesis to understand responses to capture stressors among fish discarded from commercial fisheries and options for mitigating their severity. Fish and Fisheries, 20(1):25–43, January 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hossain:2019:THS

[829] M. Shahadat Hossain, S. M. Sharifuzzaman, Muhammad Abdur Rouf, Robert S. Pomeroy, Md Delwer Hossain, Sayedur R. Chowdhury, and Sheikh AftabUddin. Tropical hilsa shad (Tenualosa ilisha): Biology, fishery and management. Fish and Fisheries, 20(1):44–65, January 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Fulton:2019:ESG

[830] Elizabeth A. Fulton, André E. Punt, Catherine M. Dichmont, Chris J. Harvey, and Rebecca Gorton. Ecosystems say good management pays off. Fish and Fisheries, 20(1):66–96, January 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Belschner:2019:EFS

[831] Tobias Belschner, Johanna Ferretti, Harry v. Strehlow, Sarah B. M. Kraak, Ralf Döring, Gerd Kraus, Alexander Kempf, and Christopher Zimmermann. Evaluating fisheries systems: a comprehensive analytical framework and its application to the EU’s Common Fisheries Policy. Fish and Fisheries, 20(1):97–109, January 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Siple:2019:FFF

[832] Margaret C. Siple, Timothy E. Essington, and Éva E. Plagányi. Forage fish fisheries management requires a tailored approach to balance tradeoffs. Fish and Fisheries, 20(1):110–124, January 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Takasuka:2019:DDT

[833] Akinori Takasuka, Michio Yoneda, and Yoshioki Oozeki. Density dependence in total egg production per spawner for marine fish. Fish and Fisheries, 20(1):125–137, January 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Goulding:2019:EBM

[834] Michael Goulding, Eduardo Venticinque, Mauro L. de B. Ribeiro, Ronaldo B. Barthem, Rosseval G. Leite, Bruce Forsberg, Paulo Petry, Urbano Lopes da Silva-Júnior, Polliana Santos Ferraz, and Carlos Cañas. Ecosystem-based management of Amazon fisheries and wetlands. Fish and Fisheries, 20(1):138–158, January 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Thorson:2019:FSP

[835] James T. Thorson. Forecast skill for predicting distribution shifts: a retrospective experiment for marine fishes in the Eastern Bering Sea. Fish and Fisheries, 20(1):159–173, January 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Dowling:2019:GSD

[836] Natalie A. Dowling, Anthony D. M. Smith, David C. Smith, Ana M. Parma, Catherine M. Dichmont, Keith Sainsbury, Jono R. Wilson, Dawn T. Dougherty, and Jason M. Cope. Generic solutions for datalimited fishery assessments are not so simple. Fish and Fisheries, 20 (1):174–188, January 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Marandel:2019:EEP

[837] Florianne Marandel, Pascal Lorance, Olivier Berthelé, Verena M. Trenkel, Robin S. Waples, and Jean-Baptiste Lamy. Estimating effective population size of large marine populations, is it feasible? Fish and Fisheries, 20(1):189–198, January 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Thurstan:2019:FSH

[838] Ruth H. Thurstan. Fishing and shipwreck heritage. Marine archaeology’s greatest threat? Fish and Fisheries, 20(1):199–200, January 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2019:IIb

[839] Anonymous. Issue information. Fish and Fisheries, 20(2):i–ii, March 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Makris:2019:IAP

[840] Nicholas C. Makris, Olav Rune Godø, Dong Hoon Yi, Gavin J. Macaulay, Ankita D. Jain, Byunggu Cho, Zheng Gong, Josef Michael Jech, and Purnima Ratilal. Instantaneous areal population density of entire Atlantic cod and herring spawning groups and group size distribution relative to total spawning population. Fish and Fisheries, 20(2):201–213, March 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Giron-Nava:2019:MMS

[841] Alfredo Giron-Nava, Andrew F. Johnson, Andrés M. Cisneros-Montemayor, and Octavio Aburto-Oropeza. Managing at Maximum Sustainable Yield does not ensure economic well-being for artisanal fishers. Fish and Fisheries, 20(2):214–223, March 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Prince:2019:WDW

[842] Jeremy Prince and Adrian Hordyk. What to do when you have almost nothing: a simple quantitative prescription for managing extremely datapoor fisheries. Fish and Fisheries, 20(2):224–238, March 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Stewart:2019:UMS

[843] Kelly R. Stewart, Erin L. LaCasella, Michael P. Jensen, Sheryan P. Epperly, Heather L. Haas, Lesley W. Stokes, and Peter H. Dutton. Using mixed stock analysis to assess source populations for at-sea bycaught juvenile and adult loggerhead turtles (Caretta caretta) in the north-west Atlantic. Fish and Fisheries, 20(2):239–254, March 2019. CODEN FFI-IAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

MacKeracher:2019:SRM

[844] Tracy MacKeracher, Amy Diedrich, and Colin A. Simpfendorfer. Sharks, rays and marine protected areas: a critical evaluation of current perspectives. Fish and Fisheries, 20(2):255–267, March 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anderson:2019:HCF

[845] Christopher M. Anderson, Melissa J. Krigbaum, Martin C. Arostegui, Megan L. Feddern, John Zachary Koehn, Peter T. Kuriyama, Christina Morrisett, Caitlin I. Allen Akselrud, Melanie J. Davis, Courtney Fiamengo, Ava Fuller, Qi Lee, Katherine N. McElroy, Maite Pons, and Jessica Sanders. How commercial fishing effort is managed. Fish and Fisheries, 20(2):268–285, March 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Siskey:2019:STI

[846] Matthew R. Siskey, Oliver N. Shipley, and Michael G. Frisk. Skating on thin ice: Identifying the need for species-specific data and defined migration ecology of Rajidae spp. Fish and Fisheries, 20(2):286–302, March 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Harford:2019:CIG

[847] William J. Harford, Skyler R. Sagarese, and Mandy Karnauskas. Coping with information gaps in stock productivity for rebuilding and achieving maximum sustainable yield for grouper–snapper fisheries. Fish and Fisheries, 20(2):303–321, March 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Olmos:2019:ESC

[848] Maxime Olmos, Félix Massiot-Granier, Etienne Prévost, Gérald Chaput, Ian R. Bradbury, Marie Nevoux, and Etienne Rivot. Evidence for spatial coherence in time trends of marine life history traits of Atlantic salmon in the North Atlantic. Fish and Fisheries, 20(2):322–342, March 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Afflerbach:2019:QUW

[849] Jamie C. Afflerbach, Melanie Frazier, Halley E. Froehlich, Sean C. Anderson, and Benjamin S. Halpern. Quantifying uncertainty in the wildcaught fisheries goal of the Ocean Health Index. Fish and Fisheries, 20 (2):343–354, March 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Rios:2019:LSB

[850] María F. Ríos, Leonardo A. Venerus, Paraskevi K. Karachle, William D. K. Reid, Karim Erzini, Konstantinos I. Stergiou, and David E. Galván. Linking size-based trophodynamics and morphological traits in marine fishes. Fish and Fisheries, 20(2):355–367, March 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Clavelle:2019:IMF

[851] Tyler Clavelle, Sarah E. Lester, Rebecca Gentry, and Halley E. Froehlich. Interactions and management for the future of marine aquaculture and capture fisheries. Fish and Fisheries, 20(2):368–388, March 2019. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hart:2019:FLA

[852] Paul J. B. Hart. Fishing lessons. Artisanal fisheries and the future of our oceans. Fish and Fisheries, 20(2):389, March 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

McQuatters-Gollop:2019:MEF

[853] Abigail McQuatters-Gollop. Marine Ecological Field Methods. Fish and Fisheries, 20(2):390, March 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Martinsohn:2019:DAM

[854] Jann Th. Martinsohn, Paul Raymond, Trey Knott, Kevin A. Glover, Einar Eg Nielsen, Lars Bonde Eriksen, Rob Ogden, John Casey, and Jordi Guillen. DNA-analysis to monitor fisheries and aquaculture: Too costly? Fish and Fisheries, 20(2):391–401, March 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2019:IIc

[855] Anonymous. Issue information. Fish and Fisheries, 20(3):i–ii, May 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Harrison:2019:TEP

[856] Philip M. Harrison, Eduardo G. Martins, Dirk A. Algera, Trina Rytwinski, Brent Mossop, Alf J. Leake, Michael Power, and Steven J. Cooke. Turbine entrainment and passage of potadromous fish through hydropower dams: Developing conceptual frameworks and metrics for moving beyond turbine passage mortality. Fish and Fisheries, 20(3): 403–418, May 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Grisel:2019:MFC

[857] Florian Grisel. Managing the fishery commons at Marseille: How a medieval institution failed to accommodate change in an age of globalization. Fish and Fisheries, 20(3):419–433, May 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Planque:2019:PSM

[858] Benjamin Planque, Christian Mullon, Per Arneberg, Arne Eide, Jean-Marc Fromentin, Johanna Jacomina Heymans, Alf Håkon Hoel, Susa Niiranen, Geir Ottersen, Anne Britt Sandø, Martin Sommerkorn, Olivier Thébaud, and Thorbjørn Thorvik. A participatory scenario method to explore the future of marine social-ecological systems. Fish and Fisheries, 20(3):434–451, May 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Travaille:2019:KAR

[859] Kendra L. Thomas Travaille, Larry B. Crowder, Gary A. Kendrick, and Julian Clifton. Key attributes related to fishery improvement project (FIP) effectiveness in promoting improvements towards sustainability. Fish and Fisheries, 20(3):452–465, May 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Musyl:2019:MAP

[860] Michael K. Musyl and Eric L. Gilman. Meta-analysis of post-release fishing mortality in apex predatory pelagic sharks and white marlin. Fish and Fisheries, 20(3):466–500, May 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Catalan:2019:CEK

[861] Ignacio A. Catalán, Dominik Auch, Pauline Kamermans, Beatriz Morales-Nin, Natalie V. Angelopoulos, Patricia Reglero, Tina Sandersfeld, and Myron A. Peck. Critically examining the knowledge base required to mechanistically project climate impacts: a case study of Europe’s fish and shellfish. Fish and Fisheries, 20(3):501–517, May 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Zimmermann:2019:CTR

[862] Fabian Zimmermann, Marion Claireaux, and Katja Enberg. Common trends in recruitment dynamics of north-east Atlantic fish stocks and their links to environment, ecology and management. Fish and Fisheries, 20(3):518–536, May 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Estrella-Martinez:2019:RAH

[863] Juan Estrella-Martínez, Bernd R. Schöne, Ruth H. Thurstan, Elisa Capuzzo, James D. Scourse, and Paul G. Butler. Reconstruction of Atlantic herring (Clupea harengus) recruitment in the North Sea for the past 455 years based on the δ ^{13}C from annual shell increments of the ocean quahog (Arctica islandica). Fish and Fisheries, 20(3):537–551, May 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ladich:2019:ESC

[864] Friedrich Ladich. Ecology of sound communication in fishes. Fish and Fisheries, 20(3):552–563, May 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bradley:2019:OIF

[865] Darcy Bradley, Matt Merrifield, Karly M. Miller, Serena Lomonico, Jono R. Wilson, and Mary G. Gleason. Opportunities to improve fisheries management through innovative technology and advanced data systems. Fish and Fisheries, 20(3):564–583, May 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Nahuelhual:2019:TBT

[866] Laura Nahuelhual, Omar Defeo, Ximena Vergara, Gustavo Blanco, Sandra L. Marín, and Fabio Bozzeda. Is there a blue transition underway? Fish and Fisheries, 20(3):584–595, May 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cook:2019:FPD

[867] Robin Cook. Fish population dynamics, monitoring and management sustainable fisheries in the eternal ocean. Fish and Fisheries, 20(3): 596, May 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Dimoff:2019:BEP

[868] Sean Dimoff, Kelsie A. Murchy, Niallan C. O’Brien, Jessica Qualley, and Francis Juanes. The Behavior and Ecology of Pacific Salmon and Trout: Second Edition. Fish and Fisheries, 20(3):597–598, May 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Everson:2019:BRC

[869] Inigo Everson. Book review: The Curious Life of Krill. A conservation story from the bottom of the world, Nicol S., Island Press, Washington, DC, Covelo, London, 2018. Price: USD 30; GBP 22.99. ISBN: 978-1-61091-853-4. Fish and Fisheries, 20(3):599–600, May 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2019:IId

[870] Anonymous. Issue information. Fish and Fisheries, 20(4):i–ii, July 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Beaudreau:2019:TYC

[871] Anne H. Beaudreau, Eric J. Ward, Richard E. Brenner, Andrew O. Shelton, Jordan T. Watson, Jennifer C. Womack, Sean C. Anderson, Alan C. Haynie, Kristin N. Marshall, and Benjamin C. Williams. Thirty years of change and the future of Alaskan fisheries: Shifts in fishing participation and diversification in response to environmental, regulatory and economic pressures. Fish and Fisheries, 20(4):601–619, July 2019. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bezerra:2019:NMA

[872] Luis Artur Valões Bezerra, Matheus Oliveira Freitas, Vanessa Salete Daga, Thiago Vinicius Trento Occhi, Larissa Faria, Ana Paula Lula Costa, André Andrian Padial, Viviane Prodocimo, and Jean Ricardo Simões Vitule. A network meta-analysis of threats to South American fish biodiversity. Fish and Fisheries, 20(4):620–639, July 2019. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Soares:2019:UFM

[873] Marta C. Soares, Jo Cable, Monica G. Lima-Maximino, Caio Maximino, and Raquel Xavier. Using fish models to investigate the links between microbiome and social behaviour: the next step for translational microbiome research? Fish and Fisheries, 20(4):640–652, July 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Slabbekoorn:2019:PLC

[874] Hans Slabbekoorn, John Dalen, Dick de Haan, Hendrik V. Winter, Craig Radford, Michael A. Ainslie, Kevin D. Heaney, Tobias van Kooten, Len Thomas, and John Harwood. Population-level consequences of seismic surveys on fishes: an interdisciplinary challenge. Fish and Fisheries, 20(4):653–685, July 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Toresen:2019:SCL

[875] Reidar Toresen, Hein Rune Skjoldal, Frode Vikebø, and Monica Bente Martinussen. Sudden change in long-term ocean climate fluctuations corresponds with ecosystem alterations and reduced recruitment in Norwegian spring-spawning herring (Clupea harengus, Clupeidae). Fish and Fisheries, 20(4):686–696, July 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Steenbergen:2019:GIS

[876] Dirk J. Steenbergen, Michael Fabinyi, Kate Barclay, Andrew M. Song, Philippa J. Cohen, Hampus Eriksson, and David J. Mills. Governance interactions in small-scale fisheries market chains: Examples from the Asia–Pacific. Fish and Fisheries, 20(4):697–714, July 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bellanger:2019:IEJ

[877] Manuel Bellanger, Daniel S. Holland, Christopher M. Anderson, and Olivier Guyader. Incentive effect of joint and several liability in fishery cooperatives on regulatory compliance. Fish and Fisheries, 20(4):715– 728, July 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Radinger:2019:EMF

[878] Johannes Radinger, J. Robert Britton, Stephanie M. Carlson, Anne E. Magurran, Juan Diego Alcaraz-Hernández, Ana Almodóvar, Lluís Benejam, Carlos Fernández-Delgado, Graciela G. Nicola, Francisco J. Oliva-Paterna, Mar Torralva, and Emili García-Berthou. Effective monitoring of freshwater fish. Fish and Fisheries, 20(4):729–747, July 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Lyons:2019:DQQ

[879] Timothy J. Lyons, Quenton M. Tuckett, and Jeffrey E. Hill. Data quality and quantity for invasive species: a case study of the lionfishes. Fish and Fisheries, 20(4):748–759, July 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kotwicki:2019:ERD

[880] Stan Kotwicki and Kotaro Ono. The effect of random and densitydependent variation in sampling efficiency on variance of abundance estimates from fishery surveys. Fish and Fisheries, 20(4):760–774, July 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Arostegui:2019:RLS

[881] Martin C. Arostegui and Thomas P. Quinn. Reliance on lakes by salmon, trout and charr (Oncorhynchus, Salmo and Salvelinus): an evaluation of spawning habitats, rearing strategies and trophic polymorphisms. Fish and Fisheries, 20(4):775–794, July 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Moody:2019:TGG

[882] Eric K. Moody, Ellen Albright, Katie Cope, Rachel Fleck, Haley Grigel, David Ortiz, and Grace M. Wilkinson. Taxonomic and geographic gaps in understanding the functional effects of imperilled fishes on freshwater ecosystems. Fish and Fisheries, 20(4):795–801, July 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Heim:2019:GMT

[883] Kurt C. Heim, Thomas E. McMahon, Leonardo Calle, Mark S. Wipfli, and Jeffrey A. Falke. A general model of temporary aquatic habitat use: Water phenology as a life history filter. Fish and Fisheries, 20(4):802– 816, July 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2019:IIe

[884] Anonymous. Issue information. Fish and Fisheries, 20(5):i–ii, September 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Samia:2019:MAF

[885] Diogo S. M. Samia, Eduardo Bessa, Daniel T. Blumstein, José A. C. C. Nunes, Ernesto Azzurro, Lorenzo Morroni, Valerio Sbragaglia, Fraser A. Januchowski-Hartley, and Benjamin Geffroy. A meta-analysis of fish behavioural reaction to underwater human presence. Fish and Fisheries, 20 (5):817–829, September 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Syed:2019:MGN

[886] Shaheen Syed, Lia ní Aodha, Callum Scougal, and Marco Spruit. Mapping the global network of fisheries science collaboration. Fish and Fisheries, 20(5):830–856, September 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Guo:2019:BFF

[887] Chuanbo Guo, Yushun Chen, Rodolphe E. Gozlan, Zhongjie Li, Thomas Mehner, Sovan Lek, and Craig P. Paukert. Biogeographic freshwater fish pattern legacy revealed despite rapid socio-economic changes in China. Fish and Fisheries, 20(5):857–869, September 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Takasuka:2019:DDD

[888] Akinori Takasuka, Michio Yoneda, and Yoshioki Oozeki. Disentangling density-dependent effects on egg production and survival from egg to recruitment in fish. Fish and Fisheries, 20(5):870–887, September 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hordyk:2019:MSA

[889] Adrian R. Hordyk, Quang C. Huynh, and Thomas R. Carruthers. Misspecification in stock assessments: Common uncertainties and asymmetric risks. Fish and Fisheries, 20(5):888–902, September 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Sambrook:2019:BRW

[890] Katie Sambrook, Andrew S. Hoey, Serge Andréfouët, Graeme S. Cumming, Stephanie Duce, and Mary C. Bonin. Beyond the reef: the widespread use of non-reef habitats by coral reef fishes. Fish and Fisheries, 20(5):903–920, September 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Paijmans:2019:TUE

[891] Kai C. Paijmans, David J. Booth, and Marian Y. L. Wong. Towards an ultimate explanation for mixed-species shoaling. Fish and Fisheries, 20 (5):921–933, September 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Losee:2019:CSA

[892] James P. Losee, Neala W. Kendall, and Aaron Dufault. Changing salmon: an analysis of body mass, abundance, survival, and productivity trends across 45 years in Puget Sound. Fish and Fisheries, 20(5):934–951, September 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Saterberg:2019:ESE

[893] Torbjörn Säterberg, Michele Casini, and Anna Gårdmark. Ecologically sustainable exploitation rates — a multispecies approach for fisheries management. Fish and Fisheries, 20(5):952–961, September 2019. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Stige:2019:DSD

[894] Leif Christian Stige, Lauren A. Rogers, Anna B. Neuheimer, Mary E. Hunsicker, Natalia A. Yaragina, Geir Ottersen, Lorenzo Ciannelli, Øystein Langangen, and Joël M. Durant. Density- and size-dependent mortality in fish early life stages. Fish and Fisheries, 20(5):962–976, September 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Townhill:2019:MRF

[895] Bryony L. Townhill, Zachary Radford, Gretta Pecl, Ingrid van Putten, John K. Pinnegar, and Kieran Hyder. Marine recreational fishing and the implications of climate change. Fish and Fisheries, 20(5):977–992, September 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Silver:2019:HDC

[896] Jennifer J. Silver and Joshua S. Stoll. How do commercial fishing licences relate to access? Fish and Fisheries, 20(5):993–1004, September 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

ODea:2019:DTA

[897] Rose E. O’Dea, Malgorzata Lagisz, Andrew P. Hendry, and Shinichi Nakagawa. Developmental temperature affects phenotypic means and variability: a meta-analysis of fish data. Fish and Fisheries, 20(5):1005–1022, September 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Mason:2019:GPF

[898] Debby Mason. Gyotaku prints of fish and crustaceans of Southeast Alaska. Fish and Fisheries, 20(5):1023, September 2019. CODEN FFI-IAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Lowe:2019:NRR

[899] Benjamin S. Lowe, Susan K. Jacobson, Happiness Anold, Athanasio S. Mbonde, and Kai Lorenzen. The neglected role of religion in fisheries management. Fish and Fisheries, 20(5):1024–1033, September 2019. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Barnett:2019:RPT

[900] Lewis A. K. Barnett, Nis S. Jacobsen, James T. Thorson, and Jason M. Cope. Realizing the potential of trait-based approaches to advance fisheries science. Fish and Fisheries, 20(5):1034–1050, September 2019. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2019:IIf

[901] Anonymous. Issue information. Fish and Fisheries, 20(6):i–ii, November 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Nevoux:2019:EIL

[902] Marie Nevoux, Bengt Finstad, Jan Grimsrud Davidsen, Ross Finlay, Quentin Josset, Russell Poole, Johan Höjesjö, Kim Aarestrup, Lo Persson, Oula Tolvanen, and Bror Jonsson. Environmental influences on life history strategies in partially anadromous brown trout (Salmo trutta, Salmonidae). Fish and Fisheries, 20(6):1051–1082, November 2019. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Thorson:2019:STM

[903] James T. Thorson, Grant Adams, and Kirstin Holsman. Spatio-temporal models of intermediate complexity for ecosystem assessments: a new tool for spatial fisheries management. Fish and Fisheries, 20(6):1083–1099, November 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Luiz:2019:TBE

[904] Osmar J. Luiz, Julian D. Olden, Mark J. Kennard, David A. Crook, Michael M. Douglas, Thor M. Saunders, and Alison J. King. Traitbased ecology of fishes: a quantitative assessment of literature trends and knowledge gaps using topic modelling. Fish and Fisheries, 20(6): 1100–1110, November 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

McClenachan:2019:SPR

[905] Loren McClenachan, Jonathan H. Grabowski, Madison Marra, C. Seabird McKeon, Benjamin P. Neal, Nicholas R. Record, and Steven B. Scyphers.

Shifting perceptions of rapid temperature changes’ effects on marine fisheries, 1945–2017. Fish and Fisheries, 20(6):1111–1123, November 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Fossile:2019:PCF

[906] Thiago Fossile, Jessica Ferreira, Dione da Rocha Bandeira, Levy Figuti, Sérgio Dias da Silva, Niklas Hausmann, Harry K. Robson, David Orton, and André Carlo Colonese. Pre-Columbian fisheries catch reconstruction for a subtropical estuary in South America. Fish and Fisheries, 20(6): 1124–1137, November 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Larouche:2019:CAA

[907] Olivier Larouche, Miriam L. Zelditch, and Richard Cloutier. A critical appraisal of appendage disparity and homology in fishes. Fish and Fisheries, 20(6):1138–1175, November 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Funge-Smith:2019:FLI

[908] Simon Funge-Smith and Abigail Bennett. A fresh look at inland fisheries and their role in food security and livelihoods. Fish and Fisheries, 20(6): 1176–1195, November 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Stevens:2019:LHD

[909] Molly H. Stevens, Steven G. Smith, and Jerald S. Ault. Life history demographic parameter synthesis for exploited Florida and Caribbean coral reef fishes. Fish and Fisheries, 20(6):1196–1217, November 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Richardson:2019:EFG

[910] Kelsey Richardson, Britta Denise Hardesty, and Chris Wilcox. Estimates of fishing gear loss rates at a global scale: a literature review and metaanalysis. Fish and Fisheries, 20(6):1218–1231, November 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cook:2019:IDS

[911] Robin M. Cook. Inclusion of discards in stock assessment models. Fish and Fisheries, 20(6):1232–1245, November 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

DeMartini:2019:HMD

[912] Edward E. DeMartini. Hazards of managing disparate species as a pooled complex: a general problem illustrated by two contrasting examples from Hawaii. Fish and Fisheries, 20(6):1246–1259, November 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Guo:2019:TAM

[913] Chuanbo Guo and Rodolphe Elie Gozlan. Treatise on aquaculture in the modern era. Fish and Fisheries, 20(6):1260–1261, November 2019. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2020:IIa

[914] Anonymous. Issue information. Fish and Fisheries, 21(1):i–ii, January 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Okamoto:2020:ASS

[915] Daniel K. Okamoto, Melissa R. Poe, Tessa B. Francis, André E. Punt, Phillip S. Levin, Andrew O. Shelton, Derek R. Armitage, Jaclyn S. Cleary, Sherri C. Dressell, Russ Jones, Harvey Kitka, Lynn C. Lee, Alec D. MacCall, Jim A. McIsaac, Steve Reifenstuhl, Jennifer J. Silver, Jörn O. Schmidt, Thomas F. Thornton, Rüdiger Voss, and John Woodruff. Attending to spatial social–ecological sensitivities to improve trade-off analysis in natural resource management. Fish and Fisheries, 21(1):1–12, January 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Friedman:2020:ICP

[916] Kim Friedman, Matias Braccini, Matthew Bjerregaard-Walsh, Ramón Bonfil, Corey J. A. Bradshaw, Stephen Brouwer, Ian Campbell, Rui Coelho, Enric Cortés, Wetjens Dimmlich, Michael G. Frisk, Irene Kingma, Sophy R. McCully Phillips, Colman O’Criodain, Denham Parker, Samuel Shephard, Javier Tovar-Ávila, and Kotaro Yokawa. Informing CITES parties: Strengthening science-based decision-making when listing marine species. Fish and Fisheries, 21(1):13–31, January 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Lauchlan:2020:SRS

[917] Shannon S. Lauchlan and Ivan Nagelkerken. Species range shifts along multistressor mosaics in estuarine environments under future climate. Fish and Fisheries, 21(1):32–46, January 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Zhang:2020:FFC

[918] Wenbo Zhang, Min Liu, Yvonne Sadovy de Mitcheson, Ling Cao, Duncan Leadbitter, Richard Newton, David C. Little, Songlin Li, Yi Yang, Xiao Chen, and Wei Zhou. Fishing for feed in China: Facts, impacts and implications. Fish and Fisheries, 21(1):47–62, January 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Fall:2020:NRD

[919] Johanna Fall and Øyvind Fiksen. No room for dessert: a mechanistic model of prey selection in gut-limited predatory fish. Fish and Fisheries, 21(1):63–79, January 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Janes:2020:SII

[920] Holger Jänes, Peter I. Macreadie, Emily Nicholson, Daniel Ierodiaconou, Simon Reeves, Matthew D. Taylor, and Paul E. Carnell. Stable isotopes infer the value of Australia’s coastal vegetated ecosystems from fisheries. Fish and Fisheries, 21(1):80–90, January 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Potts:2020:WCE

[921] Warren M. Potts, Nicola Downey-Breedt, Pablo Obregon, Kieran Hyder, Roy Bealey, and Warwick H. H. Sauer. What constitutes effective governance of recreational fisheries? — A global review. Fish and Fisheries, 21(1):91–103, January 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Lawrence:2020:ISU

[922] Karl P. Lawrence, Antony R. Young, Brian L. Diffey, and Mary Norval. The impact of solar ultraviolet radiation on fish: Immunomodulation and photoprotective strategies. Fish and Fisheries, 21(1):104–119, January 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Alexander:2020:STE

[923] Steven M. Alexander, Phillip P. A. Staniczenko, and Örjan Bodin. Social ties explain catch portfolios of small-scale fishers in the Caribbean. Fish and Fisheries, 21(1):120–131, January 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Sherley:2020:SCF

[924] Richard B. Sherley, Hannah Ladd-Jones, Stefan Garthe, Olivia Stevenson, and Stephen C. Votier. Scavenger communities and fisheries waste:

North Sea discards support 3 million seabirds, 2 million fewer than in 1990. Fish and Fisheries, 21(1):132–145, January 2020. CODEN FFI-IAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Steins:2020:ICR

[925] Nathalie A. Steins, Marloes L. Kraan, Karin J. van der Reijden, Floor J. Quirijns, Wouter van Broekhoven, and Jan Jaap Poos. Integrating collaborative research in marine science: Recommendations from an evaluation of evolving science-industry partnerships in Dutch demersal fisheries. Fish and Fisheries, 21(1):146–161, January 2020. CODEN FFI-IAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

vanHelmond:2020:EMF

[926] Aloysius T. M. van Helmond, Lars O. Mortensen, Kristian S. Plet-Hansen, Clara Ulrich, Coby L. Needle, Daniel Oesterwind, Lotte Kindt-Larsen, Thomas Catchpole, Stephen Mangi, Christopher Zimmermann, Hans Jakob Olesen, Nick Bailey, Heidrikur Bergsson, Jørgen Dalskov, Jon Elson, Malo Hosken, Lisa Peterson, Howard McElderry, Jon Ruiz, Johanna P. Pierre, Claude Dykstra, and Jan Jaap Poos. Electronic monitoring in fisheries: Lessons from global experiences and future opportunities. Fish and Fisheries, 21(1):162–189, January 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Lynch:2020:EPT

[927] Abigail J. Lynch, Devin M. Bartley, Thomas Douglas Beard, Jr., Ian G. Cowx, Simon Funge-Smith, William W. Taylor, and Steven J. Cooke. Examining progress towards achieving the Ten Steps of the Rome Declaration on Responsible Inland Fisheries. Fish and Fisheries, 21(1):190– 203, January 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hartill:2020:DCM

[928] Bruce W. Hartill, Stephen M. Taylor, Krystle Keller, and Marc Simon Weltersbach. Digital camera monitoring of recreational fishing effort: Applications and challenges. Fish and Fisheries, 21(1):204–215, January 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Lidstrom:2020:EBF

[929] Susanna Lidström and Andrew F. Johnson. Ecosystem-based fisheries management: a perspective on the critique and development of the concept. Fish and Fisheries, 21(1):216–222, January 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2020:IIb

[930] Anonymous. Issue information. Fish and Fisheries, 21(2):i–ii, March 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Coulthard:2020:TAD

[931] Sarah Coulthard, Carole White, Nasheera Paranamana, K. P. G. L. Sandaruwan, R. Manimohan, and R. Maya. Tackling alcoholism and domestic violence in fisheries — a new opportunity to improve well-being for the most vulnerable people in global fisheries. Fish and Fisheries, 21(2):223–236, March 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Thorson:2020:PRD

[932] James T. Thorson. Predicting recruitment density dependence and intrinsic growth rate for all fishes worldwide using a data-integrated lifehistory model. Fish and Fisheries, 21(2):237–251, March 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Melli:2020:POC

[933] Valentina Melli, Bent Herrmann, Junita Diana Karlsen, Jordan Paul Feekings, and Ludvig Ahm Krag. Predicting optimal combinations of by-catch reduction devices in trawl gears: a meta-analytical approach. Fish and Fisheries, 21(2):252–268, March 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Booth:2020:MHS

[934] Hollie Booth, Dale Squires, and Eleanor Jane Milner-Gulland. The mitigation hierarchy for sharks: a risk-based framework for reconciling tradeoffs between shark conservation and fisheries objectives. Fish and Fisheries, 21(2):269–289, March 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Luo:2020:MMA

[935] Jiangang Luo, Jerald S. Ault, Bruce T. Ungar, Steven G. Smith, Michael F. Larkin, Thomas N. Davidson, David R. Bryan, Nicholas A. Farmer, Scott A. Holt, A. Scott Alford, Aaron J. Adams, Robert Humston, Adam S. Marton, David Mangum, Russell Kleppinger, Angel Requejo, and Julian Robertson. Migrations and movements of Atlantic tarpon revealed by two decades of satellite tagging. Fish and Fisheries, 21(2):290–318, March 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

McConnaughey:2020:CBP

[936] Robert A. McConnaughey, Jan G. Hiddink, Simon Jennings, C. Roland Pitcher, Michel J. Kaiser, Petri Suuronen, Marija Sciberras, Adriaan D. Rijnsdorp, Jeremy S. Collie, Tessa Mazor, Ricardo O. Amoroso, Ana M. Parma, and Ray Hilborn. Choosing best practices for managing impacts of trawl fishing on seabed habitats and biota. Fish and Fisheries, 21 (2):319–337, March 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Moro:2020:ADW

[937] Stefano Moro, Giovanna Jona-Lasinio, Barbara Block, Fiorenza Micheli, Giulio De Leo, Fabrizio Serena, Massimiliano Bottaro, Umberto Scacco, and Francesco Ferretti. Abundance and distribution of the white shark in the Mediterranean Sea. Fish and Fisheries, 21(2):338–349, March 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cao:2020:NSS

[938] Jie Cao, James T. Thorson, André E. Punt, and Cody Szuwalski. A novel spatiotemporal stock assessment framework to better address fine-scale species distributions: Development and simulation testing. Fish and Fisheries, 21(2):350–367, March 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hill:2020:RPP

[939] Simeon L. Hill, Jefferson Hinke, Sophie Bertrand, Lowell Fritz, Robert W. Furness, James N. Ianelli, Matthew Murphy, Ricardo Oliveros-Ramos, Lorien Pichegru, Rowland Sharp, Richard A. Stillman, Peter J. Wright, and Norman Ratcliffe. Reference points for predators will progress ecosystem-based management of fisheries. Fish and Fisheries, 21(2):368–378, March 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Belhabib:2020:CIF

[940] Dyhia Belhabib, William W. L. Cheung, David Kroodsma, Vicky W. Y. Lam, Philip J. Underwood, and John Virdin. Catching industrial fishing incursions into inshore waters of Africa from space. Fish and Fisheries, 21(2):379–392, March 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Link:2020:CHW

[941] Jason S. Link, Geir Huse, Sarah Gaichas, and Anthony R. Marshak. Changing how we approach fisheries: a first attempt at an operational framework for ecosystem approaches to fisheries management. Fish and Fisheries, 21(2):393–434, March 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Su:2020:EMF

[942] Shu Su, Yi Tang, Bowen Chang, Wenbin Zhu, and Yong Chen. Evolution of marine fisheries management in China from 1949 to 2019: How did China get here and where does China go next? Fish and Fisheries, 21 (2):435–452, March 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hutchings:2020:IFI

[943] Jeffrey A. Hutchings and Anna Kuparinen. Implications of fisheriesinduced evolution for population recovery: Refocusing the science and refining its communication. Fish and Fisheries, 21(2):453–464, March 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Thurstan:2020:VFS

[944] Ruth H. Thurstan. Vanishing fish: Shifting baselines and the future of global fisheries. Fish and Fisheries, 21(2):465–466, March 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Penca:2020:TCP

[945] Jerneja Penca. Towards constructing a positive narrative for fisheries: Report from International Symposium on Fisheries Sustainability: Strengthening the Science-Policy Nexus. Fish and Fisheries, 21(2):467– 469, March 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2020:IIc

[946] Anonymous. Issue information. Fish and Fisheries, 21(3):i–ii, May 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Teh:2020:FPG

[947] Lydia C. L. Teh, Yoshitaka Ota, Andrés M. Cisneros-Montemayor, Lucy Harrington, and Wilf Swartz. Are fishers poor? Getting to the bottom of marine fisheries income statistics. Fish and Fisheries, 21(3):471–482, May 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ahrens:2020:SLF

[948] Robert N. M. Ahrens, Micheal S. Allen, Carl Walters, and Robert Arlinghaus. Saving large fish through harvest slots outperforms the classical minimum-length limit when the aim is to achieve multiple harvest and catch-related fisheries objectives. Fish and Fisheries, 21(3):483–510, May 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic). See correction [1252].

Rubio:2020:SRD

[949] Iratxe Rubio, Unai Ganzedo, Alistair J. Hobday, and Elena Ojea. Southward re-distribution of tropical tuna fisheries activity can be explained by technological and management change. Fish and Fisheries, 21(3):511– 521, May 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Staudinger:2020:RSL

[950] Michelle D. Staudinger, Holly Goyert, Justin J. Suca, Kaycee Coleman, Linda Welch, Joel K. Llopiz, David Wiley, Irit Altman, Andew Applegate, Peter Auster, Hannes Baumann, Julia Beaty, Deirdre Boelke, Les Kaufman, Pam Loring, Jerry Moxley, Suzanne Paton, Kevin Powers, David Richardson, Jooke Robbins, Jeffrey Runge, Brian Smith, Caleb Spiegel, and Halley Steinmetz. The role of sand lances (Ammodytes sp.) in the Northwest Atlantic Ecosystem: a synthesis of current knowledge with implications for conservation and management. Fish and Fisheries, 21(3):522–556, May 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Jones:2020:UKB

[951] Matthew Jon Jones and Robin Hale. Using knowledge of behaviour and optic physiology to improve fish passage through culverts. Fish and Fisheries, 21(3):557–569, May 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Burgess:2020:OAB

[952] Matthew G. Burgess, Ernesto Carrella, Michael Drexler, Robert L. Axtell, Richard M. Bailey, James R. Watson, Reniel B. Cabral, Michaela Clemence, Christopher Costello, Chris Dorsett, Steven D. Gaines, Emily S. Klein, Philipp Koralus, George Leonard, Simon A. Levin, Lorne Richard Little, John Lynham, Jens Koed Madsen, Andreas Merkl, Brandon Owashi, Steven E. Saul, Ingrid E. van Putten, and Sharon Wilcox. Opportunities for agent-based modelling in human dimensions of fisheries. Fish and Fisheries, 21(3):570–587, May 2020. CODEN FFI-IAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Matte:2020:DDG

[953] Jean-Michel Matte, Dylan J. Fraser, and James W. A. Grant. Densitydependent growth and survival in salmonids: Quantifying biological mechanisms and methodological biases. Fish and Fisheries, 21(3):588– 600, May 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Zhang:2020:RCY

[954] Hui Zhang, Myounghee Kang, Li Shen, Jinming Wu, Junyi Li, Hao Du, Chengyou Wang, Haile Yang, Qiong Zhou, Zhigang Liu, Harry Gorfine, and Qiwei Wei. Rapid change in Yangtze fisheries and its implications for global freshwater ecosystem management. Fish and Fisheries, 21(3):601– 620, May 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Vasilakopoulos:2020:SMF

[955] Paraskevas Vasilakopoulos, Ernesto Jardim, Christoph Konrad, Dominic Rihan, Alessandro Mannini, Cecilia Pinto, John Casey, Iago Mosqueira, and Finbarr G. O’Neill. Selectivity metrics for fisheries management and advice. Fish and Fisheries, 21(3):621–638, May 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Trochta:2020:HLH

[956] John T. Trochta, Trevor A. Branch, Andrew O. Shelton, and Douglas E. Hay. The highs and lows of herring: a meta-analysis of patterns and factors in herring collapse and recovery. Fish and Fisheries, 21(3):639– 662, May 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Huynh:2020:IMP

[957] Quang C. Huynh, Adrian R. Hordyk, Robyn E. Forrest, Clay E. Porch, Sean C. Anderson, and Thomas R. Carruthers. The interim management procedure approach for assessed stocks: Responsive management advice and lower assessment frequency. Fish and Fisheries, 21(3):663–679, May 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hart:2020:ACC

[958] Paul J. B. Hart. Alaska codfish chronicle. A history of the Pacific cod fishery in Alaska. Fish and Fisheries, 21(3):680, May 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Engelhard:2020:SJH

[959] Georg H. Engelhard and John K. Pinnegar. Sidney J. Holt, eminent fisheries scientist and conservationist (28 February 1926–22 December 2019). Fish and Fisheries, 21(3):681–682, May 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2020:IId

[960] Anonymous. Issue information. Fish and Fisheries, 21(4):i–ii, July 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

McClanahan:2020:ISY

[961] Timothy R. McClanahan and Maxwell Kodia Azali. Improving sustainable yield estimates for tropical reef fisheries. Fish and Fisheries, 21(4): 683–699, July 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Fulton:2020:MMH

[962] Christopher J. Fulton, Charlotte Berkström, Shaun K. Wilson, Rene A. Abesamis, Michael Bradley, Carolina Åkerlund, Luke T. Barrett, Abner A. Bucol, Dinorah H. Chacin, Karen M. Chong-Seng, Darren J. Coker, Martial Depczynski, Linda Eggertsen, Maria Eggertsen, David Ellis, Richard D. Evans, Nicholas A. J. Graham, Andrew S. Hoey, Thomas H. Holmes, Michel Kulbicki, Priscilla T. Y. Leung, Paul K. S. Lam, Joshua van Lier, Paloma A. Matis, Mae M. Noble, Alejandro Pérez-Matus, Camilla Piggott, Ben T. Radford, Stina Tano, and Paul Tinkler. Macroalgal meadow habitats support fish and fisheries in diverse tropical seascapes. Fish and Fisheries, 21(4):700–717, July 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Gruss:2020:STA

[963] Arnaud Grüss, James T. Thorson, Gemma Carroll, Elizabeth L. Ng, Kirstin K. Holsman, Kerim Aydin, Stan Kotwicki, Hem N. Morzaria-Luna, Cameron H. Ainsworth, and Kevin A. Thompson. Spatio-temporal analyses of marine predator diets from data-rich and data-limited systems. Fish and Fisheries, 21(4):718–739, July 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Willis:2020:TTO

[964] Ciara Willis and Megan Bailey. Tuna trade-offs: Balancing profit and social benefits in one of the world’s largest fisheries. Fish and Fisheries, 21(4):740–759, July 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Zhou:2020:ISB

[965] Shijie Zhou, André E. Punt, Yeming Lei, Roy Aijun Deng, and Simon D. Hoyle. Identifying spawner biomass per-recruit reference points from life-history parameters. Fish and Fisheries, 21(4):760–773, July 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Caswell:2020:SOS

[966] Bryony A. Caswell, Emily S. Klein, Heidi K. Alleway, Johnathan E. Ball, Julián Botero, Massimiliano Cardinale, Margit Eero, Georg H. Engelhard, Tomaso Fortibuoni, Ana-Judith Giraldo, Jonas Hentati-Sundberg, Peter Jones, John N. Kittinger, Gesche Krause, Dmitry L. Lajus, Julia Lajus, Sally C. Y. Lau, Ann-Katrien Lescrauwaet, Brian R. MacKenzie, Matthew McKenzie, Henn Ojaveer, John M. Pandolfi, Saša Raicevich, Bayden D. Russell, Andreas Sundelöf, Robert B. Thorpe, Philine S. E. zu Ermgassen, and Ruth H. Thurstan. Something old, something new: Historical perspectives provide lessons for blue growth agendas. Fish and Fisheries, 21(4):774–796, July 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Pihlajamaki:2020:CFA

[967] Mia Pihlajamäki, Inari Helle, Päivi Haapasaari, Simo Sarkki, Sakari Kuikka, and Annukka Lehikoinen. Catching the future: Applying Bayesian belief networks to exploratory scenario storylines to assess longterm changes in Baltic herring (Clupea harengus membras, Clupeidae) and salmon (Salmo salar, Salmonidae) fisheries. Fish and Fisheries, 21 (4):797–812, July 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Durante:2020:STA

[968] Leonardo Maia Durante, Michael Peter Beentjes, and Stephen Richard Wing. Shifting trophic architecture of marine fisheries in New Zealand: Implications for guiding effective ecosystem-based management. Fish and Fisheries, 21(4):813–830, July 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Song:2020:CDS

[969] Andrew M. Song, Joeri Scholtens, Kate Barclay, Simon R. Bush, Michael Fabinyi, Dedi S. Adhuri, and Milton Haughton. Collateral damage? Small-scale fisheries in the global fight against IUU fishing. Fish and Fisheries, 21(4):831–843, July 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Farmery:2020:MMC

[970] Anna K. Farmery, Ingrid E. van Putten, Michelle Phillipov, and Alistair McIlgorm. Are media messages to consume more under-utilized seafood species reliable? Fish and Fisheries, 21(4):844–855, July 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2020:IIe

[971] Anonymous. Issue information. Fish and Fisheries, 21(5):i–ii, September 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Canales:2020:RFS

[972] T. Mariella Canales, Gustav W. Delius, and Richard Law. Regulation of fish stocks without stock–recruitment relationships: the case of small pelagic fish. Fish and Fisheries, 21(5):857–871, September 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Wijermans:2020:BDF

[973] Nanda Wijermans, Wiebren J. Boonstra, Kirill Orach, Jonas Hentati-Sundberg, and Maja Schlüter. Behavioural diversity in fishing — Towards a next generation of fishery models. Fish and Fisheries, 21(5): 872–890, September 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Smith:2020:MFR

[974] Brian E. Smith and Laurel A. Smith. Multispecies functional responses reveal reduced predation at high prey densities and varied responses among and within trophic groups. Fish and Fisheries, 21(5):891–905, September 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

DiLorenzo:2020:ASM

[975] Manfredi Di Lorenzo, Paolo Guidetti, Antonio Di Franco, Antonio Calò, and Joachim Claudet. Assessing spillover from marine protected areas and its drivers: a meta-analytical approach. Fish and Fisheries, 21(5): 906–915, September 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kuparinen:2020:ACR

[976] Anna Kuparinen and Silva Uusi-Heikkilä. Atlantic cod recovery from the Allee effect zone: contrasting ecological and evolutionary rescue. Fish and Fisheries, 21(5):916–926, September 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Wegscheider:2020:MMF

[977] Bernhard Wegscheider, Tommi Linnansaari, and R. Allen Curry. Mesohabitat modelling in fish ecology: a global synthesis. Fish and Fisheries, 21(5):927–939, September 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Sharma:2020:OMD

[978] Rishi Sharma, Polina Levontin, Toshihide Kitakado, Laurence Kell, Iago Mosqueira, Ai Kimoto, Rob Scott, Carolina Minte-Vera, Paul De Bruyn, Yimin Ye, Jana Kleineberg, Jo Lindsay Walton, Shana Miller, and Arni Magnusson. Operating model design in tuna regional fishery management organizations: Current practice, issues and implications. Fish and Fisheries, 21(5):940–961, September 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

vandeWolfshaar:2020:FWF

[979] Karen E. van de Wolfshaar, P. Daniël van Denderen, Tim Schellekens, and Tobias van Kooten. Food web feedbacks drive the response of benthic macrofauna to bottom trawling. Fish and Fisheries, 21(5):962–972, September 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Solomon:2020:FMS

[980] Christopher T. Solomon, Colin J. Dassow, Carolyn M. Iwicki, Olaf P. Jensen, Stuart E. Jones, Greg G. Sass, Ashley Trudeau, Brett T. van Poorten, and Dane Whittaker. Frontiers in modelling social–ecological dynamics of recreational fisheries: a review and synthesis. Fish and Fisheries, 21(5):973–991, September 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Belhabib:2020:NFG

[981] Dyhia Belhabib, Philippe Le Billon, and David J. Wrathall. Narco-fish: Global fisheries and drug trafficking. Fish and Fisheries, 21(5):992–1007, September 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Gullestad:2020:MTS

[982] Peter Gullestad, Svein Sundby, and Olav Sigurd Kjesbu. Management of transboundary and straddling fish stocks in the Northeast Atlantic in view of climate-induced shifts in spatial distribution. Fish and Fisheries, 21(5):1008–1026, September 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anderson:2020:PES

[983] David M. Anderson and James F. Gillooly. Predicting egg size across temperatures in marine teleost fishes. Fish and Fisheries, 21(5):1027– 1033, September 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Watson:2020:ASI

[984] Joseph W. Watson, Kieran Hyder, Robin Boyd, Robert Thorpe, Marc Simon Weltersbach, Keno Ferter, Steven J. Cooke, Shovonlal Roy, and Richard M. Sibly. Assessing the sublethal impacts of anthropogenic stressors on fish: an energy-budget approach. Fish and Fisheries, 21(5): 1034–1045, September 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kennelly:2020:BBM

[985] Steven J. Kennelly. Bycatch beknown: Methodology for jurisdictional reporting of fisheries discards — using Australia as a case study. Fish and Fisheries, 21(5):1046–1066, September 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Sheaves:2020:CPC

[986] Marcus Sheaves, Kátya Abrantes, Adam Barnett, Claudia Benham, Patricia Dale, Carlo Mattone, Alison Sheaves, Nathan Waltham, and Michael Bradley. The consequences of paradigm change and poorly validated science: the example of the value of mangroves to fisheries. Fish and Fisheries, 21(5):1067–1075, September 2020. CODEN FFI-IAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2020:IIf

[987] Anonymous. Issue information. Fish and Fisheries, 21(6):i–ii, October 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Glover:2020:FLL

[988] Kevin A. Glover, Vidar Wennevik, Kjetil Hindar, Øystein Skaala, Peder Fiske, Monica F. Solberg, Ola H. Diserud, Terje Svåsand, Sten Karlsson, Lasse B. Andersen, and Ellen S. Grefsrud. The future looks like the past: Introgression of domesticated Atlantic salmon escapees in a risk assessment framework. Fish and Fisheries, 21(6):1077–1091, October 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Holland:2020:LPT

[989] Matthew M. Holland, James A. Smith, Jason D. Everett, Adriana Vergés, and Iain M. Suthers. Latitudinal patterns in trophic structure of temperate reef-associated fishes and predicted consequences of climate change. Fish and Fisheries, 21(6):1092–1108, October 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Medina:2020:RAB

[990] Antonio Medina. Reproduction of Atlantic bluefin tuna. Fish and Fisheries, 21(6):1109–1119, October 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Oyanedel:2020:SNC

[991] Rodrigo Oyanedel, Stefan Gelcich, and E. J. Milner-Gulland. A synthesis of (non-)compliance theories with applications to small-scale fisheries research and practice. Fish and Fisheries, 21(6):1120–1134, October 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kindsvater:2020:CSS

[992] Holly K. Kindsvater, Kim Tallaksen Halvorsen, Tonje Knutsen Sørdalen, and Suzanne H. Alonzo. The consequences of size-selective fishing mortality for larval production and sustainable yield in species with obligate male care. Fish and Fisheries, 21(6):1135–1149, October 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hasan:2020:MAG

[993] Md. Mehedi Hasan, Rachel L. Tulloch, Peter C. Thomson, Herman W. Raadsma, and Mehar S. Khatkar. Meta-analysis of genetic parameters of production traits in cultured shrimp species. Fish and Fisheries, 21 (6):1150–1174, October 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Pleizier:2020:MAG

[994] Naomi K. Pleizier, Dirk Algera, Steven J. Cooke, and Colin J. Brauner. A meta-analysis of gas bubble trauma in fish. Fish and Fisheries, 21(6): 1175–1194, October 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Chollett:2020:CBD

[995] Iliana Chollett and D. Ross Robertson. Comparing biodiversity databases: Greater Caribbean reef fishes as a case study. Fish and Fisheries, 21(6):1195–1212, October 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Nevatte:2020:SRC

[996] Ryan J. Nevatte and Jane E. Williamson. The sawshark redemption: Current knowledge and future directions for sawsharks (Pristiophoridae). Fish and Fisheries, 21(6):1213–1237, October 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bolgan:2020:UPL

[997] Marta Bolgan and Eric Parmentier. The unexploited potential of listening to deep-sea fish. Fish and Fisheries, 21(6):1238–1252, October 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cucherousset:2020:DFF

[998] Julien Cucherousset and Julian D. Olden. Are domesticated freshwater fish an underappreciated culprit of ecosystem change? Fish and Fisheries, 21(6):1253–1258, October 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Sheaves:2020:OVS

[999] Marcus Sheaves, Michael Bradley, Cesar Herrera, Carlo Mattone, Caitlin Lennard, Janine Sheaves, and Dmitry A. Konovalov. Optimizing video sampling for juvenile fish surveys: Using deep learning and evaluation of assumptions to produce critical fisheries parameters. Fish and Fisheries, 21(6):1259–1276, October 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Radkhah:2020:FPT

[1000] Ali Reza Radkhah and Soheil Eagderi. Fish protection technologies and fish ways for downstream migration. Fish and Fisheries, 21(6):1277– 1278, October 2020. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2021:IIa

[1001] Anonymous. Issue information. Fish and Fisheries, 22(1):i–ii, January 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ma:2021:CIN

[1002] Shuyang Ma, Yongjun Tian, Caihong Fu, Haiqing Yu, Jianchao Li, Yang Liu, Jiahua Cheng, Rong Wan, and Yoshiro Watanabe. Climate-induced nonlinearity in pelagic communities and non-stationary relationships with physical drivers in the Kuroshio ecosystem. Fish and Fisheries, 22(1):1–17, January 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

vanDeurs:2021:BLR

[1003] Mikael van Deurs, Mollie E. Brooks, Martin Lindegren, Ole Henriksen, and Anna Rindorf. Biomass limit reference points are sensitive to estimation method, time-series length and stock development. Fish and Fisheries, 22(1):18–30, January 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Tixier:2021:WLM

[1004] Paul Tixier, Mary-Anne Lea, Mark A. Hindell, Dirk Welsford, Camille Mazé, Sophie Gourguet, and John P. Y. Arnould. When large marine predators feed on fisheries catches: Global patterns of the depredation conflict and directions for coexistence. Fish and Fisheries, 22(1):31–53, January 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

France:2021:HAF

[1005] Robert L. France. Historical anecdotes of fishing pressure: Misconstrued “sea serpent” sightings provide evidence for antecedent entanglement of marine biota in the British Isles. Fish and Fisheries, 22(1):54–71, January 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Mazor:2021:TFI

[1006] Tessa Mazor, C. Roland Pitcher, Wayne Rochester, Michel J. Kaiser, Jan G. Hiddink, Simon Jennings, Ricardo Amoroso, Robert A. Mc-Connaughey, Adriaan D. Rijnsdorp, Ana M. Parma, Petri Suuronen, Jeremy Collie, Marija Sciberras, Lara Atkinson, Deon Durholtz, Jim R. Ellis, Stefan G. Bolam, Michaela Schratzberger, Elena Couce, Jacqueline Eggleton, Clement Garcia, Paulus Kainge, Sarah Paulus, Johannes N. Kathena, Mayya Gogina, P. Daniël van Denderen, Aimee A. Keller, Beth H. Horness, and Ray Hilborn. Trawl fishing impacts on the status of seabed fauna in diverse regions of the globe. Fish and Fisheries, 22(1):72–86, January 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

MagalhaesdaSilvaFreitas:2021:QSK

[1007] Tiago Magalhães da Silva Freitas, Juliana Stropp, Bárbara Borges Calegari, Joaquín Calatayud, Paulo De Marco, Jr., Luciano Fogaça de Assis Montag, and Joaquín Hortal. Quantifying shortfalls in the knowledge on neotropical Auchenipteridae fishes. Fish and Fisheries, 22(1):87–104, January 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Flowers:2021:BSU

[1008] Kathryn I. Flowers, Michael R. Heithaus, and Yannis P. Papastamatiou. Buried in the sand: Uncovering the ecological roles and importance of rays. Fish and Fisheries, 22(1):105–127, January 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Miranda:2021:BEV

[1009] Leandro E. Miranda, Giancarlo Coppola, Hunter R. Hatcher, Matthew B. Jargowsky, Zachary S. Moran, and Michael C. Rhodes. A bird’s-eye view of reservoirs in the Mississippi Basin tips a need for large-scale coordination. Fish and Fisheries, 22(1):128–140, January 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Goethel:2021:WDY

[1010] Daniel R. Goethel, Katelyn M. Bosley, Brian J. Langseth, Jonathan J. Deroba, Aaron M. Berger, Dana H. Hanselman, and Amy M. Schueller. Where do you think you’re going? Accounting for ontogenetic and climate-induced movement in spatially stratified integrated population assessment models. Fish and Fisheries, 22(1):141–160, January 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Post:2021:SGT

[1011] Søren Post, Karl Michael Werner, Ismael Núñez-Riboni, Léon Chafik, Hjálmar Hátún, and Teunis Jansen. Subpolar gyre and temperature drive boreal fish abundance in Greenland waters. Fish and Fisheries, 22 (1):161–174, January 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hamza:2021:CIL

[1012] Faseela Hamza, Vinu Valsala, Anju Mallissery, and Grinson George. Climate impacts on the landings of Indian oil sardine over the south-eastern Arabian Sea. Fish and Fisheries, 22(1):175–193, January 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Welch:2021:SCW

[1013] David Warren Welch, Aswea Dawn Porter, and Erin Leanne Rechisky. A synthesis of the coast-wide decline in survival of West Coast Chinook salmon (Oncorhynchus tshawytscha, Salmonidae). Fish and Fisheries, 22 (1):194–211, January 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Schickele:2021:ESP

[1014] Alexandre Schickele, Eric Goberville, Boris Leroy, Gregory Beaugrand, Tarek Hattab, Patrice Francour, and Virginie Raybaud. European small pelagic fish distribution under global change scenarios. Fish and Fisheries, 22(1):212–225, January 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Haas:2021:RFM

[1015] Bianca Haas, Marcus Haward, Jeffrey McGee, and Aysha Fleming. Regional fisheries management organizations and the new biodiversity agreement: Challenge or opportunity? Fish and Fisheries, 22(1):226– 231, January 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

White:2021:EEC

[1016] Easton R. White, Halley E. Froehlich, Jessica A. Gephart, Richard S. Cottrell, Trevor A. Branch, Rahul Agrawal Bejarano, and Julia K. Baum. Early effects of COVID-19 on US fisheries and seafood consumption. Fish and Fisheries, 22(1):232–239, January 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hart:2021:ORS

[1017] Paul J. B. Hart. Ocean recovery. a sustainable future for global fisheries? Fish and Fisheries, 22(1):240, January 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2021:IIb

[1018] Anonymous. Issue information. Fish and Fisheries, 22(2):i–ii, March 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hart:2021:RHB

[1019] Paul J. B. Hart and Robert Arlinghaus. Representing human behaviour in ecosystem models. Fish and Fisheries, 22(2):241–242, March 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Reid:2021:TES

[1020] Andrea J. Reid, Lauren E. Eckert, John-Francis Lane, Nathan Young, Scott G. Hinch, Chris T. Darimont, Steven J. Cooke, Natalie C. Ban, and Albert Marshall. “Two-eyed seeing”: an indigenous framework to transform fisheries research and management. Fish and Fisheries, 22 (2):243–261, March 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Saraux:2021:SIN

[1021] Claire Saraux, William J. Sydeman, John F. Piatt, Tycho Anker-Nilssen, Jonas Hentati-Sundberg, Sophie Bertrand, Philippe M. Cury, Robert W. Furness, James A. Mills, Henrik Österblom, Giannina Passuni, Jean-Paul Roux, Lynne J. Shannon, and Robert J. M. Crawford. Seabird-induced natural mortality of forage fish varies with fish abundance: Evidence from five ecosystems. Fish and Fisheries, 22(2):262–279, March 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Frawley:2021:CSF

[1022] Timothy H. Frawley, Barbara A. Muhling, Stephanie Brodie, Mary C. Fisher, Desiree Tommasi, Gwendal Le Fol, Elliott L. Hazen, Stephen S. Stohs, Elena M. Finkbeiner, and Michael G. Jacox. Changes to the structure and function of an albacore fishery reveal shifting social-ecological realities for Pacific northwest fishermen. Fish and Fisheries, 22(2):280– 297, March 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cochrane:2021:RSE

[1023] Kevern L. Cochrane. Reconciling sustainability, economic efficiency and equity in marine fisheries: Has there been progress in the last 20 years? Fish and Fisheries, 22(2):298–323, March 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

McArley:2021:FHC

[1024] Tristan J. McArley, Erik Sandblom, and Neill A. Herbert. Fish and hyperoxia — from cardiorespiratory and biochemical adjustments to aquaculture and ecophysiology implications. Fish and Fisheries, 22(2):324– 355, March 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

LaMesa:2021:PCR

[1025] Mario La Mesa, Facundo Llompart, Emilio Riginella, and Joseph T. Eastman. Parental care and reproductive strategies in notothenioid fishes. Fish and Fisheries, 22(2):356–376, March 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Dias:2021:FET

[1026] Rosa Maria Dias, Anielly Galego de Oliveira, Matheus Tenório Baumgartner, Mirtha Amanda Angulo-Valencia, and Angelo Antonio Agostinho. Functional erosion and trait loss in fish assemblages from neotropical reservoirs: The man beyond the environment. Fish and Fisheries, 22(2):377–390, March 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Finucci:2021:GDB

[1027] Brittany Finucci, Jessica Cheok, David A. Ebert, Katelyn Herman, Peter M. Kyne, and Nicholas K. Dulvy. Ghosts of the deep — biodiversity, fisheries, and extinction risk of ghost sharks. Fish and Fisheries, 22 (2):391–412, March 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Mion:2021:MCF

[1028] Monica Mion, Stefanie Haase, Jakob Hemmer-Hansen, Annelie Hilvarsson, Karin Hüssy, Maria Krüger-Johnsen, Uwe Krumme, Kate McQueen, Maris Plikshs, Krzysztof Radtke, Franziska Maria Schade, Francesca Vitale, and Michele Casini. Multidecadal changes in fish growth rates estimated from tagging data: a case study from the Eastern Baltic cod (Gadus morhua, Gadidae). Fish and Fisheries, 22(2):413–427, March 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Fulton:2021:OIE

[1029] Elizabeth Ann Fulton. Opportunities to improve ecosystem-based fisheries management by recognizing and overcoming path dependency and cognitive bias. Fish and Fisheries, 22(2):428–448, March 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Jamwal:2021:MOH

[1030] Ankur Jamwal and Vikas Phulia. Multisectoral one health approach to make aquaculture and fisheries resilient to a future pandemic-like situation. Fish and Fisheries, 22(2):449–463, March 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2021:IIc

[1031] Anonymous. Issue information. Fish and Fisheries, 22(3):i–ii, May 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hart:2021:PPR

[1032] Paul J. B. Hart and Robert Arlinghaus. Plastic pollution in rivers and lakes — an indicator of an even bigger consequence of global change? Fish and Fisheries, 22(3):465–466, May 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Parker:2021:MFF

[1033] Ben Parker, Demetra Andreou, Iain D. Green, and J. Robert Britton. Microplastics in freshwater fishes: Occurrence, impacts and future perspectives. Fish and Fisheries, 22(3):467–488, May 2021. CODEN FFI-IAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Andrews:2021:FBC

[1034] Evan J. Andrews, Jeremy Pittman, and Derek R. Armitage. Fisher behaviour in coastal and marine fisheries. Fish and Fisheries, 22(3):489– 502, May 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Shelton:2021:RSP

[1035] Andrew Olaf Shelton, Genoa H. Sullaway, Eric J. Ward, Blake E. Feist, Kayleigh A. Somers, Vanessa J. Tuttle, Jordan T. Watson, and William H. Satterthwaite. Redistribution of salmon populations in the northeast Pacific Ocean in response to climate. Fish and Fisheries, 22 (3):503–517, May 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Vince:2021:PCE

[1036] Joanna Vince, Britta Denise Hardesty, and Chris Wilcox. Progress and challenges in eliminating illegal fishing. Fish and Fisheries, 22(3):518– 531, May 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Johnson:2021:EUF

[1037] Andrew Frederick Johnson, Susanna Lidström, Ingrid Kelling, Chris Williams, Simone Niederműller, Katrin Vilhelm Poulsen, Samantha Burgess, Rhona Kent, and William Davies. The European Union’s fishing activity outside of European waters and the Sustainable Development Goals. Fish and Fisheries, 22(3):532–545, May 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Zak:2021:SLF

[1038] Jakub Žák, Milan Vrtílek, Matej Polačik, Radim Blažek, and Martin Reichard. Short-lived fishes: Annual and multivoltine strategies. Fish and Fisheries, 22(3):546–561, May 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Pettitt-Wade:2021:CME

[1039] Harri Pettitt-Wade, Lisa L. Loseto, Andrew Majewski, and Nigel E. Hussey. Cod movement ecology in a warming world: Circumpolar Arctico gadids. Fish and Fisheries, 22(3):562–591, May 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Schiller:2021:RIE

[1040] Laurenne Schiller and Megan Bailey. Rapidly increasing eco-certification coverage transforming management of world’s tuna fisheries. Fish and Fisheries, 22(3):592–604, May 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ovegaard:2021:CPE

[1041] Maria K. Ovegård, Niels Jepsen, Mikaela Bergenius Nord, and Erik Petersson. Cormorant predation effects on fish populations: a global metaanalysis. Fish and Fisheries, 22(3):605–622, May 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Pacheco:2021:MEU

[1042] Carlos Pacheco, Carlos Bustamante, and Miguel Araya. Mass-effect: Understanding the relationship between age and otolith weight in fishes. Fish and Fisheries, 22(3):623–633, May 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Miller:2021:SSS

[1043] Timothy J. Miller and Elizabeth N. Brooks. Steepness is a slippery slope. Fish and Fisheries, 22(3):634–645, May 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Karnauskas:2021:EEQ

[1044] Mandy Karnauskas, John F. Walter III, Christopher R. Kelble, Matthew McPherson, Skyler R. Sagarese, J. Kevin Craig, Adyan Rios, William J. Harford, Seann Regan, Steven D. Giordano, and Morgan Kilgour. To EBFM or not to EBFM? that is not the question. Fish and Fisheries, 22(3):646–651, May 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Halpern:2021:LNP

[1045] Benjamin S. Halpern, Jason Maier, Heather J. Lahr, Gordon Blasco, Christopher Costello, Richard S. Cottrell, Olivier Deschenes, Danielle M. Ferraro, Halley E. Froehlich, Gavin G. McDonald, Katherine D. Millage, and Michael J. Weir. The long and narrow path for novel cell-based seafood to reduce fishing pressure for marine ecosystem recovery. Fish and Fisheries, 22(3):652–664, May 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2021:IId

[1046] Anonymous. Issue information. Fish and Fisheries, 22(4):i–ii, July 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Coscia:2021:WUE

[1047] Ilaria Coscia and Allan D. McDevitt. Are we using eDNA metabarcoding to its full potential? Fish and Fisheries, 22(4):665–666, July 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Spence:2021:QUD

[1048] Michael A. Spence, Robert B. Thorpe, Paul G. Blackwell, Finlay Scott, Richard Southwell, and Julia L. Blanchard. Quantifying uncertainty and dynamical changes in multi-species fishing mortality rates, catches and biomass by combining state-space and size-based multi-species models. Fish and Fisheries, 22(4):667–681, July 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Birdsong:2021:RAS

[1049] Max Birdsong, Len M. Hunt, and Robert Arlinghaus. Recreational angler satisfaction: What drives it? Fish and Fisheries, 22(4):682–706, July 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Borland:2021:IST

[1050] Hayden P. Borland, Ben L. Gilby, Christopher J. Henderson, Javier X. Leon, Thomas A. Schlacher, Rod M. Connolly, Simon J. Pittman, Marcus Sheaves, and Andrew D. Olds. The influence of seafloor terrain on fish and fisheries: a global synthesis. Fish and Fisheries, 22(4):707–734, July 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hershey:2021:UCF

[1051] Henry Hershey. Updating the consensus on fishway efficiency: a metaanalysis. Fish and Fisheries, 22(4):735–748, July 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Paxton:2021:FDR

[1052] Avery B. Paxton, Stacey L. Harter, Steve W. Ross, Christina M. Schobernd, Brendan J. Runde, Paul J. Rudershausen, Katrina H. Johnson, Kyle W. Shertzer, Nathan M. Bacheler, Jeffrey A. Buckel, G. Todd Kellison, and J. Christopher Taylor. Four decades of reef observations illuminate deep-water grouper hotspots. Fish and Fisheries, 22(4):749– 761, July 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Righton:2021:IQP

[1053] David Righton, Adam Piper, Kim Aarestrup, Elsa Amilhat, Claude Belpaire, John Casselman, Martin Castonguay, Estibaliz Díaz, Hendrik Dörner, Elisabeth Faliex, Eric Feunteun, Nobuto Fukuda, Reinhold Hanel, Celine Hanzen, Don Jellyman, Kenzo Kaifu, Kieran McCarthy, Michael J. Miller, Thomas Pratt, Pierre Sasal, Robert Schabetsberger, Hiromi Shiraishi, Gaël Simon, Niklas Sjöberg, Kristen Steele, Katsumi Tsukamoto, Alan Walker, Håkan Westerberg, Kazuki Yokouchi, and Matthew Gollock. Important questions to progress science and sustainable management of anguillid eels. Fish and Fisheries, 22(4):762–788, July 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Zhang:2021:LBT

[1054] Fan Zhang, Kevin B. Reid, and Thomas D. Nudds. The longer the better? Trade-offs in fisheries stock assessment in dynamic ecosystems. Fish and Fisheries, 22(4):789–797, July 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Su:2021:CFO

[1055] Shu Su, Yi Tang, Jessica Chen, Bowen Chang, and Yong Chen. A comprehensive framework for operating science-based fisheries management: a checklist for using the best available science. Fish and Fisheries, 22(4): 798–811, July 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Giron-Nava:2021:SFE

[1056] Alfredo Giron-Nava, Vicky W. Y. Lam, Octavio Aburto-Oropeza, William W. L. Cheung, Benjamin S. Halpern, U. Rashid Sumaila, and Andrés M. Cisneros-Montemayor. Sustainable fisheries are essential but not enough to ensure well-being for the world’s fishers. Fish and Fisheries, 22(4):812–821, July 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

DAlessandro:2021:SED

[1057] Simone D’Alessandro and Stefano Mariani. Sifting environmental DNA metabarcoding data sets for rapid reconstruction of marine food webs. Fish and Fisheries, 22(4):822–833, July 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Spijkers:2021:IPI

[1058] Jessica Spijkers, Gerald G. Singh, Colette C. C. Wabnitz, Henrik Österblom, Graeme S. Cumming, and Tiffany H. Morrison. Identifying predictors of international fisheries conflict. Fish and Fisheries, 22 (4):834–850, July 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Li:2021:TIF

[1059] Mi-Ling Li, Yoshitaka Ota, Philip J. Underwood, Gabriel Reygondeau, Katherine Seto, Vicky W. Y. Lam, David Kroodsma, and William W. L. Cheung. Tracking industrial fishing activities in African waters from space. Fish and Fisheries, 22(4):851–864, July 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Rogan:2021:EPE

[1060] Emer Rogan, Andrew J. Read, and Per Berggren. Empty promises: The European Union is failing to protect dolphins and porpoises from fisheries by-catch. Fish and Fisheries, 22(4):865–869, July 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2021:IIe

[1061] Anonymous. Issue information. Fish and Fisheries, 22(5):i–ii, September 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Arlinghaus:2021:MHF

[1062] Robert Arlinghaus and Paul J. B. Hart. Mental health and fisheries — an understudied topic of global relevance. Fish and Fisheries, 22(5): 871–873, September 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Harrison:2021:GAF

[1063] Trevor D. Harrison and Alan K. Whitfield. A global assessment of fish estuary associations: a numerical approach to assessing estuary-associated fish functional guilds. Fish and Fisheries, 22(5):874–898, September 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Woodill:2021:OSP

[1064] A. John Woodill, Maria Kavanaugh, Michael Harte, and James R. Watson. Ocean seascapes predict distant-water fishing vessel incursions into exclusive economic zones. Fish and Fisheries, 22(5):899–910, September 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hilborn:2021:GSG

[1065] Ray Hilborn, Daniel J. Hively, Nicole Baker Loke, Carryn L. de Moor, Hiroyuki Kurota, Johannes N. Kathena, Pamela M. Mace, Cóilín Minto, Ana M. Parma, Juan-Carlos Quiroz, and Michael C. Melnychuk. Global status of groundfish stocks. Fish and Fisheries, 22(5):911–928, September 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Grewelle:2021:RRD

[1066] Richard E. Grewelle, Elizabeth Mansfield, Fiorenza Micheli, and Giulio De Leo. Redefining risk in data-poor fisheries. Fish and Fisheries, 22(5): 929–940, September 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Beveren:2021:FFR

[1067] Elisabeth Van Beveren, Hugues P. Benoît, and Daniel E. Duplisea. Forecasting fish recruitment in age-structured population models. Fish and Fisheries, 22(5):941–954, September 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Nunez-Riboni:2021:EDS

[1068] Ismael Núñez-Riboni, Anna Akimova, and Anne F. Sell. Effect of data spatial scale on the performance of fish habitat models. Fish and Fisheries, 22(5):955–973, September 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Takasuka:2021:RSR

[1069] Akinori Takasuka, Haruka Nishikawa, Sho Furuichi, and Ryuji Yukami. Revisiting sardine recruitment hypotheses: Egg-production-based survival index improves understanding of recruitment mechanisms of fish under climate variability. Fish and Fisheries, 22(5):974–986, September 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ledee:2021:CSA

[1070] Elodie J. I. Lédée, Michelle R. Heupel, Matthew D. Taylor, Robert G. Harcourt, Fabrice R. A. Jaine, Charlie Huveneers, Vinay Udyawer, Hamish A. Campbell, Russell C. Babcock, Xavier Hoenner, Adam Barnett, Matias Braccini, Stephanie Brodie, Paul A. Butcher, Gwenael Cadiou, Ross G. Dwyer, Mario Espinoza, Luciana C. Ferreira, Lachlan Fetterplace, Anthony Fowler, Alastair R. Harborne, Nathan A. Knott, Michael Lowry, Jaime McAllister, Rory McAuley, Mark Meekan, Kade Mills, Victor M. Peddemors, Richard Pillans, Jayson Semmens, Amy F. Smoothey, Conrad Speed, Kilian Stehfest, Dylan van der Meulen, and Colin A. Simpfendorfer. Continental-scale acoustic telemetry and network analysis reveal new insights into stock structure. Fish and Fisheries, 22(5):987–1005, September 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cottrell:2021:SBT

[1071] Richard S. Cottrell, Danielle M. Ferraro, Gordon D. Blasco, Benjamin S. Halpern, and Halley E. Froehlich. The search for blue transitions in aquaculture-dominant countries. Fish and Fisheries, 22(5):1006–1023, September 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Moore:2021:CRP

[1072] Jonathan W. Moore, Brendan M. Connors, and Emma E. Hodgson. Conservation risks and portfolio effects in mixed-stock fisheries. Fish and Fisheries, 22(5):1024–1040, September 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Gatti:2021:RTV

[1073] Paul Gatti, Jonathan A. D. Fisher, Frédéric Cyr, Peter S. Galbraith, Dominique Robert, and Arnault Le Bris. A review and tests of validation and sensitivity of geolocation models for marine fish tracking. Fish and Fisheries, 22(5):1041–1066, September 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bolin:2021:ASC

[1074] Jessica A. Bolin, David S. Schoeman, Karen J. Evans, Scott F. Cummins, and Kylie L. Scales. Achieving sustainable and climate-resilient fisheries requires marine ecosystem forecasts to include fish condition. Fish and Fisheries, 22(5):1067–1084, September 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

McClanahan:2021:VME

[1075] Timothy R. McClanahan, Emily S. Darling, Sangeeta Mangubhai, Georgina G. Gurney, W. Peni Lestari, Margaret Fox, Stacy D. Jupiter, Dhila A. Yulistianti, Nyawira A. Muthiga, and Stephanie D’Agata. Views of management effectiveness in tropical reef fisheries. Fish and Fisheries, 22(5):1085–1104, September 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Walker:2021:EVC

[1076] Terence I. Walker, Robert W. Day, Cynthia A. Awruch, Justin D. Bell, Juan Matias Braccini, Derek R. Dapp, Licia Finotto, Lorenz H. Frick, Karla C. Garcés-García, Leonardo Guida, Charlie Huveneers, Camila L. Martins, Bastien E. A. Rochowski, Javier Tovar-Ávila, Fabian I. Trinnie, and Richard D. Reina. Ecological vulnerability of the chondrichthyan fauna of southern Australia to the stressors of climate change, fishing and other anthropogenic hazards. Fish and Fisheries, 22(5):1105–1135, September 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

King:2021:MHC

[1077] Tanya J. King, Rachel Turner, Vincent Versace, Kirsten Abernethy, Sue Kilpatrick, and Susan Brumby. Mental health in the commercial fishing industry: Modern uncertainties and traditional risks. Fish and Fisheries, 22(5):1136–1149, September 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Thorpe:2021:RFE

[1078] Robert Thorpe. Review of “Fish Ecology, Evolution, and Exploitation: a new theoretical synthesis” by Ken Andersen. Fish and Fisheries, 22(5): 1150–1151, September 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2021:C

[1079] Anonymous. Corrigendum. Fish and Fisheries, 22(5):1152, September 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2021:IIf

[1080] Anonymous. Issue information. Fish and Fisheries, 22(6):i–ii, November 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

vanPutten:2021:TBB

[1081] Ingrid van Putten. Tricky business: Blue crimes in small scale fisheries. Fish and Fisheries, 22(6):1153–1154, November 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Temple:2021:CRT

[1082] Andrew J. Temple, Ethan Westmerland, and Per Berggren. By-catch risk for toothed whales in global small-scale fisheries. Fish and Fisheries, 22(6):1155–1159, November 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Satizabal:2021:ECR

[1083] Paula Satizábal, Philippe Le Billon, Dyhia Belhabib, Lina M. Saavedra-Díaz, Isabela Figueroa, Gina Noriega, and Nathan J. Bennett. Ethical considerations for research on small-scale fisheries and blue crimes. Fish and Fisheries, 22(6):1160–1166, November 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Siple:2021:CMS

[1084] Margaret C. Siple, Laura E. Koehn, Kelli F. Johnson, André E. Punt, T. Mariella Canales, Piera Carpi, Carryn L. de Moor, José A. A. De Oliveira, Jin Gao, Nis S. Jacobsen, Mimi E. Lam, Roberto Licandeo, Martin Lindegren, Shuyang Ma, Gudmundur J. Óskarsson, Sonia Sanchez-Maroño, Szymon Smoliński, Szymon Surma, Yongjun Tian, Desiree Tommasi, Mariano Gutiérrez T., Verena Trenkel, Stephani G. Zador, and Fabian Zimmermann. Considerations for management strategy evaluation for small pelagic fishes. Fish and Fisheries, 22(6):1167– 1186, November 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Downie:2021:IHA

[1085] Adam T. Downie, Jeffrey M. Leis, Peter F. Cowman, Mark I. McCormick, and Jodie L. Rummer. The influence of habitat association on swimming performance in marine teleost fish larvae. Fish and Fisheries, 22(6): 1187–1212, November 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Robinson:2021:USA

[1086] Lucy M. Robinson, Ingrid van Putten, Blake S. Cavve, Catherine Longo, Matt Watson, Lynda Bellchambers, Emily Fisher, and Fabio Boschetti. Understanding societal approval of the fishing industry and the influence of third-party sustainability certification. Fish and Fisheries, 22(6):1213– 1226, November 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Druon:2021:MPF

[1087] Jean-Noël Druon, Didier Gascuel, Maurizio Gibin, Antonella Zanzi, Jean-Marc Fromentin, Francesco Colloca, Pierre Hélaouët, Marta Coll, Alessandro Mannini, Joanna K. Bluemel, Chiara Piroddi, Francois Bastardie, Diego Macias-Moy, Paraskevas Vasilakopoulos, Henning Winker, Natalia Serpetti, Jordi Guillen, Andreas Palialexis, Michaël Gras, Zeynep Hekim, Laurent Dubroca, Cecilia Pinto, Jeroen Steenbeek, and Jann Martinsohn. Mesoscale productivity fronts and local fishing opportunities in the European Seas. Fish and Fisheries, 22(6):1227–1247, November 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bryndum-Buchholz:2021:SCC

[1088] Andrea Bryndum-Buchholz, Derek P. Tittensor, and Heike K. Lotze. The status of climate change adaptation in fisheries management: Policy, legislation and implementation. Fish and Fisheries, 22(6):1248–1273, November 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Gilbey:2021:EMD

[1089] John Gilbey, Kjell Rong Utne, Vidar Wennevik, Alexander Christian Beck, Kyrre Kausrud, Kjetil Hindar, Carlos Garcia de Leaniz, Corrine Cherbonnel, Jamie Coughlan, Tom F. Cross, Eileen Dillane, Dennis Ensing, Eva García-Vázquez, Lars R. Hole, Marianne Holm, Jens Christian Holst, Jan Arge Jacobsen, Arne J. Jensen, Sten Karlsson, Niall Ó Maoiléidigh, Kjell Arne Mork, Einar Eg Nielsen, Leif Nøttestad, Craig R. Primmer, Paulo Prodöhl, Sergey Prusov, Jamie R. Stevens, Katie Thomas, Ken Whelan, Philip McGinnity, and Eric Verspoor. The early marine distribution of Atlantic salmon in the North-east Atlantic: a genetically informed stock-specific synthesis. Fish and Fisheries, 22(6): 1274–1306, November 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

DeFilippo:2021:SRA

[1090] Lukas B. DeFilippo and Jan Ohlberger. Stochastic recruitment alters the frequencies of alternative life histories in age-structured populations. Fish and Fisheries, 22(6):1307–1320, November 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Lowerre-Barbieri:2021:MEE

[1091] Susan K. Lowerre-Barbieri, Claudia Friess, Lucas P. Griffin, Danielle Morley, Gregory B. Skomal, Joel W. Bickford, Neil Hammerschlag, Mitchell J. Rider, Matthew J. Smukall, Maurits P. M. van Zinnicq Bergmann, Tristan L. Guttridge, Andrea M. Kroetz, R. Dean Grubbs, Carissa L. Gervasi, Jennifer S. Rehage, Gregg R. Poulakis, Kim Bassos-Hull, Jayne M. Gardiner, Grace A. Casselberry, Joy Young, Matt Perkinson, Debra L. Abercrombie, Dustin T. Addis, Barbara A. Block, Alejandro Acosta, Aaron J. Adams, Andy J. Danylchuk, Steven J. Cooke, Frederick G. Whoriskey, and Jacob W. Brownscombe. Movescapes and eco-evolutionary movement strategies in marine fish: Assessing a connectivity hotspot. Fish and Fisheries, 22(6):1321–1344, November 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Silvar-Viladomiu:2021:MRP

[1092] Paula Silvar-Viladomiu, Cóilín Minto, Ghassen Halouani, Luke Batts, Deirdre Brophy, Colm Lordan, and David G. Reid. Moving reference point goalposts and implications for fisheries sustainability. Fish and Fisheries, 22(6):1345–1358, November 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Thorson:2021:EFS

[1093] James T. Thorson, Steven J. Barbeaux, Daniel R. Goethel, Kelly A. Kearney, Edward A. Laman, Julie K. Nielsen, Matthew R. Siskey, Kevin Siwicke, and Grant G. Thompson. Estimating fine-scale movement rates and habitat preferences using multiple data sources. Fish and Fisheries, 22(6):1359–1376, November 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ovando:2021:IES

[1094] Daniel Ovando, Ray Hilborn, Cole Monnahan, Merrill Rudd, Rishi Sharma, James T. Thorson, Yannick Rousseau, and Yimin Ye. Improving estimates of the state of global fisheries depends on better data. Fish and Fisheries, 22(6):1377–1391, November 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ramirez-Valdez:2021:AAI

[1095] Arturo Ramírez-Valdez, Timothy J. Rowell, Katherine E. Dale, Matthew T. Craig, Larry G. Allen, Juan Carlos Villaseñor-Derbez, Andrés M. Cisneros-Montemayor, Arturo Hernández-Velasco, Jorge Torre, Jennifer Hofmeister, and Brad E. Erisman. Asymmetry across international borders: Research, fishery and management trends and economic value of the giant sea bass (Stereolepis gigas). Fish and Fisheries, 22(6):1392–1411, November 2021. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2022:IIa

[1096] Anonymous. Issue information. Fish and Fisheries, 23(1):i–ii, January 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hart:2022:E

[1097] Paul J. B. Hart, Gary Carvalho, and Robert Arlinghaus. Editorial. Fish and Fisheries, 23(1):1–3, January 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Galappaththi:2022:CCA

[1098] Eranga K. Galappaththi, Vasantha B. Susarla, Samantha J. T. Loutet, Stephanie T. Ichien, Amanda A. Hyman, and James D. Ford. Climate change adaptation in fisheries. Fish and Fisheries, 23(1):4–21, January 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kuczenski:2022:PGL

[1099] Brandon Kuczenski, Camila Vargas Poulsen, Eric L. Gilman, Michael Musyl, Roland Geyer, and Jono Wilson. Plastic gear loss estimates from remote observation of industrial fishing activity. Fish and Fisheries, 23 (1):22–33, January 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Simmance:2022:NFA

[1100] Fiona A. Simmance, Philippa J. Cohen, Cindy Huchery, Sarah Sutcliffe, Sharon K. Suri, Xavier Tezzo, Shakuntala H. Thilsted, Peter Oosterveer, Cynthia McDougall, Molly Ahern, Sarah Freed, Kendra A. Byrd, Joshua Wesana, Ian G. Cowx, David J. Mills, Michael Akester, Chin Yee Chan, Joseph Nagoli, Jillian T. Wate, and Michael J. Phillips. Nudging fisheries and aquaculture research towards food systems. Fish and Fisheries, 23 (1):34–53, January 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Holm:2022:AEN

[1101] Poul Holm, John Nicholls, Patrick W. Hayes, Josh Ivinson, and Bernard Allaire. Accelerated extractions of North Atlantic cod and herring, 1520– 1790. Fish and Fisheries, 23(1):54–72, January 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Mildenberger:2022:IPA

[1102] Tobias K. Mildenberger, Casper W. Berg, Alexandros Kokkalis, Adrian R. Hordyk, Chantel Wetzel, Nis S. Jacobsen, André E. Punt, and J. Rasmus Nielsen. Implementing the precautionary approach into fisheries management: Biomass reference points and uncertainty buffers. Fish and Fisheries, 23(1):73–92, January 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Czegledi:2022:TDF

[1103] István Czeglédi, András Specziár, and Tibor Erős. Temporal dynamics of freshwater fish assemblages, their background and methods of quantifications — a synthesis. Fish and Fisheries, 23(1):93–108, January 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Arthur:2022:SSF

[1104] Robert I. Arthur, Daniel J. Skerritt, Anna Schuhbauer, Naazia Ebrahim, Richard M. Friend, and U. Rashid Sumaila. Small-scale fisheries and local food systems: Transformations, threats and opportunities. Fish and Fisheries, 23(1):109–124, January 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Koehn:2022:FHD

[1105] John Zachary Koehn, Edward H. Allison, Karen Villeda, Zelin Chen, Marisa Nixon, Emily Crigler, Lily Zhao, Michelle Chow, Bapu Vaitla, Shakuntala H. Thilsted, Joeri Scholtens, Christina C. Hicks, and Neil Andrew. Fishing for health: Do the world’s national policies for fisheries and aquaculture align with those for nutrition? Fish and Fisheries, 23 (1):125–142, January 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Fryxell:2022:SSC

[1106] David C. Fryxell, Emma R. Moffett, Michael T. Kinnison, Kevin S. Simon, and Eric P. Palkovacs. From southern swamps to cosmopolitan model: Humanity’s unfinished history with mosquitofish. Fish and Fisheries, 23(1):143–161, January 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Stoffers:2022:ERH

[1107] Twan Stoffers, Antonie D. Buijse, Johan A. J. Verreth, and Leopold A. J. Nagelkerke. Environmental requirements and heterogeneity of rheophilic fish nursery habitats in European lowland rivers: Current insights and future challenges. Fish and Fisheries, 23(1):162–182, January 2022. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Carvalho:2022:GRC

[1108] Paul G. Carvalho and Austin Humphries. Gear restrictions create conservation and fisheries trade-offs for management. Fish and Fisheries, 23 (1):183–194, January 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Nieman:2022:SSC

[1109] Chelsey L. Nieman and Christopher T. Solomon. Slow social change: Implications for open access recreational fisheries. Fish and Fisheries, 23 (1):195–201, January 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Sinan:2022:CDR

[1110] Hussain Sinan, Megan Bailey, Quentin Hanich, and Kamal Azmi. Common but differentiated rights and responsibilities in tuna fisheries management. Fish and Fisheries, 23(1):202–212, January 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Sydeman:2022:ISD

[1111] William J. Sydeman, Sarah Ann Thompson, John F. Piatt, Stephani G. Zador, and Martin W. Dorn. Integrating seabird dietary and groundfish stock assessment data: Can puffins predict pollock spawning stock biomass in the North Pacific? Fish and Fisheries, 23(1):213–226, January 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ladich:2022:SSL

[1112] Friedrich Ladich. Shut up or shout loudly: Predation threat and sound production in fishes. Fish and Fisheries, 23(1):227–238, January 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

English:2022:CCV

[1113] Philina A. English, Eric J. Ward, Christopher N. Rooper, Robyn E. Forrest, Luke A. Rogers, Karen L. Hunter, Andrew M. Edwards, Brendan M. Connors, and Sean C. Anderson. Contrasting climate velocity impacts in warm and cool locations show that effects of marine warming are worse in already warmer temperate waters. Fish and Fisheries, 23 (1):239–255, January 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Nielsen:2022:CSS

[1114] Max Nielsen, Peder Andersen, Frank Asche, Hans Ellefsen, Cecilia Hammarlund, Ayoe Hoff, Dadi Mar Kristofersson, Rasmus Nielsen, Heri á Rógvi, Kristin Roll, Hordur Sævaldsson, Jarno Virtanen, and Staffan Waldo. Can small-scale fisheries survive market-based management?

Nordic evidence. Fish and Fisheries, 23(1):256–272, January 2022. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2022:IIb

[1115] Anonymous. Issue information. Fish and Fisheries, 23(2):i–ii, March 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Waples:2022:CKM

[1116] Robin S. Waples and Pierre Feutry. Close-kin methods to estimate census size and effective population size. Fish and Fisheries, 23(2):273–293, March 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bosley:2022:FPM

[1117] Katelyn M. Bosley, Amy M. Schueller, Daniel R. Goethel, Dana H. Hanselman, Kari H. Fenske, Aaron M. Berger, Jonathan J. Deroba, and Brian J. Langseth. Finding the perfect mismatch: Evaluating misspecification of population structure within spatially explicit integrated population models. Fish and Fisheries, 23(2):294–315, March 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Marshall:2022:HDS

[1118] Dustin J. Marshall, Diego R. Barneche, and Craig R. White. How does spawning frequency scale with body size in marine fishes? Fish and Fisheries, 23(2):316–323, March 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hogrenning:2022:ECB

[1119] Egil Hogrenning and Arne Eide. Environmental constraints, biological growth and fleet dynamics of a developing fishery: A model study of the Barents Sea snow crab (Chionoecetes opilio, Majidae) fishery. Fish and Fisheries, 23(2):324–341, March 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ciannelli:2022:OSC

[1120] Lorenzo Ciannelli, Anna B. Neuheimer, Leif C. Stige, Kenneth T. Frank, Joel M. Durant, Mary Hunsicker, Lauren A. Rogers, Steve Porter, Geir Ottersen, and Natalia A. Yaragina. Ontogenetic spatial constraints of sub-arctic marine fish species. Fish and Fisheries, 23(2):342–357, March 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Durif:2022:UHS

[1121] Caroline M. F. Durif, Hans Hagen Stockhausen, Anne Berit Skiftesvik, Alessandro Cresci, Daniel Nyqvist, and Howard I. Browman. A unifying hypothesis for the spawning migrations of temperate anguillid eels. Fish and Fisheries, 23(2):358–375, March 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Buba:2022:SDF

[1122] Yehezkel Buba, John P. DeLong, and Jonathan Belmaker. Synthesizing drivers of fish functional responses across species. Fish and Fisheries, 23(2):376–391, March 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Tirronen:2022:TAE

[1123] Maria Tirronen, Tommi Perälä, and Anna Kuparinen. Temporary Allee effects among non-stationary recruitment dynamics in depleted gadid and flatfish populations. Fish and Fisheries, 23(2):392–406, March 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Franco:2022:GDS

[1124] Ana Clara Sampaio Franco, Ana Cristina Petry, Marcela Rosa Tavares, Taís de Fátima Ramos Guimarães, and Luciano Neves dos Santos. Global distribution of the South American peacock basses Cichla spp. follows human interference. Fish and Fisheries, 23(2):407–421, March 2022. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

McClanahan:2022:BPF

[1125] Timothy R. McClanahan, Alan M. Friedlander, Laurent Wantiez, Nicholas A. J. Graham, J. Henrich Bruggemann, Pascale Chabanet, and Remy M. Oddenyo. Best-practice fisheries management associated with reduced stocks and changes in life histories. Fish and Fisheries, 23(2): 422–444, March 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ammar:2022:QSE

[1126] Yosr Ammar, Rudi Voss, Susa Niiranen, and Thorsten Blenckner. Quantifying socio-economic novelty in fisheries social-ecological systems. Fish and Fisheries, 23(2):445–461, March 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Zhou:2022:EME

[1127] Shijie Zhou, Roy A. Deng, Matthew R. Dunn, Simon D. Hoyle, Yeming Lei, and Ashley J. Williams. Evaluating methods for estimating shark natural mortality rate and management reference points using life-history parameters. Fish and Fisheries, 23(2):462–477, March 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Watz:2022:ASR

[1128] Johan Watz, David Aldvén, Patrik Andreasson, Khadija Aziz, Marco Blixt, Olle Calles, Kristine Lund Bjørnås, Ivan Olsson, Martin Österling, Sanna Stålhammar, Johan Tielman, and John J. Piccolo. Atlantic salmon in regulated rivers: Understanding river management through the ecosystem services lens. Fish and Fisheries, 23(2):478–491, March 2022. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hilborn:2022:ABM

[1129] Ray Hilborn, Vera N. Agostini, Milani Chaloupka, Serge M. Garcia, Leah R. Gerber, Eric Gilman, Quentin Hanich, Amber Himes-Cornell, Alistair J. Hobday, David Itano, Michel J. Kaiser, Hilario Murua, Daniel Ovando, Graham M. Pilling, Jake C. Rice, Rishi Sharma, Kurt M. Schaefer, Craig J. Severance, Nathan G. Taylor, and Mark Fitchett. Areabased management of blue water fisheries: Current knowledge and research needs. Fish and Fisheries, 23(2):492–518, March 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2022:IIc

[1130] Anonymous. Issue information. Fish and Fisheries, 23(3):i–ii, May 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Oomen:2022:NBN

[1131] Rebekah A. Oomen, Halvor Knutsen, Even Moland, Esben Moland Olsen, Jan Atle Knutsen, Louis Bernatchez, Nils Chr. Stenseth, and Paul Hart. Nothing so boundless, nothing so patient: Obituary for Professor Jeffrey Alexander Hutchings (1958–2022). Fish and Fisheries, 23(3):519– 521, May 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Mason:2022:ACR

[1132] Julia G. Mason, Jacob G. Eurich, Jacqueline D. Lau, Willow Battista, Christopher M. Free, Katherine E. Mills, Kanae Tokunaga, Lily Z. Zhao, Mark Dickey-Collas, Mireia Valle, Gretta T. Pecl, Joshua E. Cinner, Tim R. McClanahan, Edward H. Allison, Whitney R. Friedman, Claudio Silva, Eleuterio Yáñez, María Á. Barbieri, and Kristin M. Kleisner. Attributes of climate resilience in fisheries: From theory to practice. Fish and Fisheries, 23(3):522–544, May 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Poisson:2022:NTI

[1133] François Poisson, Pierre Budan, Sylvain Coudray, Eric Gilman, Takahito Kojima, Michael Musyl, and Tsutomu Takagi. New technologies to improve bycatch mitigation in industrial tuna fisheries. Fish and Fisheries, 23(3):545–563, May 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bessa:2022:IAD

[1134] Eduardo Bessa, Manuela Lombardi Brandão, and Eliane Gonçalves de Freitas. Integrative approach on the diversity of nesting behaviour in fishes. Fish and Fisheries, 23(3):564–583, May 2022. CODEN FFI-IAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Barclay:2022:TWB

[1135] Kate M. Barclay, Arlene N. Satapornvanit, Victoria M. Syddall, and Meryl J. Williams. Tuna is women’s business too: Applying a gender lens to four cases in the Western and Central Pacific. Fish and Fisheries, 23(3):584–600, May 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kjesbu:2022:HMI

[1136] Olav Sigurd Kjesbu, Svein Sundby, Anne Britt Sandø, Maud Alix, Solfrid Sætre Hjøllo, Maik Tiedemann, Mette Skern-Mauritzen, Claudia Junge, Maria Fossheim, Cecilie Thorsen Broms, Guldborg Søvik, Fabian Zimmermann, Kjell Nedreaas, Elena Eriksen, Hannes Höffle, Ann Merete Hjelset, Cecilie Kvamme, Yves Reecht, Halvor Knutsen, Asgeir Aglen, Ole Thomas Albert, Erik Berg, Bjarte Bogstad, Caroline Durif, Kim Tallaksen Halvorsen, Åge Høines, Carsten Hvingel, Edda Johannesen, Espen Johnsen, Even Moland, Mari Skuggedal Myksvoll, Leif Nøttestad, Erik Olsen, Georg Skaret, Jon Egil Skjæraasen, Aril Slotte, Arved Staby, Erling Kåre Stenevik, Jan Erik Stiansen, Martina Stiasny, Jan H. Sundet, Frode Vikebø, and Geir Huse. Highly mixed impacts of near-future climate change on stock productivity proxies in the North East Atlantic. Fish and Fisheries, 23(3):601–615, May 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ovando:2022:HEC

[1137] Daniel Ovando, Christopher M. Free, Olaf P. Jensen, and Ray Hilborn. A history and evaluation of catch-only stock assessment models. Fish and Fisheries, 23(3):616–630, May 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bassi:2022:EAL

[1138] Arshpreet Bassi, Oliver P. Love, Steven J. Cooke, Theresa R. Warriner, Christopher M. Harris, and Christine L. Madliger. Effects of artificial light at night on fishes: a synthesis with future research priorities. Fish and Fisheries, 23(3):631–647, May 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Clark:2022:RCS

[1139] Timothy P. Clark. Racial capitalism and the sea: Development and change in Black maritime labour, and what it means for fisheries and a blue economy. Fish and Fisheries, 23(3):648–662, May 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Horbowy:2022:AFL

[1140] Jan Horbowy and Kristiina Hommik. Analysis of F_{msy} in light of lifehistory traits — effects on its proxies and length-based indicators. Fish and Fisheries, 23(3):663–679, May 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Stoffels:2022:FLC

[1141] Rick J. Stoffels, Paul Humphries, Nick R. Bond, and Amina E. Price. Fragmentation of lateral connectivity and fish population dynamics in large rivers. Fish and Fisheries, 23(3):680–696, May 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hodgdon:2022:GCS

[1142] Cameron T. Hodgdon, Noah S. Khalsa, Yunzhou Li, Ming Sun, Robert Boenish, and Yong Chen. Global crustacean stock assessment modelling: Reconciling available data and complexity. Fish and Fisheries, 23(3):697– 707, May 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Canty:2022:DSS

[1143] Steven W. J. Canty and Jessica L. Deichmann. Do small-scale fisheries have the capacity to provide food security to coastal populations? Fish and Fisheries, 23(3):708–718, May 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Puley:2022:DCM

[1144] Melina Puley and Anthony Charles. Dissecting co-management: Fisher participation across management components and implications for governance. Fish and Fisheries, 23(3):719–732, May 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kerry:2022:MGF

[1145] Christopher R. Kerry, Owen M. Exeter, and Matthew J. Witt. Monitoring global fishing activity in proximity to seamounts using automatic identification systems. Fish and Fisheries, 23(3):733–749, May 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Suthers:2022:MGR

[1146] Iain M. Suthers, Zoe White, Charles Hinchliffe, Daniel S. Falster, Anthony J. Richardson, and Jason D. Everett. The Mortality/Growth ratio of larval fish and the slope of the zooplankton size-spectrum. Fish and Fisheries, 23(3):750–757, May 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Schuch:2022:TIJ

[1147] Esther Schuch and Andries Richter. Tracing intuitive judgement of experts in fish stock assessment data. Fish and Fisheries, 23(3):758–767, May 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2022:C

[1148] Anonymous. Corrigendum. Fish and Fisheries, 23(3):768–769, May 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2022:IId

[1149] Anonymous. Issue information. Fish and Fisheries, 23(4):i–ii, July 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Arlinghaus:2022:OWR

[1150] Robert Arlinghaus. The ocean’s whistleblower: The remarkable life and work of Daniel Pauly. Fish and Fisheries, 23(4):771–772, July 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Grati:2022:MSS

[1151] Fabio Grati, Ernesto Azzurro, Martina Scanu, Anna Nora Tassetti, Luca Bolognini, Stefano Guicciardi, Sergio Vitale, Danilo Scannella, Pierluigi Carbonara, Branko Dragičević, Zdravko Ikica, Arian Palluqi, Bojan Marčeta, Hisham Ghmati, Akram Turki, Mourad Cherif, Marouene Bdioui, Othman Jarboui, Nader Benhadjhamida, Jurgen Mifsud, Nicoletta Milone, Luca Ceriola, and Enrico Arneri. Mapping small-scale fisheries through a coordinated participatory strategy. Fish and Fisheries, 23(4):773–785, July 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Wosnick:2022:GAS

[1152] Natascha Wosnick, Renata D. Leite, Eloísa P. Giareta, Danny Morick, and Michael Musyl. Global assessment of shark strandings. Fish and Fisheries, 23(4):786–799, July 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Robinson:2022:MFM

[1153] James P. W. Robinson, Kirsty L. Nash, Julia L. Blanchard, Nis S. Jacobsen, Eva Maire, Nicholas A. J. Graham, M. Aaron MacNeil, Jessica Zamborain-Mason, Edward H. Allison, and Christina C. Hicks. Managing fisheries for maximum nutrient yield. Fish and Fisheries, 23(4): 800–811, July 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Rindorf:2022:SCD

[1154] Anna Rindorf, Mikael van Deurs, Daniel Howell, Eider Andonegi, Aaron Berger, Bjarte Bogstad, Noel Cadigan, Bjarki Tór Elvarsson, Niels Hintzen, Marie Savina Roland, Marc Taylor, Vanessa Trijoulet, Tobias van Kooten, Fan Zhang, and Jeremy Collie. Strength and consistency of density dependence in marine fish productivity. Fish and Fisheries, 23(4):812–828, July 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

MacKenzie:2022:NFD

[1155] Brian R. MacKenzie, Piero Addis, Pietro Battaglia, Pierpaolo Consoli, Franco Andaloro, Gianluca Sarà, Anders Nielsen, and Teresa Romeo. Neglected fishery data sources as indicators of pre-industrial ecological properties of Mediterranean swordfish (Xiphias gladius, Xiphiidae). Fish and Fisheries, 23(4):829–846, July 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Wilson:2022:CMA

[1156] Shaun K. Wilson, Christopher J. Fulton, Nicholas A. J. Graham, Rene A. Abesamis, Charlotte Berkström, Darren J. Coker, Martial Depczynski, Richard D. Evans, Rebecca Fisher, Jordan Goetze, Andrew Hoey, Thomas H. Holmes, Michel Kulbicki, Mae Noble, James P. W. Robinson, Michael Bradley, Carolina Åkerlund, Luke T. Barrett, Abner A. Bucol, Matthew J. Birt, Dinorah H. Chacin, Karen M. Chong-Seng, Linda Eggertsen, Maria Eggertsen, David Ellis, Priscilla T. Y. Leung, Paul K. S. Lam, Joshua van Lier, Paloma A. Matis, Alejandro Pérez-Matus, Camilla V. H. Piggott, Ben T. Radford, Stina Tano, and Paul Tinkler. The contribution of macroalgae-associated fishes to small-scale tropical reef fisheries. Fish and Fisheries, 23(4):847–861, July 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Llorente-Rodriguez:2022:EHF

[1157] Laura Llorente-Rodríguez, Oliver E. Craig, André Carlo Colonese, Matthew von Tersch, Eufrasia Roselló-Izquierdo, Eduardo González Gómez de Agüero, Carlos Fernández-Rodríguez, Juan Antonio Quirós-Castillo, Begoña López-Arias, Ricard Marlasca-Martín, James Nottingham, and Arturo Morales Muñiz. Elucidating historical fisheries’ networks in the Iberian Peninsula using stable isotopes. Fish and Fisheries, 23(4):862– 873, July 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Zimmer:2022:PAR

[1158] Alex M. Zimmer and Steve F. Perry. Physiology and aquaculture: a review of ion and acid-base regulation by the gills of fishes. Fish and Fisheries, 23(4):874–898, July 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Garlock:2022:GIM

[1159] Taryn Garlock, James L. Anderson, Frank Asche, Martin D. Smith, Edward Camp, Jingjie Chu, Kai Lorenzen, and Stefania Vannuccini. Global insights on managing fishery systems for the three pillars of sustainability. Fish and Fisheries, 23(4):899–909, July 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Roscher:2022:SDO

[1160] Matthew B. Roscher, Edward H. Allison, David J. Mills, Hampus Eriksson, Denis Hellebrandt, and Neil L. Andrew. Sustainable development outcomes of livelihood diversification in small-scale fisheries. Fish and Fisheries, 23(4):910–925, July 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Lennox:2022:DFD

[1161] Robert J. Lennox, Valerio Sbragaglia, Knut Wiik Vollset, Lene K. Sortland, Loren McClenachan, Ivan Jarić, Meaghan L. Guckian, Keno Ferter, Andy J. Danylchuk, Steven J. Cooke, Robert Arlinghaus, and William M. Twardek. Digital fisheries data in the Internet age: Emerging tools for research and monitoring using online data in recreational fisheries. Fish and Fisheries, 23(4):926–940, July 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Belton:2022:DFI

[1162] Ben Belton, Derek S. Johnson, Eric Thrift, Jonah Olsen, Mostafa Ali Reza Hossain, and Shakuntala Haraksingh Thilsted. Dried fish at the intersection of food science, economy, and culture: a global survey. Fish and Fisheries, 23(4):941–962, July 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Amos:2022:CRS

[1163] Holly Amos, Alfredo Giron-Nava, Tu Nguyen, Andrés M. Cisneros-Montemayor, Mathieu Colléter, Pedro C. González-Espinosa, and Wilf Swartz. Collapse and recovery of seafood wholesale prices in time of COVID-19. Fish and Fisheries, 23(4):963–976, July 2022. CODEN FFI-IAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Saleh:2022:CVD

[1164] Alzayat Saleh, Marcus Sheaves, and Mostafa Rahimi Azghadi. Computer vision and deep learning for fish classification in underwater habitats: a survey. Fish and Fisheries, 23(4):977–999, July 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Murphy:2022:GPI

[1165] Robert Murphy Jr, Michael Downs, Nathan Wolf, and Bradley P. Harris. Guiding principles for integrating stakeholder-based data into marine fisheries decision-making with a focus on USA fisheries management. Fish and Fisheries, 23(4):1000–1008, July 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Jin:2022:BSA

[1166] Binsong Jin, Kirk O. Winemiller, Wenwei Ren, Dave Tickner, Xiaohua Wei, Li Guo, Qin Li, Hui Zhang, Paulo Santos Pompeu, Marc Goichot, Sikai Wang, Ruichang Shen, and Xiaoping Wu. Basin-scale approach needed for Yangtze River fisheries restoration. Fish and Fisheries, 23 (4):1009–1015, July 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Sohns:2022:EPI

[1167] Antonia Sohns, Gordon M. Hickey, and Owen Temby. Exploring the potential impacts of machine learning on trust in fishery management. Fish and Fisheries, 23(4):1016–1023, July 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2022:E

[1168] Anonymous. Erratum. Fish and Fisheries, 23(4):1024, July 2022. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2022:IIe

[1169] Anonymous. Issue information. Fish and Fisheries, 23(5):i–ii, September 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hamza:2022:PAS

[1170] Faseela Hamza, Vinu Valsala, and Hamza Varikoden. The puzzle of the anchovy–sardine inverse fishery at the south-eastern coast of the Arabian Sea and climate variability. Fish and Fisheries, 23(5):1025–1038, September 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Willer:2022:LDF

[1171] David F. Willer, Joshua I. Brian, Christina J. Derrick, Marcus Hicks, Alerick Pacay, Arlie H. McCarthy, Sophie Benbow, Holly Brooks, Carolina Hazin, Nibedita Mukherjee, Chris J. McOwen, Jessica Walker, and Daniel Steadman. ‘Destructive fishing’ — a ubiquitously used but vague term? Usage and impacts across academic research, media and policy. Fish and Fisheries, 23(5):1039–1054, September 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Watson:2022:DIM

[1172] James R. Watson and A. John Woodill. Detecting illegal maritime activities from anomalous multiscale fleet behaviours. Fish and Fisheries, 23(5):1055–1069, September 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Virdin:2022:SEB

[1173] John Virdin, Tibor Vegh, Stephen Akester, Jingjie Chu, Andrew Baio, Jill Hamilton, Javier Macias, Steinar Matthiasson, Sheck Sherif, and Dana I. Grieco. A snapshot of the economic benefits from foreign bottom trawling in coastal West Africa: a mutually-beneficial trade in services, no winners or extractivism? Fish and Fisheries, 23(5):1070–1082, September 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Nilsen:2022:ERN

[1174] Ina Nilsen, Cecilie Hansen, Isaac Kaplan, Elizabeth Holmes, and Øystein Langangen. Exploring the role of Northeast Atlantic cod in the Barents Sea food web using a multi-model approach. Fish and Fisheries, 23(5): 1083–1098, September 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Galappaththi:2022:WEI

[1175] Madu Galappaththi, Derek Armitage, and Andrea M. Collins. Women’s experiences in influencing and shaping small-scale fisheries governance. Fish and Fisheries, 23(5):1099–1120, September 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Zhang:2022:ALS

[1176] Fan Zhang and Noel G. Cadigan. An age- and length-structured statistical catch-at-length model for hard-to-age fisheries stocks. Fish and Fisheries, 23(5):1121–1135, September 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Chollett:2022:TBU

[1177] Iliana Chollett, Larry Perruso, and Shay O’Farrell. Toward a better use of fisheries data in spatial planning. Fish and Fisheries, 23(5):1136–1149, September 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Talwar:2022:ERR

[1178] Brendan S. Talwar, Brooke Anderson, Cristopher G. Avalos-Castillo, María del Pilar Blanco-Parra, Alejandra Briones, Diego Cardeñosa, John K. Carlson, Patricia Charvet, Charles F. Cotton, Zoe Crysler, Danielle H. Derrick, Michael R. Heithaus, Katelyn B. Herman, Olga Koubrak, David W. Kulka, Peter M. Kyne, Oscar M. Lasso-Alcalá, Paola A. Mejía-Falla, Jorge Manuel Morales-Saldaña, Beatriz Naranjo-Elizondo, Andrés F. Navia, Nathan Pacoureau, Juan C. Peréz-Jiménez, Riley A. Pollom, Cassandra L. Rigby, Eric V. C. Schneider, Nikola Simpson, and Nicholas K. Dulvy. Extinction risk, reconstructed catches and management of chondrichthyan fishes in the Western Central Atlantic Ocean. Fish and Fisheries, 23(5):1150–1179, September 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Roberson:2022:SER

[1179] Leslie Roberson, Chris Wilcox, Germain Boussarie, Emma Dugan, Cristina Garilao, Kristofer Gonzalez, Madeline Green, Salit Kark, Kristin Kaschner, Carissa J. Klein, Yannick Rousseau, Dan Vallentyne, James E. M. Watson, and Jeremy J. Kiszka. Spatially explicit risk assessment of marine megafauna vulnerability to Indian Ocean tuna fisheries. Fish and Fisheries, 23(5):1180–1201, September 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Lindkvist:2022:USE

[1180] Emilie Lindkvist, Kara E. Pellowe, Steven M. Alexander, Elizabeth Drury O’Neill, Elena M. Finkbeiner, Alfredo Girón-Nava, Blanca González-Mon, Andrew F. Johnson, Jeremy Pittman, Caroline Schill, Nanda Wijermans, Örjan Bodin, Stefan Gelcich, and Marion Glaser. Untangling social–ecological interactions: a methods portfolio approach to tackling contemporary sustainability challenges in fisheries. Fish and Fisheries, 23(5):1202–1220, September 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Li:2022:TDO

[1181] Lingbo Li, Anne B. Hollowed, Edward D. Cokelet, Michelle M. McClure, Aimee A. Keller, Steve J. Barbeaux, and Wayne A. Palsson. Threedimensional ontogenetic shifts of groundfish in the Northeast Pacific. Fish and Fisheries, 23(5):1221–1239, September 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ford:2022:ICB

[1182] Jessica H. Ford, Chris Wold, Duncan Currie, and Chris Wilcox. Incentivising change to beneficial ownership and open registers — holding flag states responsible for their fleets and costs of illegal fishing. Fish and Fisheries, 23(5):1240–1248, September 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2022:IIf

[1183] Anonymous. Issue information. Fish and Fisheries, 23(6):i–ii, November 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Thurstan:2022:BRL

[1184] Ruth H. Thurstan. Book review: The limits of private governance: Norms and rules in a Mediterranean fishery. Grisel, F. Hart Publishing, 2021, 174 p. Fish and Fisheries, 23(6):1249–1250, November 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Clavero:2022:KAD

[1185] Miguel Clavero. The King’s aquatic desires: 16th-century fish and crayfish introductions into Spain. Fish and Fisheries, 23(6):1251–1263, November 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Heidrich:2022:APD

[1186] Kristina N. Heidrich, Maria José Juan-Jordá, Hilario Murua, Christopher D. H. Thompson, Jessica J. Meeuwig, and Dirk Zeller. Assessing progress in data reporting by tuna Regional Fisheries Management Organizations. Fish and Fisheries, 23(6):1264–1281, November 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Nyboer:2022:OSF

[1187] Elizabeth A. Nyboer, Holly S. Embke, Ashley M. Robertson, Robert Arlinghaus, Shannon Bower, Claudio Baigun, Douglas Beard, Steven J. Cooke, Ian G. Cowx, John D. Koehn, Roman Lyach, Marco Milardi, Warren Potts, and Abigail J. Lynch. Overturning stereotypes: The fuzzy boundary between recreational and subsistence inland fisheries. Fish and Fisheries, 23(6):1282–1298, November 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Calo:2022:MSS

[1188] Antonio Calò, Antonio Di Franco, Federico Quattrocchi, Charalampos Dimitriadis, Patricia Ventura, Marco Milazzo, and Paolo Guidetti. Multispecific small-scale fisheries rely on few, locally essential, species: Evidence from a multi-area study in the Mediterranean. Fish and Fisheries, 23(6):1299–1312, November 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hilborn:2022:RTA

[1189] Ray Hilborn, Claudio C. Buratti, Erich Díaz Acuña, Daniel Hively, Jeppe Kolding, Hiroyuki Kurota, Nicole Baker, Pamela M. Mace, Carryn L. de Moor, Soyoka Muko, Giacomo Chato Osio, Ana M. Parma, Juan-Carlos Quiroz, and Michael C. Melnychuk. Recent trends in abundance and fishing pressure of agency-assessed small pelagic fish stocks. Fish and Fisheries, 23(6):1313–1331, November 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Pepin:2022:IKC

[1190] Pierre Pepin, Jacquelyne King, Carrie Holt, Helen-Gurney Smith, Nancy Shackell, Kevin Hedges, and Alida Bundy. Incorporating knowledge of changes in climatic, oceanographic and ecological conditions in Canadian stock assessments. Fish and Fisheries, 23(6):1332–1346, November 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Wei:2022:MFU

[1191] Yaoguang Wei, Yunhong Duan, and Dong An. Monitoring fish using imaging sonar: Capacity, challenges and future perspective. Fish and Fisheries, 23(6):1347–1370, November 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Clarke:2022:LOC

[1192] Tayler M. Clarke, Colette C. C. Wabnitz, Thomas L. Frölicher, Gabriel Reygondeau, Daniel Pauly, and William W. L. Cheung. Linking observed changes in pelagic catches to temperature and oxygen in the Eastern Tropical Pacific. Fish and Fisheries, 23(6):1371–1382, November 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Sherman:2022:MRF

[1193] C. Samantha Sherman, Glenn Sant, Colin A. Simpfendorfer, Eric D. Digel, Patrick Zubick, Grant Johnson, Michael Usher, and Nicholas K. Dulvy. M-Risk: A framework for assessing global fisheries management efficacy of sharks, rays and chimaeras. Fish and Fisheries, 23(6):1383– 1399, November 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Holder:2022:WCU

[1194] Peter Elliot Holder, Chris M. Wood, Michael J. Lawrence, Timothy D. Clark, Cory D. Suski, Jean-Michel Weber, Andy J. Danylchuk, and Steven J. Cooke. Are we any closer to understanding why fish can die after severe exercise? Fish and Fisheries, 23(6):1400–1417, November 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Golden:2022:FWM

[1195] Abigail S. Golden, Brett van Poorten, and Olaf P. Jensen. Focusing on what matters most: Evaluating multiple challenges to stability in recreational fisheries. Fish and Fisheries, 23(6):1418–1438, November 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hossain:2022:GMA

[1196] Mostafa A. R. Hossain and Julian D. Olden. Global meta-analysis reveals diverse effects of microplastics on freshwater and marine fishes. Fish and Fisheries, 23(6):1439–1454, November 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Martino:2022:UIO

[1197] Jasmin C. Martino, Clive N. Trueman, Debashish Mazumder, Jagoda Crawford, and Zoë A. Doubleday. The universal imprint of oxygen isotopes can track the origins of seafood. Fish and Fisheries, 23(6):1455– 1468, November 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Oostdijk:2022:WAF

[1198] Maartje Oostdijk and Griffin Carpenter. Which attributes of fishing opportunities are linked to sustainable fishing? Fish and Fisheries, 23(6): 1469–1484, November 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Malone:2022:BET

[1199] Margaret A. Malone and Carlos M. Polivka. The behavioural ecology toolkit for fish management and conservation. Fish and Fisheries, 23(6): 1485–1506, November 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Prince:2022:SHT

[1200] Jeremy Prince, William J. Harford, Brett M. Taylor, and Steven J. Lindfield. Standard histological techniques systematically under-estimate the size fish start spawning. Fish and Fisheries, 23(6):1507–1516, November 2022. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2023:IIa

[1201] Anonymous. Issue information. Fish and Fisheries, 24(1):i–ii, January 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hart:2023:OEM

[1202] Paul J. B. Hart. Ocean ecology. Marine life in the age of humans. Fish and Fisheries, 24(1):1–2, January 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bakun:2023:VAC

[1203] Andrew Bakun. Vertical ambush corridors: Intriguing multi-mechanism ecological structures embedded in the kinetic fluid architectures of ocean living resource production systems. Fish and Fisheries, 24(1):3–20, January 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Law:2023:FBB

[1204] Richard Law and Michael J. Plank. Fishing for biodiversity by balanced harvesting. Fish and Fisheries, 24(1):21–39, January 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ma:2023:NSE

[1205] Shuyang Ma, Caihong Fu, Jianchao Li, Peng Sun, Yang Liu, Zhenjiang Ye, Yoshiro Watanabe, and Yongjun Tian. Non-stationary effects of multiple drivers on the dynamics of Japanese sardine (Sardinops melanostictus, Clupeidae). Fish and Fisheries, 24(1):40–55, January 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cronin:2023:PTG

[1206] Melissa R. Cronin, Julia E. Amaral, Alexis M. Jackson, Jennifer Jacquet, Katherine L. Seto, and Donald A. Croll. Policy and transparency gaps for oceanic shark and rays in high seas tuna fisheries. Fish and Fisheries, 24(1):56–70, January 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Karp:2023:PSD

[1207] Melissa A. Karp, Stephanie Brodie, James A. Smith, Kate Richerson, Rebecca L. Selden, Owen R. Liu, Barbara A. Muhling, Jameal F. Samhouri, Lewis A. K. Barnett, Elliott L. Hazen, Daniel Ovando, Jerome Fiechter, Michael G. Jacox, and Mercedes Pozo Buil. Projecting species distributions using fishery-dependent data. Fish and Fisheries, 24(1):71–92, January 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Abbott:2023:SST

[1208] Joshua K. Abbott, Yutaro Sakai, and Daniel S. Holland. Species, space and time: a quarter century of fishers’ diversification strategies on the US West Coast. Fish and Fisheries, 24(1):93–110, January 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Spijkers:2023:DGF

[1209] Jessica Spijkers, Mary Mackay, Jemma Turner, Asha McNeill, Kendra Travaille, and Chris Wilcox. Diversity of global fisheries governance: Types and contexts. Fish and Fisheries, 24(1):111–125, January 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bi:2023:HCA

[1210] Rujia Bi, Chip Collier, Roger Mann, Katherine E. Mills, Vincent Saba, John Wiedenmann, and Olaf P. Jensen. How consistent is the advice from stock assessments? Empirical estimates of inter-assessment bias and uncertainty for marine fish and invertebrate stocks. Fish and Fisheries, 24(1):126–141, January 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Sun:2023:UUF

[1211] Ming Sun, Yunzhou Li, and Yong Chen. Unveiling unselective fishing in China: a nationwide meta-analysis of multispecies fisheries. Fish and Fisheries, 24(1):142–158, January 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Rodrigues:2023:MDM

[1212] Lucas dos Santos Rodrigues, Maria Grazia Pennino, David Conesa, Eidi Kikuchi, Paul Gerhard Kinas, Fabiana Gonçalves Barbosa, and Luís Gustavo Cardoso. Modelling the distribution of marine fishery resources: Where are we? Fish and Fisheries, 24(1):159–175, January 2023. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Langangen:2023:EEL

[1213] Øystein Langangen, Jan Ohlberger, Leif Christian Stige, Rémi Patin, Lucie Buttay, Nils Christian Stenseth, Kotaro Ono, and Joël M. Durant. Effects of early life mass mortality events on fish populations. Fish and Fisheries, 24(1):176–186, January 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Amat-Trigo:2023:BTC

[1214] Fatima Amat-Trigo, Demetra Andreou, Phillipa K. Gillingham, and J. Robert Britton. Behavioural thermoregulation in cold-water freshwater fish: Innate resilience to climate warming? Fish and Fisheries, 24 (1):187–195, January 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2023:IIb

[1215] Anonymous. Issue information. Fish and Fisheries, 24(2):i–ii, March 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Carvalho:2023:RLF

[1216] Gary Carvalho, Robert Arlinghaus, Bronwyn Gillanders, and Anna Kuparinen. Re-launch of Fish and Fisheries special issues and call for proposals. Fish and Fisheries, 24(2):197, March 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Johnson:2023:FED

[1217] Andrew F. Johnson. Fishery ecosystem dynamics. Fish and Fisheries, 24(2):198, March 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ruiz-Diaz:2023:UEL

[1218] Raquel Ruiz-Díaz. Using an EBFM lens to guide the management of marine biological resources under changing conditions. Fish and Fisheries, 24(2):199–211, March 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Dowling:2023:FAF

[1219] Natalie A. Dowling, Jono R. Wilson, Jason M. Cope, Dawn T. Dougherty, Serena Lomonico, Carmen Revenga, Brian J. Snouffer, Natalio Godoy Salinas, Felipe Torres-Cañete, Rowan C. Chick, Ashley M. Fowler, and Ana M. Parma. The FishPath approach for fisheries management in a data- and capacity-limited world. Fish and Fisheries, 24(2):212–230, March 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Fischer:2023:RED

[1220] Simon H. Fischer, José A. A. De Oliveira, John D. Mumford, and Laurence T. Kell. Risk equivalence in data-limited and data-rich fisheries management: An example based on the ICES advice framework. Fish and Fisheries, 24(2):231–247, March 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Free:2023:HCR

[1221] Christopher M. Free, Tracey Mangin, John Wiedenmann, Conner Smith, Halley McVeigh, and Steven D. Gaines. Harvest control rules used in US federal fisheries management and implications for climate resilience.

Fish and Fisheries, 24(2):248–262, March 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ainsworth:2023:PFI

[1222] Rachel F. Ainsworth, Ian G. Cowx, and Simon J. Funge-Smith. Putting the fish into inland fisheries — a global allocation of historic inland fish catch. Fish and Fisheries, 24(2):263–278, March 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Carruthers:2023:MER

[1223] Thomas R. Carruthers, Quang C. Huynh, Adrian R. Hordyk, David Newman, Anthony D. M. Smith, Keith J. Sainsbury, Kevin Stokes, Alexander Morison, David Agnew, Ana Parma, Ignacio Sobrino, and Catherine Longo. Method evaluation and risk assessment: a framework for evaluating management strategies for data-limited fisheries. Fish and Fisheries, 24(2):279–296, March 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kjesbu:2023:LDS

[1224] Olav Sigurd Kjesbu, Maud Alix, Anne Britt Sandø, Espen Strand, Peter J. Wright, David G. Johns, Anders Thorsen, C. Tara Marshall, Kjell Gunnar Bakkeplass, Frode B. Vikebø, Mari Skuggedal Myksvoll, Geir Ottersen, Bridie J. M. Allan, Maria Fossheim, Jan Erik Stiansen, Geir Huse, and Svein Sundby. Latitudinally distinct stocks of Atlantic cod face fundamentally different biophysical challenges under on-going climate change. Fish and Fisheries, 24(2):297–320, March 2023. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cooper:2023:AES

[1225] Katja Cooper, Hugh Breakey, Melea Lewis, Rebecca Marshallsay, Alex Naraniecki, and Charles Sampford. Aquaculture ethics: a systematic quantitative review and critical analysis of aquaculture ethics scholarship. Fish and Fisheries, 24(2):321–338, March 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Shephard:2023:RUP

[1226] Samuel Shephard, Charles J. List, and Robert Arlinghaus. Reviving the unique potential of recreational fishers as environmental stewards of aquatic ecosystems. Fish and Fisheries, 24(2):339–351, March 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2023:IIc

[1227] Anonymous. Issue information. Fish and Fisheries, 24(3):i–ii, May 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Mackinson:2023:BRA

[1228] Steven Mackinson. Book review: Alaska Herring History: The story of Alaska’s herring fisheries and industry, James Mackovjak. Fish and Fisheries, 24(3):353, May 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bigman:2023:RCP

[1229] Jennifer S. Bigman, Nicholas C. Wegner, and Nicholas K. Dulvy. Revisiting a central prediction of the Gill Oxygen Limitation Theory: Gill area index and growth performance. Fish and Fisheries, 24(3):354–366, May 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Felix:2023:MAR

[1230] Luís Félix, Raquel Vieira, Sandra M. Monteiro, and Carlos Venâncio. A meta-analytic review of monoterpene for fish anaesthesia. Fish and Fisheries, 24(3):367–380, May 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Craig:2023:IPT

[1231] J. Kevin Craig and Jason S. Link. It is past time to use ecosystem models tactically to support ecosystem-based fisheries management: Case studies using Ecopath with Ecosim in an operational management context. Fish and Fisheries, 24(3):381–406, May 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Leurs:2023:WLM

[1232] Guido Leurs, Brian O. Nieuwenhuis, Thije J. Zuidewind, Nadia Hijner, Han Olff, and Laura L. Govers. Where land meets sea: Intertidal areas as key-habitats for sharks and rays. Fish and Fisheries, 24(3):407–426, May 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Croll:2023:CDD

[1233] Jasper C. Croll, Tobias van Kooten, and André M. de Roos. The consequences of density-dependent individual growth for sustainable harvesting and management of fish stocks. Fish and Fisheries, 24(3):427–438, May 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Szuwalski:2023:UCC

[1234] Cody S. Szuwalski, Anne B. Hollowed, Kirstin K. Holsman, James N. Ianelli, Christopher M. Legault, Michael C. Melnychuk, Dan Ovando, and Andre E. Punt. Unintended consequences of climate-adaptive fisheries management targets. Fish and Fisheries, 24(3):439–453, May 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Malick:2023:FTC

[1235] Michael J. Malick, James P. Losee, Gary Marston, Mickey Agha, Barry A. Berejikian, Brian R. Beckman, and Matthew Cooper. Fecundity trends of Chinook salmon in the Pacific Northwest. Fish and Fisheries, 24(3):454–465, May 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic). See corrigendum [1253].

Arevalo:2023:FLD

[1236] Elorri Arevalo, Henrique N. Cabral, Bertrand Villeneuve, Carl Possémé, and Mario Lepage. Fish larvae dynamics in temperate estuaries: a review on processes, patterns and factors that determine recruitment. Fish and Fisheries, 24(3):466–487, May 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Whitfield:2023:MIL

[1237] Alan K. Whitfield, Ian C. Potter, Francisco J. Neira, and Edward D. Houde. Modes of ingress by larvae and juveniles of marine fishes into estuaries: From microtidal to macrotidal systems. Fish and Fisheries, 24(3):488–503, May 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Nahuelhual:2023:RIF

[1238] Laura Nahuelhual, Tomás Vallejos, Gonzalo Campos, Ximena Vergara, Stefan Gelcich, and Rodrigo Estévez. Reframing illegal fishing in smallscale fisheries as a wicked problem. Fish and Fisheries, 24(3):504–521, May 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2023:IId

[1239] Anonymous. Issue information. Fish and Fisheries, 24(4):i–ii, July 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Metcalfe:2023:BRP

[1240] Neil B. Metcalfe. Book review of: A primer of life histories — Ecology, evolution and application by Jeffrey A. Hutchings Oxford University Press, 226 pp. Fish and Fisheries, 24(4):523–524, July 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anarkhis:2023:BRP

[1241] Nuranti Anarkhis. Book review: Probiotics in Aquaculture, Brian Austin and S. M. Sharifuzzaman Cham, Switzerland: Springer. 2022. 308 pp. $126.28 ebk. ISBN: 978-3-030-98621-6. Fish and Fisheries, 24(4):525– 526, July 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Xing:2023:MEM

[1242] Qinwang Xing, Haiqing Yu, Hui Wang, Shin ichi Ito, and Fei Chai. Mesoscale eddies modulate the dynamics of human fishing activities in the global midlatitude ocean. Fish and Fisheries, 24(4):527–543, July 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Licandeo:2023:DDM

[1243] Roberto Licandeo, Santiago de la Puente, Villy Christensen, Ray Hilborn, and Carl Walters. A delay-differential model for representing small pelagic fish stock dynamics and its application for assessing alternative management strategies under environmental uncertainty. Fish and Fisheries, 24(4):544–566, July 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Haberle:2023:FCI

[1244] Ines Haberle, Lav Bavčević, and Tin Klanjscek. Fish condition as an indicator of stock status: Insights from condition index in a food-limiting environment. Fish and Fisheries, 24(4):567–581, July 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Mangubhai:2023:GBV

[1245] Sangeeta Mangubhai, Kate M. Barclay, Sarah Lawless, and Natalie Makhoul. Gender-based violence: Relevance for fisheries practitioners. Fish and Fisheries, 24(4):582–594, July 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Atlas:2023:TCS

[1246] William I. Atlas, Matthew R. Sloat, William H. Satterthwaite, Thomas W. Buehrens, Charles K. Parken, Jonathan W. Moore, Nathan J. Mantua, Jon Hart, and Anna Potapova. Trends in Chinook salmon spawner abundance and total run size highlight linkages between life history, geography and decline. Fish and Fisheries, 24(4):595–617, July 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ladich:2023:HCY

[1247] Friedrich Ladich. Hearing in catfishes: 200 years of research. Fish and Fisheries, 24(4):618–634, July 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Lawson:2023:SCA

[1248] Julia Margaret Lawson and Conner Muir Smith. A synthetic control approach to estimate the effect of total allowable catches in the high seas. Fish and Fisheries, 24(4):635–651, July 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Free:2023:IMH

[1249] Christopher M. Free, Sean C. Anderson, Elizabeth A. Hellmers, Barbara A. Muhling, Michael O. Navarro, Kate Richerson, Lauren A. Rogers, William H. Satterthwaite, Andrew R. Thompson, Jenn M. Burt, Steven D. Gaines, Kristin N. Marshall, J. Wilson White, and Lyall F. Bellquist. Impact of the 2014–2016 marine heatwave on US and Canada West Coast fisheries: Surprises and lessons from key case studies. Fish and Fisheries, 24(4):652–674, July 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic). See correction [1298].

Tian:2023:LTV

[1250] Yongjun Tian, Caihong Fu, Akihiko Yatsu, Yoshiro Watanabe, Yang Liu, Jianchao Li, Dan Liu, Yumeng Pang, Jiahua Cheng, Ching-Hsein Ho, and Shuyang Ma. Long-term variability in the fish assemblage around Japan over the last century and early warning signals of regime shifts. Fish and Fisheries, 24(4):675–694, July 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bergseth:2023:CCG

[1251] Brock J. Bergseth, Adrian Arias, Michele L. Barnes, Iain Caldwell, Amber Datta, Stefan Gelcich, Sam H. Ham, Jacqueline D. Lau, Cristina Ruano-Chamorro, Patrick Smallhorn-West, Damian Weekers, Jessica Zamborain-Mason, and Joshua E. Cinner. Closing the compliance gap in marine protected areas with human behavioural sciences. Fish and Fisheries, 24(4):695–704, July 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2023:C

[1252] Anonymous. Corrigendum. Fish and Fisheries, 24(4):705, July 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic). See [948].

Ahrens:2023:CFT

[1253] Robert N. M. Ahrens, Micheal S. Allen, Carl Walters, and Robert Arlinghaus. Corrigendum to ‘Fecundity trends of Chinook salmon in the Pacific Northwest’. Fish and Fisheries, 21(3):706, July 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic). See [1235].

Anonymous:2023:IIe

[1254] Anonymous. Issue information. Fish and Fisheries, 24(5):i–ii, September 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Mutmainnah:2023:BRM

[1255] Mutmainnah Mutmainnah, Abigail Mary Moore, and Khusnul Khatimah. Book review: Maria E. Abate and David L.G Noakes. The Behavior, Ecology, and Evolution of Cichlid Fishes. Switzerland: Springer Dordrecht. 2021. 832 pp. $158.3 ebk. ISBN 978–94–024-2080-7. Fish and Fisheries, 24(5):707–708, September 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cavanagh:2023:EBF

[1256] Rachel D. Cavanagh. Book review: Ecosystem-based fisheries management — Progress, importance and impacts in the United States. Jason S. Link and Anthony R. Marshak Oxford, UK: Oxford University Press. 2021. 712 pp. Fish and Fisheries, 24(5):709–710, September 2023. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Melnychuk:2023:CSM

[1257] Michael C. Melnychuk, Charmane E. Ashbrook, Richard J. Bell, Lyall Bellquist, Kate Kauer, Jono R. Wilson, Ray Hilborn, and Jay Odell. Characterizing state-managed and unmanaged fisheries in coastal marine states and territories of the United States. Fish and Fisheries, 24(5): 711–729, September 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bigman:2023:GGA

[1258] Jennifer S. Bigman, Nicholas C. Wegner, and Nicholas K. Dulvy. Gills, growth and activity across fishes. Fish and Fisheries, 24(5):730–743, September 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Lancker:2023:UBT

[1259] Kira Lancker, Rudi Voss, Fabian Zimmermann, and Martin F. Quaas. Using the best of two worlds: a bio-economic stock assessment (BESA) method using catch and price data. Fish and Fisheries, 24(5):744–758, September 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Lennox:2023:PFP

[1260] Robert J. Lennox, Henrik H. Berntsen, Åse Helen Garseth, Scott G. Hinch, Kjetil Hindar, Ola Ugedal, Kjell R. Utne, Knut Wiik Vollset, Frederick G. Whoriskey, and Eva B. Thorstad. Prospects for the future of pink salmon in three oceans: From the native Pacific to the novel Arctic and Atlantic. Fish and Fisheries, 24(5):759–776, September 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hasan:2023:CSP

[1261] Md Robiul Hasan, Jennifer A. Chaplin, Peter B. Spencer, and Matias Braccini. Consumption of shark products: The interface of sustainability, trade (mis)labelling, human health and human rights. Fish and Fisheries, 24(5):777–795, September 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Goodrich:2023:WDS

[1262] Harriet R. Goodrich and Timothy D. Clark. Why do some fish grow faster than others? Fish and Fisheries, 24(5):796–811, September 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Peatman:2023:ETM

[1263] Tom Peatman, Valerie Allain, Lui Bell, Berry Muller, Aurélien Panizza, Naiten B. Phillip, Graham Pilling, and Simon Nicol. Estimating trends and magnitudes of bycatch in the tuna fisheries of the Western and Central Pacific Ocean. Fish and Fisheries, 24(5):812–828, September 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kuiper:2023:ARS

[1264] Spencer Dakin Kuiper, Nicholas C. Coops, Scott G. Hinch, and Joanne C. White. Advances in remote sensing of freshwater fish habitat: a systematic review to identify current approaches, strengths and challenges. Fish and Fisheries, 24(5):829–847, September 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Townhill:2023:CCP

[1265] Bryony L. Townhill, Elena Couce, Jonathan Tinker, Susan Kay, and John K. Pinnegar. Climate change projections of commercial fish distribution and suitable habitat around north western Europe. Fish and Fisheries, 24(5):848–862, September 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Robert:2023:LFL

[1266] Dominique Robert, Jun Shoji, Pascal Sirois, Akinori Takasuka, Ignacio A. Catalán, Arild Folkvord, Stuart A. Ludsin, Myron A. Peck, Su Sponaugle, Patricia M. Ayón, Richard D. Brodeur, Emily Y. Campbell, Evan K. D’Alessandro, John F. Dower, Louis Fortier, Alberto G. García, Klaus B. Huebert, Marc Hufnagl, Shin ichi Ito, Mikimasa Joh, Francis Juanes, Mitsuo Nyuji, Yoshioki Oozeki, Guido Plaza, Motomitsu Takahashi, Yosuke Tanaka, Naoki Tojo, Shingo Watari, Naotaka Yasue, and Pierre Pepin. Life in the fast lane: Revisiting the fast growth — high survival paradigm during the early life stages of fishes. Fish and Fisheries, 24(5):863–888, September 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Wilson:2023:DFD

[1267] James A. Wilson and Jarl Giske. Does fishing dismantle fish culture and ecosystem structure? Questions about the implications of social learning among fish and fishers. Fish and Fisheries, 24(5):889–895, September 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Barclay:2023:SHC

[1268] Kate M. Barclay, Simon R. Bush, Jan Jaap Poos, Andries Richter, Paul A. M. van Zwieten, Katell G. Hamon, Eira Carballo-Cárdenas, Annet P. Pauwelussen, Rolf A. Groeneveld, Hilde M. Toonen, Amanda Schadeberg, Marloes Kraan, Megan Bailey, and Judith van Leeuwen. Social harvest control rules for sustainable fisheries. Fish and Fisheries, 24(5): 896–905, September 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2023:IIf

[1269] Anonymous. Issue information. Fish and Fisheries, 24(6):i–ii, November 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

McKenzie:2023:GFP

[1270] David J. McKenzie. Gasping fish and panting squids oxygen, temperature and the growth of aquatic animals. Fish and Fisheries, 24(6):907–908, November 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Gill:2023:FGC

[1271] Andrew B. Gill. A field guide to coastal fishes of Bermuda, Bahamas and the Caribbean Sea. Fish and Fisheries, 24(6):909, November 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Sabal:2023:WOE

[1272] Megan C. Sabal, Kate Richerson, Paul Moran, Taal Levi, Vanessa J. Tuttle, and Michael Banks. Warm oceans exacerbate Chinook salmon bycatch in the Pacific hake fishery driven by thermal and diel depth-use behaviours. Fish and Fisheries, 24(6):910–923, November 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Liu:2023:CDE

[1273] Yiwen Liu, Chongliang Zhang, Xiujin Wei, Binduo Xu, Ying Xue, Yiping Ren, and Yong Chen. Coupling dynamic energy budget and population dynamic models to inform stock enhancement in fisheries management. Fish and Fisheries, 24(6):924–939, November 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Frid:2023:RIP

[1274] Alejandro Frid, Kyle L. Wilson, Jennifer Walkus, Robyn E. Forrest, and Mike Reid. Re-imagining the precautionary approach to make collaborative fisheries management inclusive of Indigenous Knowledge Systems. Fish and Fisheries, 24(6):940–958, November 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Laurel:2023:PCA

[1275] Benjamin J. Laurel, Alisa Abookire, Steve J. Barbeaux, L. Zoe Almeida, Louise A. Copeman, Janet Duffy-Anderson, Thomas P. Hurst, Michael A. Litzow, Trond Kristiansen, Jessica A. Miller, Wayne Palsson, Sean Rooney, Hillary L. Thalmann, and Lauren A. Rogers. Pacific cod in the Anthropocene: an early life history perspective under changing thermal habitats. Fish and Fisheries, 24(6):959–978, November 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Pons:2023:BCS

[1276] Maite Pons, David Kaplan, Gala Moreno, Lauriane Escalle, Francisco Abascal, Martin Hall, Victor Restrepo, and Ray Hilborn. Benefits, concerns, and solutions of fishing for tunas with drifting fish aggregation devices. Fish and Fisheries, 24(6):979–1002, November 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic). See correction [1284].

Fowler:2023:TSH

[1277] Ashley M. Fowler, Natalie A. Dowling, Jeremy M. Lyle, Josep Alós, Leif E. Anderson, Steven J. Cooke, Andy J. Danylchuk, Keno Ferter, Heath Folpp, Clifford Hutt, Kieran Hyder, Daniel K. Lew, Michael B. Lowry, Tim P. Lynch, Nicholas Meadows, Estanis Mugerza, Kjell Nedreaas, Domingos Garrone-Neto, Faith A. Ochwada-Doyle, Warren Potts, David Records, Scott Steinback, Harry V. Strehlow, Sean R. Tracey, Michael D. Travis, Jun ichi Tsuboi, Jon Helge Vølstad, and Rowan C. Chick. Toward sustainable harvest strategies for marine fisheries that include recreational fishing. Fish and Fisheries, 24(6):1003– 1019, November 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cotter:2023:OFI

[1278] Emma Cotter and Garrett Staines. Observing fish interactions with marine energy turbines using acoustic cameras. Fish and Fisheries, 24(6): 1020–1033, November 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Froehlich:2023:BLH

[1279] Halley E. Froehlich, Jamie C. Montgomery, David R. Williams, Casey O’Hara, Caitlin D. Kuempel, and Benjamin S. Halpern. Biological lifehistory and farming scenarios of marine aquaculture to help reduce wild marine fishing pressure. Fish and Fisheries, 24(6):1034–1047, November 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Tringali:2023:RRS

[1280] Michael D. Tringali and Susan K. Lowerre-Barbieri. Reproductive resilience or sweepstakes recruitment? Assessing drivers of lifetime reproductive success in exploited marine fish. Fish and Fisheries, 24(6): 1048–1066, November 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Andrews:2023:EST

[1281] Adam J. Andrews, Christophe Pampoulie, Antonio Di Natale, Piero Addis, Darío Bernal-Casasola, Veronica Aniceti, Gabriele Carenti, Verónica Gómez-Fernández, Valerie Chosson, Alice Ughi, Matt Von Tersch, Maria Fontanals-Coll, Elisabetta Cilli, Vedat Onar, Fausto Tinti, and Michelle Alexander. Exploitation shifted trophic ecology and habitat preferences of Mediterranean and Black Sea bluefin tuna over centuries. Fish and Fisheries, 24(6):1067–1083, November 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Vazquez:2023:EIS

[1282] Diego M. Vazquez and Luis O. Lucifora. Estimating intrinsic susceptibility to extinction when little ecological information is available: the case of neotropical freshwater stingrays (Chondrichthyes: Potamotrygoninae). Fish and Fisheries, 24(6):1084–1102, November 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Mao:2023:CFG

[1283] Jiangning Mao, Jiaqi Lu, Jiong Chen, and Jinbo Xiong. Consistent features of the gut microbiota in response to diverse shrimp Litopenaeus vannamei diseases: a meta-analysis. Fish and Fisheries, 24(6):1103– 1117, November 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2023:CPM

[1284] Anonymous. Correction to ‘Pons, M., Kaplan, D., Moreno, G., Escalle, L., Abascal, F., Hall, M., Restrepo, V., & Hilborn, R. (2023). Benefits, concerns, and solutions of fishing for tunas with drifting fish aggregation devices. Fish and Fisheries. Fish and Fisheries, 24(6):1118–1119, November 2023. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic). See [1276].

Anonymous:2024:IIa

[1285] Anonymous. Issue information. Fish and Fisheries, 25(1):i–ii, January 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Gilman:2024:GGG

[1286] Eric Gilman, Milani Chaloupka, Nathan Taylor, Lauren Nelson, Kim Friedman, and Hilario Murua. Global governance guard rails for sharks: Progress towards implementing the United Nations international plan of action. Fish and Fisheries, 25(1):1–17, January 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Griffiths:2024:IOF

[1287] Christopher A. Griffiths, Henning Winker, Valerio Bartolino, Håkan Wennhage, Alessandro Orio, and Massimiliano Cardinale. Including older fish in fisheries management: a new age-based indicator and reference point for exploited fish stocks. Fish and Fisheries, 25(1):18–37, January 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Eurich:2024:DPC

[1288] Jacob G. Eurich, Whitney R. Friedman, Kristin M. Kleisner, Lily Z. Zhao, Christopher M. Free, Meghan Fletcher, Julia G. Mason, Kanae Tokunaga, Alba Aguion, Andrea Dell’Apa, Mark Dickey-Collas, Rod Fujita, Christopher D. Golden, Anne B. Hollowed, Gakushi Ishimura, Kendra A. Karr, Stephen Kasperski, Yuga Kisara, Jacqueline D. Lau, Sangeeta Mangubhai, Layla Osman, Gretta T. Pecl, Jörn O. Schmidt, Edward H. Allison, Patrick J. Sullivan, Joshua E. Cinner, Roger B. Griffis, Timothy R. McClanahan, Richard C. Stedman, and Katherine E. Mills. Diverse pathways for climate resilience in marine fishery systems. Fish and Fisheries, 25(1):38–59, January 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Frawley:2024:DHO

[1289] Timothy H. Frawley, Barbara Muhling, Stephanie Brodie, Hannah Blondin, Heather Welch, Martin C. Arostegui, Steven J. Bograd, Camrin D. Braun, Megan A. Cimino, Nima Farchadi, Elliott L. Hazen, Desiree Tommasi, and Michael Jacox. Dynamic human, oceanographic, and ecological factors mediate transboundary fishery overlap across the Pacific high seas. Fish and Fisheries, 25(1):60–81, January 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Li:2024:RIT

[1290] Lingbo Li, Anne B. Hollowed, Edward D. Cokelet, Aimee A. Keller, Steve J. Barbeaux, Michelle M. McClure, and Wayne A. Palsson. The relative influence of temperature and ontogeny on groundfish distribution varies across life stages. Fish and Fisheries, 25(1):82–96, January 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Pecoraro:2024:HOI

[1291] Carlo Pecoraro, Chiara Papetti, and Carmelo Fruciano. Hidden in the ocean: The importance of detecting hybridisation in pelagic and deepwater fishes. Fish and Fisheries, 25(1):97–113, January 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Herrera-R:2024:SDF

[1292] Guido A. Herrera-R, Sebastian A. Heilpern, Thiago B. A. Couto, Lulu Victoria-Lacy, Fabrice Duponchelle, Sandra B. Correa, Aldo Farah-Pérez, Silvia López-Casas, Carlos M. Cañas-Alva, Carolina R. C. Doria, and Elizabeth P. Anderson. A synthesis of the diversity of freshwater fish migrations in the Amazon basin. Fish and Fisheries, 25(1):114–133, January 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Conneely:2024:MAA

[1293] Ellie-Ann Conneely and Christopher J. Coates. Meta-analytic assessment of physiological markers for decapod crustacean welfare. Fish and Fisheries, 25(1):134–150, January 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Guillotreau:2024:FFE

[1294] P. Guillotreau, F. Salladarré, M. Capello, A. Dupaix, L. Floc’h, A. Tidd, M. Tolotti, and L. Dagorn. Is FAD fishing an economic trap? Effects of seasonal closures and other management measures on a purse-seine tuna fleet. Fish and Fisheries, 25(1):151–167, January 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Thorson:2024:TFN

[1295] James T. Thorson. Trees for fishes: The neglected role for phylogenetic comparative methods in fisheries science. Fish and Fisheries, 25(1):168– 179, January 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Crimmen:2024:EDF

[1296] Oliver Crimmen. [Book review:] Eponym Dictionary of Fishes by Bo Beolens, Michael Grayson and Michael Watkins 2023. Whittles Publishing. xiv, 1463 pp. Fish and Fisheries, 25(1):180–181, January 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Campana:2024:BRL

[1297] Steven E. Campana. Book review of The lives of sharks: a natural history of shark life. Fish and Fisheries, 25(1):182, January 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2024:CAI

[1298] Anonymous. Correction to article “Impact of the 2014–2016 marine heatwave on US and Canada WestCoast fisheries: Surprises and lessons from key case studies”. Fish and Fisheries, 25(1):183, January 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic). See [1249].

Anonymous:2024:IIb

[1299] Anonymous. Issue information. Fish and Fisheries, 25(2):i–ii, March 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Coulthard:2024:APF

[1300] Sarah Coulthard. Asia–Pacific fishing livelihoods. Fish and Fisheries, 25(2):185–186, March 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Panzeri:2024:IPA

[1301] Diego Panzeri, Tommaso Russo, Enrico Arneri, Roberto Carlucci, Gianpiero Cossarini, Igor Isajlović, Svjetlana Krstulović Šifner, Chiara Manfredi, Francesco Masnadi, Marco Reale, Giuseppe Scarcella, Cosimo Solidoro, Maria Teresa Spedicato, Nedo Vrgoč, Walter Zupa, and Simone Libralato. Identifying priority areas for spatial management of mixed fisheries using ensemble of multi-species distribution models. Fish and Fisheries, 25(2):187–204, March 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Singer:2024:SSA

[1302] Merrill Singer. Syndemics of the sea: Adverse disease interactions and the stressors of fisher livelihoods. Fish and Fisheries, 25(2):205–217, March 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ovando:2024:SBC

[1303] Daniel Ovando, Darcy Bradley, Echelle Burns, Lennon Thomas, and James Thorson. Simulating benefits, costs and trade-offs of spatial management in marine social-ecological systems. Fish and Fisheries, 25(2): 218–234, March 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Rasal:2024:DSS

[1304] Jennifer Rasal, Michael C. Melnychuk, Amanda Lejbowicz, Carlos Montero-Castaño, Sophie Ferber, and Catherine Longo. Drivers of success, speed and performance in fisheries moving towards Marine Stewardship Council certification. Fish and Fisheries, 25(2):235–250, March 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Karp:2024:QUM

[1305] William A. Karp, Michael C. Melnychuk, Robyn E. Forrest, Lorne Richard Little, Kristin McQuaw, Chad Demarest, Ray Hilborn, Nicole Baker, Brian Mose, Bruce Turris, and Ernesto Penas Lado. Quota use in mixedstock fisheries. Fish and Fisheries, 25(2):251–267, March 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Hardy:2024:TBA

[1306] Natasha A. Hardy, Cindy Matuch, Zachary Roote, Iris George, Barbara A. Muhling, Michael G. Jacox, Elliott L. Hazen, Steven J. Bograd, Larry B. Crowder, and Stephanie J. Green. Trait-based analyses reveal global patterns in diverse diets of albacore tuna (Thunnus alalunga). Fish and Fisheries, 25(2):268–282, March 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Mayer:2024:TTP

[1307] Noa B. Mayer, Scott G. Hinch, and Erika J. Eliason. Thermal tolerance in Pacific salmon: a systematic review of species, populations, life stages and methodologies. Fish and Fisheries, 25(2):283–302, March 2024. CO-DEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Chong:2024:GSE

[1308] Lisa Chong, Zachary A. Siders, Kai Lorenzen, Robert N. M. Ahrens, and Edward V. Camp. Global synthesis of effects and feedbacks from artificial reefs on socioecological systems in recreational fisheries. Fish and Fisheries, 25(2):303–319, March 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ma:2024:RRS

[1309] Shuyang Ma, Geir Huse, Kotaro Ono, Richard D. M. Nash, Anne Britt Sandø, Kjell Nedreaas, Solfrid Sætre Hjøllo, Svein Sundby, Tom Clegg, Jon Helge Vølstad, and Olav Sigurd Kjesbu. Recruitment regime shifts and nonstationarity are widespread phenomena in harvestable stocks experiencing pronounced climate fluctuations. Fish and Fisheries, 25(2): 320–348, March 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Gravel:2024:MPG

[1310] Sarah Gravel, Jennifer S. Bigman, Sebastián A. Pardo, Serena Wong, and Nicholas K. Dulvy. Metabolism, population growth, and the fastslow life history continuum of marine fishes. Fish and Fisheries, 25(2): 349–361, March 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Maia:2024:PAM

[1311] Caroline Marques Maia, João Luis Saraiva, and Eliane Gonçalves de Freitas. Preference, avoidance, motivation and their importance to fish welfare. Fish and Fisheries, 25(2):362–379, March 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2024:IIc

[1312] Anonymous. Issue information. Fish and Fisheries, 25(3):i–ii, May 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Dupaix:2024:CAE

[1313] Amaël Dupaix, Frédéric Ménard, John D. Filmalter, Yannick Baidai, Nathalie Bodin, Manuela Capello, Emmanuel Chassot, Hervé Demarcq, Jean-Louis Deneubourg, Alain Fonteneau, Fabien Forget, Francesca Forrestal, Daniel Gaertner, Martin Hall, Kim N. Holland, David Itano, David Michael Kaplan, Jon Lopez, Francis Marsac, Alexandra Maufroy, Gala Moreno, Jeff A. Muir, Hilario Murua, Liliana Roa-Pascuali, Géraldine Pérez, Victor Restrepo, Marianne Robert, Kurt M. Schaefer, Grégory Sempo, Marc Soria, and Laurent Dagorn. The challenge of assessing the effects of drifting fish aggregating devices on the behaviour and biology of tropical tuna. Fish and Fisheries, 25(3):381–400, May 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Rice:2024:FGR

[1314] Emma D. Rice, Edith Gondwe, Abigail E. Bennett, Patrick Asango Okanga, Nimah F. Osho-Abdulgafar, Kafayat Fakoya, Ayodele Oloko, Sarah Harper, Patrick Chimseu Kawaye, Ernest O. Chuku, and Hillary Smith. The future of gender research in small-scale fisheries: Priorities and pathways for advancing gender equity. Fish and Fisheries, 25(3): 401–408, May 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Schiller:2024:LPD

[1315] Laurenne Schiller, Gregory L. Britten, Graeme Auld, and Boris Worm. Learning from positive deviants in fisheries. Fish and Fisheries, 25(3): 409–428, May 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Foldvik:2024:OTR

[1316] Anders Foldvik, Eva Marita Ulvan, and Tor Næsje. Optimal timing of return migration in Atlantic salmon. Fish and Fisheries, 25(3):429– 440, May 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

DelSantoONeill:2024:ETF

[1317] T. J. Del Santo O’Neill, A. G. Rossberg, and R. B. Thorpe. An efficient tool to find multispecies MSY for interacting fish stocks. Fish and Fisheries, 25(3):441–454, May 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Lin:2024:FWR

[1318] Zhen Lin and Shin ichi Ito. Fish weight reduction in response to intraand interspecies competition under climate change. Fish and Fisheries, 25(3):455–470, May 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Goethel:2024:LDY

[1319] Daniel R. Goethel, Aaron M. Berger, Simon D. Hoyle, Patrick D. Lynch, Caren Barceló, Jonathan Deroba, Nicholas D. Ducharme-Barth, Alistair Dunn, Dan Fu, Francisco Izquierdo, Craig Marsh, Haikun Xu, Giancarlo M. Correa, Brian J. Langseth, Mark N. Maunder, Jeremy McKenzie, Richard D. Methot, Matthew T. Vincent, Teresa A’mar, Massimiliano Cardinale, Marta Cousido-Rocha, Nick Davies, John Hampton, Carolina Minte-Vera, and Agurtzane Urtizberea. ‘Drivin’ with your eyes closed’: Results from an international, blinded simulation experiment to evaluate spatial stock assessments. Fish and Fisheries, 25(3):471–490, May 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Robertson:2024:TMI

[1320] Matthew D. Robertson, Noel G. Cadigan, Paul M. Regular, Mariano Koen-Alonso, Frédéric Cyr, Fan Zhang, and Tyler D. Eddy. Testing models of increasing complexity to develop ecosystem-informed fisheries advice. Fish and Fisheries, 25(3):491–507, May 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Champagnat:2024:RMF

[1321] J. Champagnat, E. Rivot, and O. Le Pape. The response of marine fish population productivity to juvenile habitat modifications depends upon life histories. Fish and Fisheries, 25(3):508–522, May 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bradley:2024:FML

[1322] Michael Bradley, Alexia Dubuc, Camilla V. H. Piggott, Katie Sambrook, Andrew S. Hoey, Martial Depczynski, Tim J. Langlois, Monica Gagliano, Shaun K. Wilson, Katherine Cure, Thomas H. Holmes, Glenn I. Moore, Michael Travers, Ronald Baker, Ivan Nagelkerken, and Marcus Sheaves. The fish–mangrove link is context dependent: Tidal regime and reef proximity determine the ecological role of tropical mangroves. Fish and Fisheries, 25(3):523–541, May 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Haas:2024:PPR

[1323] Bianca Haas, Kamal Azmi, Hussain Sinan, and Quentin Hanich. Policy pathways to reduce disproportionate burdens in tuna fisheries. Fish and Fisheries, 25(3):542–548, May 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cosseau:2024:BRE

[1324] Céline Cosseau and Jeremie Vidal-Dupiol. Book review: Epigenetics in aquaculture by F. Piferrer, H. P. Wang (Eds). New Jersey: John Wiley & Sons. 2023. pp. 512. ISBN: 978-1-119-82191-5. Fish and Fisheries, 25(3):549–550, May 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2024:IId

[1325] Anonymous. Issue information. Fish and Fisheries, 25(4):i–ii, July 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Langan:2024:OBB

[1326] Joseph A. Langan, Curry J. Cunningham, Jordan T. Watson, and Skip McKinnell. Opening the black box: New insights into the role of temperature in the marine distributions of Pacific salmon. Fish and Fisheries, 25(4):551–568, July 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bano:2024:SMP

[1327] Bálint Bánó, Aleksey Bolotovskiy, Boris Levin, George M. T. Mattox, Mauricio Cetra, István Czeglédi, and Péter Takács. Scale morphology is a promising, additional tool for exploring the taxonomy and ecology of freshwater fishes. Fish and Fisheries, 25(4):569–588, July 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Somerville:2024:UVM

[1328] Jasmine Somerville, Jon Blount, and Martin Stevens. The use of vision modelling to design bycatch reduction devices using light. Fish and Fisheries, 25(4):589–601, July 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Farchadi:2024:MHR

[1329] Nima Farchadi, Heather Welch, Camrin D. Braun, Andrew J. Allyn, Steven J. Bograd, Stephanie Brodie, Elliott L. Hazen, Alex Kerney, Nerea Lezama-Ochoa, Katherine E. Mills, Dylan Pugh, Riley Young-Morse, and Rebecca L. Lewison. Marine heatwaves redistribute pelagic fishing fleets. Fish and Fisheries, 25(4):602–618, July 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Bavinck:2024:MSE

[1330] Maarten Bavinck, Joeri Scholtens, and Michael Fabinyi. Maximum sustainable employment: Adding to the beacons of wild fisheries governance. Fish and Fisheries, 25(4):619–629, July 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Vigo:2024:DMS

[1331] Maria Vigo, Virgilio Hermoso, Joan Navarro, Joan Sala-Coromina, Joan B. Company, and Sylvaine Giakoumi. Dynamic marine spatial planning for conservation and fisheries benefits. Fish and Fisheries, 25(4): 630–646, July 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Jones:2024:NDI

[1332] Benjamin L. H. Jones, Rolando O. Santos, W. Ryan James, Sophia V. Costa, Aaron J. Adams, Ross E. Boucek, Lucy Coals, Leanne C. Cullen-Unsworth, Samuel Shephard, and Jennifer S. Rehage. New directions for indigenous and local knowledge research and application in fisheries science: Lessons from a systematic review. Fish and Fisheries, 25(4): 647–671, July 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Cano-Barbacil:2024:MDR

[1333] Carlos Cano-Barbacil, Julian D. Olden, and Emili García-Berthou. Myers’ divisions revisited: Contemporary evidence for distinct trait differences among global inland fishes. Fish and Fisheries, 25(4):672–685, July 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Ma:2024:NAF

[1334] Shuyang Ma, Geir Huse, Kotaro Ono, Richard D. M. Nash, Jon Helge Vølstad, and Olav Sigurd Kjesbu. Northeast Atlantic fish stock productivity hindcasts and forecasts from a Bayesian framework reveal pronounced climate-induced dynamics. Fish and Fisheries, 25(4):686–710, July 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Sando:2024:MSA

[1335] Anne Britt Sandø, Solfrid S. Hjøllo, Cecilie Hansen, Morten D. Skogen, Robinson Hordoir, and Svein Sundby. A multi-scenario analysis of climate impacts on plankton and fish stocks in northern seas. Fish and Fisheries, 25(4):711–732, July 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Howard:2024:PMF

[1336] Rebecca A. Howard, Lauren A. Rogers, Kelly A. Kearney, Laura L. Vary, and Lorenzo Ciannelli. Projecting marine fish distributions during early life stages under future climate scenarios. Fish and Fisheries, 25(4):733– 749, July 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Harrison:2024:CRH

[1337] Hannah L. Harrison and Valerie Berseth. A commentary on the role of hatcheries and stocking programs in salmon conservation and adapting ourselves to less-than-wild futures. Fish and Fisheries, 25(4):750–760, July 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Anonymous:2024:IIe

[1338] Anonymous. Issue information. Fish and Fisheries, 25(5):i–ii, September 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Arlinghaus:2024:CUR

[1339] Robert Arlinghaus. Catch uncertainty and recreational fishing attraction: Propositions and future research directions. Fish and Fisheries, 25(5): 761–780, September 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Mangubhai:2024:VCA

[1340] Sangeeta Mangubhai, Margaret Fox, Yashika Nand, and Natalie Mason. Value chain analysis of a women-dominated wild-caught mud crab fishery. Fish and Fisheries, 25(5):781–792, September 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Flink:2024:RFS

[1341] Henrik Flink, Göran Sundblad, Juha Merilä, and Petter Tibblin. Recreational fisheries selectively capture and harvest large predators. Fish and Fisheries, 25(5):793–805, September 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Drymon:2024:DOC

[1342] James Marcus Drymon, Amanda E. Jargowsky, Evan G. Prasky, Edward V. Camp, Ashley Oliphant, Sean P. Powers, and Steven B. Scyphers. Depredation: An old conflict with the sea. Fish and Fisheries, 25(5):806–810, September 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Dierickx:2024:FRP

[1343] Katrien Dierickx, Peter Schauer, Jennifer Harland, Alan Pipe, Tarek Oueslati, Alexander Lehouck, Anton Ervynck, Wim Wouters, Matthew Von Tersch, David Orton, and Michelle Alexander. Finding the right plaice at the right time: Multi-molecular analysis of flatfish reveals historical catch habitats. Fish and Fisheries, 25(5):811–822, September 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Roa-Ureta:2024:POT

[1344] Ruben H. Roa-Ureta, Patrícia Amorim, and Susana Segurado. Probability ogives for trends in stock biomass and fishing mortality from landings time series. Fish and Fisheries, 25(5):823–836, September 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Campbell:2024:CDC

[1345] Brooke Campbell, Dirk Steenbergen, Owen Li, Abel Sami, Beia Nikiari, Aurélie Delisle, Pita Neihapi, Tarateiti Uriam, and Neil Andrew. Characterising a diversity of coastal community fisheries in Kiribati and Vanuatu. Fish and Fisheries, 25(5):837–857, September 2024. CODEN FFI-IAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Kangosjarvi:2024:EDF

[1346] Henna Kangosjärvi, Per-Arne Amundsen, Pär Byström, Anders G. Finstad, Michael Power, Javier Sánchez-Hernández, and Antti P. Eloranta. Environmental drivers of food webs in charr and trout-dominated coldwater lakes. Fish and Fisheries, 25(5):858–875, September 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).

Moore:2024:CSL

[1347] Alec B. M. Moore, Keith Brander, Shaun Evans, Poul Holm, and Jan Geert Hiddink. Century-scale loss and change in the fishes and fisheries of a temperate marine ecosystem revealed by qualitative historical sources. Fish and Fisheries, 25(5):876–894, September 2024. CODEN FFIIAK. ISSN 1467-2960 (print), 1467-2979 (electronic).